A Selected Bibliography of Publications by, and about, Niels Bohr

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

27 March 2024 Version 1.326

Title word cross-reference

+ [VIR^{+}08]. $1 [Duf46]. $1.00 [N.38, Bal39]. $105.95 [Dor79]. $11.95 [Bus20]. $12.00 [Kra07, Lan08]. $189 [Tan09]. $21.95 [Hub14]. $24.95 [RS07]. $29.95 [Gor17]. $32.00 [RS07]. $35.00 [Par06]. $45 [Cuf24]. $47.50 [Kri91]. $6.95 [Sha67]. $61 [Kra16b]. $9 [Jam67]. ^{−} [VIR^{+}08]. ^{238} [Tur46a, Tur46b]. ^{∘} [Fra55]. _{2} [Som18]. β [Gau14]. c [Dar92d, Gam39]. G [Gam39]. h [Gam39]. q [Dar92d]. × [wB90].

-numbers [Dar92d].

/Hasse [KZN^{+}88]. /Rath [GRE^{+}01].

0 [wB90, Hub14, Tur06]. 0-19-852049-2 [Ano93a, Red93, Seg93]. 0-19-853977-0 [Hub14]. 0-521-35366-1 [Kri91]. 0-674-01519-3 [Tur06]. 0-85224-458-4 [Hen86a]. 0-9672617-2-4 [Kra07, Lan08].

1.5 [GRE^{+}01]. 100-året [BR^{+}85]. 100th [BR^{+}85, KRW05, Sch13, vM02]. 110th [Rub97a]. 121 [Boh87a]. 153 [MP97]. 16 [SE13]. 17 [Boh55a, KRBR62]. 175 [Bad83]. 18.11.1962 [Hei63a]. 1911 [Meh75]. 1915 [SE13]. 1915/16 [SE13, SE13]. 1918 [Boh21a]. 1920s [PP16]. 1922 [Boh22a]. 1923 [Ros18]. 1925 [Cla13, Bor13, Jan17, Sho13]. 1927 [Ano28]. 1929 [HEB^{+}80, HvMW79, Pye81]. 1930 [Lin81, Whe81]. 1930/41 [Fer68, Fer71]. 1930s [Aas85b, Stu79]. 1933 [CCJ^{+}34]. 1933b [CS79g]. 1935 [Hei35]. 1937 [ABH^{+}37b, ABH^{+}37d, ABH^{+}38a]. 1939 [ABH^{+}39a, HP85, OSWR86, vMHW85, Rec86]. 1940 [Dah43]. 1941 [Aas02f]. 1943 [SO93]. 1944 [Sch21]. 1945 [CS79d, Dah14, DF14, Ros45, Ros61, Ros79b]. 1945e [CS79d]. 1946 [CSW97b]. 1949 [PvM93, Wal08, vM93, Bro95, Kra96]. 1950 [Boh50a, Boh50d]. 1950a [CS79c]. 1952 [vM96]. 1954 [vM99]. 1954/Mladjenovic [ARK^{+}99]. 1955 [Boh55a]. 1955b [CS79f]. 1956 [vM01]. 1957 [Ano57b]. 1958 [Boh61c, KMP59, vM05]. 1958-1962 [Boh66b]. 1958/1962 [Lin64, Nie65]. 1960s [ARK^{+}99, Mla98]. 1960s/Frenkel [ARK^{+}99]. 1961 [Boh85e]. 1962 [Boh66a, Coc63, KRBR62, Mal70]. 1963c [CS79b]. 1963h [CS79e]. 1967 [RB67]. 1968 [Wei68]. 1983 [K^{+}84]. 1984 [Wes85]. 1985 [CW85, DDU86, LM85]. 1996 [Nie97]. 1998 [Che99a, How99]. 1999 [Rot00]. 1er [LRdB^{+}23]. 1S [Har87]. 1st [LRdB^{+}23].

20 [Hen86a]. 2000 [Log00]. 2010 [KLR13]. 208 [Nie55]. 20th [APB96, ABP98, Bre97, Gin01]. 22 [CCJ^{+}34]. 23.50 [Che85a]. 24 [Ano28]. 298.00 [Hen86b]. 2nd [Dar91a, Ros79b]. 2P [Har87].

3 [wB90, Gor17, Tur06]. 3-540-13609-6 [Hen86b]. 3-540-19334-0 [wB90]. 3-540-90642-8 [Hen86a]. 3-540-90667-3 [Hen86a]. 3-540-90674-6 [Hen86a]. 3-540-90675-4 [Hen86a]. 3-540-90680-0 [Hen86a]. 35 [Hel98b]. 3D [Har87].

4 [Stu84]. 4. [How99]. 4/Pauli [OSWR86]. 49/Atmanspacher/Primas [MGG^{+}95]. 4F [Har87].

5. [Che99a]. 565 [Hen92].

60th [Ber15, Pau45, Pau94]. 63 [Che93]. 6th [LRdB^{+}23].

7 [Cla13, Hei35]. 70th [B^{+}55, BM55, CS79f, Fra55, Gus55, HH55, Her55, KP55, Las55, LH55, Møl55, Nie55, PRW55, Pih55, RM55, Ros79d]. 75th [Gam60].

8/Richardson [KZN^{+}88]. 80th [CS79b, Ros79a]. 81 [Nie55]. 8th [Rot00].

9 [Boh50a, Boh50d]. 9.40 [Sch64c]. 90-277-1648-X [Che85a]. 978 [Cla13, Gor17]. 978-0-19-023299-3 [Gor17]. 978-0-19-881926-4 [Bus20]. 978-0-19-965498-7 [Bor13, Cla13]. 978-1-4020-5253-8 [Tan09]. 9O [Bro86]. 9O-277-1648-X [Bro86].

= [BD92, MR02].

A. [Boh62b, Som20c, PHvMW93]. A./Kraus [GRE^{+}01]. a.o [BDR86, Hen86b, HvMW79, vMHW85, vM93, vM96, vM99, vM01, vM05]. Åbent [Boh50a, Boh64d]. aabnedes [Ano29]. Aage [Boh63a, Cha09, KNA01]. Aar [Dah43, Boh64b, Boh64a, Kra13a]. årene [Boh64e, Cas64]. året [BR^{+}85]. års [Cou64]. Aaserud [Bor16, Far13a, Kra16b, Kri91, Ren15, Kub09, Was92, For91]. Abb [Sch64c]. Abhandlungen [Boh21a, Bor63]. Abich [Kuh67]. Abides [Gla00]. ablaze [Hei15]. Abraham [Ano93a, Ber91, Dar92a, Kle93, Red93, Stu92, Hen92, Seg93]. Absence [Red63, RR00]. absolute [BL18b, BL63b]. absorption [Kon00, LR23]. abstract [Pai48]. Abteilung [BDR86]. Academic [SDG15]. Academy [Ano57b, FMO88, Mai93, Ped64, BR^{+}85]. Acceleration [HEB^{+}80]. Acceleration/Brosche [HEB^{+}80]. accept [Bar89]. According [KH25b, Pat85, Som15]. Account [Hof47, Hof59, Hof14, Rad82]. accounts [Sha87]. act [Stu13]. acting [Dje07, POC^{+}10]. Action [Nia98, Tib13, Boh29a, Boh38e, Boh39f, Ish15, Ish17, Nak15, PB17]. action-angle [Nak15]. Actions [Boh77]. active [HS39a]. actors [Kai94a]. actually [Bel86]. ad [Seg86a]. Adam [DT18]. adiabatic [PV15, Ehr23]. Adiabatische [Ehr23]. Adolf [CL83]. advanced [Jon14]. Advantages [Mar54]. adventure [MB11b]. af [Boh37e, Boh39c, Boh55a, BR64, KvG85, Kle64, Koc64, RRK^{+}64, Sch68]. After [Den99, Hee95, RY13, Wyd84, Aas20, And89, Jon85b, Kae48, LSW12, Pai64, Plo94]. Aftermath [Wal08]. again [Boh63e]. against [Mas04]. Age [Opp50, Pur18, Stu18, Dje07]. ago [Bar99, Bec57]. ai [Seg83]. aims [Sch15]. AIP [Ano68a]. Akademisk [Bad83]. aktiver [HS39a]. akzeptieren [Bar89]. al [Car97]. Alain [Hub14]. Alamos [Bet91]. alarm [Yor75]. Albert [Adl03, Kon15, BR55, Kle10, Kon15, Meh87, Ros18, Sch49a, Sch49b, VW86]. Alexandra [Fre90]. Alfred [Adl03, Sha67]. Algebraic [OK12]. algebras [HLS12]. aliens [A^{+}09]. Alkali [Lin24]. alkaline [HS39b]. allemandes [Kar07]. allen [Sch64c]. allgemeine [Som15]. allgemeinen [Pau33]. Allgemeinverständliche [BDR86]. almenfatteligt [HK22, HKK29]. almost [Büh98a]. Alpha [Nia98, Fra55]. alpha- [Fra55]. alsidighed [Dir64]. also [Pai48]. always [Gra15]. am [Hei35]. Amateur [Har01]. Ambition [Jon08]. America [Stu85a]. American [FMO88, Hei96, Nie97, Val06]. americana [Val06]. Among [Ano22, CS79f, Ros79d]. Amount [Ano39a]. Amsterdam [Dor79, Jam67]. Amy [Dje07]. analogies [AI15]. Analogy [Dar92d, Dar88]. Analyses [Bou49a, Bou49b, Che93, Dar87, Dar90, Dar91a, Dar92a, Dar92b, MP97, dB86a]. Analysis [Bra61, Kat11d, PB17, vFH76, Bub89, Car04, CDH^{+}06, Urt06]. Ancient [Jon15, Ryn83, Sac81]. Andrew [Hen86a]. Andwendung [Boh31b]. angewandte [Som20b]. angle [Møl55, Nak15]. Angst [Ano02a]. annihilation [SO93]. Anniversary [Ano93b, Pau01, Sch13, Ano85e, BR^{+}85, Jac85, Rub97a]. Announced [Ano39d]. anomalies [Mas04]. anomaly [Gau14]. Anschauung [Hei10b]. answers [Som20c]. Anti [Che93, CHR96, Fay91, Fol96, Hen93, Plo94]. anti-epistemology [Plo94]. Anti-Realism [CHR96, Fol96]. Anti-Realist [Hen93, Che93, Fay91]. antiphonies [KvG85]. antiquity [Los11]. Anton [Fis10a, Fis12a, Fis10a, Fis12a]. Antwort [BL19]. Antworten [Som20c]. Anwendung [Boh23h, Boh23i]. Any [Lau33]. aos [Seg87]. Apfel [Büh98b]. appeal [BBCN86]. apple [Büh98b]. Application [Boh75b, Mil24, Smi76, Boh16, Boh23h, Boh23i, Boh24d, FF78b, RH00, Boh23a]. Applications [HEB^{+}80, Har23, Har26]. Applications/Lehmann [HEB^{+}80]. Applied [Som20b, Wer23]. Appreciation [Pei63, Har01]. apprentice [Jak13]. Approach [Tan04a]. approdo [A^{+}09]. appropriations [HF15]. approximate [Har23]. approximation [Kle78]. April [Wei68, LRdB^{+}23]. Arbeiten [ARK^{+}99, BL19, Hou30, Ber15]. Archimedes [Sch68]. architecture [Kra97]. Archival [Hei96, DS09, Nie97]. Archive [Aas00b, Aas02d]. Archives [Cra02]. archivio [DS09]. Area [SWW88, KZN^{+}88]. arena [Boh05]. Argument [HH94, Rös95, BO09, Dar85, Kat98]. arguments [Hon81, Pat85]. Argyris [MGG^{+}95]. Argyris/Faust [MGG^{+}95]. Arising [Bas71]. Aristole [vW02b]. Aristoteles [vW02b]. Arkady [Tan09]. Arkimedes [Sch68]. armaments [BBCN86]. arme [Kar07]. Armin [Rec86]. Arms [She85a, Far13c, She86]. Arnold [Eck13a, Eck13b, Set10]. Arthur [Adl03, Duf46, dR92, Coh88, Coh91, Coh92]. Articles [Boh64e, Boh32c]. Artikler [Boh64e]. Artistic [Dow09]. Arts [FMO88]. Ascent [Bro73a]. Ask [Jen03]. Aspect [Hub14, Lau33]. aspects [Des48, Meh75]. Assembly [Som20b]. Assessment [Whe57, Whi04]. assistance [All56]. Association [AG81, Boh38b]. assured [Byr10]. astounding [Haw11]. Astronomy [RS07, Dea13, FF91]. Astrophysics [CO07]. Atlantis [Che94b]. Atom [AH13, Ano21, Ano23c, Ano39b, Ano39a, Ano39d, Ano62c, ABH^{+}36a, ABH^{+}37d, ABH^{+}38a, Bec57, Beh41, Beh43b, Boh15b, Boh26a, Boh62d, Boh65a, Boh01, BL63b, Bor13, Bor16, Cla13, CT65, Dar92c, Duf46, Far13a, Gam63, Gar63a, Har28, HK69, Hei77, Hei85b, Hei13, Her72, HN70, HLA^{+}10, HK26, Hug90, Jan17, Kae45, Kra11c, Kra12a, KN13, Kra16b, KH23, KH25a, Lau51, Nia98, Nis67, Pei55, Pod10b, Ren15, Rod19, Roq17, Sch13, Sho13, Vic08, Wer23, vM02, AK15, Ano93a, AI15, Bag10, Boh15c, Boh23f, Boh45b, Boh85h, Boh85j, Bro73b, CDH^{+}06, Eck14, El’85, FM12, Gan17, Hei81, Hei85c, Hei85d, Hei15, Het95, HG15, KN11, Kra15c, KH25b, MB24, Nau13, Nav15, Nie15, Pet14, Pol60, Red93, Sie13, Ter08, VIR^{+}08, Wyk08, Jon14]. Atom-Institute [Ano21]. Atom-Kommission [ABH^{+}36a, ABH^{+}37d, ABH^{+}38a]. Atom-Model [Wer23]. Atombau [Boh21a, Boh23h, Boh23i, Boh22d, Boh23c, Boh24b]. atombomben [Jak13]. Atome [Boh22b, Rec09, Rec10, Boh23b, Ano23a, Boh23g, Boh24f, Boh25c]. Atomen [Boh25d, BH23, Som20a, Som20c, Som20b]. Atomernes [Boh23d, Aas02e, Boh21f]. Atomes [LRdB^{+}23, Pia24a, Pia24b]. Atomfizika [Boh84]. Atomforskning [Boh24c]. Atomfysik [Boh58f, Boh59a, Boh64e, BR^{+}85]. Atomi [Pet88, ABH^{+}36b, Dar90, Goo92]. Atomic [Ano22, Ano39a, Ano43, Ano95b, ARK^{+}99, Ayl26, Ayl27, Beh43a, Beh43c, Boh14a, Boh21a, Boh21b, Boh21c, Boh23b, Boh25a, Boh26b, Boh32a, Boh32b, Boh31d, Boh34, Boh36e, BK37, Boh37f, Boh39b, Boh47, Boh49, Boh58a, Boh58b, Boh60a, Boh61a, Boh62a, Boh63b, Boh65b, Boh75b, Boh75a, Boh78, Boh81, Boh85b, Boh87b, BB87, Boh10, Boh11b, Bor58, BBSR69, Bor13, Bør12, Bur18, Can00, Cla13, Com56, Dar97, Duf46, FR13b, Fri54, Gam29, Gla00, Hei64a, HN64, Hof47, Hof59, Hof14, Hoy73, HnB74, Jan17, Kel83a, Kow53, Kra79, Kra10b, Kra12a, Kra12b, Kra23a, Lan94, Lau51, Lin64, McW73, Mil24, MB11a, Nie65, Pal00, Pei97d, Pia23, Kam64, Pla23, Rad82, Rei84, San24, Sch64b, Sho13, SH59, Som20b, Som24a, Som24d, Som24c, Som24b, SM08, Tur46a, Tur46b, Yag72, Aas02e]. atomic [Ano97, Ano13, Bai13, Bar89, Bec57, BZ85, Ber89, Bet91, Boh20b, Boh22a, Boh22d, Boh22e, Boh22f, Boh22g, Boh23h, Boh23i, Boh23c, Boh23a, Boh24d, Boh24b, Boh24c, Boh24e, Boh28b, Boh28c, Boh28d, Boh28e, Boh29c, Boh35c, Boh36d, Boh37d, Boh37e, Boh38a, Boh38e, Boh39f, Boh39c, Boh40c, Boh41e, Boh42a, Boh45c, Boh46c, Boh48b, Boh55a, Boh60c, Boh63d, BH64, Boh64b, Boh64a, Boh66b, BR^{+}85, Boh85m, Boh87c, Boh87d, Boh87a, BL18b, BL18a, BL19, BL63a, CM85, CCJ^{+}34, Cos22, Dam85, Dar55, DNT08, Eck13a, Eck15, El’85, Föp15, Fri64, HK22, HKK29, Hou30, Ina16, Jak13, Jea19, Kel07, Kel09, Kos23, Kra97, Kra10a, Kra11a, Kra13a, Kra14a, LR23, MW46a, MW07, Mel00, Pla24a, Pla24b, Pol60, RR00, Ree15a, She72, Sie13, SE13, Som13b, Som13a, Tuc97, Van03, Wea88, Yag64, dBdB21, Ano21]. Atomic [Boh21f, Boh24a, Boh25b, Boh26c, Boh29b, Boh30a, Boh30b, Boh31a, Boh32c, Boh58c, Boh58f, Boh58d, Boh59a, Boh61b, Boh63c, Boh64e, Boh66a, Boh72, Boh85e, Boh88, Boh93, Bro36, Som31, Gla60, Lan59, McC34, Pol58, Rie60, Ste58]. atómica [Boh88]. atomique [Boh32c, Boh61b, Boh72, Boh93, Dar92b]. atomiques [Boh23a, Boh39f, CCJ^{+}34]. atomism [Boh28a]. Atomistik [Boh28a]. atomistique [Boh28b]. Atomkern [Boh38e]. atomkerne [Boh55a, Boh36d]. atomkerne-energien [Boh55a]. atomkernen [Fri64]. Atomkerner [Boh37e]. Atomkernernereaktioner [Boh37d]. Atomkernernes [Boh35c, Boh39c, Boh41e, Boh42a, Boh45c, Boh46c]. Atomkerners [Boh40c]. Atomkerns [Hou30]. Atommodel [Ano13]. Atommodell [Bar89, Pol60, Som13b, BH64, BL18a, BL19, BL63a, SE13]. Atommodelle [BL18b, BL63b]. Atommodelles [Som24c, Föp15, Som24a, Som24b]. Atommodells [SE13, Som13a, Pla24a, Pla24b, Som24d]. átomo [KH25b]. Atomowa [Kar06, Boh63c]. Atomphysik [Boh58c, Boh66a, Boh85e, Sch64c, Sch64b, Mal70]. Atomphysiker [Eck13a].

Atomphysikerin [KZN^{+}88]. Atomphysikers [Dar87]. Atoms [Ano23d, ABH^{+}37a, ABH^{+}37c, ABH^{+}38b, ABH^{+}39b, ABH^{+}40a, ABH^{+}40b, Boh13a, Boh13b, Boh13c, Boh23g, Boh24f, Boh58e, Boh62c, Boh63f, Boh77, Faf58, Hei70, Lin24, Pol60, Sal19, Som20a, Ano23a, Boh21d, Boh22b, Boh23d, Boh25c, Boh25d, BH23, Gao15, JVK22, Rec10, Ree05, Sim14, Som20c, Urt06, Ano57a, Ano57b, ABH^{+}36a, ABH^{+}36b, ABH^{+}37d, ABH^{+}38a, Bec57, Kra23a, LRdB^{+}23, Pet88, Mac95]. Atomteoretiske [Boh24a]. Atomteori [Boh58d, HK22, HKK29, Kra13a, Boh25b, Boh29b]. Atomteorien [Boh30a]. Atomteoriens [Boh29c]. Atomtheorie [Boh30b, Kos23, SE13, Boh26c, Boh28e, Boh31a, Hoy73, HnB74, LR23, Pla23, Som31, Hei75]. atomvåbnenes [Boh64b, Boh64a]. atomvåbnet [CM85]. Atoommodel [Bur18]. Atoomtheorie [Cos22]. attempts [Kar07, Nav06, Yor75]. Attributed [Lau33]. attributes [Mof06]. Audience [Ano23d]. Aufbruch [MGG^{+}95]. Aufl [Sch64c]. Aufsätze [Boh22d, Boh24b, Boh31a, Boh66a, Boh85e, Mal70]. Augenblick [Büh98a]. August [K^{+}84]. Ausarbeitung [Boh22a]. Ausbau [Som13a]. Ausgewählte [Bor63]. auspices [Ano28, CCJ^{+}34]. author [Dur00]. authority [Dar85]. Autobiographical [BDR86]. Autobiographisches [BDR86]. autoradiography [LH55]. Autour [Rhe91]. aux [Boh23a, Pat85, Seg84]. Aver [Cho02]. avril [LRdB^{+}23]. Award [Ano57b, Ano62c, Bec57].

B [Kuh67, vM05, Bet37]. B. [Hei48]. back [Kai94a, Mei64a, Mei64b]. Background [Che93, HW07b, Fav92, Mai93, Ree15b]. Baconian [Per13]. bag [Jak13]. Balance [Dow09]. Ball [Ano39a]. Balmer [RR00]. Balmer/Rydberg/Bohr [RR00]. Band [ARK^{+}99, BDR86, Eis00, HvMW79, MGG^{+}95, PvM93, vMHW85, vM93, vM96, vM99, vM01, vM05, Kra23b]. Bandenspektren [Kra23b]. bands [SWW88]. Banknotes [Ano05]. Bar [UM92]. Bar-Hillel [UM92]. Barad [FJ21a]. Barbara [Kuh67]. Bare [Ano39a]. barium [HS39a]. Bariumisotope [HS39a]. Barndom [Adl64, WS09]. Barr [Dje07]. Barth [LN96]. Barton [Hub14, Hof94]. Based [Boh61c, Boh62e]. Basic [Aas85b, Los61, Boh23h]. Basics [Som20c]. basis [Boh55a]. Bass [OSWR86]. basso [KvG85]. Baths [Ano22]. Battle [Kae45]. Bau [Ano23a, Boh22b, Boh23g, Boh24f, Boh25c, Boh36d]. Bauer [Dar92b, Bar23]. Baues [HK26, KH25a]. Bd [Dar87, HEB^{+}80, HP85]. be [Bar89, BO09, Boh35a, EPR35, Enz02a, Jak13, Mor95, Nik08]. bear [Dom06]. became [Ree15a]. Becker [DT18]. becomes [Gla00]. Bedeutung [Ish15, MGG^{+}95]. Before [Kra10a, Wyd84, Kon02, Nik08, Nik12, Pet11]. Began [FW67, Kae48, Fri54, RS07]. Beginning [Lem31]. Beginnings [Hei64a]. behalf [How99]. Behandlung [Ehr23]. behavior [UG25]. behind [Jak13]. Beijing [CHR96]. Being [Bad71]. Beitrag [Pau01]. Beiträge [FH60, KRW05]. Beitragen [Kuh67, MA65]. bekannt [Cas35]. Bell [WB06, CW12, HP01]. Bemerkungen [BL19]. Ben [KNA01]. Berechnung [BL18b, BL63b]. Bericht [ABH^{+}36a, ABH^{+}37d, ABH^{+}38a]. Berichte [KRW05]. Berkeley [Kra07, Lan08]. Berlin [wB90, Eis00, Hen86a, Hen86b, KLR13, Tan09]. Berlin/Heidelberg/New [Eis00]. Bern [Pau01]. Berner [Pau01]. berømte [Ano29, WS09]. Bertrand [Bal39, Coh42, N.38, Wei39]. berücksichtigt [CSW97a]. beskrivelse [Pih55]. besøk [Aas02h]. bestätigt [Kon15]. Bestätigung [Cas35]. bestemte [Jak13]. Bestrahlung [HS39b]. Beta [Jen00, Fra55, LH55]. beta-brass [Fra55]. Beth [Bor58]. Bethe [Lee07]. Betragtninger [Boh41c]. Between [AA89, AA90, Ang10, Lan08, Stu18, VW86, Bal85, Boh20a, Boh22c, Cam08, Che92, Che99a, CS79b, DS09, Dör05, HP01, Kra07, Kra13b, Kra25, Lau07, Oku01, Ros79a, Val06]. Bewußtseinswandel [vW88, vW91]. Beyond [Bag04, Bas71, Bok08b, Cas09, Dje07, Fav99, Rod19, Bro00a]. Beziehungen [Kos23]. bezüglich [UG25]. Bibliographie [Ano85b, Ano85a]. Bibliography [Ano85b, Ano85a]. Biennial [AG81]. Bifurkation [MGG^{+}95]. Big [Ano39d, Hür22]. Bildern [wB90]. Binding [AA89, AA90, Boh13a, PKWF08]. biografi [Blæ85]. Biografie [Eck13a]. Biographical [BDR86]. Biographie [Rec09, Rec10]. Biographies [SS98, Sim14]. Biographisches [BDR86]. Biography [Gam61b, Gam61a, Gam62b, Gam62a, Blæ85, Bod16, Bus20, Eck13a, Enz02a, Ken85b, Rec10, Sho13]. Biology [Boh38a, HH94, JFEH15, Ste89, RH00, Vol88, ZDSF11]. Birth [BH83a, BH83b, Bel92a, BD92, BR^{+}85, BH82, Jac85, Kel07, Kel09]. Birthday [CS79b, CS79d, CS79f, Gam60, KRW05, Pau45, Pau94, Ros79a, Ros79b, Ros79d, vM02, Ano85e, Ari35, Ber15, B^{+}55, Boh35b, BM55, Fra55, Gus55, HH55, Hei35, Her55, KP55, Las55, LH55, Møl55, Nie55, PRW55, Pih55, RM55, Ros45, Ros61, Mor56]. Birthplace [Rob15]. bis [Boh85e, Büh98b, Fis10a, Fis12a, Rec09, Rec10, Sch54, vW02b]. Bit [Ano39e]. Blaedel [wB90, Aas90a, Fre89]. Blaise [Hof94, Ruh92]. Blast [Ano39e, Ano43]. Blegdamsvej [PSV01, Smo93]. Blood [Hol80]. Bloomfield [Coh42]. Blow [Ano39c]. blue [Mer85]. Boats [LDN99]. Bohm [Lee06, Nik08]. Bohmsche [GRE^{+}01]. Bohr [Aas85b, Aas88, Aas00b, Aas02d, Adl03, All56, Ano23a, Ano29, Ano39a, Ano43, Ano68a, Ano68b, Ano71, Ano93a, Ano93b, Ano02a, Ano09, Ano10, Arr06a, ARK^{+}99, Bad83, Bal39, Bar23, Bel86, Ber91, Bis04, BAV01, Bor58, Bor13, Bor16, BG24, Bou49a, Bou49b, wB90, Bro36, Bro86, Bro95, Bro99, Bro00a, Bub90a, Büh98a, Bus12, Bus20, Cao90, Car97, Che85a, Che93, Che99a, Cla13, Coc63, Coh42, Cuf24, Cus94a, Dar87, Dar90, Dar91a, Dar92a, Dar92b, DDU86, Dor79, Duf46, Eis00, EFM01, FT19, Far13a, Far13d, Far13e, Far13f, Far13g, Fay90, Fay07, Fer07, FMO88, Fis10b, Fis12b, For91, For09, Fre90, Fri67, Gam60, Gam66, Goo92, GRE^{+}01, HEB^{+}80, HH55, Hei96, Hei86a, Hei35, Hei63a, Hei86b, Hel98a, Hel85, Hel98b, Hen86a, Hen86b]. Bohr [Hen92, Hen93, HvMW79, HP85, How86, How99, Jam67, Jan17, Jon14, Kat86, Kle93, KZN^{+}88, Kra96, Kra07, Kra16b, Kri91, Kub09, Kuh67, Lak96, Lan72, Lan08, Lau07, Len48a, Len48b, Lin81, Mac86, Mac95, Mai93, MGG^{+}95, MP97, Mil24, Mor56, Mot85, Mur90a, N.38, Nat99, Nie65, Nie68, OSWR86, Par06, PHvMW93, PvM93, Paw02, Ped64, Pia24a, Pia24b, Pye81, Rec86, Red93, Ren15, RS07, Rob79, Rob15, Rös92, Ryn83, Sch64c, Sch64b, Sch88, Sch11, Sch21, Seg93, Ser58, SvMS85, Sha67, Shi85a, Shi85b, Sho13, Smo93, Som20c, Som31, SE13, Som13a, Stu84, Stu92, Sud11, Tan09, Tur06, VGEK87, Was92, Wei39, Wes85, Whe81, Wil92, YRBD09, Zin19, dB86a, vMHW85, vM93, vM96, vM99, vM01, vM02, vM05, vW02a]. Bohr [Aas85a, Aas90b, Aas99, Aas00a, Aas02a, Aas02b, Aas02d, Aas02c, Aas02e, Aas02g, Aas02f, Aas02h, Aas03, Aas07, AH13, Aas15a, AK15, Aas20, Ach93, Alb92, AMH06, Alt82, AF17, Amu14, AA89, AA90, Ang10, Ano22, Ano23c, Ano23d, Ano23b, Ano23a, Ano29, Ano57a, Ano57b, Ano62a, Ano62b, Ano62c, Ano63a, Ano63b, Ano65, Ano85c, Ano85d, Ano85e, Ano85f, Ano93b, Ano94b, Ano95a, Ano95b, Anoxx, Ano02a, Ano04a, Ano13, Deg89, Ant96, AI15, Ari35, Arr06a, Ayl26, Ayl27, Bac15, Bal85, Ban00, Bar89, BO09, Bec57, Beh41, Beh43b, Beh43c, Bel92a, Bel92b, Bel93, BF94, BZ85, BA06, BV85, Ben00, Ber95, Ber89, Bet85, BGS95, BRB12, BH11, Blæ85, Blæ88, Bla85, Blo63, B^{+}55, Boh71a, Boh85a, BS50, BM55, Boh63a]. Bohr [BH64, BR64, Boh85f, BR^{+}85, BBCN86, BS87, BB87, BB^{+}89, Boh05, BA05, Boh07b, BHK^{+}08, Boh11a, BB05, Bok08a, Bol04, Bon10, BL18b, BL18a, BL19, BH23, BL63a, BG24, Bri58, Bro03, Bro09, Bro71, Bro73b, Bub73, Bub89, Bub90b, Büh98a, Bun79, Bun85, Bun88, Bun92, Bur18, Cam08, Cam07, CS15, Car83, Car04, Cas35, Cas85, CW12, Cav14, Cha09, CDH^{+}06, Che92, Che94a, Che94b, Che97, Che99b, Chi60, CM85, Cla88, CS79b, CS79c, CS79e, CS79d, CS79f, CS79g, Con49, Cos22, CG14, D’A01a, Dah02a, Dah02b, Dai96, Dam85, Dan89, Dar91b, Dar92c, Dar97, Dav86, Dav16, DS09, De 12, De 14, Des48, Dir24, Dir28, Dir64, Dir67, Dol95, Dom06, DG11, Dör05, Dot08, DJ15, DNT08, Eck14, Eck15, Ehr23, El’85, Eng03, Enz02b, Eps24, Erl72, Eva23]. Bohr [Faf58, Fal98, FG21, Fav91, Fav92, Fav94, Fav05, Fav09, Fay88, Fay91, FF94, FF17a, FJ21a, FJ21b, FM12, FR13a, Fol85, Fol94, Fol95, Fol96, Föp15, Fox80, Fra63, Fra55, FK85, Fre86, Fre89, Fri67, FF78b, Fuc85, Gam60, Gam63, Gao15, Gar72, GS04, Gar63a, Gol94, Gom06, GLSC01, Got02, Gow85, Gra15, Gre00, Gru14, Gus55, Hal12, Ham05, HH55, Har83, Har87, Har23, Har26, Harxx, HF15, Haw94, HK69, Hei77, Hei81, Hei85b, Hei85c, Hei85d, Hei13, Hei15, Hei20, Hei28, Hei35, Hei62, Hei86b, Hei10a, Hel98a, Hen81, Hen84, HP85, Her55, Her23, Her72, HP01, Het95, HN64, HN70, Hoc83, HLA^{+}10, Hol86, HK22, HK26, HKK29, Hol88a, Hol05, Hol80, HG15, Hon81, Hon82a, Hon82b, Hon88, Hoo91, Hoo94, How79]. Bohr [How18, Hoy73, HnB74, HH94, Hug90, Hun85, Jac85, JVK22, Jam74a, JFEH15, Jon86, Kai94a, Kai94b, KvG85, KP55, Kal16, Kan96, Kat98, Kat11a, Kat11b, Kat11c, Kat15, Kel58, Kel83b, Kle64, Kle70, Kle10, KN12, Kom70, Kon15, Kon83, Kon00, Kon02, KK03, Kos23, Kra79, KNA01, Kra09, Kra10a, Kra10b, Kra11a, Kra11b, Kra11c, KN11, Kra12a, Kra13a, Kra13b, Kra13c, KN13, Kra15c, Kra15a, Kra16a, Kra17, KH23, Kra24, KH25a, KH25b, Kra14a, Kra14c, Kri94, Kub08, KRBR62, Kum10, Kum11, LR23, Lan59, Lan72, Lan06, Las55, Lat83, Lau33, Lau13, Lee06, LH55, Lib75, Lin07, Lin08, Lin24, LN96, Los61, Mac85, Mac94, MGG09, MA13, Mal15, Mar88, Mas04, McC34, McE72, McK05, MR82e, MR82f, Meh87, Mer04, MWL^{+}08, Mig85, MW84, MB24, Mlo14]. Bohr [Mof06, Møl55, MB11a, Moo66, Moo67, Moo69, Moo85b, Moo85a, Mor85b, Mot08, Mur87, Mur89, Mur90b, Mur94, Nak15, Nau13, Nau15, Nav15, New09, Nie55, Nie65, Nie15, Nie63, Nik08, Nik12, Nis37, Nis67, Oku01, Oli85, Opp63c, Opp63a, Opp63b, Opp64, OBxx, OP04, OPH^{+}09, OK12, Pai48, Pai85a, Pai91, Pal15, PRKH09, Pas03, Pat85, Pau45, PRW55, PB58, PHvMW93, PvM93, Pau94, Pau01, Paw02, Ped64, Pei40, Pei55, Pei63, Pei85, Pei88, Pei97a, Pei97b, Pei97c, Pei08, Pei10, Pel47b, Pel47a, Per85, PV15, PP16, Per06, Per13, Per21, PKWF08, Pet63, Pet85, Pet88, Pih55, oMP64, Kam64, Pih64, POC^{+}10, Pla23, Pla24a, Pla24b, Plo94, Plo06b, Plo10, Plo12, Pod10b, Pol60, Pow02, Rad82, RM55, Red63, RR00, Ree05, Rhe91, Rio17]. Bohr [RC90, RH00, Roq17, Rös92, Rös94, Rös95, Ros45, Ros61, Ros63, RRK^{+}64, Ros79a, Ros79c, Ros79b, Ros79d, Rot87, Roz67, Roz85, Roz98, Rub97a, Rub97b, Ruh92, Sac81, Sac88, San24, Sce94, Sch15, SWW88, Sch68, Sch54, Sch88, Sch22, Sch24, Sch13, Sch00b, Sch02, Sch77, SvMS85, SK75, She72, She85a, She86, Shi83, Shi86, Sho08, Sie13, Sim14, Smi76, Som15, Som18, Som20a, Som20d, Som20b, Som24a, Som24d, Som24c, Som24b, SE13, Som13b, Som13a, SM95, SM08, Sta09a, Ste11, Ste89, Ste86a, Ste86b, Ste87, Ste88a, Ste88b, Stu85b, Suv85, SB72, SSH05, SSH14a, SSH14b, Swe02, Szu05, Tal79, Tan02, Tan04a, Tan04b, Tel69, Tel80, Ter94, Ter08, TM13, Tuc97, UPG^{+}12, Urt06, Vic08]. Bohr [VIR^{+}08, VW86, Was81, Wea88, Wei14, Wei63, Wei64, Wei85a, Wei85b, Wei91, Wer23, Whe63, Whe85b, Whi96, Whi04, Whi06, Wil92, Wil85, Wyk08, Yam02, Zin01, Zin15, Zin16, dB86b, dBdB21, dB24, dB63, del81, vCK17, vD15, vFH76, vH23, vM02, Aas88, Aas90a, Amo78, Gla60, Hof94, Hub14, Lin64, Mal70, Pia23, Pol58, Rie60, Ste58, Mer89]. Bohr-Dokumente [Paw02]. Bohr-Festschrift [Ano93b, Pau01]. Bohr-Heisenberg [Dah02b]. Bohr-like [MWL^{+}08]. Bohr-Sommerfeldschen [vM02]. Bohr/Einstein [Zin01]. Bohrende [Paw02]. Bohrian [BL63b, RC14]. Bohrification [HLS12]. Bohrs [BR^{+}85, Car83, Fra63, Hei10a, Ano13, Dir64, Hei62, HK22, HKK29, Kra13a, MA65, oMP64, Sch24, Kuh67]. Bohr’sche [Cas35, BH64, HK26, KH25a, Pla23, Pol60, Sch22, Som13b, Her23]. Bohr’schen [Som20a, Som20c, Som24a, Som24d, Som24b, Bar89, BH23, Föp15, Hei75, Hoy73, HnB74, Kos23, LR23, Pla24a, Pla24b, Som15, Som18, Som20d, Som20b, Som24c, SE13]. Bohrscher [BL18b, BL63b]. Bohrsches [BL18a, BL19, BL63a]. bolt [Mer85]. Bomb [Ano02a, Ber95, BL03, Cho02, Duf46, Far13d, Far13e, Gla00, Kae45, Luc11, MD67, Rei84, Bag10, Cas09, Far13b, Far13c, FP40a, FP40b, Jak13, Kar05b, Kar07, Kel07, Kel09, MW46a, MW07, Mel00, Ree15a, She72, Van03, Zan02, Duf46]. bomba [Kar06]. Bombarded [Ano23d]. Bombarding [Ano23d]. Bombardment [Fri39, MF39b]. Bombe [Ano02a, Kar05b, Kar07]. bomben [Jak13]. Bombs [Can00]. bonding [WB06]. Book [Aas88, Aas90a, Amo78, And89, Ano93a, Ano09, Bad83, Bal39, Bel86, Bis04, BDR86, Bor58, Bor13, Bor16, wB90, Bro86, Bro95, Bro99, Bro00b, Bro00a, Bub90a, Bus20, Cao90, Car97, Che85a, Cla13, Coh42, Cus94a, DT18, Dje07, Dor79, FT19, Far13a, Fay90, Fay07, Fer07, For91, For09, Fre89, Fre90, Gla60, Goo92, Gor17, Hei96, Hei86a, Hei75, Hel85, Hen86a, Hen86b, Hen92, Hen93, Hof94, Hol00, How86, Hub14, Jam67, Jan17, Kle93, Kra96, Kra07, KN13, Kra16b, Kri91, Lan59, Lan08, Lau07, Len48a, Len48b, Lev95, Lin64, Lin81, Mac86, Mac95, Mai93, Mal70, McC34, Mer89, Mil24, Mor56, Mur90a, N.38, Nat99, Nie65, Nie68, Par06, Pia23, Pol58, Pye81, Rec86, Red93, Ren15, Rie60, RS07, Ryn83, Sch64c, Sch64b, Seg93]. Book [Sha67, Shi85a, Shi85b, Ste58, Stu84, Stu92, Tan09, Tur06, Wal08, Was92, Wei39, Wes85, Whe81, Zin19, dB13, Dea13, Wil92]. Books [All56, Fri67, AFG90, Nie15]. Bor [K^{+}86, Moo69, PB58]. Bora [Ter94]. borders [Bei13, Ben13]. born [Nik08, CS79b, FL10, Nik08, Ros79a]. Borsche [vH23]. Boru [PB58]. boson [Kra14b]. Boston [Che85a, FMO88, Wes85, CW85, MP97]. Bottles [Bro09]. Boulder [Wei68]. bound [FF78a, Kle78, Nie15, Pai86]. boundary [Gus55, POC^{+}10]. Bragg [Pia24a, Pia24b]. Brandmüller [MGG^{+}95]. brass [Fra55]. brauchen [Aas00c]. Braunschweig [Sch64c]. Breakdown [BAV01, EFM01, Mel00]. breakthrough [oER64]. breakthroughs [Lig05]. Breit [FR13a]. brev [Boh50a, Boh64d]. Brief [Kat11a, KZN^{+}88, Enz02a, How99]. Briefe [KRW05]. Briefwechsel [ARK^{+}99, Car97, Eis00, GRE^{+}01, HEB^{+}80, Hen86b, HvMW79, HP85, Kra96, MGG^{+}95, OSWR86, PHvMW93, PvM93, Pye81, vMHW85, vM93, vM96, vM99, vM01, vM05]. Brigitte [Hel98b]. Brillouin [Pia24a, Pia24b]. bring [Yor75]. Bringing [Dom06, Kai94a, Stu85a]. Brintspektret [Boh14c]. Britain [Far13c]. British [Boh38b, Jon78, Kra11c]. Broadway [Sch00b]. Brock [Bis04]. Brod [MGG^{+}95]. Broglie [Bou49a, Bou49b, Pia24a, Pia24b, Ano62a, Len48a, Len48b, Pel47b, Pel47a, Ser58]. Broken [Cas55, Bro15, Dje07]. bronia [Kar06]. Brosche [HEB^{+}80]. Brown [KZN^{+}88]. Bruchstücke [HS39a]. brudstykker [KvG85]. Brussels [Ano28, CCJ^{+}34, LRdB^{+}23]. Bruxelles [Ano28, CCJ^{+}34, Far01, LRdB^{+}23, Pel47a, Bou49a, Bou49b]. Buchbesprechungen [BDR86, MGG^{+}95, OSWR86]. Bücher [ARK^{+}99, GRE^{+}01, HEB^{+}80]. Bühne [Aas00c]. Builders [MD67]. Built [Pei55]. bulk [OBxx, Rabxx]. Bulletin [FR98]. Bunde [MGG^{+}95]. Bunsen [Som20b]. Bunsen-Gesellschaft [Som20b]. Byen [Ano29]. Bygning [Boh21f, Aas02e, Boh23d].

C [Adl03, BDR86, dB13, BDR86, Fra55]. C. [SR99]. CA [Kra07]. Cadmium [Alt82]. Calculation [BL63b, BL18b, Bor23]. Calculus [Fey33, San24]. California [Lan08, Ano22]. called [Che99a]. Cambridge [FMO88, Kri91, Par06, Tur06]. Campaign [Bør12]. Can [Boh35a, EPR35, Lau33, Mas04, Bar89]. Canada [Inf78]. Cancer [Kae39]. Candid [HH04a]. Cannon [Ano39a]. Cannonball [Ano39d]. Capture [Ano36a, Boh36c, BL54, Faf58, Boh36d, Las55]. Carbon [Faf58]. Cards [Ghi05]. Carl [Ano02a, GRE^{+}01, Ano02a]. Carnap [Bal39, Coh42, N.38, Wei39, FJ21b]. carpenter [Jak13]. Carrington [RS07]. Case [Eng03, Joh13, PSV01, Gau14, Het95]. Cassidy [MGG^{+}95]. Cat [Sta09f, Ste87, Tri80, Vil86]. catch [MMS^{+}19]. Catherine [Dar92b]. Cathode [Nia98]. Causality [Boh37a, Boh39b, Boh48a, Boh50c, Boh51, Boh85b, Boh91, BFF98, Bor58, Los11, Boh36b]. caused [UPG^{+}12]. Cavendish [Nav06]. Celebration [KRW05]. Celia [FB00]. Centenary [DDU86, Hei86a, FK85]. Centennial [FMO88]. Central [Har28, K^{+}84]. Centuries [Dor79, SDG15]. Century [ABP98, Den99, FT19, OP04, Zin19, APB96, Bra09, Cas83, D’A01b, FF17a, GL03, Jac12, LSW12, Lig05, OPH^{+}09, Rio17, SSH14a, VGEK87, Wei77, Bre97, D’A00, FW60, Gin01, Bro00b]. CERN [Kra14b, Lin55]. chain [And73]. Challenge [Boh45a, Rös94, Sta09b, Boh45b, Boh85h, BBCN86, BB^{+}89, CS15]. champions [Lan06]. Chance [Hof94, Hub14, Ruh92]. change [Gau14, vW88, vW91]. changed [GL03, Moo66, Moo85b, Hel85, Gam66, Kuh67, Sha67]. changer [Ree15a]. Changing [Aas85b]. Chaos [MGG^{+}95]. Chaos/Plaschko/Brod [MGG^{+}95]. Chaotische [MGG^{+}95]. Charge [BR50, BR79a, CS79c, Pai48, Pau55]. charged [BO09]. charges [Las55]. Charles [Bal39, Coh42, Hof94, Hub14, N.38, Wei39]. chelovek [Moo69]. Chemical [Nis67, Boh21f, Boh22b, Kos23, WB06]. Chemie [ABH^{+}37d, ABH^{+}38a, Ber15, MGG^{+}95, Som20b, ABH^{+}36a]. Chemie/Dictionary [MGG^{+}95]. chemischen [Boh22b]. chemischer [Kos23]. Chemistry [Boh32b, Joh13, Nia98, NM12, Sch90, Som20b, dB13, Ber15, Kra13b, Sce94, ABH^{+}36a, ABH^{+}36b, ABH^{+}37a, ABH^{+}37c, ABH^{+}37d, ABH^{+}38a, ABH^{+}38b, ABH^{+}39b, ABH^{+}40a, ABH^{+}40b, MGG^{+}95, Sto97]. Chemists [Joh13]. Chevalley [Dar92b]. Chicago [Cuf24]. childhood [WS09, Adl64]. chimica [ABH^{+}36b]. Chimie [Bar23]. Chimie-Physique [Bar23]. China [Kan96]. Chinese [DC96]. choice [Ken85a, MW84]. Chronological [Ano64]. Churchill [Far13b, Far13c]. cinquième [Ano28]. circle [BR64]. Civilization [Boh45a, Boh45d, Boh46a, Boh07a]. Clarendon [Ano93a, Red93, Seg93]. Clarify [Sch02]. Clarity [Pei86a]. Clark [RS07]. Clarke [Sch90]. Classical [BB05, Bok08b, Bon10, Dar92d, Dar97, How18, Plo11, Amu14, CS15, Gru14, Kav14, Mlo14, SSH14b, Zin16, Pih55]. classical/quantum [Zin16]. Cline [Kuh67]. clockwork [DNT08]. closing [Fre85]. Closure [Vic08]. cloth [Cuf24, Nie15, Par06]. club [Ber96]. cm [wB90]. co [RRK^{+}64]. co-workers [RRK^{+}64]. codiscoverer [CSW97b]. Coefficient [Fra55]. Cohen [GSW95]. Cohérence [Dar91b]. Coins [Ano05]. Cold [Sal19]. Collaboration [KN12, BLW19, Koc64, Wei64]. Colleagues [Jam67, Kle10, BR64, Roz67, Roz85, Nie68]. Collected [BDR86, BRH^{+}99, Boh81, Bro99, Bro00a, Dor79, KZN^{+}88, Kub09, Bel87, Boh05, Boh07b, BHK^{+}08, Pei08, Stu84]. collection [Ano29]. Collections [Hei96, Sch64b, Nie97]. Collective [HEB^{+}80, Gus55, KNA01]. collects [Fri64]. collide [Wal08]. collision [Her23]. Collisions [Boh77]. Colloquium [Bas71, CW85, UM92]. Colorado [Wei68]. Columbia [Ano39a, Ano39d, Lau51]. comme [Dar85]. Commemorations [Bel99a]. Comment [BAV01, EFM01, YRBD09, Hel98b]. commentary [DP97, PD97]. Comments [dR92]. Commission [ABH^{+}36a, ABH^{+}36b, ABH^{+}37d, ABH^{+}38a]. commissione [ABH^{+}36b]. Committee [ABH^{+}37a, ABH^{+}37c, ABH^{+}38b, ABH^{+}39b, ABH^{+}40a, ABH^{+}40b]. Common [FG21, KN12]. Communicating [vD15]. communication [Hei28, Kat15]. communicazione [Hei28]. Community [GSW95, Hab69]. Como [And89, Boh85g, DS09]. Company [Bro86, Che85a]. Comparison [Los61]. Compendium [GHW09]. complaint [Ros35]. complementarità [DS09]. Complémentarité [Che85b, Ano85b, Ano85a, BV85, Bla85, Dar85, Des48, Pat85]. Complementarity [Ban00, Bel92b, Bel93, Boh37a, Boh48a, Boh50c, Boh51, BFF98, Bro09, Bro00a, Che85b, Cla88, D’A01a, Fav99, Fol94, Gra85, Hol86, Hol70, How79, Jam74b, Jam74c, Jon86, Jon85a, Kal16, Kat11b, Kat11c, Kat11d, Kat11e, Kat86, Lau33, Mac96, Ott03, Pet11, Plo94, Plo06a, RC14, Rös95, Sta09c, Ste86a, Ste86b, Ste87, Ste88a, Ste88b, Ste58, Tel69, Vol88, Ano85b, Ano85a, Bel92a, BV85, Boh36b, Boh91, Cam07, De 12, De 14, Des48, Fol85, Hel98b, Kat98, Kot85, Pat85, Per13, Plo12, RH00, Ros64, Tan02, vD15, Dar85, DS09, Meh87, Shi85a, How86, Mac86, Shi85b]. Complementary [Bus12, Ste11, Wei35]. Complete [Boh35a, EPR35]. Completeness [Hoo91]. Completion [MR00]. complétude [Dar91b]. composition [KvG85]. Compounds [Chi60]. Compton [Duf46]. computation [Mer04]. Computer [Hol80, Wal08]. concept [Ano85b, Ano85a, Pel47b]. ConceptDraw [GRE^{+}01]. Concepts [Dar97, GHW09, Mot08, CS15, Ste86a]. Conceptual [Jam89, Kai94b, Kra11a, Kuh67, Jam66]. concernant [Ano85b]. Concerning [Fuc85, Ano85b, Ano85a, Boh71a, Boh85a, Mar58, HS39b]. Condensed [ARK^{+}99, KZN^{+}88]. condition [Dir24, FF78b, KK03, Sch22, TM13]. Conditions [Aas85b, FF78a, Hoy73, Kel58, Pei39, POC^{+}10]. conductivity [UPG^{+}12]. Conference [KLR13, Roq17, Ros79e, AK15, CHR96, Far01, Pau01, BV09, Bor16]. Conferences [Anoxx, Meh75, Boh23b]. confinement [UPG^{+}12]. Conflict [CS79b, Ros79a, Che92]. conflit [Che92]. Confronts [New09]. Congratulations [Hei86b]. Congress [Ano28]. connaissance [Boh61b, Boh72, Dar92b]. Conrad [MGG^{+}95]. conscience [Wil70]. consciousness [Wal08, vW88, vW91]. conseil [Ano28, CCJ^{+}34, LRdB^{+}23]. Consensus [Jen00]. Conservation [Boh32a, Boh36a, GLR14]. consideration [CSW97a]. Considerations [Nis67, Boh20b]. Considered [Boh35a, EPR35]. Consistency [Vic08, BO09]. consolidated [Ros64]. conspicuous [MB24]. Conspiracy [Wer02]. Constant [Ayl26, Ayl27, SB72]. Constituting [BKP09]. Constitution [Ano36a, Boh13a, Boh13b, Boh13c, Boh21d, Boh32b, Boh36c, Boh36e, Boh62c, Boh63f, Boh65b, HN64, HG15, Boh22b, Boh22g, Boh24e, Som18, Pia23]. Constructing [Hen86a]. Construction [HK26, Mac95, Che97, Fol95, KH25a, Pet88]. Constructions [Set03]. Containing [Boh13b, Boh13c, Pai48]. Contamination [Ano91]. contemporaine [Ano62a]. contemporanea [Seg76, Seg83]. contemporanee [Cas14]. Contemporary [Aas02b, FF94, Kai94b, Ste89, Ano62a, Cas14, Che94a, MP97, Seg76]. Content [Vic08, Kom70]. Context [KN13, Bel92a, Kai94a]. Contingency [Lev95, Cus94b]. continued [dR92]. Continues [BSMB99]. Continuing [Mur90a, Sac88]. Continuous [BPP39, BPP97]. Contrast [Mor85b]. Contribution [CS79e, Ros79c, Pau01, Ros63]. Contributions [KRW05, FH60, oMP64]. Control [She85a, Her55, She86]. Controversies [Kra76, Mur90a, Sac88]. Controversy [Jen00]. Conundrums [FM12]. conversations [HH04a]. Cooperation [Boh47, Boh60a, Sal19]. Copenhagen [And89, Bad83, Che99a, How04, Kra07, Lan08, Lev95, Mai93, Smo93, Wal08, Aas00a, Aas00c, Aas02f, Ano29, Ano71, Ano02b, Bel99a, BJN01, BSMB99, Cas00, Cas02, Che85b, Cus94b, Dör05, The00, Fra98, Fra00, FB01, Gom06, Hei85a, HB63, Kar05a, Lau07, MA13, Mer04, Paw02, PA01, Pow03, Ros79e, Seg07b, SvMS85, S^{+}06, Tan02, Wal19, Whe85a, vW71, Dje07, BW00a, BW00b, Hol00, Ros00, Sch00b]. Copenhagen. [Log00]. Copenhague [Che85b]. Copper [Alt82]. corpusculaires [dBdB21]. Correct [New09]. Corrected [CS79d, Kel58]. Correspondence [AF17, BS50, Bro95, Car97, D’A01a, Dah02a, Dah02b, Dor79, Eis00, Hel98b, Hen86b, HvMW79, Pet14, Rec86, vMHW85, vM93, vM96, vM99, vM01, vM05, Boh33c, BAV01, DNT08, EFM01, El’85, Harxx, HP85, Jäh15, Kav14, Lee07, Lee09, Lib75, PHvMW93, PvM93, Ryn15, Tan02, Kom70, Kra23a]. correspondences [PP16]. Cosmic [Faf58, MGG^{+}95]. Cosmological [ARK^{+}99]. Cosmology [Rei72]. costruzione [Dar90, Goo92, Pet88]. Could [Ano39e, CW12]. Coulomb [Har28]. courage [Gol87, Hei15]. covering [Dea13]. Cradle [Kel83a]. Crafting [Set10]. Crazy [Arr06b]. Creating [The00, ZDSF11]. Creation [Bel86, Bro86, Che85a, Hei13, Pur18, Hen84, Stu13, Wes85, dB86a]. Creative [Rot87]. creativity [Aas15a]. Creator [Dje07]. creators [Kel07, Kel09, Str11]. Credited [Ano39b]. Crisis [Hen09a, Hoo94]. Critical [Pei39, Sta09b, Pop82]. Criticism [Rad82, Car83]. Critique [Kal16, Kri94, Per06, Som20d]. Crocodile [Mac97]. cross [Las55]. cross-sections [Las55]. Crossing [Bei13, Ben13]. crossroads [BV09, ET07]. Crusade [Aas03, Aas99]. Crystal [BL63b, BL18b, BL19, BL18a, BL63a]. cuántica [MR02, HBS96]. cuatro [Boh88]. Cultural [Ang10, Eck13a, KN12]. culture [Boh41c, Cas14, Dah43, LSW12]. Cultures [Boh39e, Cas14]. Curies [Kae48]. curious [Gau14]. Curtiss [Mil24]. Cushing [Lev95]. Cyclotrons [Row09].

D [Ano43, Bro86, Che85a, Dje07, KP55, Kha89, KvG85]. D-flat [KvG85]. d-Moll [KvG85]. dêbat [Kum11]. d’action [Boh39f]. Dag [Ano21]. dagegen [CSW97a]. dai [Seg83]. Dalton [Kra14c]. Dan [Che93]. Dane [Ano43, Ano62c, Jak13]. Danes [IH07, WS09]. Danish [CM85, Mai93, Ped64, Aas02e, Adl64, And64, Ano21, Ano29, Ano64, Ano65, Ano13, Blæ85, Boh11, Boh14c, Boh21f, Boh23d, Boh24a, Boh24c, Boh25b, Boh29b, Boh29c, Boh30a, Boh31e, Boh35c, Boh37d, Boh37e, Boh39c, Boh40c, Boh41c, Boh41e, Boh42a, Boh42b, Boh45c, Boh46c, Boh50a, Boh55a, Boh58f, Boh58d, Boh64d, Boh64e, Boh64b, Boh64a, BR64, Boh64c, BR^{+}85, Boh11a, Cas64, Cou64, Dah43, Dir64, Fav05, Fav09, Fri64, Gam42, Gol87, HH55, Hei64b, HK22, HKK29, IH07, Jak13, Kal64, KvG85, Kie07, Kle64, Koc64, KNA01, Kra13a, Kra15a, Kra25, Lev00, LDN99, MW46b, Mor85b, NN01, Pai64, Ped64, Pih55, oMP64, Kam64, Pih64, Ros64, RRK^{+}64, Roz64a, Roz64b, oER64, Sch64a, Sch68, S^{+}06, SO93, Teu02, Wei64, Wer02, WS09]. Danmarks [Dah43]. dans [BV85, Che85b, Che97, Dem15, Pat85]. dansk [Teu02, Boh41c]. Danske [Ped64, NN01, Pih64]. danskere [IH07]. Danskeren [Jak13]. danskeres [WS09]. Darkness [Lev00]. Darstellung [Mor11]. Data [Cuf24, Per21]. d’atome [dBdB21]. Daughter [Pal00]. d’autorité [Dar85]. David [Ano57b, Ano63a, Bro00a, Che93, Coc63, OBxx, Wal08, Fer07, For09, Mai93, RS07]. Davies [KZN^{+}88]. dawn [Wei14]. Day [Lau51, Wyd84, HHW99]. Days [Bar87, Pop82]. DC [Ano57b]. Dear [Coh88, Coh91, Coh92, dR92]. Debate [Ano39c, Ano09, Arr06a, BSMB99, FG21, Fol94, Fol96, Hel98a, Jam74a, Kra07, Lan08, Lau07, Sud11, Zin01, Bok08a, Bun79, Cam08, Dör05, Kai94a, Kum10, Kum11, Lan06, Oku01]. Debates [Mac94]. Debatte [Hel98a, Arr06a]. Decay [Jen00, Gau14]. December [Sch21]. Decency [Wer02]. decidability [HP01]. decided [Jak13]. decisive [Hoc83]. déclare [Ano62a]. decoherence [CS15]. Deconstruction [Kat11b]. découvertes [Seg84]. Decrease [Boh13d]. Dedicated [CS79b, CS79d, CS79f, Mor56, Ros79a, Ros79b, Ros79d, B^{+}55, BM55, BBCN86, BB^{+}89, Fra55, Gus55, Her55, KP55, Las55, LH55, Møl55, Nie55, PRW55, Pih55, RM55, Ros45, Ros61]. Defect [HEB^{+}80, PKWF08]. Defects [MGG^{+}95]. Defects/Weiss [MGG^{+}95]. Defend [Ach93]. defiant [Bol04]. Defining [ARK^{+}99, Mla98, De 12, De 14]. Definite [Amo78, Sch77]. deformations [PRKH09]. Degenerating [Rad82]. degli [ABH^{+}36b]. del [Wei77]. Delayed [Ken85a, MW84]. Delayed-choice [Ken85a]. Delbrück [Dom06, JFEH15, McK05, RH00]. della [ABH^{+}36b, Wei77]. dell’indeterminazione [Cav14]. demonstrative [Mas04]. Denmark [Smo93, Ano05, BJ03, Dah43, Lev00, Sch11, SM95]. dens [Hei64b]. deny [Bør12]. Depth [Hoo94]. derived [TM13]. Derrida [Plo94]. descobertas [Seg87]. descripción [Boh88]. Description [Boh30a, Boh31a, Boh34, Boh35a, Boh61a, Boh78, Boh88, Boh11b, Bro36, Cao90, McC34, Boh29b, Boh29a, Boh30b, Boh32c, Boh58d, Boh87b, BB87, Boh93, EPR35, Hon88, Pih55, Som31]. deserve [KNA01]. destruction [Byr10]. detect [Nav06]. Determination [Boh09a, Boh09b, Boh10, ES74, Che92]. determiner [SWW88]. deuterium [ES74]. Deuteron [Boh41a, Boh41d]. Deuteron-Induced [Boh41a, Boh41d]. Deutsch [GRE^{+}01]. Deutsch-Englisch [GRE^{+}01]. Deutsch-Englisch/ConceptDraw [GRE^{+}01]. Deutschen [Som20b, Kar05b]. Deutschland [MGG^{+}95]. Deutung [HB63, Kos23]. develop [Jak13]. Development [All56, Beh41, Beh43b, Boh75a, Bra61, Hen86a, Jam89, Kat11a, Kuh67, Mor56, Mot08, Nis67, Sch90, Was81, Yag72, Boh28c, Boh28d, Boh87a, El’85, Hei55, Hoy73, Jam66, Meh75, MR82d, Pai85a, PRW55, PB58]. Developments [Beh43c, Boh61c, Boh62e, Pai48, Boh28b, Boh28a, Boh29c]. développement [Boh28b]. Deviant [Gar72]. Dewey [Bal39, Coh42, N.38, Wei39]. Dexer [Duf46]. DF1 [Che85a]. Diagram [Dow09]. dialectic [LN96]. Dialectical [FL10]. Dialog [Aas00c, MGG^{+}95]. Dialogue [Bel92b, Che85a, Kle70, RW96, dB86a, Aas00c, Bel99b, Boh85f, Car04, Hen84, Kra13c, Bro86, Wes85, Bel86]. dialogues [Bel92a]. diaries [Ros18]. dibattito [Cam08]. dice [MB11b]. Dictionary [MGG^{+}95]. Did [Bac15, Ber95, BGS95, Rös95, CSW97b, Das15, Wil92]. Didn’t [Wei94, Dav86]. Dielectric [Ayl26, Ayl27]. Dies [Ano62c, Cha09]. difference [Boh22c, Che99a]. Different [Ano23d]. difficult [DG11]. Difficulties [Hen86a, MR82e, MR82f]. difficulty [MB24]. diffraction [BO09]. Diffusion [ARK^{+}99]. Dilemma [Whi06, Whi96]. dimensional [FF78a, Urt06]. d’incomplétude [Pat85]. Dinner [Ano23b]. Diplomatic [Aas20]. Dirac [Lak96, Bal85, Bok08a, Moy81a, Moy81b, Moy81c, TM13]. Direct [PKWF08]. Disadvantages [Mar54]. disappears [Ano62a]. disciple [Kal64]. disciples [Kal64]. discontinuity [Kra13c]. Discoveries [Kra76, OP04, Bra09, GL03, Lig05, OPH^{+}09, Seg76, Seg80, Seg84, Seg87, Seg07a]. Discovery [Ano39a, BHT86, Boh62a, FW67, Gar63b, Gra64, Hei64a, Hen86a, Kon02, Mei62, Sim00, Boh63d, Dar55, FW85, Gou71, GA71, Kae48, MR82a, Rob12]. Discreteness [BH11]. Discusses [Ano23c, Ano23d]. Discussion [Boh49, Dir28, Gam29, VW86, dB86b, Som20c]. Discussione [Hei28]. Discussions [Bas71, Boh38b, Meh87, Ano28, CCJ^{+}34, LRdB^{+}23, Boh21a, Hei28]. Disintegration [Boh39d, MF39a]. Disintegrations [Boh39a]. Diskussion [Som20c]. disparait [Ano62a]. Dispersion [Kra24, LR23, Dav16, Kon00, Som15, Som18]. Dispersionsformel [Som15]. Dispersionstheorie [Som18]. Dispute [Mur94]. Dissidents [Fre15]. Distance [Aas02a]. disunities [Fal98, Hel98b]. Ditfurth [vW02a]. Divergent [Hoo91, Cam07, Dom06]. divide [Zin16]. divine [DNT08]. Division [Fri39]. DM [wB90, Hen86a, Hen86b, Sch64c]. do [Che99b, Gao15, Kra11a, Mas04, Wil92]. Dobzhansky [Adl03]. Doctor [Dje07]. doctrine [CS15]. documentary [GA71]. documentation [Aas02h]. documenti [DS09]. Documents [Kra07, Lan08, Lau07, Aas02d, Aas02g, Ano02a, CSW97b, DS09, Dör05, Paw02]. Does [CS15, MB11b]. dokumentasjon [Aas02h]. Dokumente [Ano02a, Paw02]. donor [PRKH09].

Doppelspalt [Kon15]. Doppler [Dir24, Sch22]. Dopplerprinzip [Sch22]. Dopros [Ter94]. Dordrecht [Che85a]. Dordrecht/Boston/Lancaster [Che85a]. Dörries [Kra07, Lan08, Lau07]. double [Kon15]. Doubleday [RS07]. Doubtful [Kae39]. Douglas [Harxx]. d’ouvrages [Bou49a, Bou49b, Che93, Dar87, Dar90, Dar91a, Dar92a, Dar92b, MP97, dB86a]. doxographie [Che97]. doxography [Che97]. Dr. [Ano23c, Ano23d, Ano23b, Ano43, Ano62b]. Draft [Ano62c]. Drama [Aas02h, PSV01, Gla00]. Drawing [Ber03]. dream [A^{+}09]. dreams [Haw11]. Drei [Boh22d, Boh24b]. Dritter [ABH^{+}38a]. driven [Bro15]. drogi [Inf33]. Drømmeland [Gam42]. Drop [Ano94a, Stu94]. dropped [Jak13]. Drudesche [Som18]. Drudian [Som18]. Dual [Lau33]. Duality [AMH06, Kal16, NM12, dB13]. Dubna [Ano97]. Dugald [Bub90a, Dar91a, Fay90]. d’une [Boh32c]. durch [Ehr23, HS39a, UG25]. During [Wer02, Aas99, Aas03, Aas20, KN12, Mel00]. Dürr [GRE^{+}01]. Dutch [Bur18, Cos22]. Dyby [LDN99]. Dynamics [Bet37, Bon10]. Dynamik [MGG^{+}95].

E-P-R [Mer85]. E. [Bar23, Boh62b]. Each [Ano39d, Rös95, UG25]. earliest [Hei85a]. Early [Bai13, Bar87, Beh41, Beh43b, Beh43c, Hen09a, Her72, Jam74b, Kra10b, Kra11c, Lin81, Nav06, Whe81, Bad83, Bar89, Dar88, Eck15, Kat98, Pal15, PP16, Rob79, Rob15, Rot00, ZDSF11, Sal19]. Earth [Eva96, FF17b, HEB^{+}80, HS39b, Ano39e]. East [Ros18]. eastern [Kot85]. easy [HK22, HKK29]. Ed [Jam67, Lan08, Dar91a, Kra07]. Eddington [Adl03]. Edinburgh [Hen86a]. Edited [Dor79, Duf46, Eis00, Sch64b, All56, Fri67, Kub09, Mor56]. Edition [CS79d, Ros79b]. Editor [Ben13, Pow02]. Editorial [Wei85a]. editors [And89]. Edmond [Dar92b]. Edmund [Nau13]. Eds [Zin19, Bor16]. Education [PSV01, Mof06]. educational [Mof06]. Edward [vW02a]. Effect [Alt82, Boh14b, Boh15a, Chi60, Hol80, Red63, Boh23e, Boh25d, DJ15, Jäh15]. Effectiveness [Plo11]. Effects [Boh38c, Boh38d]. Efforts [Gla00]. efterkrigstiden [Pai64]. Efterskrift [Aas02e]. Egenskaber [Boh35c, Boh21f, Boh31e]. ego [PB58]. Ehrenfest [Kle10, PV15, PP16, Pia24a, Pia24b]. Eigenschaften [Boh22b]. eigentlich [Bel86]. Einheit [Lan72]. einleitenden [Boh31a]. einmal [Boh63e]. Einstein [Adl03, Ano09, ARK^{+}99, Bal85, Bro95, Car97, Eis00, Fay07, Fer07, For09, GRE^{+}01, HEB^{+}80, Hel98a, Hen86a, Hen86b, HvMW79, HP85, Kon15, Kra96, LM85, MGG^{+}95, OSWR86, Par06, PHvMW93, PvM93, Pye81, Rec86, RS07, Sud11, Tur06, vMHW85, vM93, vM96, vM99, vM01, vM05, Acz01, Alb92, Deg89, Arr06a, Bel64, Bel93, Bod16, Boh49, BR55, Boh85f, Bok08a, Bol04, Bun79, Cam08, Car04, CW12, CS79b, CG14, Dan89, Dav86, Erl72, FG21, Fin86, Fol96, Fuc85, Gam88, Hel98a, HP85, HP01, HW96, Hol88b, Hol05, HW07a, HW07b, Hoo91, Jam74a, Kai94a, Kle70, Kle10, Kon15, Kum10, Kum11, Lan06, Lau13, Lin07, Lin08, Mar88, MR82e, MR82f, Meh87, Mur94, New09, Nik12, Oku01, Pai79, Pai94, Pat85, PHvMW93, PvM93, Per85, Ros79a, Ros18, Rot87]. Einstein [Sac81, Sac88, Sch49a, Sch49b, Sha87, Sta09d, VW86, Whi96, Whi06, Zin01, vCK17, vD15, Ryn83, Mur90a]. Einstein-Debatte [Hel98a, Arr06a]. Einzeldarst [Sch64c]. einzelnen [UG25]. Electric [Boh14b, Boh15a, Boh23e]. Electrified [Boh13d]. electrodynamical [Dar88]. Electrodynamics [BR50, BR79a, CS79c, CS79f, Ros79d, Tal79]. Electromagnetic [BR79b, CS79g, BO09, BR33]. Electron [Ano23d, Ano39d, BL54, Hei64a, Ric14, Ric16, Tho26, Boh11, Boh12, Boh33c, Gou71, Her23, JVK22, Kra97, Las55, Moy81a, Moy81b, Moy81c, Pai48, PRKH09, Ree05, UG25]. Electronic [Dje07]. Electrons [Ano23d, Eps24, OSWR86, UG26, UGB84, Boh13a, Boh26e, Gao15, Kon00, Kra11a, LRdB^{+}23, Sim14, Ano28, Pia24a, Pia24b, Ano28, LRdB^{+}23]. elektromagnetischen [BR33]. Elektronenstoß [Her23]. Elektrons [UG25]. Elektronteori [Boh11]. Element [Sto97, Rig02, SKST69]. Elementarteilchen [Hei62]. Elementary [BH83a, HEB^{+}80, Hei63b, KH23, BH82, Hei62, MW84]. elementary-particle [BH82]. Elemente [Boh20c, Boh22b, BC23]. Elements [AA89, AA90, Ano23c, Ano39a, Eva96, Jaf72, Kra76, Sim00, Boh20c, Boh22b, Boh22e, BC23]. Elephant [Mac97]. Elettrica [MSB^{+}37]. ellipsoidal [Gus55]. elliptic [DNT08, Nau15]. Elster [Ber15, Ber15]. Embattled [Sal19]. emberi [Boh84]. Emergence [Per13, Pol60, Tan02]. Emission [Som20d, Kon00]. Emits [Ano23d]. Emmet [Los61]. Empirical [Hoo94]. Empiricists [Rös95]. Employed [Ano23c]. emulsions [B^{+}55, LH55]. Encyclopedia [Coh42, CF98, NCM38, NBD^{+}38, Wei39, NCM55, NCM70, N.38, Bal39]. ends [Zan02]. Energia [MSB^{+}37]. energien [Boh55a]. energies [Har83, Har87]. Energiesatz [Sch24]. Energy [Ano36b, Ano39a, Ano39d, BPP39, Boh45b, Boh85h, BPP97, GLR14, Mei62, Kam64, Seg85, Tur46a, Tur46b, Boh55a, Fri54, Kra17, Oku01, PKWF08, Sch24]. Engine [Dje07]. Englisch/ConceptDraw [GRE^{+}01]. English [Cas55, Hei74]. enhed [Blæ85]. Enriched [Bro09]. ensayos [Boh88]. Entanglement [Acz01, Bar07, TWD03]. enter [Das15]. Entering [Den99]. Enterprise [Aas85b, UM92]. Entgegnung [Car83]. entitled [Boh23e]. entre [Bal85, Che92]. entstehenden [HS39b]. Entstehung [HS39a, Pol60]. Entwicklung [Boh28a, Hoy73]. envision [Gau14]. epic [Rob12]. epilepsy [RR00]. Episodes [Eva96, Bra09]. Epistemological [Boh49, CS79b, GRZ99, Ros79a, Kom70]. Epistemology [CS79e, Plo06a, Plo10, Ros79c, Che99a, Plo94, Ros63]. Epoch [Ano39d]. eponymous [KZN^{+}88]. EPR [Bac15, BF94, Bub89, Bub90b, Nik12, OK12, Whi04]. Equal [Ano23d]. Equation [ARK^{+}99, KK03]. Equation/Gaponenko [ARK^{+}99]. Equations [Hen86a, MR82c]. Era [MGG^{+}95]. Erdalkalimetalle [HS39b]. Erforschung [Fis10a, Fis12a, MGG^{+}95]. Ergebnisse [Kon15, KZN^{+}88]. erhielt [CSW97a]. Eric [Gor17]. Erindringer [Cas64]. Erinnerungen [Gam60]. erkendelse [Boh58f, Boh64e]. Erkenntnis [Boh66a, Boh85e, Mal70, Sch64c, Sch64b, Boh58c]. Erkenntnisfragen [Boh59b, Boh59c]. Erklärung [GRE^{+}01]. Ermakov [TM13]. Ermakov-type [TM13]. Ernest [Boh26d, Coh42, dR92, Bad71, Coh88, Coh91, Coh92]. erneut [Kon15]. Ernst [FS06]. Errata [Ano94a]. Ersetzung [UG25]. Ersten [Ano23a, Büh98a]. Erster [ABH^{+}36a]. Erweiterung [SE13]. Erwin [Adl03, Bro36, BD92, MR87, Meh87, RY13]. erzählt [FS06]. especially [DG11]. espionage [Ano95a]. esposizione [DS09]. essai [RC90]. essais [Dar92b]. Essay [CS79d, Hoo94, Kub09, Ros79b, Oku01, RC90, Ros45, Ros61]. Essays [APB96, ABP98, Bab02, Bas71, Boh63b, Boh66a, Boh66b, Boh78, Boh85e, Boh87c, Boh87d, Boh11b, CS79f, CL83, D’A00, GSW95, Kra07, Lan08, Lau07, Mor56, Pai94, Ros79d, Boh22g, Boh24e, Boh34, Boh61a, BBCN86, BS87, Boh88, BB^{+}89, D’A01b, Dör05, PRW55, Nie65, Lin64]. essential [Rig02]. est [Dem15]. Establishment [Kam64]. estructura [KH25b]. Ethics [Log00]. Ettore [GR08]. Études [Rhe91]. Eugene [HH04b]. Europe [Fer68, Fer71, LSW12, Rod19, SDG15]. Evaluations [Moy81a]. Eve [dR92, dR92, Coh88, Coh91, Coh92]. Even [Wei94]. events [B^{+}55]. Everett [Byr10, Whe57]. Evert [Bor58]. Everything [KZN^{+}88]. Evidence [Fri39, Kon83, HS39a, HS39a]. evidential [Fra04]. Evolution [CT65, BV85, Pel47b]. example [Che92]. Exchange [BL03]. Exclusion [Pau55]. exemple [Che92, Dar91b]. exile [Hei84]. existence [HS39b]. existing [Gau14]. expanding [Dea13]. Expands [Aas00b]. expansion [SE13, Som13a]. Expelled [MP01]. Expense [Ros00]. Experiment [Ano39c, CS15, Kon15, Nik12]. Experimental [Kal16, BG24, Cas35, GRZ99, Per06, KZN^{+}88]. Experimentalists [Gea14]. experimentelle [Cas35]. experimentellen [BG24]. Experiments [Ban00, CW12, Gea14, GHW09, Ken85a, MW84, Sha87, vD15]. Explain [Ano23c]. Explained [Gam63]. explanation [Boh39c]. Explanations [Kra06]. Explorers [JFEH15]. explores [FM12]. Explosion [Ano39b]. exposition [Nie15]. exposure [DS09]. extended [Eck14, Ros64]. Extending [Eck15]. extensions [Ano85b, Ano85a]. Extraordinary [Gib19]. eye [Tuc97]. Eyes [Bab02]. eyewitnesses [Kel07, Kel09].

F [Coh42, Mil24, Pye81, Rec86]. F.R.S [Boh37c]. fa [Seg86b]. faces [Kra15c]. Fachwörterbuch [GRE^{+}01]. facing [Ano62b]. facsimiles [Bey49]. Factor [SB72]. Factory [Ano91]. facts [Ant96]. Fail [GLR14]. fails [Tuc97]. Falkenburg [Hel98b]. Fall [Dje07, MGG^{+}95]. family [Byr10]. famous [Ano29, HH04a, WS09]. Fantastic [WD06]. Far [Ano39a, Ros18, Boh64c]. Faraday [Ano68a, Boh32b, Boh65b]. Farm [BL08, Ber96, Luc07]. Fashioned [Duf46]. fast [Büh98a]. faster [POC^{+}10]. Father [Boh64c]. Faust [MGG^{+}95, PSV01, Seg07b, Wal08]. Faustu [Dje07]. Favrholdt [Mai93, Che93, Bro00a]. Faye [Che93, FT19, MP97, Zin19, Cus94a, Hen93].

Fear [Ano02a]. feature [JVK22]. February [Sch21]. Feldgrößen [BR33]. fem [BR^{+}85]. Fermi [Ano39a, GR08, Lau13, Wei94]. Feshbach [Fre90]. Festival [BDR86, HH55]. Festrede [BDR86]. Festschrift [Ano93b, Ber15, KRW05, Ber15, KMP59, Pau01]. Festskrift [HH55]. few [Tel85]. Feynman [Hal17]. fiction [Ant96]. Field [BR50, BR79a, BR79b, CS79c, CS79g, Har28, BO09, BR33, HF15, KP55, Pai48]. Fields [Boh14b, Boh15a, Ber15, Boh23e, Lan55, Sch64b]. Fifth [ABH^{+}40a, ABH^{+}40b, Ano28]. fifties [Roz64b]. fiftieth [Ari35, Hei35]. Fifty [Kae48, Cou64]. Fighter [Kie07]. figure [Ano62a, Cas14, Moo85a]. figures [Cas14]. File [Les94, McM94]. Filme [HEB^{+}80]. Filosoffen [Fav09]. filters [Her55]. Final [Sto97]. Finally [Sto97]. Find [Che99b]. Finding [FG21]. Finland [LM85]. Finn [Far13a, For91, Kub09, Was92, Bor16, Kra16b, Kri91, Ren15]. First [ABH^{+}36b, ABH^{+}37a, Can00, FT19, Hei85b, Kle70, Kra06, MGG^{+}95, Ros79e, She85a, Zin19, Ano23a, Ano57b, Bag10, Bec57, Büh98a, Far13c, FF17a, Hei85c, Rio17, ABH^{+}36a, LSW12, Stu18]. firsthand [Sha87]. Fischer [FS06]. Fisica [MR02, Pet88, Seg76, Seg83, Wei77, HBS96, Dar90, Goo92]. fisicos [Seg87]. fisikaim [Seg86a]. Fission [Ano62c, BHT86, Bar87, BW39a, BW39b, Boh39a, Boh40a, Boh40b, BBBL40, Boh41a, Boh41b, FW67, Gar63b, Gra64, Mei39, MF39c, NBDG40, Sim00, Sim12, Stu85a, Stu94, Tur40, Boh37e, Boh39c, Boh41d, BW85, CSW97b, FW85, GA71, HS39a, Las55, MF39b, Stu13, Whe09, Ano94a, Boh39c, CSW97a, Fav05]. five [BR^{+}85, CW85]. Fixer [MGG^{+}95]. Fixer/Wenske [MGG^{+}95]. fizikata [K^{+}86]. fiziki [PB58]. Fizyka [Boh63c]. flash [Gar63c]. flat [KvG85]. Flavors [Bro09]. Flawed [Bør12]. flight [MMS^{+}19]. Flimsy [McM94]. flow [Gus55]. fluorescent [Boh24g]. Fluorescenzlichtes [Boh24g]. fødsel [BR^{+}85]. Folio [Dar92b]. Folio/essais [Dar92b]. follow [Gra15]. following [Dir28]. Folse [Shi85b, Zin19, FT19, How86, Mac86, MP97, Shi85a]. Fölsing [MGG^{+}95]. FONTANUS [dR92]. Force [Ano39d, UG25]. Forces [Bri65, Pai86]. Ford [Ano62c]. Forderung [UG25]. foredrag [Boh55a, BR^{+}85]. Forenede [Boh64d, Boh50a]. Foreword [Boh46a, Boh60b, Boh07a, Duf46, Hub14, Pei86b, Roz64a]. Forgetting [Sim12]. Forgotten [GR08]. Forholdet [Kal64]. Forklaring [Boh39c]. Forlag [Bad83]. Formalism [Tan04a]. Format [wB90]. Formation [HN64, Gau14, HS39a]. Formed [Boh10, Faf58]. Formula [Ano23c, Cas14, vM02, Som15, Cas14]. Formulation [Eve57, Hen86a, MR82b, Whe57]. Forord [Boh42b, Boh46b, Roz64a]. Forscher [Kon15]. forsker [Kle64]. fortalt [BR64, RRK^{+}64]. Forties [Roz64b]. fortjent [KNA01]. fortolkning [Hei64b]. Forty [Bec57]. Forty-four [Bec57]. Found [Ano39a, Kra14b]. foundation [Boh24c, Das15, MR82e, MR82f]. Foundations [And89, Bey49, BKR85, Bro99, Fre15, GL23, HH76, Hen86a, K^{+}84, LM85, Neu44, NCM70, Tal79, Wei39, Hon81, Mac85, Kal85, Kal96]. Founder [Boh61c, Boh62e]. Four [Ada72, Boh31a, Boh32c, Boh34, Boh61a, Boh88, Boh11b, Bec57, Boh78]. Fourth [ABH^{+}39b, dB86b]. Fra [Sch68, Boh64e, Cas64, Jak13, KvG85, Pai64, Sch64a, Val06, WS09]. Fractals [MGG^{+}95]. Frage [Boh21e, BR33]. Fragen [Paw02]. Fragments [Boh40a, BBBL40, Boh41b, Pal00, HS39a, Las55]. Framework [Fol94, How86, Mac86, Shi85a, Shi85b, Fol85, Mof06]. Franck [Gea14, Jäh15]. Frayn [Kra07, Lan08, Lau07, Log00, Dör05, Ros00]. Free [Dar27, Kra11a]. Freed [Ano39d]. Frees [Ano39b]. freigegeben [Paw02]. fremstillet [HK22, HKK29]. French [Hei86a, Hub14, Ano28, Ano62a, Ano85b, Ano85a, Bal85, BV85, Boh23a, Boh23b, Boh28b, Boh32c, Boh37b, BS38, Boh39f, Boh61b, Boh72, Boh93, Che85b, Che92, Che97, CCJ^{+}34, Dar85, Dem15, Des48, HP81, Kar07, Kum11, LRdB^{+}23, Pat85, Pel47b, RC90, Ros35, Seg84, dBdB21, dB24, dB25, dB63]. Frenkel [ARK^{+}99, Fre96]. frequencies [RR00]. frequency [Dir24, Sch22]. Frequenzbedingung [Sch22]. Fresnel [SWW88]. Friction [HEB^{+}80]. Friedr [Sch64c]. Friedrich [Ano02a, ARK^{+}99, GRE^{+}01, Boh35b]. Friends [Jam67, Kle10, Nie68, BR64, RRK^{+}64, Roz67, Roz85]. Friendship [Gla00, Cou64]. Frisch [CSW97a, CSW97a, Gar63b, Stu13]. Fritz [CSW97a, CSW97a]. fröhliche [Rec09, Rec10]. frühen [Bar89]. Fueled [BSMB99]. Full [Duf46, MW46a, MW07]. function [PRKH09, Zin16]. Fund [Aas85a, Ano62c]. Fundamental [Hen86a, Mil24, Bey49, Boh24d, MR82c]. fundamentalism [Zin15]. Fundamentals [Som24a, Som24b, Som20a, Som24d, Som24c]. fünfzigsten [Hei35]. Further [Beh43c, MSB^{+}37, HS39a, Som13a]. fusion [Rei72]. future [DG11]. fylder [Kra13a]. Fyrrerne [Roz64b]. Fys [Che93]. fysiker [Mor85b, Ano29]. Fysiker-Kongres [Ano29]. Fysikere [Ano29]. fysikken [oMP64]. fysikkens [Sch68]. fysiske [Boh21f, Boh55a].

G [Adl03, Hei74, Hof94, Hub14]. Gaar [Ano29]. Gadamer [vW02a]. Galilei [Büh98b, MGG^{+}95]. Galileo [Büh98b, Cro01, Gam88, Sha87]. Gallery [Pai00]. Game [New09, Fin86, Ree15a]. game-changer [Ree15a]. gamma [Nie55]. gamma-rays [Nie55]. Gamow [Har01, Kuh67, Rei72, Wei68]. Gaponenko [ARK^{+}99]. gases [Las55]. gathered [Ano29]. Gauthier [Pia24a, Pia24b]. Gauthier-Villars [Pia24a, Pia24b]. geb [Sch64c]. Gebaude [Bor58]. Gebieten [Ber15, Sch64c]. Geburtstag [Gam60, vM02, Ber15, Boh35b, KRW05]. Geburtstage [Hei35]. Gedankenexperiment [Kon15, LM85]. Gegen [Lan72]. gegenwärtige [Sch35c, Sch35a, Sch35b]. geheime [Kar05b]. geheimnisvolles [Kar05a]. Geist [SvMS85]. Geistesgeschichte [Kuh67, MA65]. Geisteswelt [Sch54]. Geitel [Ber15, Ber15]. General [Boh62a, Hof47, Hof59, Hof14, NM12, dB13, BJN01, Boh18b, Boh63d, Bus12, Bus20, Jon14, Pau33, RR00, Sho13, Som15, Sud11, BK37, Møl55, San24, Som20b, Wer23]. Generalization [BB05]. generalized [FF78b]. Generation [Hür22]. Generations [Bro00b, Ada72]. Genesis [Bel92b, HK69, Her71, HN70, Kon83, Jon14]. Geneva [Boh55b].

Genius [Dje07, Pai00, Bol04, Gar63c, SM95]. Gennembrudsårene [oER64]. genre [Car04]. Genshi [Bō90, Boh08]. gentle [SM95]. genuine [Sch15]. geometric [Ham05, Pih55]. Geometrical [KZN^{+}88]. geometrisk [Pih55]. geopolitical [Ree15a]. Georg [vW02a]. George [Kuh67, Har01, Rei72, Wei68]. Gerald [Fay07, Par06, Tur06]. German [Eis00, Kar07, Som20b, Wal08, Aas00c, Ano23a, Ano93b, Ano02a, Arr06a, Arr06b, ABH^{+}36a, ABH^{+}37b, ABH^{+}37d, ABH^{+}38a, ABH^{+}39a, Bar89, Ber15, BDR86, Boh20c, Boh21a, Boh21e, Boh22b, Boh22d, Boh23h, Boh23i, Boh23g, Boh23c, BC23, Boh24f, Boh24b, Boh24g, BKS24b, Boh25c, Boh25d, Boh26c, Boh28a, Boh28e, Boh29a, Boh30b, Boh31a, Boh31b, BR33, Boh35b, Boh36b, Boh36d, Boh38e, Boh55e, Boh58c, Boh59b, Boh59c, Boh63e, BH64, Boh66a, Boh85e, BL18b, BL18a, BL19, BH23, Bor23, Bor63, BL63b, BL63a, BG24, Büh98a, Büh98b, Car83, Cas35, Cas02, CSW97a, Dah14, DF14, Eck13a, Ehr23, FH60, FS06, Fis10a, Fis12a, Föp15, Fra63, Gam60, HS39a, HS39b, Hei35, Hei62, HB63, Hei86b, Hei10b, Hei10a, Hel98a, HP85, Her23, HK26, Hou30, Hoy73, HnB74, Hun67, Ish15, Joh13, Kar05a, Kar05b, KRW05, Kon15]. German [Kos23, KZN^{+}88, Kra23a, KH25a, Kra23b, LR23, Lan72, Mor11, Pau33, PHvMW93, Pau85a, PvM93, Pau01, Paw02, Pla23, Pla24a, Pla24b, Pol60, Rec10, Rös92, Sch54, Sch88, Sch22, Sch24, Sch35c, Sch35a, Sch35b, SvMS85, Sim12, Som15, Som18, Som20a, Som20d, Som20c, Som20b, Som24a, Som24d, Som24c, Som24b, Som31, SE13, Som13b, Som13a, UG25, Wei35, vH23, vM02, vW88, vW91, vW02b]. Germany [Bør12, Gla00, Joh13, Set03]. Gershenfeld [ARK^{+}99]. Gesammelte [BDR86]. Geschichte [FH60, Hei75, HnB74, Kar05b, KZN^{+}88, MGG^{+}95, Sch88, Hun67]. Gesellschaft [Som20b]. get [Kar07]. gewesen [Bel86]. Giants [MD67]. Gift [MP01]. Gino [Wal08]. given [Rot00]. gleichnamigen [KZN^{+}88]. glimpse [Kle64, Mor85a]. Glimt [Kle64]. global [Ree15a]. Glückwunschschreiben [Hei86b]. God [Bab02, Ghi05, MB11b]. Gödel [Per85]. Goethe [KvG85]. Goethe-temaer [KvG85]. Goetzberger [MGG^{+}95]. Gogh [Bab02]. Going [Bro36, Sta09b, Dav86]. good [Ber89]. Google [Dje07]. Gospel [KN13, KN11]. Göttingen [Hun85]. Göttinger [GRE^{+}01]. Goudsmit [Lak96]. grand [Kum11]. grande [Ano62a]. Gravity [Zin01]. Great [Ano09, Cro01, Hal09, SKST69, Sha87, Sud11, Tib13, Ano62a, Bre97, Büh98b, Gam88, Kae48, Kum10, Kum11, Lig05, WS09, AFG90, IH07, vW02b]. Greatest [Acz01, Ano39a, Hür22, Bod16, Pop82]. Gregg [Wal08]. grips [Yor75]. Gross [KK03]. Grosse [ARK^{+}99, Dar87, vW02b]. Großforschung [MGG^{+}95]. Ground [FG21, Har83]. Ground-state [Har83]. Group [Far01, Pau55, RRK^{+}64, Rot00]. Growth [Hof47, Hof59, Hof14]. Grünbaum [CL83]. Grundkurs [MGG^{+}95]. grundlag [Boh55a]. Grundlage [GRE^{+}01]. Grundlagen [Som20c, Som20b, Som24a, Som24b, Som20a, Som24d, Som24c]. Grundlaget [Boh24c]. Grundpostulate [Boh23h]. Grundprincipperne [Boh30a]. Guide [Hei96, Rod19, Nie97]. Guthrie [Boh23e].

H [Adl03, Ano43, Coh42, Duf46, FT19, Mil24, Pia24a, Pia24b, Som18, VIR^{+}08, Som18, VIR^{+}08]. H. [Hei74]. Haas [HEB^{+}80, Pia24a, Pia24b]. Habermas [vW02a]. hafnium [Gan17, Sce94]. Hahn [CSW97a, GRE^{+}01, MGG^{+}95, Ano39b, CSW97a, Mor11, Sim12]. Hahn-Meitner-Instituts [MGG^{+}95]. Hahn-Meitner-Instituts/Fölsing [MGG^{+}95]. Hail [Lau33]. Hailed [Ano39a, Ano39d]. Halbleiterbauelemente [OSWR86]. Halbleiterbauelemente/Stitch [OSWR86]. half [Cas83]. Halfway [Bar07]. Hall [BL08, Ber96, Fra55, Luc07]. halvfjerdsårsdagen [HH55]. Halvtreds [Cou64]. halvtredserne [Roz64b]. Hamiltonian [Smi76]. Hamiltonians [GS04]. Handbook [OSWR86]. Hans [Ber15, Car83, vW02a, BR64, HH55, Ber15, Car83]. Hans-Georg [vW02a]. Haphazard [Cas83]. happened [Nau15]. Happens [FM12]. Happy [Rec10]. har [Ano29, KNA01]. Harald [Ang10]. Hardcover [Tan09, Bor13, Gor17, Kra16b]. Hardy [Adl03]. Harmoni [Blæ85, Mor85b]. Harmonious [Dow09]. Harmony [Blæ88, Mor85b, Blæ85, wB90, Aas90a, Fre89]. Hartree [Harxx]. Harvard [Par06, Tur06]. harvest [Bra09]. Hasse [KZN^{+}88]. Hauptversammlung [Som20b]. Havlin [MGG^{+}95]. Hawking [KZN^{+}88, Cro01]. head [Ano62b]. head-and-shoulders [Ano62b]. Heavy [Boh39d, Boh39g, BL54, Bør12, Fri39, Hol80, B^{+}55, Boh40c, VIR^{+}08]. Heelan [Bab02]. hegemony [Cus94b, Lev95]. Heidegger [vW02a]. Heidelberg [wB90, Eis00, Hen86a, Hen86b]. Heilbron [Bus20, Kra16b, Ren15, Far13a]. Heisenberg [Adl03, ARK^{+}99, BDR86, Bor58, Bro95, Car97, Dah02a, Eis00, Fay07, Fer07, For09, GRE^{+}01, HEB^{+}80, Hen86b, HvMW79, HP85, KRW05, Kra96, Kra07, Lak96, MGG^{+}95, OSWR86, Par06, PHvMW93, PvM93, Paw02, Pye81, Rec86, Rec09, Rec10, RS07, Sch00b, Tur06, vMHW85, vM93, vM96, vM99, vM01, vM05, Aas00a, Aas02a, Aas02d, Aas02g, Aas02f, Aas02h, Ano02a, Bel92b, Ber95, BL03, BDR86, Bre97, Bro71, Cam07, Cam09, Cas09, Cav14, CG14, Dah02b, Dir28, Dör05, Dur00, Fav05, GLSC01, Got02, Hei70, Hei84, HP85, Hol05, JVK22, Kle78, KRW05, Lan72, Lan08, Lin07, Lin08, Los61, Mas04, Meh87, PLS02, PHvMW93, PvM93, Paw02, Plo10, Pow02, Rec10, Rhe91, Sch00a, Shi83, Shi86, Tan04b, Yam02, Zan02, vD15, vM02, vW02b]. Heisenberg [Lau07]. Heisenbergs [vM02]. Heisenberg’schen [Lan72]. Heitjans [ARK^{+}99]. held [Ano28, BR^{+}85, CCJ^{+}34, LRdB^{+}23]. Helge [Bor13, Bor16, Bro00b, Sho13, Cla13, Jan17]. helium [Boh13e, Boh15d, Dav16, Har83, Har87]. Helmholtzian [Bis04, Bro03]. Helmut [Hen86a]. Helped [Ano62c, Jak13]. Heme [Chi60]. Hemocyanin [Red63]. Hemoglobin [Alt82, Chi60]. Hendry [Che85a, Wes85, Bro86, dB86a]. Henri [Adl03]. Henrik [Ano57b, Ano63a, Coc63, OBxx]. Henry [How86, Mac86, MP97, Shi85a, Zin19, FF17b, Shi85b]. Hentschel [Wal08]. her [Ano29, CSW97b]. Herausgeg [Sch64c]. Herausgegeben [Eis00]. Hercules [Kel83a]. Here [Ano39c, Pai94, Ano29, Bre97]. heretics [DT18].

Heritage [Che93, Cus94a, Fay91, Hen93]. Herman [Fre90]. Hermann [HEB^{+}80, Pye81, Rec86]. Hermeneutic [Bab02]. Hermeneutics [Kat11b, Dol95, Hee95]. Heroic [Pur18]. Herrn [BL19]. Hertz [Gea14]. Heuristic [Eng03]. Heutigen [Bor58]. Hevesy [Pal15]. Hexeneinmaleins [vM02]. Hidden [Rös94, Far13b]. Higgs [Kra14b]. High [Smi76]. higher [Boh22e]. Highlights [Teu02]. highly [BO09]. Hilfe [BL18b, BL63b]. Hillel [UM92]. Him [CS79d, Ros79b, Ano62a, Ros45, Ros61]. Hintertreppe [Fis10a, Fis12a]. Hiroshima [Lau51, Wyd84]. histoire [Bou49a, Bou49b, HP81, Kar07, Pel47a]. historia [Kar06]. Historian [Jen03]. historians [Kel07, Kel09]. Historical [ABP98, Cas00, Cra02, GL23, Hen86a, Jam74d, Kat11d, Kra07, KN13, Lan08, Lau07, Lev95, Pea08, PA01, Seg85, APB96, Ben00, Cus94b, Dör05, MR82d]. Historicity [Rös94]. historie [Sch68]. historien [NN01]. historier [WS09]. Histories [Pei97d]. historiography [DS09]. History [Ber03, D’A00, DC96, Dar92d, Dur00, Eva96, GSW95, GHW09, Hei64a, Hen81, Hen86a, Her72, HnB74, HR74, Hun74, Kar05b, Kat11b, Kie07, KZN^{+}88, Kra07, Lan08, Pau85b, RW96, Ros00, Sch00a, UM92, WP85, Wei77, dB86a, Ano97, Ano02b, Bag10, Bag11, Bro00b, D’A01b, ET07, Far13b, FH60, Fra04, GA71, HP81, Kar07, KLR13, KHFA67, Rob15, Sch15, Sch68, Sch88, Hun67]. Hit [Ano04b]. Hitachi [K^{+}84]. Hitler [Ber96, Kar05b, Kar07, MP01]. Hitlera [Kar06]. Hitlers [Kar05b]. Hitting [Kow53]. hjalp [Jak13]. hoax [Dur00]. Høffding [Ang10, Fav91, Fay88]. Hoimar [vW02a]. Højdepunkter [Teu02]. holdt [BR^{+}85]. hole [UPG^{+}12]. Holland [Dor79, Jam67]. Holocaust [BJ03]. Holstebro [Jak13]. Holton [Fay07, Par06, Tur06]. Home [Whe92, Wei68]. Hon [Boh37c]. Honner [Cao90]. Honor [Ano23b, Bab02, GSW95]. Honour [CL83]. hopes [McK05]. hos [Aas02h]. House [Duf46]. Houtermans [ARK^{+}99]. Hoyer [Hei75, Stu84]. HQ [KLR13]. HQ-3 [KLR13]. Hugh [Byr10]. humaine [Boh61b, Boh72, Dar92b]. Human [Boh39e, Boh58e, Boh58b, Boh63b, Boh10, Bro36, Duf46, Gla60, Lan59, Lin64, Nie65, Pol58, Sch64b, Ste58, Boh58c, Boh58f, Boh58a, Boh59a, Boh61b, Boh63c, Boh64e, Boh66a, Boh66b, Boh72, Boh85e, Boh87c, Boh87d, Kai94a, Rie60, Yor75]. Humanism [Mar88]. humanistic [SKST69]. humanitarian [Pas03]. Humor [Hal12]. Hund [Jäh15]. Hundred [Bor16, Roq17, AK15]. hundredth [Ano85e, Jac85]. Hungary [Pal15]. hvor [Jak13, KNA01]. Hyde [Lau33]. Hydrogen [Ayl26, Ayl27, Fox80, Lak96, Rig02, Boh13e, Boh14c, Boh15d, Boh15c, Boh18c, Dav16, JVK22]. hydrogen-like [JVK22]. hydrogenic [PKWF08]. hyperfine [Ina16]. Hypothese [UG25]. Hypothesis [Pau85a, PV15, Sch24, UG25, Gau14].

I. [Rabxx]. Idea [Tel69]. Ideal [Boh85c, She72]. Idealization [Tan04a]. Ideals [Hoo91, KN12]. Ideas [D’A00, Hof47, Hof59, Hof14, Sac81, Bre97, D’A01b, HF15, Ryn83]. If [Dav86, Nik08]. Ignorance [Hoo94]. ihr [CSW97a]. ihre [Cas35, Ehr23]. II [ARK^{+}99, Boh23i, Boh66a, Bro99, Eis00, Kal96, Mal70, OSWR86, Rec86, dR92, vMHW85, vM99, Aas99, Aas03, Boh13b, Boh18c, Boh23i, Boh64e, Boh66a, Boh85e, Bub90b, Coh91, Har87, Mar63b, Mel00, Sim12, Ste86b, Wer02, Yag72]. III [Bro95, Dor79, PvM93, vM93, vM01, Boh13c, Boh22e, Byr10, HH76, Ste87]. Ilich [Fre96]. ill [Seg93, Tur06]. Illies [vW02a]. illus [Jam67]. Illustrations [Sha67]. Illustrious [Fer68, Fer71]. Im [Büh98a, Bor58, Kon15]. image [FM12]. imagination [Büh98a]. imagining [HG15]. Imidazole [Chi60]. Immanent [Rad82]. immigrants [Fer68, Fer71]. Impact [BK37, Bro71, PLS02]. impacts [Boh25d]. Impinging [Ano23d]. Implications [Nia98, NM12, dB13]. Importance [Bad71]. impossibility [Büh98a]. impression [And64]. inaugurated [Ano21]. Including [SS98]. incommensurable [KvG85]. Incompatible [Gom06]. Incompleteness [Jam74c, Pat85]. ‘inconceivable’ [Far13d, Far13e]. incongruent [Kav14]. inconsistent [Bar89]. indetermination [Cav14]. index [Sha67]. Indians [Ano22]. Indifference [Kon02]. Individualitat [Kuh67, MA65]. Indivisibility [RC14]. indledende [Boh41c]. indsats [oMP64]. Induced [Boh41a, Boh41d]. induction [Mas04]. industrial [Boh55a]. industriel [Boh55a]. Indvies [Ano21]. inertia [BM55]. Infancy [Kel83a]. Inference [HH76]. infinitely [Ano85d]. influence [McK05]. Influential [Rog10]. Information [Plo06a, Wal08, Whi06]. ingen [MW46b]. initiative [Far13f, Far13g]. Injustice [CSW97b, CSW96]. inkommensurable [KvG85]. inkonsistente [Bar89]. Inner [Hei84, UG25]. Innere [MGG^{+}95]. inneren [UG25]. Innocence [Stu18]. Innovator [Gre00]. Innsbruck [Som24c]. Innsbrucker [Som24c]. Inquiry [Gib19]. ins [MGG^{+}95]. Insights [FM12]. inspirator [Dam85]. inspired [McK05]. Institut [Ano29, CCJ^{+}34]. Institute [Ano21, Ano28, Ano71, Bad83, CCJ^{+}34, Hei96, Lin81, Nie97, Whe81, Bet85, Aas85b, Ano21, Ano29, Hal12, Mot85, Rob79, Rob15, Sch11, Sch21, Smo93]. Instituts/Fölsing [MGG^{+}95]. Integral [Amo78, Sch77, FF78a]. integrity [Hol88a]. Intellectual [BLW19, Fer68, Fer71]. Intellectuals [DF14]. Intellektuelle [DF14]. Intelligence [Bør12, Jon78]. intelligenten [GRE^{+}01]. Intelligibility [Hoo91, Hoo94, Dea06]. Interaction [HEB^{+}80, Boh20a, Kra25]. Interactions [Alt82, Fox80, Ina16]. Interessen [Fri64]. interest [Bus12, Bus20, Gra15, Sho13, Sud11, Fri64]. interference [Her55]. International [Ano28, ABH^{+}37a, ABH^{+}37c, ABH^{+}38b, ABH^{+}39b, ABH^{+}40a, ABH^{+}40b, Boh47, Boh60a, CHR96, K^{+}84, KLR13, KN12, NBD^{+}38, NCM55, CCJ^{+}34, Kat15, NCM70, She72, Wei64, Ano97, ABH^{+}36a, ABH^{+}36b, ABH^{+}37b, ABH^{+}37d, ABH^{+}38a, ABH^{+}39a, Coh42, Wei39]. Internationale [ABH^{+}37b, ABH^{+}39a]. Internationalen [ABH^{+}36a, ABH^{+}37d, ABH^{+}38a]. internationalt [Wei64]. internazionale [ABH^{+}36b]. Interpretation [And89, Ano94a, Dow09, Gom06, Hei62, HB63, How04, Hug89, Jam74b, Jam74c, MA13, RC14, Stu94, Tel80, Ano02b, Cam09, Che85b, Che99a, CSW97b, Foc57, Hei55, Hei64b, Hon81, Kos23, Per06, Stu13, Tan02, vW71, Che85b]. Interpretations [Deg89, BSMB99, Jam74d, Kat11e, Mar54]. Interpreting [Lee06]. Interrogation [Ter94]. Interscience [Jam67]. Intersections [Kat11b, LSW12]. Interview [Boh62b, KRBR62, Wei68]. Interwar [KN12]. Intimidator [Gre00]. intrigue [FB01]. Intrigues [Ano95b]. introducción [Boh88]. Introduces [MB11a]. Introduction [CO07, Duf46, Kay14, Kra15b, Pei86c, Boh32c, Hei20, Nak15, Plo12, Wal08, Bus20]. Introductory [Boh11b, Boh31a, Boh32c, Boh34, Boh41c, Boh61a, Boh88]. invariant [TM13]. Invented [How04]. Invention [Ano43]. inventory [KHFA67]. Investigation [Ano95b, Kra10b, Bar89, FB00]. investigations [Boh41e, Boh42a, Boh46c, RM55]. involvement [LDN99]. Inward [Pai86]. Ion [Fox80, HEB^{+}80]. Ion-Solid [Fox80]. ionization [Ree05]. Ionizing [Den99]. Ions [BL54, Hol80]. irradiation [HS39a, HS39b]. Irreducibility [HH94]. Isador [Ano68b]. ISBN [Ano93a, Bor13, wB90, Bro86, Bus20, Che85a, Cla13, Gor17, Hen86a, Hen86b, Hub14, Kra07, Kri91, Lan08, Red93, Seg93, Tan09, Tur06]. Ishiwara [PB17]. isoelectronic [Har83]. Isotope [Ano39e, ABH^{+}37b, ABH^{+}39a]. Isotopes [ABH^{+}37b, ABH^{+}39a, NBDG40, Boh22c, HS39a, RM55]. Issues [Wes85]. Italian [ABH^{+}36b, A^{+}09, Cam08, Cas14, Cav14, DS09, Hei28, Pet88, Seg76, Val06]. IV [ARK^{+}99, Eis00, dR92, vM96, vM99, vM01, vM05, Coh92, Far01, HH04a, Ste88a]. IV/Teil [ARK^{+}99]. Izbrannye [Boh71b].

J [Amo78, Bus20, Coh42, Dor79, FT19, Hei74, Hei77, Hei85d, Hei86a, How86, Kra97, Mac86, MP97, Mil24, Pia24a, Pia24b, Ren15, Ser58, Stu84, Zin19]. J. [Hei77, Hei85d, Kra97, OBxx, PC06, RB67]. jadrowa [Kar06]. Jaeger [Wal08]. Jaenicke [KZN^{+}88]. Jagdish [Hen86a]. Jahr [Ano93b, Pau85a, Pau01]. Jahre [Ano23a]. Jahren [Boh21a, Boh66a, Boh85e, Mal70]. James [Bro36, Lev95, Ste86a, Ste86b, Ste87, Ste88a, Ste88b]. Jammer [Kuh67]. Jan [Che93, Cus94a, Hen93, MP97, Zin19]. Japan [K^{+}84, Nis37]. Jean [Bou49a, Bou49b, Len48a, Len48b]. Jeans [Mil24]. jedes [UG25]. jeg [KNA01, KNA01]. Jekyll [Lau33]. Jet [Boh09a, Boh09b]. Jewish [DF14]. Jews [Gol87, Lev00, LDN99, SO93, Wer02]. Joachim [vW02a]. Jocular [Bel99a]. Joel [Dje07]. Joensuu [LM85]. John [Bal39, Bro86, Cao90, Coh42, Far13a, Kra16b, N.38, Wei39, Wes85, dB86a, Che85a, Hal17, Sch90]. Jordan [Deg89, JFEH15]. Jørgen [Bro99]. journals [Bey49]. Journey [FR13b]. Journeys [Jon15, WD06]. jüdische [DF14]. Jugend [Rec09, Rec10]. Julius [Ber15, Ber15]. July [KLR13]. jump [MMS^{+}19]. Jumping [Eps24]. Jun [PB17]. June [Boh50a, Boh50d, KLR13, LM85]. Jung [MGG^{+}95]. junge [Gam60]. juni [Boh50a]. Jürgen [vW02a]. Just [Bar99, Rös95].

Kalckar [Bro99]. Kamerlingh [Pia24a, Pia24b]. Kann [Bar89]. Kantian [Bro09]. Kantianism [Che94b]. Kapitza [Ano65, Rub97b]. Kärger [ARK^{+}99]. Karl [Car97, Eis00, Kra96, Rec86, Bro95, Hon81]. karne [Seg86a]. kastes [Jak13]. Kathlerine [Duf46]. Kausalität [Boh36b]. keepers [Van03]. Keeping [Bal14]. kemiske [Boh21f]. Kennedy [Hei86a]. Kepler [Hol88b]. kernekollektiv [KNA01]. Kerner [KZN^{+}88]. Kernphysik [BDR86]. Kernspaltung [CSW97a]. Kernwaffenversuche [Kar05b]. kiadás [Boh84]. kijutsu [Bō90, Boh08]. King [Ber91]. Kings [RS07]. Kirche [MGG^{+}95]. Kirschner [OSWR86]. Kirsten [Dje07]. Kjøbenhavn [Ano29]. klassiske [Pih55]. Klaus [Kuh67, Wal08, vW02a]. kleinsten [Fis10a, Fis12a]. KLM [Jen03]. Knobloch [MGG^{+}95]. Knopf [Sha67]. Know [Lau33, Boh11a, KNA01]. Knowledge [Boh58e, Boh58b, Boh63b, Boh10, Hof47, Hof59, Hof14, Lau33, Nie65, Boh58c, Boh58f, Boh58a, Boh59a, Boh61b, Boh63c, Boh64e, Boh66a, Boh66b, Boh72, Boh85e, Boh87c, Boh87d, Gla60, Lin64, Pol58, Ste58, Rie60, Lan59, Nie65]. Known [CW12, Wei91, Cas35]. Knows [GLSC01]. Knud [LDN99]. København [Bad83, S^{+}06]. Københavns [Ano10]. Koch [HEB^{+}80]. Kokubunji [K^{+}84]. Kommission [ABH^{+}36a, ABH^{+}37d, ABH^{+}38a]. Komplementäre [Wei35]. Komplementaritat [Kuh67, MA65, Boh36b]. Komplementaritetssynspunktet [Ros64]. Kongelige [Ped64]. Kongres [Ano29, Ano29]. konsolideres [Ros64]. Konstitution [Som18]. Kontrast [Mor85b]. Kopenhagen [Aas00c, Paw02, Kar05a]. Kopenhagener [HB63, SvMS85]. Korrespondenz [MA65, Kuh67]. Korrespondenzprinzip [Kra23a]. Kragh [Bor13, Bor16, Cla13, Jan17, Bro00b, Sho13]. Kramers [BG24, Hen81, Kon83, Kra09, Was81, BG24]. Kraus [GRE^{+}01]. kreds [BR64, RRK^{+}64]. Krigens [Boh64b, Boh64a]. Kripke [How18]. Kristalleigenschaften [BL18b, BL63b]. Kristallgitter [BL19, BL18a, BL63a]. Kritik [Car83, Som20d]. Kronologisk [Ano64]. krypton [RM55]. Kuczyński [vW02a]. Kuhn [Boh62b]. Kultur [Boh41c, Dah43]. Kulturbote [Eck13a]. Kumar [Sud11]. Kunz [HEB^{+}80]. kurze [KZN^{+}88]. Kvantenpostulatet [Boh29c]. Kvanteteorien [Hei64b]. kvarkim [Seg86a]. kwanty [Inf33].

L [All56, BL19, Bus20, Far13a, Hei48, Kuh67, Mil24, Mor56, Pia24a, Pia24b, Ren15, Stu84]. L. [Kha89, Pai48]. Labels [Jen03]. Laboratory [K^{+}84, Nav06]. Labyrinth [Hal17]. Lahti [And89]. Lakatos [Car83, Rad82]. Lancaster [Che85a]. Landau [Kha89]. landing [A^{+}09]. Landscape [Ano39c]. Landscapes [SDG15]. landsindustrimödet [Boh55a]. Langage [RC90, Che85b]. Langenscheidts [GRE^{+}01]. Langmuir [Ano22, Eva23]. Language [OP04, Che85b, OPH^{+}09, Rec10, Bor58, RC90]. Langville [Dje07]. L’application [Boh23a]. large [Lib75]. Laser [OSWR86]. lastly [dR92]. late [MSB^{+}37]. Latency [Hol86]. Later [Gla00, Jam74c, SSH14a]. latest [Boh29c]. Latour [BSMB99]. lattice [BL18a, BL19, BL63a]. Laudationes [BDR86]. Laudations [BDR86]. Laureate [SS98]. laureates [Hoi01]. Law [Kra24, Gau14, POC^{+}10, RR00]. Laws [Boh32a, Boh36a]. Layman [Pai94]. Lead [Faf58]. Leaders [Lau51]. leading [Cro01]. leap [Fis10a, Fis12a]. leaps [Ano09]. learn [FJ21b]. learned [Mer04]. leaves [Ham05]. Leben [ARK^{+}99, Boh63e, Dar87, MGG^{+}95, Rec09, Rec10]. Leben-noch [Boh63e]. Lebensgeschichte [KZN^{+}88]. Lecons [Bou49a, Bou49b, Pel47a]. Lecture [Ano23c, Boh61c, Boh65a, Boh65b, Som24c, Boh55a, Boh85g, BR^{+}85, Boh85j, Hei70, Som20c, Boh23e, Boh32b, Bra61, And89]. Lectures [SS98, Boh22a, Boh22d, Boh22f, Boh23b, Boh24b, Boh31a, Boh32c, Boh66a, Boh85e]. Lee [KP55]. left [Ano62b, Inf78]. Legacy [Che93, Fay91, Hen93, Kai94b, Ste89, Cus94a]. Lehmann [HEB^{+}80]. Lehrbuch [HEB^{+}80]. Leninism [Gra85]. Lenzen [Coh42]. Leo [BS50]. Léon [CS79a, Jac12]. Leonard [Coh42]. Leonidovich [Ano65, Rub97b]. less [Cas35]. Lesser [Wei91]. lesson [DDU86]. Let [Kra17]. L’étrange [HP81]. Letter [Ben13, Boh50d, Boh67b, Boh85d, Gol94, Haw94, Pow02, Suv85, Wei94, Boh50a, Boh64d, Boh85k, Dem15, Pow03, Yam02, Hei48]. letterature [Cas14]. Letters [Cho02, Coh88, Coh91, Coh92, KRW05, Mai93, Sch02, Swe02, dR92, vW02a, Fre96, Hei74, Ped64]. leurs [Seg84]. Level [Bus20, Bus12, Jon14, Sho13, Sud11]. Levels [Ano36b]. L’évolution [BV85, Pel47b, Pel47a, Bou49a, Bou49b, Len48a, Len48b, Ser58]. Lewin [Mor11]. Lewis [Ano22]. li [Seg86b]. Liberate [Ano39d]. Liberated [Ano39a]. Library [Hei96, Nie97, Ano68a]. libre [Bou49a, Bou49b, Pel47a]. Licht [Boh63e]. Lichtemission [Som20d]. lidt [KNA01]. Life [Aas90a, Boh32d, wB90, Fre89, Jam67, KZN^{+}88, MGG^{+}95, Nie68, OP04, Rec10, Ste89, Blæ88, Boh33a, Boh33b, Boh63e, BR64, Boh85i, Cro01, Dom06, Eck13b, Fre86, Fre96, Fri67, Hei74, Hei84, Jon85a, McK05, OPH^{+}09, PC06, RRK^{+}64, Roz67, Roz85, Sim96, dB63]. Lifetime [Gin01]. Light [Bis04, Boh32d, Boh33a, Boh33b, Boh85i, Cas02, K^{+}84, Mal15, Som20d, Ste89, Ant15, Boh20a, Boh24g, Bro03, Kon00, Kra25, McK05, WB06, Boh63e]. like [Boh11a, JVK22, MWL^{+}08]. Lindley [Fer07, For09, RS07]. Lindner [MGG^{+}95]. Line [Boh18b, Boh18c, Boh18a, Boh22e, Boh67a, Boh23i, Boh23c]. Line-Spectra [Boh67a]. Lines [Ano23d, Boh14b, Boh15a, Boh23e]. Linienspektren [Boh23c]. l’Institut [Ano28]. Liquid [Ano94a, Stu94]. Liquid-Drop [Ano94a, Stu94]. Lise [Büh98b, CSW97a, KZN^{+}88, Mor11, Ben00, Büh98b, CSW97a, CSW97b, Gar63b, KZN^{+}88, Mei64a, Mei64b, Mor11, Sim96]. listen [Dav86]. Literature [AH13, Can00, Far13a, Kra16b, Ren15, Cas14]. little [KNA01, WS09]. liv [BR64, RRK^{+}64]. Lived [Pai94]. Liverpool [Row09]. Lives [Ano23c, Bre97]. living [Zin15]. Local [Kri94]. Locality [CW12, Sta09d]. localized [MWL^{+}08]. Location [AA89, AA90]. Logic [Ach93, HLS12]. logical [Bub89, Rös95]. Logics [Gar72]. Logik [Bor58]. London [wB90, Par06, Ano43]. Look [Ghi05]. looks [Mei64a, Mei64b]. Loosening [UPG^{+}12]. Lord [Boh37c, MSB^{+}37]. Lorentz [Pia24a, Pia24b, Pau55]. Loss [BL54]. Lösung [MGG^{+}95]. Lösung/Meÿenn [MGG^{+}95]. louder [Dje07]. Louis [Ano62a, Bou49a, Bou49b, Len48a, Len48b, Ser58, Pel47b, Pel47a]. Love [AH13, Aas15a, FF17b, CG14, Mer04, Kra16b, Ren15, Far13a]. Lovett [Kuh67]. Low [KZN^{+}88]. Ltd [K^{+}84, Log00]. ludzka [Boh63c]. lui [Ano62a]. lun [Boh64f]. l’Université [Bou49a, Bou49b]. Lyman [Kon02]. Lys [Kra25].

M [Pia24a, Pia24b, dB13]. M. [Ano62a, SR99]. MA [FMO88]. Mach [HW07b]. Macroscopic [KZN^{+}88]. made [Cli87, Haw11]. Madison [wB90]. Magic [vM02]. Magnetic [Boh14b, Boh15a, Boh23e]. Magneton [Con49]. Majorana [GR08]. Makers [AG02]. makes [Dea06]. Making [Ano39d, Bel99b, Zan02]. Man [Boh55c, Boh55d, Bro73a, Fey33, Gam63, Gam66, Hel85, Kuh67, Lau33, Sha67, Whe85b, ZDSF11, Kha89, Moo66, Moo67, Moo69, Moo85b]. Manchester [Bir62, Bir63, Kat15]. Manhattan [HHW99, Kel07, Kel09, Ree09, Ree11, Ree14, Ree15a]. Manjit [Sud11]. mänskligt [Boh59a]. many [Byr10, Kra15c]. Mapping [PRKH09]. Marcano [dB13]. March [Boh55a, Rot00]. Margenau [Hol86]. Margrethe [Aas15a]. Marks [Sch00b]. Martin [ARK^{+}99, vW02a]. marts [Boh55a]. Marxism [Gra85]. második [Boh84]. Mass [Par06, Tur06, ES74, Oku01]. Massachusetts [FMO88]. Masters [Duf46]. Mat [Che93]. Material [BK37]. materialism [FL10]. Materie [MGG^{+}95]. Materie/Brandmüller [MGG^{+}95]. materja [Inf33]. Mathematical [ARK^{+}99, CDH^{+}06, Harxx, KP55, Bad83]. Mathematician [Jon15]. Mathematicians [Acz01]. Mathematics [Ber03, Boh56, BH03, GSW95, Pau85b, Plo11, Ber15, FF91]. Mathematik [Ber15]. Matrix [Hen86a, MR82b, Sta09e, Kle78]. Matsui [Fre90]. Matter [ARK^{+}99, Bar07, Boh13d, BL54, KZN^{+}88, Ric14, Ric16, Ryn83, Sta93, Sta09g, Boh20a, Boh21f, Boh48b, Boh60c, Inf33, Inf34, Kra25, Pai86, Sac81, Sta04, Wei70]. Matter/Friedrich [ARK^{+}99]. matters [Rei72]. Matthias [Kra07, Lan08, Lau07]. Max [Adl03, Bro36, Fis10a, Fis12a, Kuh67, CS79b, Fis10a, Fis12a, FL10, KMP59, McK05, RH00, Ros79a]. Max-Planck-Festschrift [KMP59]. Maxwell [Boh31c]. May [Ano97, Far13f, Far13g]. Mayenn [Kra96]. Meaning [Bar07, Bec18, DT18, Duf46, Mar58, Bag04, Ish15, Ish17, MW46a, MW07, PB17]. Means [VW86]. meant [She85b]. Measurability [BR79b, CS79g, BR33]. measure [Bag04]. Measurement [Mar58, HF15, PKWF08, WZ83]. Measurements [BR50, BR79a, CS79c, Mar63a, Mar63b, Pai48]. Measures [Lau13]. mécanique [BD92, Dar91b, RC90, Boh37b]. Mechanical [Boh35a, Bai13, EPR35]. Mechanics [Bel86, BB05, BW00a, BW00b, Bro86, Che85a, Che93, Eve57, Fay91, Fre15, GGK02, Ghi05, GL23, Har28, Hei64a, Hel98a, Hen81, Hen86a, Hol86, Hug89, Jam74d, Jam89, K^{+}84, Kat86, Kel85, Lev95, Los61, MA13, MR82b, MR00, Plo06a, Pur18, RC14, Sta93, Sta09e, Sta09g, Tel80, Tri80, VW86, dB86a, vdW67, Amu14, Bel87, BD92, Boh25b, Boh25a, Boh26c, Boh26b, Boh35d, Boh37b, Bun79, Cam09, Cus94b, Dar27, Das15, DJ15, DNT08, ET07, FJ21a, FJ21b, Foc57, Gao15, Gru14, Hee95, Hei10b, Hen84, Hun85, Jam66, Kle78, MR82a, MR82c, MR87, Mlo14, Nau15, Pau33, Per06, Pih55, RC90, Sta04, SSH14b, Tan02, WB06, Wes85, vdW68, vdW07, Sch35c, Sch35a, Sch35b, Hen93, Kuh67]. Mechanik [GRE^{+}01, Boh26c]. Mechanism [BW39b, Boh41a, Boh41d, FW67, Whe09, BW85]. med [KvG85]. Medal [Ano65, Ano68b]. medarbejdere [BR64, RRK^{+}64]. Medd [Che93]. Medical [Boh52, Pod10a]. Medicine [Cra02]. Mediziner [HEB^{+}80]. Mediziner/Hermann [HEB^{+}80]. meet [Lan06]. Meeting [Aas02a, Aas02g, Aas02f, Ano39a, AG81, Bar07, CCJ^{+}34, Kra07, Lan08, Lau07, Aas02d, Ano29, Ano02a, Boh55a, Dör05, Gla00, Kar05a, LRdB^{+}23, Paw02]. Meetings [Jon85b]. Meets [How18, Rei84]. megismerés [Boh84]. Mehra [Hen86a]. Meitner [CSW97a, MGG^{+}95, Mor11, Ben00, Büh98b, CSW97a, CSW97b, Gar63b, KZN^{+}88, Mei64a, Mei64b, Sim96, Stu13]. Meitner/Hawking [KZN^{+}88]. mekanik [Pih55, Boh25b]. melding [Amu14, Gru14, Mlo14, SSH14b]. mellem [Kra25]. meltdown [Byr10]. members [Lin55]. memoir [Fra85, Opp63b]. Memoirs [Gam60, Nat99, Roz98, SSSS95]. Memorabilia [Ano68a]. Memorandum [Boh44, FP40a]. Memorial [Boh61c, Boh63a, Bra61, FW60, Rei72, Wei63]. Memories [McM94, Nie63, Wei85c, Tel85, Cas64, Pai64, Sch64a]. Memory [Rub97a]. Men [Cli87, SKST69, Van03]. menace [Yor75]. Mendel [Mur90a, Ryn83]. menneskelig [Boh58f, Boh64e]. menschliche [Boh58c, Boh66a, Boh85e, Mal70, Sch64c, Sch64b]. Mental [Wal08]. Mentality [Wal08]. Mercury [Ano23d]. Mermin [Wal08]. meson [B^{+}55]. Mesonic [Faf58]. Message [GLSC01]. Meßbarkeit [BR33]. messenger [Eck13a]. metafora [Cas14]. metafore [Dar90]. metafori [Goo92, Pet88]. Metal [Hol80]. Metallernes [Boh11]. metals [Boh11, HS39b]. metaphor [Cas14, Sch15]. metaphorical [Pet88]. Metaphors [Mac95]. Metaphysics [Los61]. Method [Boh09a, Boh09b, RW96, Tib13, Boh33c, Her55, Per13]. Methodological [Los61]. Methods [Har28]. Methuen [Log00]. Meyenn [Bro95, Eis00, GRE^{+}01, HEB^{+}80, Pye81, Rec86, ARK^{+}99, MGG^{+}95, Car97]. Meyer [Dje07, Kuh67]. Meyer-Abich [Kuh67]. MGM [Rei84]. Mi [Seg86a]. Mi-karne [Seg86a]. miao [Boh64f]. Michael [Kra07, Kuh67, Log00, Dör05, Lan08, Lau07]. microscope [Tan04b]. mid [MMS^{+}19]. mid-flight [MMS^{+}19]. mighty [FM12]. migration [Fer68, Fer71]. Miles [Ano39c]. Millennium [Den99]. Millikan [Pia24a, Pia24b]. Mind [Hür22, Lau33, Sta93, Sta04, Sta09g, Tuc97, WD06]. Minder [Pai64, Sch64a]. ming [Seg86b]. mirror [HW96]. Mirrors [FM12]. Missing [Joh13]. Mission [Aas20, Aas07]. missionaries [Hei85a]. mistake [Bod16]. Misunderstood [Rös95]. mittels [HS39b]. Mittelstaedt [And89]. Mix [Ano01]. Mladjenovic [ARK^{+}99]. Mobilization [Joh13]. Model [AF17, Ano94a, Bor13, Bur18, Cla13, Eps24, Gar63a, GR08, HN70, Jan17, Kra12a, Nis67, Pod10b, San24, Sch13, Sho13, Som24a, Som24d, Som24c, Som24b, Stu94, Wer23, Yag72, Amu14, Ano13, Bar89, Ber89, BH64, BL18a, BL19, BH23, BL63a, CDH^{+}06, Eck14, Gao15, Gru14, HG15, Ina16, KP55, Kon00, Kra14a, Mlo14, Pal15, Pla24a, Pla24b, Pol60, Ree05, Som15, Som18, Som13b, Som13a, SSH05, SSH14a, SSH14b, Urt06, Yag64, dBdB21]. modèle [dBdB21]. Modelers [Lak96]. Modeling [ARK^{+}99]. Modeling/Origin [ARK^{+}99]. Modell [Som15]. Modellen [BH23]. Modelles [Som18]. Models [Ano22, BL63b, Hei64a, Hug90, Joh13, Lak96, AI15, Bai13, Boh14a, BL18b, Föp15, Kra23b, Sch15]. Modern [And89, BKP09, Boh31c, Bor58, CO07, Gib19, Jon15, LM85, RS07, TWD03, dB86a, Bag04, Boh24c, BR^{+}85, Bra09, FL10, Inf34, Seg80, Seg87, Seg07a, Seg84]. moderne [Boh24c, BR^{+}85, MGG^{+}95]. modernes [Seg84]. modernim [Seg86a]. Modernizing [Joh13]. modernos [Seg87]. Modifications [Hen86a, MR82b]. Molecular [Joh13, McW73, Amu14, BRB12, Gru14, Kra23b, Mlo14, SSH05, SSH14a, SSH14b, ZDSF11]. Molecules [Boh13a, Boh13b, Boh13c, Boh62c, Boh63f, Szu05, BH23, Kon15, Urt06]. Molekeln [BH23]. Molekülen [Kon15]. Molekülmodelle [Kra23b]. Moll [KvG85]. Molluscan [Red63]. moment [CG14]. Moments [BM55, Bag11, Büh98b]. Monday [Rot00]. monographs [Sch64b]. Moore [Sha67, Fri67, Gam66, Kuh67, Hel85]. moral [Gol87]. Morris [Bal39, Coh42, N.38, Wei39]. Moseley [FF17b, Hei74, Jaf72]. Most [Rog10, Haw11]. Mother [FF17b]. Motion [Tho26, Dar27]. Motivations [Zin01]. Mottelson [KNA01]. move [Gao15]. Movement [Kie07]. Moving [Boh13d]. Mr. [BL19, Gam39, Gam42]. multiple [Byr10]. multiplication [Pei39]. Multipole [Nie55]. Münchener [BDR86]. Munich [BDR86]. Muons [Faf58]. Murdoch [Bub90a, Dar91a, Fay90]. Murphy [Bro36, Bro36]. Muse [Can00]. must [BO09]. mutual [Byr10]. My [Rot00, dR92, Coh88, Coh91, Coh92, Hei15]. Myers [KZN^{+}88]. Mysteries [Ghi05]. mysterious [Kar05a]. Mystery [Acz01, FR13b, Gla00, Sch02, FR98]. mystical [Deg89]. Mysticism [Hon82a]. Myth [Che97]. Mythe [Che97]. Mythology [How04].

N [Amo78, Ano43, wB90, Dje07, McC34, Pia24a, Pia24b, SR99, Som18, Wal08, Som18]. N. [CS79c, KH25b, Pai48, Sch77]. Nabo [NN01]. nach [Gam60, KZN^{+}88, Som15]. Nachprüfung [BG24]. Nachweis [HS39a, HS39b, HS39a]. Nagaoka [Ina16, Yag64, Yag72]. Nagel [Coh42]. Nagging [Paw02]. naissance [BD92]. Names [Sto97]. Nanocrystals [ARK^{+}99, UPG^{+}12]. Nanocrystals/Gershenfeld [ARK^{+}99]. narrative [Com56]. national [Bal14, Ano57b]. nationer [Boh50a, Boh64d]. Nations [Suv85, Boh50a, Boh50d, Boh50e, Boh64d, Boh67b, Boh85d, Boh85k]. Natur [Hel98b]. Natural [Boh39e, Boh52, Boh56, GSW95, McW73, Dol95, FH60, Sch64b, Teu02]. naturaleza [Boh88]. Naturbeschreibung [Boh31a, Boh29a, Boh30b, Som31]. naturbeskrivelse [Boh58d, Boh29b]. Naturbeskrivelsen [Boh30a]. Nature [Ano09, ARK^{+}99, Boh30a, Boh31a, Boh32c, Boh34, Boh61a, Boh78, Boh88, Boh11b, Cao90, Hon82a, Plo10, Sud11, Boh29b, Boh29a, Boh30b, Boh58d, Boh87b, BB87, Dea06, GL03, Hon88, Kum10, Kum11, Sce94, Boh87a, Bro36, McC34, Som31, Bal14].

Naturforscher [Dar87]. naturvidenskab [Teu02]. Naturwissenschaft [Sch64c, FH60]. naucnye [Boh71b]. nauki [Inf33]. Nazi [Cho02, Gla00, Joh13, MP01]. Near [Fox80]. need [Aas00c]. Negative [Faf58]. Neighbor [NN01]. nel [DS09]. nella [Seg76, Seg83]. nelle [Cas14]. Nelson [Boh37c, AF17, DNT08]. Networking [Whe88]. neue [Sch24]. neuen [Ano02a]. neuere [Boh28a, Hou30]. Neugebauer [Jon15]. Neurath [Bal39, Coh42, N.38, Wei39]. Neutrality [KN12, LSW12]. Neutrino [GLR14, Pau85a, Ros35, Nav06]. Neutrinohypothese [Pau85a]. neutrinos [Fra04]. Neutron [Ano36a, Boh36c, Bro71, Fri39, Kae39, Lau13, Boh31e, HS39a, HS39b, MF39b, Pei39, Boh36d]. Neutronen [HS39b]. Neutronenbestrahlung [HS39a]. Neutroneneinfang [Boh36d]. Neutronernes [Boh31e]. Neutrons [MF39a]. Newer [Boh41e, Boh42a, Boh46c]. News [Ano36a, Ano36b, Ant15, Stu85a, Yam02]. Newton [Büh98b]. Newtons [Büh98b]. Next [FM12]. Niaz [dB13]. nicht [CSW97a]. Nichtlineare [MGG^{+}95]. Niels [Aas88, Aas90a, Adl03, Ano23a, Ano29, Ano62b, Bad83, Bal39, B^{+}55, BM55, Bor58, wB90, Bro36, Bro00a, Büh98a, Cao90, Che93, Coh42, Cuf24, Dar90, Dar92b, Duf46, Ehr23, Far13a, FMO88, For91, Fra55, Fre89, Gam60, Gla60, Goo92, Gus55, HH55, Hei96, Hei86b, Hei10a, Hen92, Her55, Hof94, How86, How99, Hub14, KP55, Kat86, Kra07, Kri91, Kuh67, Lan59, Lan08, Las55, Lau07, Lau33, LH55, Lin64, Lin81, Mac86, Mac95, Mai93, Mal70, Mer89, MA65, Mil24, Møl55, Mor56, N.38, Nie97, Nie55, Nis37, Pia23, Pih55, Pol58, RM55, Ren15, Rie60, Sch64c, Sch64b, Sch54, Seg93, Shi85a, Shi85b, Smo93, Ste58, VGEK87, Was92, Wei39, Whe81, vW02a, Aas85a, Aas85b, Aas88, Aas90b, Aas99, Aas00b, Aas02d, Aas02c, Aas02e]. Niels [Aas02h, Aas03, Aas07, AH13, Aas15a, Aas20, Ach93, Ang10, Ano23a, Ano29, Ano57b, Ano62a, Ano62c, Ano63a, Ano63b, Ano68a, Ano68b, Ano71, Ano85c, Ano85d, Ano85e, Ano85f, Ano93a, Ano94b, Ano95b, Anoxx, Ano10, Ano13, Ari35, BO09, Bec57, BZ85, Ben00, Ber89, Bet85, Bis04, Blæ85, Blæ88, Blo63, BS50, Boh63a, BR64, BR^{+}85, BBCN86, BS87, BB87, BB^{+}89, Boh05, BA05, Boh07b, BHK^{+}08, Boh11a, BB05, Bri58, Bro03, Bro73b, Bro99, Büh98a, Bun85, Bun88, Bun92, CS15, Cas85, Che94a, Che94b, Che97, CM85, Cla13, Coc63, CS79e, CS79d, CS79f, CS79g, D’A01a, Dai96, Dam85, DDU86, Dir64, Dir67, Dol95, DG11, Dor79, Dör05, Ehr23, El’85, Eng03, Enz02b, Fav92, Fav94, Fav05, Fav09, Fay91, FF94, FF17a, FMO88, Fis10b, Fis12b]. Niels [Fol85, Fol95, Fra63, FK85, Fre86, Fri67, Fuc85, Gam60, Gam63, Gar63a, Gom06, Gow85, Gra15, HH55, Harxx, HF15, Hei15, Hei20, Hei35, Hei63a, Hei10a, Hol88a, Hon81, Hon82a, Hon82b, Hon88, How79, How18, HH94, Jac85, Jam67, Jan17, Jon86, Kai94b, KvG85, Kan96, Kat11b, Kel83b, Kle64, Kle10, KN12, KZN^{+}88, Kra79, KNA01, Kra12a, Kra13b, Kra17, Kub08, KRBR62, McE72, McK05, Meh87, Mig85, MB11a, Moo66, Moo67, Moo69, Moo85b, Moo85a, Mor85b, Mot85, Mot08, Mur87, Mur89, Mur90b, Nie65, Nie63, Opp63c, Opp63a, Opp63b, Opp64, OBxx, OP04, OPH^{+}09, Pai85a, Pai91, Pas03, Pau45, PRW55, PB58, Pau94, Ped64, Pei63, Pei08, Per13, Per21, Pet63, Pet85, Pet88, oMP64, Kam64, Pih64, Plo12, Red93, Rio17, Rob79, Rob15]. Niels [RC90, RH00, Rös92, Ros45, Ros61, Ros63, RRK^{+}64, Ros79c, Ros79b, Ros79d, Roz67, Roz85, Roz98, Rub97a, Rub97b, Ruh92, Sch68, Sch54, Sch11, Sch21, SvMS85, SK75, She72, She85a, She86, SM95, SM08, Ste89, Ste86a, Ste86b, Ste87, Ste88a, Ste88b, Stu84, Stu85b, Suv85, Tel69, Ter94, Tuc97, VW86, Was81, Wei14, Wei63, Wei64, Wei85a, Wei85b, Whe63, Whe85b, Wil85, Yam02, Zin16, Zin19, dB63, del81, vD15, vFH76, All56, Ber91, Bor13, Bro00a, Bub90a, Bus20, Che93, Cus94a, Dar87, Dar91a, Dar92a, FT19, Fay90, Fre90, Fri67, Gam66, Hei86a, Hel85, Hen93, Kle93, Kub09, Kuh67, MP97, Mor56, Nat99, Nie68, Sha67, Sho13, Stu92, Wil92]. Nielsen [Dor79]. Nielson [Stu84]. niemieckich [Kar06]. night [LDN99]. Nil [Moo69, PB58]. Nils [Fre89, K^{+}86]. Nil’sa [Ter94]. Nil’su [PB58]. nine [Van03]. nineteenth [D’A01b, D’A00, Dor79, SDG15]. No [Enz02a, Kra76, Swe02, Wei39, Yam02, Hel98b]. Nobel [NN01, Rec10, Aas02c, Ano62c, Ano93b, Boh65a, Boh85j, Cha09, CSW96, CSW97a, CSW97b, Cra02, Hoi01, NN01, SS98, SR99]. Nobelpreis [CSW97a, Rec09, Rec10, Ano93b]. Nobelpriser [NN01]. noch [Boh63e]. noget [Boh11a]. Nogle [Boh41c]. Non [Har28, Gus55, Kav14]. non-classical [Kav14]. Non-Coulomb [Har28]. non-ellipsoidal [Gus55]. Nondispersing [MGG09, YRBD09]. None [Duf46, MW46a, MW46b, MW07]. Nonlocality [CW12, Ant96]. nonseparable [Kel58]. Nørlund [Aas15a]. North [Adl03, Dor79, Jam67]. North-Holland [Dor79, Jam67]. Norwegian [Aas02h, Bør12]. Note [Boh12, Dir24, Ben00, Pop82]. Notes [RS07]. Nothing [Mor95]. notion [Bou49a, Bou49b, Des48, Len48a, Len48b, Pel47b, Pel47a, Ser58]. Notions [Boh48a, Boh50c]. nous [Ano62a]. nouveau [Boh28b]. November [FMO88, KRBR62, Far13d, Far13e]. Nowe [Inf33]. noyaux [Boh39f, CCJ^{+}34]. nucléaire [Boh37b, BS38, Kar07]. Nuclear [Aas90b, Ano36a, Ano36b, Ano94a, ARK^{+}99, BHT86, Bet37, BGS95, Boh36c, Boh38b, Boh38c, Boh38d, BW39b, BPP39, Boh39a, Boh40b, Boh61c, Boh62e, BPR86, BPP97, Bør12, Bri65, Can00, Cas02, CT65, For91, Gra64, Jen00, Kar05b, KZN^{+}88, Kri91, MGG^{+}95, MF39a, Mei62, Mla98, Mot08, NBDG40, Pei40, Pei97a, Pei08, Ros72, Seg85, Sim00, Sim12, Stu85b, Stu94, Stu18, Tur40, Was92, Whe63, Ada72, And73, Ano95a, Bet91, Bey49, BDR86, Boh37d, Boh55a, BW85, BBCN86, Byr10, CSW97b, Dar88, ES74, Far13c, Gow85, GA71, Kae48, Kar07, KNA01, LH55, Opp64, Ree15a, Stu13, Bet37, Boh37b, BS38, Stu79]. Nuclear-Fission [Sim12]. Nuclei [Boh13c, Boh36e, BK37, Boh37f, Boh39d, Boh39g, Fri39, Gam29, Boh13a, Boh35c, Boh36d, Boh37e, Boh39f, Boh39c, Boh40c, Boh41e, Boh42a, Boh45c, Boh46c, BM55, Boh85m, CCJ^{+}34]. Nucleon [GRE^{+}01]. Nucleon/Sube [GRE^{+}01]. Nucleus [Boh13b, Boh62a, FR13b, Kow53, Kra12b, Mei39, MF39c, BZ85, Boh38e, Boh63d, Fri64, Gus55, Hou30]. Number [N.38, Boh22e, Dar55, Kon00]. Numbering [Jaf72]. Numbers [Lin24, Dar92d, Lib75, Ott03]. numerical [Har23]. nuova [Dar90, Goo92, Pet88]. nur [CSW97a]. NY [Hub14]. Nyere [Boh41e, Boh42a, Boh46c].

O [Cav14, Som18, Bro95, Som18]. O. [Gar63b]. O.M [Boh26d, Boh37c]. Obituary [Ano63b, Boh37c]. Objection [Jam74c]. objections [Kra11a]. Objectives [Som20a, Som20b, Som20c]. Objectivity [BKP09, Hoo91, Mof06]. Objekten [GRE^{+}01]. Obscure [Che99b]. Observation [VIR^{+}08, UPG^{+}12]. obtenir [Kar07]. Occasion [CS79d, CS79f, Mor56, Ros79b, Ros79d, B^{+}55, BM55, Fra55, Gus55, Her55, KP55, Las55, LH55, Møl55, Nie55, PRW55, Pih55, RM55, Ros45, Ros61]. Occasional [Nie65]. Occupation [Sch21]. Occupied [Sch11]. October [Ano28, Ano57b, CCJ^{+}34, DDU86, Far01, HH55, Hei35, CS79d, Ros45, Ros61, Ros79b, SO93]. octobre [Ano28, CCJ^{+}34]. Odd [Tre99]. Odds [Lau13]. œuvre [dB63]. off [Ano09]. offentlige [BR^{+}85]. Office [Kra07, Lan08]. Oktober [Hei35, HH55]. Old [Bro09, Møl55, Jäh15, Kav14, Kra13a]. Oleson [Fre90]. Olson [HEB^{+}80]. Om [Aas02e, Boh14c, Boh64c, Jak13, Boh11a, NN01, Boh23d, Boh35c, Boh37d, Boh37e, Kra25]. Omdanndelser [Boh45c]. Omdannelser [Boh41e, Boh42a, Boh46c]. omkring [Fri64]. Omnes [Dar92b]. Oncology [Den99]. One [AK15, Gib19, Lau33, MW46a, MW07, Roq17, Sta09b, Wyd84, Ano85e, FF78a, GL03, Bor16, BG24, MW46b, Duf46]. one-dimensional [FF78a]. one-hundredth [Ano85e]. Online [CF98]. Only [Lau33, Ano39d, Boh13b, CSW97a]. Onnes [Pia24a, Pia24b]. Ontology [How79, RC14]. Opbygning [Boh35c]. Open [Boh50d, Boh50b, Boh67b, Boh85c, Boh85d, Boh85k, Suv85, Aas07, BBCN86, BB^{+}89, Boh50a, Boh64d]. opened [Ano29]. Opening [Sch00b]. Operation [Ter94]. operator [HLS12]. Operatsiia [Ter94]. opinion [Hei10b]. Oppenheimer [RB67, Haw94, OBxx, PC06, Val06]. opportunity [Boh45b, Boh85h]. Opposition [Kra11c]. Optical [ARK^{+}99, Kal16]. Optoelektronische [OSWR86]. Orbit [Faf58, Har83, Har87, PRKH09]. Orbits [Lin24, JVK22, Nau15]. order [Nie55]. Origin [Ano94a, ARK^{+}99, Stu94]. originally [Bey49]. Origins [Meh87, Moy81b, Hol88b, Nak15]. Ørsted [Kra15a]. Orthodox [Ban00]. oscillator [Kon00]. Oscillators [Hen81]. ostinato [KvG85]. Other [Rös95, MGG^{+}95, Mor85a, Rei72]. Otto [Bal39, Coh42, CSW97a, GRE^{+}01, Mor11, N.38, Wei39, CSW97a, Jon15, Sim12]. Our [Hof47, Hof59, Hof14, Hoo94, BL19]. outrageous [Büh98a]. ove [Seg86a]. ove-ad [Seg86a]. ove-tagliyotehe [Seg86a]. oversigt [Ano64]. overtook [Far13c]. Overview [Kat11c, Ano64, Boh31a, Boh32c, Wal08]. Oxford [Ano93a, Bor13, Bus20, Cla13, Gor17, Hen92, Hub14, Kra16b, Red93, Seg93]. Oxygen [Faf58, Kon15]. Oxygenation [Hol80].

P [Amo78, HH04b, Hei86a, Mer85, Pia24a, Pia24b]. P.R.S [Boh26d]. Paarläufe [FS06]. Pacifist [Sal19]. packets [MGG09, MWL^{+}08, YRBD09]. page [Ano09]. PageRank [Dje07]. pages [Eis00, Kra16b, Par06, RS07]. pair [FS06]. Pais [Ano93a, Hen92, Red93, Seg93, Ber91, Dar92a, Kle93, Stu92]. Palestine [Ros18]. Pandora [Van03]. Paper [Tri80, Was81, Whi04, ZDSF11, Nak15, Pai48]. paperback [Bus20, Hub14, Kra07]. Papers [Aas02f, BFF98, CS79a, D’A01a, Kra07, Lin55, Bel87, BV85, BS87, DP97, FB01, Haw11, OBxx, PD97, Rabxx]. Pappband [wB90]. paradossi [Dar90, Goo92, Pet88]. Paradox [Erl72, Tri80, Bel64, Mer85, Vil86, Fuc85]. Paradoxes [Joh13, Mac95, Pet88]. Parallelism [Jon86]. paranormal [Deg89]. Paris [wB90]. Park [Ano10]. Part [Car97, MR87, Mil24, vM96, vM99, vM01, vM05, Boh18b, Boh18c, Boh22e, Eis00, Boh13a, Boh13b, Boh13c, Coh91, Coh92, Har28, Mar63a, Mar63b]. Particle [AMH06, ARK^{+}99, BH83a, BH83b, Hen86a, Kai94b, Kal16, dB13, BH82, Gau14, NM12, dBdB21]. Particles [Boh13d, BK37, Hei63b, Nia98, BO09, Boh48b, Boh60c, Fis10a, Fis12a, Hei62, Sla85]. Partisans [Sal19]. parts [Hen86a]. Pascal [Hof94, Hub14, Ruh92]. Paschen [Boh35b, Boh35b]. Passing [Boh13d]. Passion [Jon08]. pathways [Inf33]. Patrick [Bab02]. Paul [OSWR86, Bok08a, Kle10]. Pauli [All56, ARK^{+}99, Bel86, Bro86, Bro95, Car97, Che85a, Eis00, GRE^{+}01, HEB^{+}80, Hen86b, HvMW79, Kra96, MGG^{+}95, Mor56, OSWR86, Pye81, Rec86, Wes85, dB86a, vMHW85, vM93, vM96, vM99, vM01, vM05, Deg89, EvM94, Enz02a, FW60, Hen84, HP85, Mas04, PHvMW93, Wei85c]. Pauli-Jung-Dialog [MGG^{+}95]. pbk [Lan08]. Peace [Ano57a, Ano57b, Bec57, Sal19, Hei70, Kra17]. Peaceful [Sal19]. Peacock [ARK^{+}99]. Peierls [DP97, Lee07, Lee09, PD97]. Peircean [Mal15]. Pekka [And89]. Pelseneer [Bou49a, Bou49b, Len48a, Len48b, Ser58]. Penetrating [BL54]. penetration [Boh48b, Boh60c]. pensée [Bou49a, Bou49b, Pel47a, RC90]. People [BA05, Boh07b]. Period [Hen09a, KN12]. Periodic [AA89, AA90, Kra06, Boh16, BC23, Boh31b, Kra85]. periodische [Boh31b]. periodisches [BC23]. Perović [Cuf24]. perpetual [Yor75]. Personaggi [Seg76, Seg83]. Personal [Kie07, Wei85c, Bet91, Com56, Fra85, Kai94a]. Personalities [Gin01, Seg76]. personality [Fra63]. Persönlichkeit [Fra63]. Perspective [Cas00, Jam74d, Pea08, Seg85, TWD03, Bet91, Gra15]. Perspectives [BKP09, FF17a, Rio17, Boh64a, ET07, GRZ99, FT19, Zin19]. perspektiver [Boh64b, Boh64a]. perturbation [Bor23]. Perturbations [San24]. Peter [And89, Bad83, Lin81, Whe81]. Petersen [Boh62b]. Petr [Rub97b]. Petruccioli [Dar90, Goo92, Mac95]. Pharmazeuten [HEB^{+}80]. Phase [Dah02b, Kle70, Rad82, BH23, Dah02a, FF78a, SWW88]. phase-integral [FF78a]. Phasenbeziehungen [BH23]. phenomena [Boh12, Boh93, Cas85]. phénomène [Bou49a, Bou49b, Len48a, Len48b, Pel47b, Pel47a, Ser58]. phénomènes [Boh32c, Boh93]. Phenomenological [Sho08]. Phenomenology [Ste88b]. Phenomenon [Ano39b, Boh39a, Dom06, MW84, Pel47b]. Philanthropy [Aas90b, For91, Kri91, Was92]. Philos [Hel98b, MP97]. Philosoph [Rös92].

Philosopher [Fol96, Kat86, Rös92, Sch49a, Sch49b, Cam09, CS15, Fav09, Kle64, Mor85b]. Philosopher-Physicist [Kat86]. Philosopher-Scientist [Sch49a, Sch49b]. Philosophical [Aas88, Boh87c, Che93, Dow09, HW07b, Hoo94, Kat11d, Mar58, Mer89, Rös94, Zin01, BB87, Dol95, Fav92, Lee06, Mai93, BDR86]. Philosophie [Wei35, BD92, Bun92, Che97, Hei10a]. Philosophy [Aas02b, Ano93a, AG81, Bab02, Ber91, Bis04, Boh39e, Boh52, Boh56, Bub90a, Bun85, Cao90, Che93, CL83, CW85, CHR96, CF98, DC96, FT19, Fay90, FF94, FF17a, GSW95, Gor17, GHW09, HH76, Hen92, How86, Jam74d, Kle93, Kra16a, Mac86, Mal15, McE72, Mur90b, Pet63, Red93, Rio17, Shi83, Shi85a, Shi85b, Shi86, Stu92, UM92, Zin19, vFH76, Bag04, Bel87, BD92, Bro03, Bun88, Che94a, Che97, Dar91a, Dar92a, EvM94, ET07, Fol85, Fol95, Hei10a, Hon81, Hon82b, Hon88, Jac12, Kot85, MP97, Mur87, Mur89, Pai91, Pet85, Plo06b, Sce16, Seg93, Wei35, Tan09]. Philosphers [Acz01]. Photo [Boh38c, Boh38d]. Photo-Effects [Boh38c, Boh38d]. photodissociation [Bon10]. Photon [Sta09a]. photons [Ano28]. Photovoltaik [MGG^{+}95]. Photovoltaik/Lindner [MGG^{+}95]. Physical [Boh35a, Boh55c, Boh55d, Fri39, GL23, Hal09, HH76, Hel98a, Hoo91, Jon86, Pau85b, PA01, Som20b, VW86, Boh21f, Boh22b, Boh35d, Boh55a, EPR35, Pai86, Pel47b, ZDSF11]. physiciens [Seg84]. Physicist [Cha09, Cho02, Kat86, Rös92, Cam09, Eck13a, Harxx, Hei74, Kha89, KZN^{+}88, Mor85b, Pas03, Rot00, Wil70]. Physicists [Ano39a, Ano39c, Hür22, Kuh67, Pod10a, Sal19, Wal08, vW02b, Ada72, Ano29, Bre97, Cli65, Cli87, Cro01, Gam88, HH04a, Seg80, Seg84, Seg87, Seg07a, dR85, Ano29]. Physics [Aas90b, Aas02b, Acz01, ABP98, And89, Ano39b, Ano93a, Anoxx, ARK^{+}99, Bar07, Bec18, Ber91, Bet37, Bis04, BKP09, Boh31c, Boh38b, Boh39b, Boh47, Boh49, Boh58b, Boh59b, Boh59c, Boh60a, Boh63b, Boh81, Boh85b, BPR86, Boh10, BBSR69, Bør12, BH83a, BH83b, Bro99, Bro00a, Can00, CL83, CSW97b, D’A00, FT19, Fav99, FF17a, Fre90, Gam61a, Gam62a, Gam85, GSW95, Gin01, Gla00, Gla60, GHW09, Hal12, Hen86a, Hen92, Hof94, How79, HH94, Hub14, Kai94b, Kal85, Kal96, Kel83a, Kle93, KZN^{+}88, Kri91, Kuh67, LM85, Lan59, Lin64, Mar58, McW73, Mla98, MB11a, Mor56, Mot08, Mur90b, Nie97, New09, Nie65, Pei91, Per21, Plo11, Pod10a, Pol58, Red93, Ree09, Rio17, SS98, Sch64b, SvMS85, Ste58, Stu85b, Stu92, Stu18, Tan09, Wes85, Whe63, Whe85a]. Physics [Zin19, vCK17, Aas15a, APB96, Ano62a, Deg89, Bag04, Bag10, Bal14, BZ85, BJN01, Ber15, Bey49, BDR86, Boh22a, Boh38a, BS38, Boh58c, Boh58f, Boh58a, Boh59a, Boh61b, Boh63c, Boh64e, Boh66a, Boh66b, Boh72, Boh85e, BKR85, BR^{+}85, Boh87c, Boh87d, Bra09, Bro03, BH82, Bro00b, Büh98b, Cas85, Cas09, Cli87, Con62, D’A01b, Dar92a, Dea13, Dom06, El’85, EvM94, ET07, FF91, FMO88, FW60, Gam61b, Gam62b, GRZ99, Haw11, Hei62, HBS96, Hei10a, Hon88, Jac12, Kae48, KLR13, Kra13b, KHFA67, LRdB^{+}23, MR02, Meh75, Mig85, Mur87, Mur89, Pai85a, Pai91, PRW55, PB58, Pei08, Pet88, oMP64, Plo06b, Ree11, Ree14, Ree15a, Rie60, Rob15, Ruh92, Sac88, Sch68, Seg93, SKST69, Seg76, Seg07b, Set03, Sha87, Sim96, Str11, Stu79, Vol88, Wal08, WP85, Wei77, Ano28]. Physics [Bor58, CCJ^{+}34, Cuf24, For91, Hal12, KZN^{+}88, All56, Bis04, Bub90a, Cao90, DT18, Fay90, Mac95, Mur90a, Wal08, Was92, Dar91a, Hei96, FL10]. Physics/Grosse [ARK^{+}99]. Physics/Peacock [ARK^{+}99]. Physics/Thorsen [KZN^{+}88]. Physik [Ber15, Büh98b, GRE^{+}01, HEB^{+}80, Hei62, Hei10a, MGG^{+}95, SvMS85]. Physik-Lehrbuch [HEB^{+}80]. Physik/Vilenkin/Shellard [MGG^{+}95]. physikalische [Hel98a, Som20b]. physikalischen [Boh22b]. Physiker [Rös92, vW02b]. Physique [Bar23, BS38, Boh61b, Boh72, Ano28, Ano62a, Bou49a, Bou49b, CCJ^{+}34, Kum11, Len48a, Len48b, LRdB^{+}23, Pel47b, Pel47a, Rhe91, Ser58, Dar92b]. Pickering [Hen86a]. picture [JVK22]. pieces [KvG85]. Pioneer [Ano62c, Bec57]. Piotr [Ano65]. Pitaevskii [KK03]. place [A^{+}09, Dol95]. Plainte [Ros35]. Planck [Adl03, Bro36, Fis10a, Fis12a, Hen86a, KMP59, Ano23c, Bro15, CG14, Fis10a, Fis12a, HN70, Kle66, MR82e, MR82f, Pet14, Rhe91]. Planck-type [Pet14]. Plaschko [MGG^{+}95]. Play [Aas00a, MB11b]. Playful [Hal12]. Plays [Ros00]. Plotnitsky [Tan09]. Plum [Nav15]. Pluralism [Bok08b]. Podolsky [Bel64, Erl72, Fuc85, LM85, Alb92]. Poincaré [Adl03]. Point [Ano23c, Ayl26, Ayl27, Lem31, Dam85]. Poland [Rot00]. polarimetric [Her55]. Polarisation [Boh21e, Boh24g]. Polarized [OSWR86]. polarons [UPG^{+}12]. Pole [Ano04b]. Polish [Boh63c, Inf33]. Political [Aas03, Aas99, Boh05, Far13f, Far13g, Kle35, Moo85a]. Politics [BA05, Hab69, Sim12, Far13b, Fav05, Inf78, Jac12, LSW12, Lev86, She72, Wal08]. politik [Fav05]. Polity [Hen92, Pai91, Seg93, Ano93a, Red93, Stu92, Ber91, Kle93, Dar92a]. Polymath [Har01]. Popular [KN13, BDR86, Nie15]. popularised [KN11]. Popularization [BA05, Boh07b]. Portrait [CW85, Pai00, Ano62b]. portraits [Far01]. Portrays [Aas00a]. Portuguese [Seg87]. Position [Boh55c, Boh55d]. positions [Dom06]. positive [Boh13a]. positivism [Lan72]. Positivismus [Lan72]. Possibilities [Kae39]. Possible [Ano22]. postgraduate [Jon14]. postscript [Aas02e]. postulat [Boh28b]. Postulate [Boh75a, Tel80, Boh28b, Boh28a, Boh28c, Boh28d, Boh29c, Boh87a, Dot08]. postulates [Boh23h, Boh24d, Mil24]. Postwar [CSW97b, Pai85b]. posvjascennyj [PB58]. potential [Gus55, Ree05]. potentials [BRB12, FF78b, FF78a]. potere [Val06]. pour [Kar07]. Power [Kae39, KZN^{+}88, Bet91, Val06]. Pp [Ano93a, Bad83, Che85a, Dor79, Hen86b, Hen92, Kra07, Kri91, Red93, Bal39, Bor13, Bro86, Bus20, Cla13, Cuf24, Duf46, Gor17, Hen86a, Hub14, Jam67, Lan08, Log00, N.38, Pia24a, Pia24b, Seg93, Sha67, Tan09, Tur06]. pp. [Mai93]. Practical [Fey33]. Practice [Sta09b, Set10]. Practising [Jäh15]. Pragmatic [Mar88]. Pragmatics [Ste86b]. Pragmatism [MA13, Mal15]. Praise [Eng03]. précédés [Boh32c]. precedidos [Boh88]. Prediction [DG11, GLR14, Sce94]. predicts [Ree05]. Preface [Aas15b, Boh42b]. Preis [CSW97a]. Preliminary [Kra10b]. Prelude [Has15]. Presence [Hol80]. Present [Hei63b, Hof47, Hof59, Hof14, Tri80, Los11, Sch35c, Sch35a, Sch35b]. Presentation [Ano57b, KH23, SS98, HK22, HKK29]. Presented [Lau33]. President [Boh44]. Press [Ano93a, Bor13, Bus20, Cla13, Cuf24, Hen86a, Hen92, Hub14, Kra16b, Kri91, Par06, Red93, RS07, Seg93, Tur06]. Price [Gor17, CSW97a]. Primacy [How18]. Primal [KZN^{+}88]. Primas [MGG^{+}95]. primer [WB06]. Primitifs [Bou49a, Bou49b, Len48a, Len48b, Pel47b, Pel47a, Ser58]. primitives [Pel47b]. Primo [ABH^{+}36b]. Princeton [RS07]. principe [Rhe91]. Principle [AF17, Bel92b, Cla88, Dor79, Kal16, Sta09c, Wer23, BAV01, Das15, Dir24, DNT08, EFM01, El’85, Jäh15, Kot85, Lib75, Pau55, Pet14, Ryn15, Ban00, Dah02a, Dah02b, Kom70, Kra23a, Sch00a, Sch22]. Principles [Los61, Set03, She85a, Boh30b, Fal98, Hel98b, Pau33, Meh87, Som20b]. Prinzipien [Boh30b, Pau33]. Prior [Kat11e]. private [Lee09]. Prize [Rec10, Ano57a, Ano93b, Boh85j, CSW97a, CSW97b, NN01, SR99]. Prizes [Cra02, Aas02c]. prób [Kar06]. Probability [HH76, Plo10, SWW88]. probably [Bre97]. Problem [Boh39b, Boh85b, Bor58, How79, Jon86, HP01]. Problemer [Boh24a]. problèmes [Boh23a]. Problems [Boh49, Boh75b, Gin01, Mar58, Boh23a, Boh24a, Kat15, Møl55, Set03]. procedures [Kos23]. Proceedings [Bor16, AK15, AG81, FMO88, Stu79, K^{+}84, Roq17, Ano97]. Process [Ano39d, Gar63b, CSW97b]. Produce [Ano23d, FM12]. Produces [Ano39d]. Production [HEB^{+}80, Her55]. Production/Olson [HEB^{+}80]. Products [Mei39, MF39c, MF39b]. Prof. [Lau33]. professées [Bou49a, Bou49b, Pel47a]. Professor [ARK^{+}99, Hei86b, B^{+}55, BM55, Fra55, Gus55, Hei70, Hei86b, Her55, KP55, Las55, LH55, Møl55, Nie55, Nis37, Pih55, RM55, Wei68]. Professors [Ano39a]. Profound [Bet91]. program [Car83]. Programm [Car83]. programmes [Het95]. progress [Gra15, Het95]. project [Ano97, Cas02, HHW99, Kel07, Kel09, Ree09, Ree11, Ree14, Ree15a, Sim12]. Projection [Hoo91, Tel80]. Proliferation [FR13a]. prolongements [Ano85b, Ano85a]. Properties [ARK^{+}99, Boh36e, BL63b, FP40a, FP40b, Boh21f, Boh22b, Boh31e, Boh35c, BL18b, CCJ^{+}34]. prophet [She86]. proposes [Kon15]. propriétés [CCJ^{+}34]. prospects [Boh64b]. Protactinium [BW39a]. Proved [Acz01]. PSA [AG81]. Psycho [Jon86]. Psycho-Physical [Jon86]. Psychoanalysis [CL83]. psychological [Mof06]. Public [Duf46, BJN01, BR^{+}85, MW46a, MW07]. Publication [GR08, Bal14, Suv85, Ano93b, Pau01]. Publications [Bar23]. published [Hub14]. Publishers [wB90]. Publishing [Bro86, Che85a, Log00, Bal14]. puddings [Nav15]. Pursuing [KN12]. Pursuit [Ach93]. puts [Dur00].

Quanta [Boh75b, Cuf24, Kle66, Per21, vCK17, Boh23a, Boh28b, HP81, Inf33, Inf34, dB25]. quantal [FF78b]. Quantenmechanik [GRE^{+}01, Hel98a, Hei10b, Sch35c, Sch35a, Sch35b].

Quantenmechanik/Thomas/Weise [GRE^{+}01]. Quantenphysik [GRE^{+}01, Boh59b, Boh59c]. Quantenpostulat [Boh28a]. Quantenspringerei [Kra11b]. Quantensprung [Fis10a, Fis12a]. Quantenstatistik [Pla24a, Pla24b]. Quantentheorie [Boh23h, Boh23i, BKS24b, Boh31b, Ehr23, Hei62, Kuh67, MA65, Pol60, Som24a, Som24d, Som24c, Som24b, Boh21e, Boh23h, Bor23, HB63, Hun67, Sch88]. Quantenwelt [Kon15, Lan72, Arr06b]. Quanthentheorie [Hou30]. quantique [BD92, Dar91b, Kum11, Rhe91, RC90]. quantisation [BRB12]. Quantitative [LH55, Har26, Kom70]. Quantities [BR79b, CS79g, BR33]. quantization [BH11, Bon10, FF78b, FF78a, GS04, Ham05, Kle35, KK03, TM13]. quantized [BO09]. Quantum [AH13, AMH06, Ano09, Arr06b, Ayl26, Ayl27, Bar07, Bas71, Bec18, Bel86, Bel99b, Bis04, Boh15b, Boh18b, Boh18c, Boh18a, Boh22e, Boh32b, Boh35a, Boh35d, Boh39f, BR50, Boh59b, Boh59c, Boh62d, Boh65b, Boh67a, Boh75a, BR79a, Bok08b, BW00a, BW00b, Bor13, Bro86, Bro99, Cao90, Che85a, Che93, Cla13, CS79c, CS79f, Cus94b, Dar92d, ET07, Eve57, Far13a, Fay91, Fre15, GGK02, Gam85, Gar72, Ghi05, GL23, GHW09, Hal12, Hal17, Hee95, Hei64a, Hei13, HB63, Hel98a, Hen81, Hen86a, Hen93, Hen09a, Her71, Hof47, Hof59, Hof14, Hol86, HW07a, How79, Hug89, HR74, Hun74, Hür22, Jam74d, Jam89, Jan17, JFEH15, Joh13, Jon08, Kal85, Kal96, Kal16, K^{+}84, Kat11a, Kat86, Kra12a, Kra16a, Kra16b, Kuh67, Kum10, Lem31, Lin24, Log00, Los61, MA13, Mar54, Mar63a]. Quantum [Mar63b, MR00, Mil24, MB11a, Nia98, OK12, Pea08, Plo06a, Plo10, Plo11, Pur18, Ren15, RC14, Ros79d, Ros79e, Sch90, Sho13, Som24a, Som24d, Som24c, Som24b, Sta93, Sta09g, Tal79, Tan04a, Tel80, TWD03, Tre99, Tri80, VW86, Was81, Wei85a, Wes85, Whe57, WZ83, Whe88, Whi06, Zin01, dB86a, vCK17, vFH76, vdW67, Amu14, Deg89, Bag04, Bag11, Bai13, Bel87, BA06, Ber91, BD92, Boh71a, Boh85a, Boh16, Boh21e, Boh23h, Boh23i, Boh23a, Boh24d, BKS24b, BKS24a, Boh28c, Boh28d, Boh29c, Boh29a, Boh31b, Boh36a, Boh38e, Boh67c, BKR85, Boh87a, Bol04, Bro03, Bub89, Bun79, Cam09, Cas09, Che99a, Cli65, Cli87, Con62, CG14, DT18, Dar88, Das15, DDU86, Dot08, Ehr23, EPR35, El’85, Fal98, FJ21a, FJ21b, Fin86, Fis10a, Fis12a]. quantum [Foc57, Gao15, GRZ99, Gru14, HF15, Haw11, Hei15, Hei55, Hei62, Hel98b, Hen84, HLS12, HG15, Hon81, Hon88, Hou30, Hun67, Hun85, Ish15, Ish17, Jäh15, Jam66, Jea19, KLR13, Kav14, Kel58, Kon15, Kon00, Kra13c, KHFA67, Lan55, Lan72, Lee06, Lib75, Mac85, MR82a, MR82c, MR82d, MR82e, MR82f, Mig85, MMS^{+}19, Mlo14, MB11b, Nie15, Pai48, Pai79, PB17, PV15, Per06, PKWF08, Pla24a, Pla24b, Pol60, Pop82, RC90, Sce94, Sch15, Sch88, Sch35c, Sch35a, Sch35b, Set10, Sie13, Sim14, Sta04, Str11, SSH14b, Tan02, UPG^{+}12, WB06, Wei14, Wei85b, Whi96, Zin15, Zin16, dB25, vdW68, vdW07, Ano09, BV09, Boh28a, Boh39f, Bor23, Hei64b, HBS96, Hei10b, Kum11, Lev95, MR02, Wal08, Bro00b, Sud11]. Quantum-Mechanical [Boh35a]. Quantum-Theoretical [Plo10]. quark [Seg83]. Quarks [Hen86a, Seg80, Seg84, Seg87, Seg07a]. Quarters [Aas00b]. Quatre [Boh32c].

Quelques [Des48]. Quest [Bec18, Bus12, DT18, KZN^{+}88, Com56, Mof06, Ste11]. Question [BR79b, CS79g, Boh21e, BR33]. Questioners [Kuh67, Cli65]. Questions [Boh59b, Boh59c, Pal00, And73, Paw02]. Qui [Dem15].

R [Gar63b, Pia24a, Pia24b, Mer85]. Rabbi [Ano04a]. Rabi [Ano68b, Rabxx]. racconti [A^{+}09]. Race [Duf46, Lau51, Bal14, Far13c]. Radder [Car83]. Radiation [Boh15b, Boh62d, Den99, HEB^{+}80, HN70, Mil24, Pod10a, Was81, BKS24b, BKS24a, Boh67c, Sch24]. Radiative [Faf58]. Radioactive [FP40a, FP40b]. Radioactivity [Kra12b, vH23]. Radioaktivität [vH23]. Radiological [dR85]. Radium [Kae48]. radius [PKWF08, UPG^{+}12]. Række [Ano29]. raggi [Seg83]. Rahner [Hon81]. raios [Seg87]. Raiser [Aas85a]. rallying [Dam85]. Raman [SR99]. Ramsauer [Jäh15]. ran [Boh64f]. Range [BBBL40, Boh41b]. Rankings [Dje07]. Ranotgen [Seg86a]. rapportant [Ano85a]. rapporto [ABH^{+}36b]. Rapports [CCJ^{+}34, LRdB^{+}23, Ano28]. Rare [Eva96, FF17b]. Rath [GRE^{+}01]. Rational [Nia98, Bar89]. Ray [Faf58, Boh14a, Bra61, BC23, Cos22]. Rayner [Harxx]. rayons [Seg84]. Rays [Kae39, Kay14, Nia98, Boh21e, Nie55, Seg80, Seg84, Seg87, Seg07a]. razvitie [PB58]. Re [Whi04]. Re-Assessment [Whi04]. Reaches [Ano43]. reacting [POC^{+}10]. Reaction [BHT86, MF39a, And73, Som13b]. Reactions [BPP39, Boh39g, BPP97, Pei40, Pei97a, Boh37d]. Reader [Hof47, Hof59, Hof14]. readership [Jon14, Sho13, Sud11]. Reading [Plo06b, Tan09]. Reaktion [Som13b]. Real [Bec18, DT18, vCK17, FM12, FR98, UPG^{+}12]. realidad [MR02]. realisiert [Kon15]. Realism [CHR96, Dai96, Dan89, Fav94, Fol94, Fol96, How79, Hug90, Kri94, Mac94, Fin86]. Realist [Hen93, Sho08, Che93, Fay91]. realitè [Rhe91, Kum11, RC90]. realities [WD06]. Reality [Boh35a, Bor58, Bus12, Dow09, Hal17, Hel98a, New09, VW86, Boh35d, Cas83, EPR35, Kum10, Kum11, MR02, RC90, Ste11, Ano09, Sud11]. Realization [Boh59b, Boh59c, MWL^{+}08, Wyk08]. realized [Kon15]. Reallism [Eng03]. really [Fra04]. realm [Kae48]. Reappraisal [HW07a]. reason [BLW19]. Reasonable [Plo11]. Reasoned [Tib13]. Rebuilding [Fre15]. Recalls [Lau51]. receive [Bec57, CSW97b]. received [CSW97a]. Recently [Aas02f, Boh10, CSW97b]. Reception [Hen86a, Kra10b, MR82c, Pal15]. Rechenberg [Hen86a]. Recherches [dB25]. Recognition [VW86]. Recollections [ABP98, Ros79e, APB96, Hei84, Kha89, Pei85]. Reconciliation [Ano22]. reconfirmed [Kon15]. reconsidered [Fay88]. reconsidering [BV09]. Reconstructing [OK12]. Reconstruction [Nia98, NM12, dB13]. recordings [Ber96]. Redirecting [Aas90b, Kri91, For91, Was92]. Redirection [Aas85b]. Reductionism [Bok08b]. Reexamining [Bok08b]. reflection [Pau55]. Reflections [Fuc85, Gin01, Shi83, Shi86, Fre85, Inf78]. Refocusing [Dot08]. Refuge [DF14, Dah14]. Refugee [Seg85]. Refutation [Fuc85]. Refutations [Kal16].

Regime [MP01]. Region [BPP39, BPP97]. Reich [BLW19]. Reichenbach [Gar72]. Reidel [Bro86, Che85a, Wes85]. Reihe [Sch64c]. relating [Aas02d, Aas02g]. Relation [AA89, AA90, BBBL40, Boh41b, Bok08b, Kal16, Oku01, dB86b]. relations [BH23, Kos23, vD15]. Relationship [Ang10, Mor11, Roz98, Fav91, Fay88, Kal64, Nat99]. Relationships [KZN^{+}88]. Relative [Eve57, Whe57]. Relativistic [JVK22, Ter08]. Relativitätskongreß [Pau01]. Relativity [Pau01, Møl55, San24, Wer23]. Release [Aas02d, Aas02g]. Released [Aas02f, CSW97b, Paw02]. Relevance [Aas02b, Per06]. Religion [RW96]. Remark [Her72]. Remarks [BK37, Boh63a, Bri58, Fav91, Boh41c, BL19]. Remind [Duf46]. reminiscence [vW85]. Reminiscences [Bar87, Blo63, Boh61c, Boh62e, Hei85e, Pai85b, Ros72]. ren [Seg86b]. renormalizable [KP55]. Renormalized [AF17]. Replacement [UG25]. Reply [OK12]. Report [Ano39a, Duf46, ABH^{+}36a, ABH^{+}36b, ABH^{+}37a, ABH^{+}37c, ABH^{+}37d, ABH^{+}38a, ABH^{+}38b, ABH^{+}39b, ABH^{+}40a, ABH^{+}40b, KHFA67, MW46a, MW07, Kae45, Mil24]. reported [Bey49]. Reports [KRW05, Ano28, Boh85l, Dir28, LRdB^{+}23, CCJ^{+}34]. Representation [Mor11]. requirement [UG25]. Rescue [Gol87, Lev00, Wer02, LDN99, SO93]. Research [Boh52, FF17b, K^{+}84, Kam64, Boh24c, Boh85l, Het95, KZN^{+}88, Mof06, dB25]. researcher [Kle64]. researchers [Kon15]. Resistance [Kie07, Lev00, LDN99]. Resisting [Kra11c]. resistivity [Fra55]. resolution [Gau14, PRKH09]. Resonance [Boh38d, Boh39a]. Response [Aas85b, Alb92, BF94, Whi04, Bub89, Bub90b, Car83, Tan04b, BL19]. Rest [Ano93c]. resulting [HS39b]. results [Kon15, KZN^{+}88]. Rethinking [Lan06]. retrait [Pat85]. retrospect [Stu79]. Retrospectives [She85b]. reveal [CSW97b]. Revealed [Kae45]. reverse [MMS^{+}19]. Review [Aas88, Aas90a, Amo78, And89, Ano09, Bad83, Bal39, Bel86, Bis04, Bor58, Bor13, Bor16, wB90, Bro36, Bro86, Bro95, Bro99, Bro00b, Bro00a, Bub90a, Bus20, Cao90, Car97, Che85a, Cla13, Coh42, Cuf24, Cus94a, DT18, Dor79, Duf46, FT19, Far13a, Fay90, Fay07, Fer07, For91, For09, Fre89, Fre90, Gam66, Gla60, Goo92, Gor17, Hei96, Hei86a, Hei75, Hel85, Hen86a, Hen86b, Hen92, Hen93, Hof94, Hol00, How86, Hub14, Jam67, Jan17, Kat11a, Kle93, Kra96, Kra07, Kra16b, Kri91, Kub09, Kuh67, Lan59, Lan08, Lau07, Len48a, Len48b, Lev95, Lin64, Lin81, Log00, Mac86, Mac95, Mai93, Mal70, McC34, Mer89, Mil24, Mor56, Mur90a, N.38, Nat99, Nie65, Nie68, Par06, Pia23, Pia24a, Pia24b, Pol58, Pye81, Rec86, Ren15, Rie60, Ryn83]. Review [Sch64c, Sch64b, Seg93, Ser58, Sha67, Shi85a, Shi85b, Ste58, Stu84, Stu92, Tan09, Tur06, Was92, Wei39, Wes85, Whe81, Wil92, Zin19, dB13, Jon14, Som31]. Reviews [Ano93a, Bar23, BDR86, Dje07, Red93, Wal08]. Revisited [AH13, AMH06, Har83, Har87, Nik12, Sim14, SSH05, Wal19, dB86b, Kra16b, Far13a, Ren15]. Revisiting [BA06, BL08, Luc07, McK05]. Reviving [Szu05]. Revolution [Kae48, Pea08, Bel99b, Bol04]. Revolutionary [Lau33, Sie13].

Revolutionized [Hal17]. Rezension [Eis00, Som31]. Rhetoric [Bel93]. rhetorical [Car04, Tuc97]. Richard [RS07, Hal17]. Richardson [KZN^{+}88]. Right [Boh37c, Mei62, Bun79]. rings [WB06]. ripples [SWW88]. riron [Bō90, Boh08]. Rise [Aas90b, Dje07, For91, Hen86a, Kri91, Was92, MR82e, MR82f, MR87]. Risø [Kam64]. Ritz [Kon02]. ro [JVK22]. ro-vibrational [JVK22]. road [Bet91]. Roads [Mei62]. Robert [RB67, GSW95]. Robertson [Bad83, Whe81, Lin81]. rock [Das15]. Rocked [Ano39e]. Roland [Dar92b]. Role [Hen81, Nis67, AI15, Dol95, Hoy73]. Roles [BW00a, BW00b]. Rolle [Hoy73]. roman [Kum11]. Röntgen [MGG^{+}95, Kay14]. Rontgenspectra [Cos22]. Röntgenspektren [BC23]. Roosevelt [Boh44, Dem15]. Roots [Hol70, Dar88, Per13]. Röseberg [Dar87]. Rosen [LM85, Alb92, Bel64, Erl72, Fuc85]. Rosen-Paradox [Fuc85]. Rosenfeld [All56, Mor56, Stu84, CS79a, Dar91b, Jac12, Pai48, Tal79]. rotating [BM55, Gus55]. Rotation [HEB^{+}80]. Rotation/Haas [HEB^{+}80]. route [Hei15]. Routledge [CF98]. Royal [Mai93, Ped64]. Rozental [Fri67, Jam67, Nat99, Nie68]. Rud [Dor79, Stu84]. Rudinger [Boh62b]. Rudolf [Bal39, Coh42, Lee09, N.38]. Rudolph [Wei39]. Ruhla [Hof94, Hub14]. Rules [Hen09b]. runs [FS06]. Rupp [vD15]. Russell [Bal39, Coh42, N.38, Wei39]. Russian [Boh71b, Moo69, PB58, Ter94]. Ruth [Fri67, Gam66, Hel85, Kuh67, Sha67, Mor11]. Rutherford [Boh26d, Boh37c, Kra14c, Lak96, Pia24a, Pia24b, Pol60, dR92, Bad71, Bir62, Bir63, Boh61c, Bra61, Bro73b, Bur18, Cam99, Coh88, Coh91, Coh92, Hei81, Kat15, Kra12b, MSB^{+}37, Oli85, Pei88, Pei97b, Pei97c, Pei10, Pod10b, Pol60, Wer23]. Rydberg [RR00, VIR^{+}08].

S [Boh62b, Bus12, Fri67, GSW95, Jam67, Nie68, Sch64c, PB58, Sal19]. S.J. [Bab02]. Sachs [Mur90a, Ryn83]. Saha [SR99]. Sakharov [Mar88]. Sallhofer [Lak96]. samarbejde [Koc64, Wei64]. samfund [Pih64]. samler [Fri64]. samlet [Ano29]. Sammlungen [Sch64c]. Sandro [Dar90, Goo92, Mac95]. Saturnian [Yag64, Yag72]. Sauerstoff [Kon15]. Sauerstoff-Molekülen [Kon15]. Saul [How18]. Say [Ano39e]. says [Ano62a]. sbornik [PB58]. Scaling [Kra23a, Urt06]. Scan [Yam02]. Scandinavia [DF14]. Scandinavian [Sch54]. Scattering [Boh40a, LR23]. Scerri [Gor17]. Schalenbau [Kra23a]. Schicksal [ARK^{+}99]. Schicksal/Kärger [ARK^{+}99]. schlägt [Kon15]. Schödinger [Dom06]. Scholder [vW02a]. Schriften [BDR86]. Schrödinger [Adl03, ARK^{+}99, Bro36, Lak96, Deg89, BD92, Che92, DJ15, Kra11b, Kra13c, MR87, Meh87, Nau15, Per06, Plo10, RY13, Shi83, Shi86, Sta09f, Ste87, Tri80, Vil86]. Schumacher [HEB^{+}80]. Schweden [Dah14]. Schweitzer [Amo78]. Science [Aas85b, Aas90b, Aas00c, AG02, Ano10, AG81, Bab02, Boh45d, Boh46a, Boh55c, Boh55d, Boh61c, Boh62e, Boh07a, wB90, Bro36, Che93, CW85, CHR96, Cra02, DC96, Den99, Dje07, Dur00, Fay07, For91, GGK02, Gam66, Gib19, Hab69, HH76, Hel85, Jon08, Jon15, Kae39, Kra07, Kri91, Kuh67, Lan08, Lev86, MGG^{+}95, NBD^{+}38, Opp50, Pai00, PSV01, Par06, PA01, Rec10, RW96, Rot87, Sch64b, Sch11, Sha67, Tur06, UM92, Wal08, Was92, Wei39, dB86a, Ano97, Ano02b, Bet91, Cas83, Dea06, Eck13b, Far13b, Fav05, FH60, FS06, FR98, GL03, HH04a, Hol88a, Hol05, Hon81, Inf33, Inf34, Inf78, Kat15, LSW12, Lig05, Lin07, Lin08, Moo66, Moo85b, NCM38, NCM55, NCM70, Nie15, Ree15b, Sce16, S^{+}06, Teu02, Tuc97, Val06, VGEK87, Bal39]. Science [Bro36, N.38, RS07, Wei39, Coh42, Fer07, Gor17, For09]. Science/Goetzberger/Voß/Knobloch [MGG^{+}95]. Sciences [Ano57b, Bad83, FMO88, GSW95, GN86, HH76, Kra15a, Mai93, Pau85b, Ped64, SDG15, Dol95, Hal09, Hee95, Neu44, BR^{+}85]. Scientific [Bro95, Car97, Eis00, GSW95, Har01, Hen86b, HvMW79, How79, Jon78, KN12, Rod19, Sim14, Tib13, UM92, vMHW85, vM93, vM96, vM99, vM01, vM05, Aas15a, Bey49, Boh71b, DP97, Enz02a, Haw11, HP85, Het95, Hol88b, Lee09, PHvMW93, PD97, Rec10, SKST69, She72, Wei64, PvM93, Rec86]. scientifique [Bou49a, Bou49b, Pel47a]. Scientist [Ano43, Kae45, Sch49a, Sch49b, Aas99, Bus12, Cam99, Moo67, Moo69, Bus20]. Scientists [Acz01, Ano39e, Ant15, Duf46, Gor17, Lau33, MP01, Rog10, Seg85, Tib13, Cas14, Hal09, Rob12, Sce16, Str11, Wil85]. scienza [Val06]. scienziati [Cas14]. Scope [Bus20, Bus12, Jon14, Sho13, Sud11]. scoperte [Seg76, Seg83]. se [Ano85a]. Search [Dje07, Gea14, Sim00, Kon00]. secolo [Wei77]. Second [ABH^{+}37c, CS79d, Kra79, ABH^{+}37d, Stu18]. Secret [Kae45, Kar05b, Bag10, Bec57, Ber96, FB00, Kar07, Mel00]. secrète [Kar07]. Secrets [BGS95]. sections [Las55]. security [Yor75]. Seems [GLR14, Mor07]. Seen [Jam67, Nie68, Gra15, Roz67, Roz85]. Segrè [Wal08]. Según [KH25b]. sein [Ano02a, Bar89]. seine [MGG^{+}95]. seinem [KRW05]. seines [HK26, KH25a]. Seiten [wB90, Eis00]. seizures [RR00]. Selected [CS79a, DP97, PD97, Lee09, Boh71b, Bor63]. Selection [ABP98, Hen09b, APB96]. Self [Whe88]. Self-Synthesized [Whe88]. Selsk [Che93]. selskab [BR^{+}85, Ped64]. Semantic [vFH76]. Semiclassical [Kel85, BA06]. Semiconductor [ARK^{+}99, PKWF08, UPG^{+}12]. semidesjatiletiem [PB58]. seneste [Boh29c]. sens [Pat85]. sense [Dea06]. Separated [NBDG40, RM55]. Sept [Che99a, How99]. September [Far13d, Far13e, Aas02f, Far13f, Far13g]. septième [CCJ^{+}34]. sequence [Har83]. Serienspektra [Boh20c]. Serienspektren [Boh23i]. Series [BDR86, Kon02, Sch64b, Boh15c, Boh20c, Boh22c, RR00, VIR^{+}08]. serious [MB24]. ses [Ano85b, Ano85a]. Session [Boh63a]. sets [Ott03]. Seven [Boh22f, Gor17, Kav14, Sce16]. seventh [Boh23e, CCJ^{+}34, Far01]. Seventieth [Mor56]. Several [Boh13c]. shaky [Fin86]. shall [Gra15]. Shaped [Gib19, OP04, OPH^{+}09]. Share [BGS95]. Shell [Jen03]. Shellard [MGG^{+}95]. Shepherd [Dje07]. Shepherd-Barr [Dje07]. shine [Kra17]. shizen [Bō90, Boh08]. Shockley [Dje07]. Shook [Gam85, Kuh67, Haw11]. short [Bus20, Hei20, Ken85b]. should [Jak13]. shoulders [Ano62b]. Shows [Ano23c, Ano68a]. shu [Boh64f]. Shurkin [Dje07]. Sicilia [A^{+}09]. Sicily [A^{+}09]. Side [Hal12, Mor85a]. siebzigsten [Boh35b]. sig [Fri64]. Signal [CW12]. Significance [Boh62a, Ste88a, Boh63d, El’85]. Silence [Ano93c]. Simes [Mor11, Mor11]. simple [Cas85, ES74, Ter08]. Simplifying [vCK17]. Simulation [Hol80]. Simultaneous [Eps24]. since [Meh75]. Single [Boh13b, Hür22, FF78b, FF78a, PKWF08]. single-well [FF78b, FF78a]. Singular [Ham05]. Sir [Adl03, Boh26d, Lee09]. sit [Ano29]. Sites [PA01]. Situation [Hei63b, Sch35c, Sch35a, Sch35b, Tri80, Tuc97]. Sixteen [Boh85l]. Sixtieth [CS79d, Ros79b, Boh35b, Ros45, Ros61]. Sixty [FR13b, Bar99]. Skandinavien [DF14]. Skandinavische [Sch54]. skulle [Jak13]. Skurrile [Arr06b]. Slater [BG24, Hen81, Kon83, Kra09, Sch90, Was81]. slept [Bre97]. slightly [Ano62b]. slit [BO09, Kon15]. Slobodan [Cuf24]. små [WS09]. small [Ano85d]. smaller [UPG^{+}12]. smallest [Fis10a, Fis12a]. Smith [KZN^{+}88]. smooth [FF78a]. Smyth [Kae45]. Sneaking [Ghi05]. so-called [Che99a]. social [Hee95, Neu44]. Société [Bar23]. Society [GGK02, Gib19, Som20b, Ano97, Pih64]. Sociological [Hen86a]. Sociology [UM92]. Soft [AA89, AA90]. Software [ARK^{+}99, GRE^{+}01]. sogno [A^{+}09]. Sohn [Sch64c]. Sokalratic [BSMB99]. Solid [Fox80]. Solids [HEB^{+}80]. Solution [Smi76]. Solvay [BV09, CCJ^{+}34, Far01, Meh75, Ano28, CW12, CCJ^{+}34]. som [Kle64, Mor85b]. Some [B^{+}55, Boh20b, Boh61c, Boh62e, Fre85, Har23, Har26, MB24, Pei85, Boh41c, Des48]. something [Boh11a]. Sommerfeld [Eck13a, Hen86a, Lak96, BRB12, BH11, Boh28e, DJ15, Eck13a, Eck13b, Eck14, Eck15, FF78b, GS04, Ham05, Kel58, KK03, MR82e, MR82f, Nau15, SWW88, Sch88, Set10, Som20c, SE13, Som13b, Som13a, TM13, vM02]. Sommerfelds [SE13, Som20c, Som13b]. Sommerfeldsche [SE13]. Sommerfeldschen [Sch88, Som13a, vM02]. Son [Cha09]. Sønderdeling [Boh40c]. Sonderforschungsbereich [KZN^{+}88]. Sonnenenergie [MGG^{+}95]. Soul [Fer07, For09, RS07, Wal08, Lin07, Lin08, Seg07b]. Sounders [Yor75]. Source [Kae39]. Sources [KHFA67, Pau85b, vdW67, vdW68, vdW07, Kuh67]. sous [Ano28, Che92, CCJ^{+}34]. sous-détermination [Che92]. Soviet [Ano97, FH60, SSSS95]. Sowjetische [FH60]. Space [Dah02b, BH11, Dah02a, Pau55, SWW88]. space-time [Pau55]. Spain [Ros18]. Spaltning [Boh37e]. Spaltningen [Fav05]. Spanish [Boh88, HBS96, KH25b]. speak [Tuc97]. Speakable [Bel87]. speaks [Dje07]. Special [Ano95b, SSSS95, KZN^{+}88]. Spectra [Ano23c, Boh18b, Boh18c, Boh18a, Boh22e, Boh67a, HLA^{+}10, Kra24, Pia23, Som20a, Som20b, UG26, UGB84, Boh13e, Boh14a, Boh15d, Boh20c, Boh22d, Boh22c, Boh22g, Boh23i, Boh23c, BC23, Boh24b, Boh24e, Boh26e, Cos22, Har23, Har26, Kra23b, dBdB21, Boh23b, Som20c]. Spectral [Ano23c, Ano23d, Boh14b, Boh15a, Boh23e, RR00]. spectres [Boh23b, dBdB21]. Spectroscopic [RM55, Mas04]. Spectrum [Ano23d, Boh14c, Boh15c, Boh18c, Gau14]. Speeches [BDR86, SS98]. Speed [Lau13, Bal14]. Spektren [Boh22d, Boh24b, Som20c, Som20a, Som20b]. Spell [Bub73, Enz02b]. Spies [Lan94]. Spin [Hen09c, Mor07, Smi76, Dur00, GS04, Gou71]. Spinning [Boh26e, Tho26, UG26, UGB84]. Spinoza [GN86, Jon86]. Spirit [Bel99a, Gib19, SvMS85, Hei85a]. Split [Ano39d]. Splits [Kae39]. splitting [Boh40c]. Spookiest [Acz01]. spooky [Kon15]. Sprache [Rec09, Rec10]. Spreading [KN11, KN13]. Springer [wB90, Eis00, Hen86a, Hen86b, Tan09]. Springer-Verlag [wB90, Hen86a]. spukhaften [Kon15]. Spur [Hei10a]. spurned [She86]. Spymaster [SSSS95]. stabilen [ABH^{+}37b, ABH^{+}39a]. Stabilität [Föp15]. Stabilitätsbetrachtungen [Hoy73]. Stability [Boh32a, Boh31d, Föp15, Hoy73]. Stable [ABH^{+}37b, ABH^{+}39a]. Stage [Aas00c, Dje07, Ano02b, Gla00, Kai94a]. staircase [Fis10a, Fis12a]. Standardization [Lin55]. Standpoint [San24, Som18]. Standpunkte [Som18]. Stark [DJ15]. stars [Kra17]. State [Eve57, Whe57, DNT08, Har83]. States [Ano23d, Mar63a, Mar63b, Smi76, FF78a, Har87, Kle78, TM13, Far13c]. statesman [Aas99]. Statistical [HH76]. statistics [Pla24a, Pla24b]. Status [VW86]. Steam [Ano22]. Steen [Bis04]. steep [FF78b]. Stefan [Nat99]. stellar [Kra17]. Stellung [Bor58]. stemning [And64]. Stenholm [Bus12]. steps [Wei70]. Sternstunden [Büh98b]. Stewart [dR92]. still [Kae48]. Stitch [OSWR86]. stochastic [DNT08]. Stockholm [WD06]. Stof [Kra25]. Stoffernes [Boh21f]. stolen [Mel00]. Stoner [Nau13]. stop [Mer04]. Stopping [Boh40a]. Store [IH07, WS09]. Storia [Wei77, Val06]. Stories [Gam39, A^{+}09, WS09]. storiogra [DS09]. Störungsrechnung [Bor23]. Story [Acz01, Gam85, Hof47, Hof59, Hof14, Jon08, Kae45, KZN^{+}88, Kuh67, MP01, Ano09, Bag11, NN01, Ree15a, Val06]. Stössen [Boh25d]. Strahlung [Boh21e, BKS24b]. Strahlungshypothese [Sch24]. Strange [Hof47, Hof59, Hof14, Log00, HP81]. stranieri [A^{+}09]. Straßmann [CSW97a, CSW97a]. strategies [Bal14]. stress [Gol87]. Strings [MGG^{+}95]. stripped [B^{+}55]. Structural [Eng03]. Structure [Boh15b, Boh21a, Boh23g, Boh24f, Boh62d, Boh65a, Bor13, Bro73b, Cla13, CCJ^{+}34, Eng03, Gam29, GRE^{+}01, Hug89, Kat11b, Kra12a, KH23, Kra23a, Mil24, Nia98, Sho13, UG26, UGB84, Aas02e, Ano23a, Bec57, Boh15c, Boh20b, Boh21f, Boh21b, Boh21c, Boh22d, Boh22f, Boh23h, Boh23i, Boh23c, Boh23d, Boh23f, Boh23b, Boh24d, Boh24b, Boh25c, Boh26e, Boh35c, Boh36d, Boh85j, Gan17, Jea19, KP55, Kra10a, KH25b, Wei70, CCJ^{+}34]. Structures [Jan17]. Struggle [Fer07, For09, RS07, Wal08, Lin07, Lin08, Seg07b]. Struggles [HW07a]. Stuart [RS07]. students [BJN01]. Studie [Kuh67, MA65]. Studier [Boh11]. Studies [Aas00c, Cra02, DC96, MP97, UM92, dB86a, Bey49, B^{+}55, BL19, Boh11]. Study [How04, Kay14, Fis10a, Fis12a, Het95]. stuff [Haw11]. suas [Seg87]. sub [PRKH09]. sub-Bohr [PRKH09]. Subatomic [Kra06]. Sube [GRE^{+}01]. Subjectivity [Ste88b]. Substance [AA89, AA90]. successes [MB24]. Succession [Jon86]. Successive [Boh40b]. Suche [KZN^{+}88]. Sudoplatov [Les94, McM94]. sulla [Hei28]. Sun [RS07, HHW99]. Sündermann [HEB^{+}80]. Super [FP40a, FP40b]. Super-bomb [FP40a, FP40b].

Superarrivals [AMH06]. Superconductivity [Con49]. superposition [Das15]. Superstrings [KZN^{+}88]. Suppl [Boh87a]. Supplement [Nie97]. Supplementary [BFF98]. Supported [Cho02]. supreme [Cam99]. Surface [Boh09a, Boh09b, Boh10]. Surface-Tension [Boh09a, Boh09b]. Surfaces [OSWR86]. Surfaces/Paul [OSWR86]. Surprises [Pei91]. surreptitiously [Das15]. Survey [Boh11b, Boh34, Boh61a, Boh88]. Suspended [OP04, OPH^{+}09]. sustained [And73]. svjazi [PB58]. Swastika [Luc11]. Sweden [Dah14]. Swedenborg [Sch54, Sch54]. Swedish [Boh59a, CSW97b, Gra15]. Sweets [Ano01]. Symposium [DDU86, FMO88, K^{+}84, Ano97, Stu79, And89, LM85, Sch00b, VGEK87]. Synchrotron [HEB^{+}80]. Synthesized [Whe88]. System [BC23, Kra06, Kra15a, Whe88, Kra85]. Systeme [MGG^{+}95, Boh31b]. Systeme/Bunde/Havlin [MGG^{+}95]. systems [Boh16, Boh31b, Kel58, Boh13b, Boh13c]. Szilard [BS50].

T [Lev95]. T. [Boh62b, KP55]. Tabelle [ABH^{+}37b, ABH^{+}39a]. Table [AA89, AA90]. Tables [ABH^{+}37b, ABH^{+}39a]. tænker [Kle64, Mor85b]. tagliyotehe [Seg86a]. Taiji [Dow09]. tajnych [Kar06]. taken [CSW97a]. Takes [Ano23c, Far13f, Far13g, Ano23c]. Tale [CSW97b, Gor17, RS07, CSW96, Sce16]. Talk [Bri58, Rot00]. Tasks [SSSS95]. taught [CG14]. Teaching [Lat83, Wea88]. Tech [wB90]. Techniques [HEB^{+}80, KZN^{+}88]. Technological [MGG^{+}95]. Technology [DC96, K^{+}84, Kra07, Lan08, Sch11]. Teil [ARK^{+}99, Eis00, OSWR86, vM96, vM99, vM01, vM05]. Teilchen [Fis10a, Fis12a]. Telephone [Ano04b]. Tell [Ano39a, Ber95, Wei94, Hol00]. Teller [vW02a]. Tells [Ano23d, FS06]. temaer [KvG85]. Temperament [Bro36]. Temperatures [KZN^{+}88]. Temperatures/Jaenicke [KZN^{+}88]. Ten [Jon08, Ano23a]. Tension [Boh09a, Boh09b, Boh10]. tentatives [Kar07]. tenu [Ano28, CCJ^{+}34, LRdB^{+}23]. teoretiske [Boh39c]. teoría [KH25b, Boh88]. Terms [Yam02]. Terrible [Duf46]. Testing [Ach93]. Tests [Ano23d, Kar05b]. Tetsuo [Fre90]. Teubner [OSWR86]. text [FL10]. Textbooks [Nia98, NM12, dB13]. textes [BV85]. Textual [Nie15]. Thales [Lak96]. theater [FR98]. Their [HHW99, Ehr23, Seg80, Seg84, Seg87, Seg07a]. them [Nak15]. Thematic [Hol88b]. themes [KvG85]. Theodosius [Adl03]. Theology [Los61, Hon81]. theorem [HP01, Sch24, dB24]. théorème [dB24]. Theoretical [Anoxx, ARK^{+}99, Bet37, Bis04, Boh31c, BK37, D’A00, FW60, Hal12, Pei91, Plo10, Boh39c, Bro03, CSW97b, D’A01b, Des48, Set03]. Theoretische [MGG^{+}95]. Theorie [Ano23a, Bar89, Boh23i, BG24, Cas35, Her23, HK26, KH25a, Som20a, Som20d, Som20c, Som20b, vH23, Bar23, Boh23a, Boh32c, Boh93, dB25]. Theories [FR13a, HH76, Hei85b, Kra10a, Los11, Che92, Hei85c, Jea19, Che92]. théoriques [Des48]. Theorist [SM08, Dam85]. Theory [Ach93, Acz01, Ano23c, Ano62c, Ayl26, Ayl27, Bas71, Beh43a, Beh43c, Boh13d, Boh15b, Boh18b, Boh18c, Boh18a, Boh22e, Boh30a, Boh31a, Boh32b, Boh32c, Boh34, Boh61a, Boh62d, Boh65b, Boh67a, Boh75b, Boh75a, Boh78, Boh88, Boh11b, Bro36, Chi60, Dar92d, Dar97, Gam85, Gar72, Gea14, HH76, HB63, Hei63b, Hen81, Hen86a, Hen09a, Her71, HN64, HN70, HK26, HW07a, Hoy73, HnB74, HR74, Hun74, Kat11a, Kon83, Kon02, KZN^{+}88, Kra79, Kra09, Kra10b, KH23, Kra24, Lau33, Lem31, Mar54, McC34, Mil24, Møl55, OK12, Pei40, Pei97a, Pia23, Pla23, Plo06a, Rad82, Ric14, Ric16, Ros79e, San24, Som20a, Som20d, Som20b, Som24a, Som24d, Som24c, Som24b, Tan04a, Vic08, Was81, Whe57, Whi06, vFH76, vH23, Ano23a, BV09, Bag04, Bar89, BA06]. theory [Boh71a, Boh85a, Boh11, Boh12, Boh16, Boh21e, Boh22f, Boh22g, Boh23h, Boh23i, Boh23a, Boh24a, Boh24d, BKS24b, BKS24a, Boh24e, Boh25b, Boh25a, Boh26c, Boh26b, Boh28b, Boh28c, Boh28d, Boh28e, Boh29b, Boh29c, Boh30b, Boh31b, Boh33c, Boh36a, Boh58d, Boh67c, Boh87b, BB87, Boh87a, Boh93, Bor23, BG24, CS15, Cas35, Che99a, Cli65, Cli87, Cos22, Dar88, Dav16, DDU86, DJ15, Eck15, Ehr23, El’85, FM12, Fin86, Har83, Har87, Har23, Har26, Hei55, Hei62, Hei64b, Her23, Het95, HK22, HKK29, HG15, Hon81, Hou30, Hun67, Jäh15, Jea19, KP55, Kav14, Kon00, Kos23, Kra85, Kra11a, Kra13a, KH25a, KH25b, LR23, Lan55, Lat83, Lee06, Mac85, MR82d, MR82e, MR82f, Nie15, Pai48, Pai79, PV15, Pet14, Pol60, Pop82, Sce94, Sch15, Sch88, Set10, Sie13, Som18, Som20c, Som31]. theory [SE13, Wea88, Wei14, WZ83, dB25, Boh23i, Har28, Mil24, Kuh67]. there [Fra04]. Thermodynamics [Kle66]. thermoelectric [Boh12]. thickness [Her55]. Things [Lau33]. Thinking [Plo10, Rot87]. thinks [Far13d, Far13e]. Third [ABH^{+}38b, Den99, KLR13, ABH^{+}38a, BLW19]. thirteen [Bey49, NN01]. Thirty [Gam85, Yor75, Kuh67]. Thomas [GR08]. Thomsen [AA89, AA90]. Thomson [Kra14c, Lak96, Hei77, Hei85d, Kra97]. Thomsons [Sim14]. Thorium [Boh39a, HS39a, Mei39, MF39b]. Thorsen [KZN^{+}88]. Thought [Hoo94, Rös94, Ste86b, Hol88b, Kon15, Nik12, Sch54]. thoughts [RC90]. Threat [Duf46, Mas04]. Three [And73, Boh23b, Wei70, ZDSF11, Boh22g, Boh24e, Boh22d, Boh24b]. Three-Man [ZDSF11]. Thus [Ano39a]. Tidal [HEB^{+}80]. til [Ano29, Boh50a, Boh64d, HH55, Jak13, Kal64, NN01, Sch68]. Tillich [Los61]. Time [Hal17, KZN^{+}88, Lau33, Rog10, Ste88a, Enz02a, Pau55]. Times [Ano93a, Ber91, Hen92, Kle93, MGG^{+}95, Red93, Ryn83, Seg93, Stu92, Bre97, Cro01, Dar92a, Eck13b, Opp63a, Pai91, Sac81, Dea13, Wil92]. Tissues [AA89, AA90]. Tisvilde [Sch64a]. Tl [Nie55]. today [Ano21]. Tokyo [wB90, K^{+}84]. Told [CW12, BR64, RRK^{+}64]. Tompkins [Gam39, Gam42, Gam42]. tømrerlærlingen [Jak13]. tonedigt [KvG85]. Too [Hür22, Gol94, Mor95]. Topological [MGG^{+}95]. Total [Las55]. Tourist [PA01]. town [Ano29]. tra [Cam08, DS09]. track [Hei10a, LH55]. Tradition [Bis04, Bro03, Lee06]. Traditions [KLR13]. Træk [Koc64, Sch68]. Tragedy [Jon08, RS07]. Traits [Sch68]. Transcendental [BKP09, Hon81, Hon82b]. Transformation [Eva23]. Transformationen [Ehr23]. Transformations [Boh40b, Jon15, Ehr23, KLR13, PP16]. transition [SWW88]. Transitions [Faf58]. translated [Hub14, Mil24]. Translation [Tri80]. Transmutation [BK37, Boh41e, Boh42a, Boh45c, Boh46c]. Transmutations [Boh37f, Boh85m]. Transuranic [Sto97]. Transuranium [Sim00]. travel [Ros18]. Travelling [Rod19]. treason [BLW19]. treasury [FF91]. treatment [Ehr23]. Treffen [Ano02a, Kar05a, Paw02]. tretten [NN01]. Tribute [Jac85, Kub08, Wei63, MSB^{+}37]. Tributes [MSB^{+}37]. Trilogy [AH13, Far13a, Kra16b, Ren15, Boh85n]. trip [WD06]. Trois [Boh23b]. trudy [Boh71b]. True [MP01, Mor95]. Truth [Pei86a]. trying [Hol00]. tumult [Pat85]. tumulte [Pat85]. Tunge [Boh40c]. turbulent [Eck13b]. Turn [Kuh67]. turns [Sie13]. Twentieth [D’A00, Dor79, SDG15, Bro00b, D’A01b, Jac12, LSW12, Lig05, Wei77, FW60]. twentieth-century [LSW12, Lig05]. Twenty [CW85, FT19, FF17a, Rio17, Zin19, GL03]. Twenty-First [FT19]. Twenty-First-Century [Zin19, FF17a, Rio17]. Twenty-five [CW85]. twenty-one [GL03]. Two [Ano39c, BO09, Bus12, Eps24, Gar72, Gom06, Lau51, Hen86a, Ree05, Ste11, Aas20]. two-electron [Ree05]. Two-slit [BO09]. Type [AA90, MF39a, Pet14, TM13]. types [Sch15]. typescript [BS87].

U [GRE^{+}01, ARK^{+}99, HEB^{+}80, KZN^{+}88, MGG^{+}95, OSWR86, PHvMW93, Tur46a, Tur46b]. U. [Hei75]. u.a [Car97, Hen86b, HvMW79, HP85, Kra96, PvM93, Pye81, vMHW85, vM93, vM96, vM99, vM01, vM05]. u.a. [BDR86, Eis00]. U.P [Gor17]. Übersicht [Boh31a]. ucheny [Moo69]. udbygges [Ros64]. udnyttelse [Boh55a]. Udvikling [Boh29c]. Ulrich [Dar87, Stu84]. Ultimate [Dow09]. unbearable [Büh98a]. unbounded [Kae48]. Uncertain [Yam02]. Uncertainty [Lin07, Lin08, Meh87, Swe02, Tan04b, Zan02, Cas09, CG14, Pet11, dB86b, vD15, Fer07, RS07, For09]. Uncovered [Hür22]. underdetermination [Che92, Mas04]. undergraduate [Jon14]. Underlying [Hof47, Hof59, Hof14]. Underøgelser [Boh41e, Boh42a, Boh46c]. understand [Bac15]. Understanding [FG21, Sch64b, Dol95]. unerträgliche [Büh98a]. Unexpected [RS07]. Unfinished [Bec18, DT18]. ungdom [Adl64]. ungeheuerliche [Büh98a]. Unified [Bal39, Coh42, N.38, NBD^{+}38, Wei39, NCM38, NCM55, NCM70]. unifying [Fal98, Hel98b]. Union [ABH^{+}36a, ABH^{+}36b, ABH^{+}37a, ABH^{+}37c, ABH^{+}37d, ABH^{+}38a, ABH^{+}38b, ABH^{+}39b, ABH^{+}40a, ABH^{+}40b]. Unione [ABH^{+}36b]. United [Boh64d, Suv85, Boh50a, Boh50d, Boh50e, Boh67b, Boh85d, Boh85k, Far13c]. Units [McW73, SH59]. Unity [Aas90a, wB90, Fre89, Nie65, Wei39, Blæ85, Blæ88, NCM70, Lan72]. universal [Ish15, Ish17, PB17]. Universe [Bar07, Whe92, Dea13, KZN^{+}88]. Universe/ [KZN^{+}88]. Universelle [Ish15]. universes [Byr10]. Université [Pel47a]. Universitet [Ano10]. Universitetets [Ano21]. Universitetsforlaget [Bad83]. Universities [SDG15]. University [Bor13, Bus20, Cla13, Cuf24, Hen86a, Hen92, Hub14, Kra15a, Kra16b, Kri91, Lan08, Par06, RS07, Tur06, Ano21, Lau51]. Universums [KZN^{+}88]. Universums/Davies [KZN^{+}88]. Unlikely [Acz01]. unlocked [Bec57]. unmechanical [UG25]. unmechanischen [UG25]. Unpredictability [CW12]. unpublished [FL10]. Unraveling [Ghi05]. Unravelling [FR13b]. Unreasonable [Plo11]. Unreliable [Les94]. unseren [BL19]. unspeakable [Bel87]. Untersuchung [Bar89]. until [Gra15]. Unveiling [Ano93a, Red93]. Unwanted [SSSS95]. unyielding [Bol04]. update [Ant15]. Uran [HS39a]. Uranium [Ano39a, Ano39d, Ano39e, Boh39a, Kae39, MF39a, MF39c, NBDG40, Ber96, HS39a, HS39b, MF39b]. Urans [HS39b]. Uranspaltung [GRE^{+}01, HS39a]. Urkraft [KZN^{+}88]. US$62.99 [Bor13]. USA [FMO88]. Use [PSV01, Boh31b, Boh55a]. Uses [Ano39d, HF15, Sch15]. Using [BL63b, BL18b].

V [All56, Mor56, Pye81, SR99, Som20b, PB58, Sch64c, Ste88b]. V. [Som20b]. valid [RR00]. Vapor [Ano23d]. var [WS09]. variables [Nak15]. Various [Mar54]. Vast [Ano39d]. ved [Dah43, KNA01]. Vegard [BL19, BL19]. vek [K^{+}86]. vekselsange [KvG85]. Vekselvirkningen [Kra25]. Velocity [Boh13d, BBBL40, Boh41b, Fox80]. Velocity-Range [BBBL40, Boh41b]. venner [BR64, RRK^{+}64]. venskab [Cou64]. Venturing [JFEH15]. Verantwortung [GRE^{+}01]. Verden [MW46b, Jak13]. Verhalten [HS39b]. Verhaltens [UG25]. Verhältnisses [Mor11]. verification [BG24, Cas35]. Verlag [wB90, Hen86a, Sch64c]. versatility [Dir64, Dir67]. Version [GRE^{+}01]. Versions [Jam74b, Jam74c]. Versus [Duf46, Bol04, Fol96, Hol86, Mur90a, Pet14, Sac88]. Vertrieb [ARK^{+}99]. very [Bus20, DG11, FF78b, Hei20]. vetande [Boh59a]. Vexation [Fay07, Par06, Tur06, Hol05]. vi [Bor13, Cla13, Hen86a]. Vibration [Boh09a, Boh09b]. vibrational [JVK22]. Victor [Coh42, Rec86]. Victory [Hol05, Fay07, Par06, Tur06]. Vid [Che93]. vide [Boh11a]. videnskab [Fav05]. videnskabeligt [Wei64]. Videnskabernes [BR^{+}85, Ped64, S^{+}06]. vie [dB63]. Vienna [MR87]. Vier [Boh31a]. View [Ayl26, Ayl27, Che93, Fay91, Got02, Hen93, Lem31, Ros64, Koc64]. Viewed [Dah02b, Dah02a, Møl55]. Vieweg [Sch64c]. Viewpoint [Bro09]. viewpoints [Bet91]. Views [Ano36a, Ano36b, Ant15, Boh71a, Boh85a, Cam07, HP01, Ste11, Bus12]. vii [Bal39, Hen86a, N.38, Pia24a, Pia24b, Sha67]. viii [Hen86a, Kra07, Lan08]. Vil [Boh11a]. Vilenkin [MGG^{+}95]. Villars [Pia24a, Pia24b]. Vindications [Moy81c]. virke [BR64, RRK^{+}64]. Virtual [Hen81]. Vision [Ano39e, Cuf24, Per21, Bro15]. Visit [Nis37, Aas02h]. visualization [Tuc97]. Vol [Bro95, KZN^{+}88, Kra07, OSWR86, Pye81, Wei39, Che93, MP97, Boh81, Boh05, Boh07b]. Volt [Ano39d]. Volts [Ano39b, Ano39d]. Volume [BDR86, Bro99, Bro00a, Car97, Dor79, Eis00, HH76, Hen86a, HvMW79, Kra96, PvM93, Rec86, Sta09b, UM92, vMHW85, vM93, vM96, vM99, vM01, vM05, BH11, FW60, FK85, HH55, HP85, Rei72, N.38, Pei08, Hei86a].

Volumes [Kub09, Stu84]. vom [Som18, UG25, vM02]. vor [Ano02a]. Vorgänge [Kos23]. Vorstellungsvermögen [Büh98a]. Vortrag [Som20c, Som24c]. Vorträge [Boh66a, Boh85e, Mal70]. Voß [MGG^{+}95]. vs [Kra11b, Lee06].

W [All56, Bal39, Coh42, Mor56, N.38, Pia24a, Pia24b, Sch64c, Sch64b, Wei39]. Waddington [Adl03]. Waerden [Hei48, Kuh67]. walls [FF78b]. wanted [FJ21b]. War [Aas99, Aas03, Aas20, Jon78, Mel00, Sim12, Wer02, Bag10, Boh64b, Boh64a, BH03, Bro15, Far13b, Kat15, Pai64, LSW12, Sal19]. warm [Gra15]. Wars [Stu18]. wartime [CSW96, Jon85b]. Warum [CSW97a]. Was [Ano93c, New09, Bel86, Bun79, Mel00]. Washington [Ano57b, Anoxx]. Water [Boh09a, Boh09b, Boh10, Bør12]. Water-Surface [Boh10]. Wave [AMH06, Har28, Kal16, NM12, dB13, Dar27, DJ15, Hei10b, MGG09, MR87, MWL^{+}08, Nau15, PRKH09, Pau33, YRBD09, Zin16]. Wave-particle [NM12]. Waves [Sla85]. Way [Duf46, BG24, De 12, De 14, Jon85a]. weapon [CM85, Kar07]. weapons [Bet91, Boh64b, Boh64a, Gow85, Opp64]. Week [Kae39, Bec57]. Weg [BG24, Ban00]. Weinberg [MGG^{+}95]. Weird [Str11]. Weise [GRE^{+}01]. Weiss [MGG^{+}95, Pia24a, Pia24b]. Weisskopf [HEB^{+}80, All56, Pye81, Rec86, Mor56]. weitere [Som13a]. weiterer [HS39a]. Weizäcker [Ano02a, Ano02a]. Weizsäcker [GRE^{+}01]. Welcher [Ban00]. Well [Ros00, Cas35, FF78b, FF78a, PKWF08, RR00, She85b]. well-known [Cas35]. Wellenmechanik [Hei10b, Pau33]. weniger [Cas35]. Wenske [MGG^{+}95]. went [Sie13]. Wer [Boh55e]. werden [CSW97a]. were [Bey49, Nik08, WS09]. Werk [Dar87, MGG^{+}95]. Werk/Weinberg [MGG^{+}95]. Werke [BDR86]. Werner [Adl03, Bor58, GRE^{+}01, Kra07, Lan08, Lau07, MGG^{+}95, vM02, vW02b, Aas02h, BDR86, Dör05, Fav05, Hei84, KRW05, Meh87, PLS02, Rec09, Rec10, Yam02, vM02, vW02b]. Western [AFG90]. Westphal [Sch64c, Sch64b]. Wheeler [Hal17]. where [Jak13, Bro36]. Whether [Ano39c]. which [Ano21]. Whitehead [Adl03]. Whittlesey [Duf46]. Who [Cha09, Duf46, Gam63, GLSC01, How04, Bun79, Cli87, Dem15, Jak13, Boh55e, Dem15]. Widely [Ano39a]. Wie [Bel86]. wiedza [Boh63c]. Wife [Wei94]. Wigner [FR13a, HH04b, SWW88]. wildly [Hei15]. Wiley [Jam67]. Wilhelm [MGG^{+}95]. Will [Ano39c, Bec57, Kra11a]. Willem [Bor58]. William [Dje07, Ste86a, Ste86b, Ste87, Ste88a, Ste88b]. Windows [GRE^{+}01]. Wine [Bro09]. Winner [Ano62c]. winners [NN01]. Wins [Ano57a]. Wir [Aas00c]. Wirken [Rec09, Rec10]. Wirklichkeit [Hel98a]. Wirkung [Boh25d]. Wirkungsquantum [Boh29a, Boh38e]. Wirkungsquantums [Ish15]. Wissenschaft [MGG^{+}95, Rec09, Rec10, Sch64c, Sch64b, Bor58, FS06]. Wissenschaft/Cassidy [MGG^{+}95]. Wissenschaftlers [GRE^{+}01]. Wissenschaftlers/Dürr [GRE^{+}01]. wissenschaftliche [Rec09, Rec10]. Wissenschaftlicher [ARK^{+}99, Car97, Eis00, GRE^{+}01, HEB^{+}80, Hen86b, HvMW79, MGG^{+}95, OSWR86, Pye81, vMHW85, vM93, vM96, vM99, vM01, vM05, HP85, PHvMW93, PvM93, Kra96]. Wissenschaftshistoriker [Aas00c]. wit [Wei35]. withdrawal [Pat85]. Without [Ach93]. Witness [Les94, SSSS95]. Wittgenstein [Bus12, Mof06, Ste11]. Witzes [Wei35]. Wizard [Jon78]. WKB [Kle78]. Wolfgang [ARK^{+}99, Car97, Eis00, HEB^{+}80, Kra96, MGG^{+}95, Pye81, Rec86, Bro95, Eis00, EvM94, Enz02a, FW60, Hen86b, HvMW79, HP85, PHvMW93, vMHW85, vM93, vM96, vM99, vM01, vM05, Car97, Pye81]. women [HHW99]. Won [Ano62c, Cha09, New09]. wonderful [Mor95]. Wonderland [Gam39, Gam42]. Woodger [Coh42]. Wootters [Kal16]. Words [Che94b, Dje07, Kel09]. Work [Boh61c, Boh62e, Jam67, Kle66, Lau51, Nie68, Rec10, Was81, El’85, Fre86, Fre96, Hou30, Roz67, Roz85, SKST69, Boh81, Fri67]. workers [RRK^{+}64]. Working [Nat99, Roz98]. Works [Ano85f, BDR86, Bro99, Bro00a, Dor79, BR64, Boh71b, BRH^{+}99, Boh81, Boh05, Boh07b, BHK^{+}08, Bor63, PLS02, Pei08, RRK^{+}64, dB63, Ber15, KZN^{+}88, Stu84, Kub09]. World [Arr06b, Boh50b, Boh85c, Duf46, Gam66, Hel85, Hür22, Kuh67, Lem31, LSW12, Sha67, Stu18, Wei85a, Whe88, Aas07, BBCN86, BB^{+}89, Dea06, FF91, FMO88, GL03, Haw11, Inf34, Kon15, Lan72, MW46a, MW46b, MW07, Moo66, Moo85b, Pai86, Sch54, Wei85b, Yor75, Zin15, Aas99, Aas03, Aas20, Fre90, Mel00, Sim12, Wer02, AFG90]. worlds [Byr10]. worrying [Mer04]. Wörterbuch [MGG^{+}95]. Would [Nik08, Boh11a]. Wreck [Ano39e]. Writings [Aas88, ABP98, EvM94, Mer89, APB96, BDR86, Boh87c, BB87, Dol95]. Wrong [Gol94, Mei62]. Wu [Seg86b].

X [Bar23, Bro86, Che85a, Kae39, Duf46, Kra16b, Boh14a, BC23, Bra61, Cos22, Kay14, Seg80, Seg83, Seg84, Seg87, Seg07a]. X-ray [BC23, Cos22, Boh14a, Bra61]. X-Rays [Kae39, Kay14, Seg80, Seg84, Seg87, Seg07a]. xi [Bro86, Hen86a, Hub14]. xian [Seg86b]. XII [wB90, Dor79, RS07, Boh14c]. XIII [Sch64c, Kri91]. xiv [Par06, Tur06]. xlvii [Hen86a]. xvi [Sha67]. xvii [Ano93a, Hen92, Red93, Seg93]. XX [VGEK87, K^{+}86, Wei77]. xxix [Hen86b]. XXV [Som20b]. XXV. [Som20b].

Yakov [Fre96]. Yale [Ano23c, Ano23d]. Year [Ano93b, FR13b, Pau85a, Pau01, Dah43, SR99]. Years [And89, Ano71, ARK^{+}99, Boh21a, Bor16, CW85, Gam85, Gla00, Kuh67, LM85, Lin81, Mla98, PLS02, RY13, Roq17, Whe81, AK15, Ano23a, Bad83, Bar99, Bec57, Boh64e, Boh64b, Boh64a, Boh66a, Boh85e, Cas64, Con62, Cou64, Dea13, Hoc83, Hoi01, Kae48, Kav14, Kra13a, Pai85b, Rob79, Rot00, oER64, Yor75]. yesterday [Ano29]. yngste [Kal64]. York [wB90, Dea13, Duf46, Eis00, Gor17, Hen86a, Hen86b, Hub14, Jam67, RS07, Sha67, Ano39e]. Young [Sch90, Gam60, Wil85]. youngest [Kal64]. youth [Adl64, Rec10]. Yuan [Boh64f]. Yukawa [Dar88, FR13a]. z [Kar06]. zehn [Ano23a]. Zeichnen [GRE^{+}01]. Zeilinger [Fis10a, Fis12a]. Zeit [KZN^{+}88]. zelloph [wB90]. Zerstreuung [LR23]. zi [Boh64f]. Ziele [Som20a, Som20c, Som20b]. Zinc [Alt82]. zones [SWW88]. Zuflucht [DF14]. Zufluchtsland [Dah14]. Zum [Gam60, vM02, Bar89, Ber15, Boh35b, Fis10a, Fis12a, Hei35, Paw02, Som20c]. Zumutung. [Büh98a]. zur [BG24, FH60, GRE^{+}01, Kos23, Kuh67, MA65, Pau01, Boh21e, Boh24g, BR33, Pla24a, Pla24b, Sch88, Som20d]. Zurek [Kal16]. Zurich [MR87]. Zwang [UG25]. Zweiter [ABH^{+}37d].

References

Attanasio:2009:SAR

[A^{+}09] Maria Attanasio et al., editors. Il sogno e l’approdo: racconti di stranieri in Sicilia. (Italian) [The dream and the landing place: stories of aliens in Sicily], volume 22 of Il contesto. Sellerio, Palermo, Italy, 2009. ISBN 88-389-2354-X. 213 pp. LCCN PQ4253.2 .S58 2009.

Ando:1989:RBL

[AA89] A. Ando and I. Ando. Relation between the location of elements in the Thomsen–Bohr periodic table and the binding substance in soft tissues. Journal of Radioanalytical and Nuclear Chemistry, 132(2):387–396, ???? 1989. CODEN JRNCDM. ISSN 0236-5731 (print), 1588-2780 (electronic).

Ando:1990:RBL

[AA90] A. Ando and I. Ando. Relation between the location of elements in the Thomsen–Bohr type periodic table and the binding substance in soft tissues. Journal of Pharmacobio-Dynamics, 13(1):S24, ???? 1990. ISSN 0386-846X.

Aaserud:1985:NBF

[Aas85a] Finn Aaserud. Niels Bohr as fund raiser. Physics Today, 38(10):38–46, October 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Aaserud:1985:RNB

[Aas85b] Finn Aaserud. The Redirection of the Niels Bohr Institute in the 1930s: Response to Changing Conditions for Basic Science Enterprise. Ph.D dissertation, The Johns Hopkins University, Baltimore, MD, USA, September 1985. 407 pp.

Aaserud:1988:BRN

[Aas88] Finn Aaserud. Book review: Niels Bohr: The Philosophical Writings of Niels Bohr. Isis, 79(2):351–352, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/233676.

Aaserud:1990:BRB

[Aas90a] Finn Aaserud. Book review: Harmony and Unity: The Life of Niels Bohr by Niels Blædel. Isis, 81(1):130–131, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/234132.

Aaserud:1990:RSN

[Aas90b] Finn Aaserud. Redirecting Science: Niels Bohr, Philanthropy, and the Rise of Nuclear Physics. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-35366-1. xiii + 356 pp. LCCN QC789.2.D4 A27 1990.

URL http://www.loc.gov/catdir/description/cam024/89048317.
html;   http://www.loc.gov/catdir/toc/cam028/89048317.html.

Aaserud:1999:SSN

[Aas99] Finn Aaserud. The scientist and the statesman: Niels Bohr’s political crusade during World War II. Historical Studies in the Physical and Biological Sciences, 30(1):1–47, ???? 1999. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.jstor.

org/stable/27757819.

Aaserud:2000:CPP

[Aas00a] Finn Aaserud. “Copenhagen” play portrays Bohr and Heisenberg. AIP History Newsletter, 32(1):1–2, ???? 2000. CODEN ???? ISSN 1048-1338.

Aaserud:2000:NBA

[Aas00b] Finn Aaserud. Niels Bohr Archive expands in new quarters. AIP History Newsletter, 31(1):1–2, ???? 2000. CODEN ???? ISSN 1048-1338.

Aaserud:2000:WBD

[Aas00c] Finn Aaserud. Wir brauchen einen Dialog in “Kopenhagen” — Wissenschaftshistoriker auf der Bühne. (German) [We need a dialogue in “Copenhagen” — science studies on the stage]. Max-Plack-Institute für Wissenschaftsgeschichte, ??(??):33–36, ???? 2000.

Aaserud:2002:BHM

[Aas02a] Finn Aaserud. The Bohr–Heisenberg meeting from a distance.

In Anonymous [Ano02b]. URL http://web.gc.cuny.edu/ashp/nml/

artsci/aaserud.htm. Three videocassetes recorded at the Graduate Center of the City University of New York, Baird Auditorium, March 2, 2002. Apparently not published in print.

Aaserud:2002:BRP

[Aas02b] Finn Aaserud. Bohr’s relevance to philosophy and contemporary physics. In Mataix and Rivadulla Rodríguez [MR02], pages 155–170. ISBN 84-7491-640-2. LCCN QC174.12 .F57 2002.

Aaserud:2002:NBN

[Aas02c] Finn Aaserud. Niels Bohr and the Nobel prizes. In Crawford [Cra02], pages 39–64. ISBN 4-946443-69-X. ISSN 0282-1036. LCCN AS911.N9 .H57 2002.

Aaserud:2002:NBA

[Aas02d] Finn Aaserud. Niels Bohr Archive: Release of documents relating to 1941 Bohr–Heisenberg meeting. Web document, February 6, 2002. URL

http://www.nba.nbi.dk/release.html.

Aaserud:2002:NBO

[Aas02e] Finn Aaserud. Niels Bohr, om atomernes bygning — efterskrift. (Danish) [Niels Bohr, on atomic structure — postscript]. In Teuber [Teu02], pages 162–169. ISBN 87-12-03847-4. LCCN Q127.D4 H65 2002.

Aaserud:2002:RRP

[Aas02f] Finn Aaserud. Recently released papers on the Bohr–Heisenberg meeting in Copenhagen in September 1941. Web document, March 8, 2002.

URL http://web.gc.cuny.edu/sciart/copenhagen/bhpapers.htm.

Aaserud:2002:RDR

[Aas02g] Finn Aaserud. Release of documents relating to 1941 Bohr–Heisenberg meeting. Naturens Verden, 84(8–9):1–40, ???? 2002. CODEN

NAVDAL. ISSN 0028-0895. URL http://www.nba.nbi.dk/papers/
introduction.htm;   http://www.nba.nbi.dk/release.html.

Aaserud:2002:WHB

[Aas02h] Finn Aaserud. Werner Heisenberg’s besøk hos Niels Bohr i 1941: Drama og dokumentasjon. (Norwegian) [Werner Heisenberg’s visit with Niels Bohr in 1941: drama and documentation]. Forskningspolitikk, 3 (??):14–15, 17, ???? 2002. ISSN 0333-0273.

Aaserud:2003:NBP

[Aas03] Finn Aaserud. Niels Bohr’s political crusade during World War II. In Booss and Høyrup [BH03], pages 299–311. ISBN 3-7643-1634-9, 0-8176-1634-9. LCCN QA10.8 .M38 2003. URL http://mmf.ruc.dk/~booss/

mathwar/BookFiles/Booss299-311.pdf. Based on [Aas99].

Aaserud:2007:NBM

[Aas07] Finn Aaserud. Niels Bohr’s mission for an ‘open world’. In Michał Kokowski, editor, The Global and the Local: The History of Science and the Cultural Integration of Europe. Proceedings of the 2nd ICESHS (Cracow, Poland, September 6–9, 2006), pages 706–709. Wydawnictwo Polskiej Akademii Umiejñetności, Kraków, Poland, 2007. ISBN 83-60183-

42-2. LCCN ???? URL http://www.2iceshs.cyfronet.pl/2ICESHS_
Proceedings/Chapter_25/R-17_Aaserud.pdf.

Aaserud:2015:LPM

[Aas15a] Finn Aaserud. Love and physics: Margrethe Nørlund and Niels Bohr’s scientific creativity, 1910–1913. In Aaserud and Kragh [AK15], pages 61–74. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/2015/
06/Bohr-Book.pdf.

Aaserud:2015:P

[Aas15b] Finn Aaserud. Preface. In Aaserud and Kragh [AK15], pages 9–12. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/2015/
06/Bohr-Book.pdf.

Aaserud:2020:NBD

[Aas20] Finn Aaserud. Niels Bohr’s diplomatic mission during and after World War Two. Berichte zur Wissenschaftsgeschichte, 43(4):493–520, December 2020. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Aston:1936:EBA

[ABH^{+}36a] F. W. Aston, N. Bohr, O. Hahn, W. D. Harkins, and G. Urbain. Erster Bericht der “Atom-Kommission” der Internationalen Union für Chemie. (German) [First report of the Commission on Atoms, International Union of Chemistry]. Berichte der deutschen chemischen Gesellschaft (Abteilung A: Vereins-Nachrichten), 69(7): A171–A174, July 8, 1936. ISSN 1099-0682 (print), 1434-1948 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/cber.
19360690747/abstract.

Aston:1936:UIC

[ABH^{+}36b] F. W. Aston, N. Bohr, O. Hahn, W. D. Harkins, and G. Urbain. Unione internazionale di chimica: Primo rapporto della commissione degli atomi. (Italian) [International Union of Chemistry: First report of the Commission on Atoms]. Il Nuovo Cimento (8), 13(8):369–372, ???? 1936. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Aston:1937:FRC

[ABH^{+}37a] F. W. Aston, N. Bohr, O. Hahn, W. D. Harkins, and G. Urbain. First report of the Committee on Atoms of the International Union of Chemistry. Journal of the Chemical Society, ??(??):1910–1912, ???? 1937. CODEN JCSOA9. ISSN 0368-1769 (print), 2050-5574 (electronic).

Aston:1937:ITS

[ABH^{+}37b] F. W. Aston, N. Bohr, O. Hahn, W. D. Harkins, and G. Urbain. Internationale Tabelle der stabilen Isotope für 1937. (German) [International tables of stable isotopes for 1937]. Zeitschrift f”ur analytische Chemie, 112(9–10):342, September 1937. CODEN ABCNBP. ISSN 1618-2642 (print), 1618-2650 (electronic).

Aston:1937:SRC

[ABH^{+}37c] F. W. Aston, N. Bohr, O. Hahn, W. D. Harkins, and G. Urbain. Second report of the Committee on Atoms of the International Union of Chemistry. The Analyst, 62(??):874, ???? 1937. CODEN ANALAO. ISSN 0003-2654 (print), 1364-5528 (electronic).

Aston:1937:ZBA

[ABH^{+}37d] F. W. Aston, N. Bohr, O. Hahn, W. D. Harkins, and G. Urbain. Zweiter Bericht der Atom-Kommission der Internationalen Union für Chemie 1937. (German) [Second report of the Commission on Atoms, International Union of Chemistry]. Berichte der deutschen chemischen Gesellschaft (Abteilung A: Vereins-Nachrichten), 70(12):A159–

A160, December 1, 1937. URL http://onlinelibrary.wiley.com/
doi/10.1002/cber.19370701242/abstract.

Aston:1938:DBA

[ABH^{+}38a] F. W. Aston, N. Bohr, O. Hahn, W. D. Harkins, and G. Urbain. Dritter Bericht der Atom-Kommission der Internationalen Union für Chemie 1937. (German) [Third report of the Commission on Atoms, International Union of Chemistry]. Berichte der deutschen chemischen Gesellschaft (Abteilung A: Vereins-Nachrichten), 71(7):A159–A162, Oc-

tober 5, 1938. URL http://onlinelibrary.wiley.com/doi/10.1002/
cber.19380710736/abstract.

Aston:1938:TRC

[ABH^{+}38b] F. W. Aston, Niels Bohr, O. Hahn, W. D. Harkins, and G. Urbain. Third report of the Committee on Atoms of the International Union of Chemistry. Journal of the Chemical Society, ??(??):1110–1112, ???? 1938. CODEN JCSOA9. ISSN 0368-1769 (print), 2050-5574 (electronic).

Aston:1939:ITS

[ABH^{+}39a] F. W. Aston, N. Bohr, O. Hahn, W. D. Harkins, and G. Urbain. Internationale Tabelle der stabilen Isotope für 1939. (German) [International tables of stable isotopes for 1939]. Zeitschrift f”ur analytische Chemie, 118(3–4):16–107, March 1939. CODEN ABCNBP. ISSN 1618-2642 (print), 1618-2650 (electronic).

Aston:1939:FRC

[ABH^{+}39b] F. W. Aston, Niels Bohr, O. Hahn, W. D. Harkins, F. Joliot, R. S. Mulliken, and M. L. Oliphant. Fourth report of the Committee on Atoms of the International Union of Chemistry. The Analyst, 64(??):887, ???? 1939. CODEN ANALAO. ISSN 0003-2654 (print), 1364-5528 (electronic).

Aston:1940:FRCa

[ABH^{+}40a] F. W. Aston, Niels Bohr, O. Hahn, W. D. Harkins, F. Joliot, R. S. Mulliken, and M. L. Oliphant. Fifth report of the Committee on Atoms of the International Union of Chemistry. The Analyst, 65(??):568b– 568b, ???? 1940. CODEN ANALAO. ISSN 0003-2654 (print), 1364-5528 (electronic).

Aston:1940:FRCb

[ABH^{+}40b] F. W. Aston, Niels Bohr, O. Hahn, W. D. Harkins, F. Joliot, R. S. Mulliken, and M. L. Oliphant. Fifth report of the Committee on Atoms of the International Union of Chemistry. Journal of the Chemical Society, ??(??):1416–1417, ???? 1940. CODEN JCSOA9. ISSN 0368-1769 (print), 2050-5574 (electronic).

Amaldi:1998:CPE

[ABP98] Edoardo Amaldi, Giovanni Battimelli, and Giovanni Paoloni, editors. 20th Century Physics: Essays and Recollections: a Selection of Historical Writings, volume 3 of Edoardo Amaldi Foundation series. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-2369-2. xiii + 747 pp. LCCN Q127.I8

A53 1998. URL http://www.worldscientific.com/worldscibooks/
10.1142/2852.

Achinstein:1993:HDT

[Ach93] Peter Achinstein. How to defend a theory without testing it: Niels Bohr and the “logic of pursuit”. Midwest Studies In Philosophy, 18(1): 90–120, September 1993. ISSN 0363-6550 (print), 1475-4975 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1111/j.1475-4975.
1993.tb00259.x/abstract.

Aczel:2001:EGM

[Acz01] Amir D. Aczel. Entanglement: The Greatest Mystery in Physics: The Unlikely Story of How Scientists, Mathematicians, and Philosphers Proved Einstein’s Spookiest Theory. Four Walls Eight Windows, New York, NY, USA, 2001. ISBN 1-56858-232-3. xviii + 284 pp. LCCN QC174.12 .A29 2001. URL http://www.loc.gov/catdir/

enhancements/fy0831/2002069338-d.html.

Adams:1972:FGN

[Ada72] J. B. Adams, F.R.S. Four generations of nuclear physicists. Notes and Records of the Royal Society of London, 27(1):75–94, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Adler:1964:BOU

[Adl64] David Jens Adler. Barndom og ungdom. (Danish) [Childhood and youth]. In Bohr and Rozental [BR64], pages 9–35. LCCN QC16.B63 N5. English translation in [Roz67].

Adler:2003:HPM

[Adl03] Mortimer J. Adler, editor. Henri Poincaré, Max Planck, Alfred North Whitehead, Albert Einstein, Sir Arthur Eddington, Niels Bohr, G. H. Hardy, Werner Heisenberg, Erwin Schrödinger, Theodosius Dobzhansky, C. H. Waddington, volume 56 of Great books of the Western world. Encyclopaedia Britannica, Chicago, IL, USA, 2003. ISBN 0-85229-531-6. xvi + 749 pp. LCCN AC1 .G72 1990.

Ammari:2017:BCP

[AF17] Zied Ammari and Marco Falconi. Bohr’s correspondence principle for the renormalized Nelson model. SIAM Journal on Mathematical Analysis, 49(6):5031–5095, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adler:1990:GBW

[AFG90] Mortimer Jerome Adler, Clifton Fadiman, and Philip W. Goetz, editors. Great books of the Western world. Encyclopaedia Britannica, Chicago, IL, USA, second edition, 1990. ISBN 0-85229-531-6 (set). 749 pp. LCCN AC1 .G72 1990.

Asquith:1981:PPB

[AG81] Peter D. Asquith and Ronald N. (Ronald Nelson) Giere, editors. PSA 1980: proceedings of the 1980 Biennial Meeting of the Philosophy of Science Association, volume 2. Philosophy of Science Association, East Lansing, MI, USA, 1981. ISBN 0-917586-16-6. LCCN ????

Allaby:2002:MS

[AG02] Michael Allaby and Derek Gjertsen. Makers of Science. Makers of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-521680-6 (set). 96 (vol. 1) pp. LCCN Q141 .A44 2002.

URL http://www.gbv.de/dms/goettingen/360423353.pdf;   http:
//www.loc.gov/catdir/enhancements/fy0611/2001048396-d.html.

Aaserud:2013:LLQ

[AH13] Finn Aaserud and J. L. Heilbron. Love, Literature, and the Quantum Atom: Niels Bohr’s 1913 Trilogy Revisited. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-968028-0 (hardcover). viii + 284 pp. LCCN QC774.B64 A19 2013.

Arabatzis:2015:RMA

[AI15] Theodore Arabatzis and Despina Ioannidou. The role of models and analogies in the Bohr atom. In Aaserud and Kragh [AK15], pages 360–376. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/2015/
06/Bohr-Book.pdf.

Aaserud:2015:OHY

[AK15] Finn Aaserud and Helge Kragh, editors. One hundred years of the Bohr atom: proceedings from a conference, volume 1 of Scientia Danica. Series M, Mathematica et physica. Det Kongelige Danske Videnskabernes Selskab, Copenhagen, Denmark, 2015. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Albert:1992:BRE

[Alb92] David Z. Albert. Bohr’s response to Einstein, Podolsky, and Rosen. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 17, pages 269–272. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346. LCCN Q174 .S354 1992. URL https://link.

springer.com/chapter/10.1007/978-94-011-2688-5_17.

Allison:1956:BBB

[All56] Samuel K. Allison. Books: Niels Bohr and the Development of Physics, edited by W. Pauli with the assistance of L. Rosenfeld and V. Weisskopf. Bulletin of the Atomic Scientists, 12(2):61, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Altermatt:1982:ICC

[Alt82] Robert Eugene Altermatt. Interactions of Cadmium, Copper and Zinc with Hemoglobin and the Bohr Effect. Ph.D dissertation, California State University, Long Beach, Long Beach, CA, USA, 1982. 78 pp.

Ali:2006:QSB

[AMH06] Md. Manirul Ali, A. S. Majumdar, and Dipankar Home. Quantum superarrivals: Bohr’s wave–particle duality revisited. Foundations of Physics Letters, 19(2):179–186, April 2006. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Amos:1978:BRB

[Amo78] D. E. Amos. Book review: A Definite Integral of N. Bohr (P. J. Schweitzer). SIAM Review, 20(1):188–190, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Amusia:2014:BMM

[Amu14] M. Y. Amusia. Bohr’s molecular model and the melding of classical and quantum mechanics. Physics Today, 67(8):8, August 2014. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [SSH14a].

Andersen:1964:SDI

[And64] Mogens Andersen. En stemning. (Danish) [An impression]. In Bohr and Rozental [BR64], pages 311–314. LCCN QC16.B63 N5. English translation in [Roz67].

Anderson:1973:TQA

[And73] Herbert Anderson. Three questions about the sustained nuclear chain reaction. The University of Chicago Magazine, 65(5):3–7, March/April

1973. ISSN 0041-9508. URL http://library.ucsd.edu/dc/object/
bb3960604v.

Andersen:1989:BRP

[And89] Per H. Andersen. Book review: Pekka Lahti and Peter Mittelstaedt (editors), Symposium on the Foundations of Modern Physics 1987: The Copenhagen Interpretation 60 Years after the Como Lecture. Physics Today, 42(9):92, 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/v42/i9/
p92/s1.

Angeloni:2010:CRB

[Ang10] Roberto Angeloni. On the cultural relationship between Niels Bohr and Harald Høffding. Nuncius, 25(2):317–356, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/10.1163/18253911-
88000015.

Anonymous:1921:UAI

[Ano21] Anonymous. Universitetets Atom-Institute, der Indvies i Dag. (Danish) [The University’s Atomic Institute, which is inaugurated today]. Politiken, ??(??):??, March 3, 1921. Partially reproduced in [Kra85, page 52].

Anonymous:1922:SBA

[Ano22] Anonymous. Steam baths among the California Indians, a possible reconciliation of the atomic models of Bohr and of Lewis and Langmuir, and more. Scientific American, 127(5):327, November 1922. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v127/
n5/pdf/scientificamerican1122-327.pdf.

Anonymous:1923:EZJ

[Ano23a] Anonymous. Die Ersten zehn Jahre der Theorie von Niels Bohr über den Bau der Atome. (German) [The first ten years of Niels Bohr’s theory of the structure of atoms]. Die Naturwissenschaften, 11(27):536–624, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Anonymous:1923:DHD

[Ano23b] Anonymous. Dinner in honor of Dr. Bohr. Science (New Series), 58 (1513):533–534, December 28, 1923. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1923:DAN

[Ano23c] Anonymous. Discusses atom from new point: Dr. Bohr, in Yale lecture takes up the theory of spectral lives. Takes up Planck formula. He shows how it is employed to explain the spectra of the elements. New York Times, ??(??):17, November 14, 1923. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/103174188/.

Anonymous:1923:DBD

[Ano23d] Anonymous. Dr. Bohr discusses bombarding atoms: Tells his Yale audience how impinging electrons produce different states: Tests in mercury vapor: Bombarded by electrons, it emits spectral lines equal to spectrum of electron. New York Times, ??(??):17, November 9, 1923. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/103145693/fulltextPDF/13F38D78EA15C2D5127.

Anonymous:1928:EPR

[Ano28] Anonymous, editor. Électrons et photons: rapports et discussions du cinquième conseil de physique tenu à Bruxelles du 24 au 29 octobre 1927 sous les auspices de l’Institut International de Physique Solvay. (French) [Electrons and photons: reports and discussions of the Fifth Congress of Physics held in Brussels from 24 to 29 October 1927 under the auspices of the Solvay International Institute of Physics]. Gauthier-Villars et cie, Paris, France, 1928. viii + 287 pp. LCCN ????

Anonymous:1929:FKK

[Ano29] Anonymous. Fysiker-Kongres i Kjøbenhavn: Niels Bohr har paa sit Institut samlet en Række berømte Fysikere til Kongres, der aabnedes her i Byen i Gaar. (Danish) [Physicists meeting in Copenhagen: Niels Bohr has gathered at his Institute a collection of famous physicists to a meeting that opened here in town yesterday]. Politiken, ??(??):??, April 9, 1929. Reproduced in [Wei85b, page 21].

Anonymous:1936:NVNa

[Ano36a] Anonymous. News and views: Neutron capture and nuclear constitution. Nature, 137(3461):351, February 29, 1936. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

www.nature.com/nature/journal/v137/n3461/pdf/137351a0.pdf.

Anonymous:1936:NVNb

[Ano36b] Anonymous. News and views: Nuclear energy levels. Nature, 137 (3461):351, February 29, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v137/n3461/pdf/137351b0.pdf.

Anonymous:1939:EFU

[Ano39a] Anonymous. 6 elements found in uranium atom: Physicists bare discovery of greatest amount of energy liberated thus far: Report widely hailed: Professors Bohr and Fermi, at Columbia meeting, tell of atomic ‘cannon ball’. New York Times, ??(??):17, February 25, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:/

/search.proquest.com/hnpnewyorktimes/docview/102952660/.

Anonymous:1939:AEF

[Ano39b] Anonymous. Atom explosion frees 200,000,000 volts; new physics phenomenon credited to Hahn. New York Times, ??(??):2, January 29, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL http://search.proquest.com/hnpnewyorktimes/

docview/102763891. From the article: “American scientists heard today of a new phenomenon in physics — explosion of atoms with a discharge of 200,000,000 volts of energy. … Dr. Enrico Fermi of the University of Rome told yesterday that this had been accomplished by Dr. G. [sic] Hahn of Berlin. … Scientists at the meeting said the discovery was comparable in significance to the original discovery of radioactivity thirty years ago.”.

Anonymous:1939:PHD

[Ano39c] Anonymous. Physicists here debate whether experiment will blow up two miles of the landscape. Washington Post, ??(??):??, April 29, 1939. ISSN 0190-8286.

Anonymous:1939:VEF

[Ano39d] Anonymous. Vast energy freed by uranium atom: Split, it produces 2 ‘cannonballs’, each of 100,000,000 electron volts: Hailed as epoch making; new process, announced at Columbia, uses only 1–30 volt to liberate big force. New York Times, ??(??):18, January 31, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:/

/search.proquest.com/hnpnewyorktimes/docview/102759255/.

Anonymous:1939:VER

[Ano39e] Anonymous. Vision Earth rocked by isotope blast: Scientists say bit of uranium could wreck New York. New York Times, ??(??), April 30, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/102790674/fulltextPDF.

Anonymous:1943:SRL

[Ano43] Anonymous. Scientist reaches London: Dr. N. H. D. Bohr, Dane, has a new atomic blast invention. New York Times, ??(??):3, October 9, 1943. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/106622843/19C4D8047E2421EPQ.

Anonymous:1957:BWA

[Ano57a] Anonymous. Bohr wins Atoms for Peace Prize. Physics Today, 10(5): ??, May 1957. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1957:PFA

[Ano57b] Anonymous. The Presentation of the first Atoms for Peace Award to Niels Henrik David Bohr at the National Academy of Sciences, Washington, DC, October 24, 1957. National Academy of Sciences, Washington, DC, USA, 1957. 33 pp. LCCN QC792.7 .P73 1957.

Anonymous:1962:ALD

[Ano62a] Anonymous. Avec lui disparait une grande figure de la physique contemporaine nous déclare M. Louis de Broglie. (French) [With him [Niels Bohr] disappears a great figure of contemporary physics, says M. Louis de Broglie]. Le Monde [Paris], ??(??):??, November 20,

1962. ISSN 0395-2037. URL http://www.lemonde.fr/archives/
article/1962/11/20/avec-lui-disparait-une-grande-figure-de-
la-physique-contemporaine-nous-declare-m-louis-de-broglie_
2360411_1819218.html.

Anonymous:1962:DNB

[Ano62b] Anonymous. Dr. Niels Bohr, head-and-shoulders portrait, facing

slightly left, 1962. URL http://hdl.loc.gov/loc.pnp/cph.3c12063.

1 photographic print.

Anonymous:1962:NBA

[Ano62c] Anonymous. Niels Bohr, atom pioneer and Nobel winner, dies. Dane, 77, helped to draft theory of fission — won Ford Fund Award. New York Times, pages 1, 35, November 19, 1962. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/docview/116246757?.

Anonymous:1963:NHD

[Ano63a] Anonymous. Niels Henrik David Bohr (1885–1962). Physics Today, 16(10):21, October 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1963:ONB

[Ano63b] Anonymous. Obituary: Niels Bohr. Physics Today, 16(1):??, January 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1964:KOD

[Ano64] Anonymous. Kronologisk oversigt. (Danish) [Chronological overview]. In Bohr and Rozental [BR64], page 341. LCCN QC16.B63 N5. English translation in [Roz67].

Anonymous:1965:DBM

[Ano65] Anonymous. Danish Bohr medal to Piotr Leonidovich Kapitza. Physics Today, 18(8):??, August 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1968:NBL

[Ano68a] Anonymous. Niels Bohr Library at AIP shows Faraday memorabilia. Physics Today, 21(5):??, May 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1968:NBM

[Ano68b] Anonymous. Niels Bohr Medal to Isador I. Rabi. Physics Today, 21 (1):163–??, January 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1971:YNB

[Ano71] Anonymous. 50 years Niels Bohr Institute, Copenhagen. Europhysics News, 2(7):3–4, ???? 1971. CODEN EUPNAS. ISSN 0531-7479 (print),

1432-1092 (electronic). URL https://www.europhysicsnews.org/
articles/epn/abs/1971/07/epn19710207p3/epn19710207p3.html.

Anonymous:1985:BSR

[Ano85a] Anonyme. Bibliographie se rapportant au concept de complémentarité et à ses prolongements. (French) [Bibliography concerning the concept of complementarity and its extensions]. Revue d’Histoire des Sciences, 38(3–4):353–363, juillet–décembre 1985. CODEN RHSAAM.

ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632503;   http:
//www.persee.fr/web/revues/home/prescript/article/rhs_0151-
4105_1985_num_38_3_4012.

Anonymous:1985:BCC

[Ano85b] Anonymous. Bibliographie concernant le concept de complémentarité et ses prolongements. (French) [Bibliography concerning the concept of complementarity and its extensions]. Revue d’Histoire des Sciences, 38(3–4):353–363, 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0048-7996, 0151-4105, 1969-6582 (electronic).

URL http://www.persee.fr/web/revues/home/prescript/article/
rhs_0151-4105_1985_num_38_3_4012.

Anonymous:1985:NB

[Ano85c] Anonymous. Niels Bohr. Physics Today, 38(10):??, October 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1985:NBI

[Ano85d] Anonymous. Niels Bohr and the infinitely small, volume 137 (v. 35, no. 1) of Impact of science on society. Unesco, Paris, France, 1985. ISSN 0019-2872. 81 + 4 pp. LCCN QC16.B63 N5 1985.

Anonymous:1985:OHA

[Ano85e] Anonymous. On the one-hundredth anniversary of the birthday of Niels Bohr. Soviet Physics. Uspekhi, 28(10):853, October 31, 1985. CO-

DEN SOPUAP. ISSN 0038-5670. URL http://stacks.iop.org/0038-
5670/28/i=10/a=E01.

Anonymous:1985:WNB

[Ano85f] Anonymous. Works by Niels Bohr. In French and Kennedy [FK85], pages 385–391. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Anonymous:1991:CF

[Ano91] Anonymous. The contamination factory. Bulletin of the Atomic Scientists, 47(8):34–39, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BRU

[Ano93a] Anonymous. Book reviews: Unveiling the atom [Abraham Pais, Niels Bohr’s Times, in Physics, Philosophy and Polity. Clarendon Press, Oxford, 1991. Pp. xvii + 565, £25.00. ISBN 0-19-852049-2]. Notes and Records of the Royal Society of London, 47(1):152–154, July 1993. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1993:JBF

[Ano93b] Anonymous. Das Jahr 1945 Die Bohr-Festschrift und der Nobelpreis (German) [The year 1945: The Bohr Anniversary Publication and the Nobel Prize]. In von Meyenn [vM93], pages 253–336. ISBN 3-540-54911-0. ISSN 0172-6315. LCCN QC16.P37 A34. URL http://www.springer.

com/physics/book/978-3-540-54911-6.

Anonymous:1993:RWS

[Ano93c] Anonymous. The rest was silence. New York Times, ??(??):BR3, March 14, 1993. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/109207251/.

Anonymous:1994:EOL

[Ano94a] Anonymous. Errata: The Origin of the Liquid-Drop Model and the Interpretation of Nuclear Fission. Perspectives on Science, 2(2):254, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Correction of printers error in bottom three equations from page 90.

Anonymous:1994:NBH

[Ano94b] Anonymous. Niels Bohr, 1994. 1 videocassette (58 minutes).

Anonymous:1995:BNE

[Ano95a] Anonymous. Bohr and nuclear espionage. Scientific American, 273 (3):10–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SIA

[Ano95b] Anonymous. Special investigation: The atomic intrigues of Niels Bohr. Scientific American, 272(5):83, May 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n5/pdf/scientificamerican0595-83.pdf.

Anonymous:1997:ISD

[Ano97] Anonymous, editor. International symposium Dubna 14–18 May 1996: science and society, history of the Soviet atomic project (40’s–50’s), proceedings. Dubna, Moscow, Russia, 1997. ISBN 5-86656-073-9 (hardcover). LCCN ????

Anonymous:19xx:NB

[Anoxx] Anonymous. Niels Bohr, 19xx. URL http://hdl.loc.gov/loc.pnp/ cph.3b33671. 1 photographic print. Year unknown.

Anonymous:2001:HMS

[Ano01] Anonymous. How to mix sweets. Physics in Perspective (PIP), 3(1): 132, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2002:AVB

[Ano02a] Anonymous. Angst vor der Bombe: Carl Friedrich von Weizäcker über die neuen Dokumente und sein Treffen mit Bohr und Heisenberg. (German) [Fear of the bomb: Carl Friedrich von Weizäcker on the new documents and his meeting with Bohr and Heisenberg]. Süddeutsche Zeitung, ??(??):??, February 8, 2002.

Anonymous:2002:CIS

[Ano02b] Anonymous, editor. The Copenhagen interpretation: science and history on stage. Graduate Center of the City University of New York and Smithsonian Institute, New York, NY, USA, 2002. Three videocassetes recorded at the Graduate Center of the City University of New York, Baird Auditorium, March 2, 2002. Apparently not published in print.

Anonymous:2004:BR

[Ano04a] Anonymous. Bohr and the rabbi. Physics in Perspective (PIP), 6 (3):367, September 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2004:HHT

[Ano04b] Anonymous. How to hit a telephone pole. Physics in Perspective (PIP), 6(4):486, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2005:CBD

[Ano05] Anonymous. The Coins and Banknotes of Denmark. Danmarks Nationalbank, København, Danmark, 2005. ISBN 87-87251-55-8, 87-87251-56-6. 28 pp. LCCN ???? URL http://www.

gbv.de/dms/zbw/724844228.pdf;   http://www.nationalbanken.
dk/C1256BE900406EF3/sysOakFil/Danmarks_penge_2005_ENG/%24File/
Coins_Banknotes.pdf;   http://www.nationalbanken.dk/DNUK/NotesAndCoins.
nsf/side/500-krone!OpenDocument. See pages 1, 12, 20–21 for color

pictures of the DKR 500 banknote with Niels Bohr’s portrait.

Anonymous:2009:BRQ

[Ano09] Anonymous. Book review: Quantum story leaps off the page: Quantum: Einstein, Bohr and the Great Debate About the Nature of Reality. Physics Education, 44(1):100, 2009. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://stacks.iop.org/0031-

9120/44/i=1/a=M04.

Anonymous:2010:NBS

[Ano10] Anonymous. Niels Bohr Science Park, Københavns Universitet. Universitets- og Bygningsstyrelsen, København, Danmark, 2010. ISBN ???? 40 pp. LCCN ????

Anonymous:2013:NBA

[Ano13] Anonymous. Niels Bohrs atommodel 1913–2013. (Danish) [Niels Bohr’s atomic model 1913–2013]. Web site at the Department of Physics, Aarhus University, Denmark for the 100th anniversary, 2013. URL http:

//phys.au.dk/niels-bohrs-atommodel-1913-2013/.

Anonymous:20xx:WCT

[Anoxx] Anonymous. Washington conferences on theoretical physics. Web

document, 20xx. URL http://home.gwu.edu/~kargaltsev/HEA/

washington-conferences.html. Undated. The page includes a photograph of a plaque with the preface “The most famous event at this 5th Washington Conference on Theoretical Physics came from the announcement by Niels Bohr at the 1939 conference, in the Hall of Government, Room 209, that the nucleus of uranium had been split by bombardment with neutrons, with significant energy released. This was the dawn of the atomic age.” and the engraving: “In this room, January 26, 1939, Niels Bohr made the first pubic announcement of the successful disintegration of uranium into barium with the attendant release of approximately two hundred million electron volts of energy per disintegration. This announcement was heard by the physicists listed below who where attending the fifth of the conferences on theoretical physics which are sponsored jointly by the Carnegie Institution of Washington and The George Washington University.” The participant listed on the plaque are: L. H. Adams; Donald Hatch Andrews; Ferdinand G. Brickwedde; Gerhard Heinrich Dieke; George A. Gamow; Maria Goeppert-Mayer; M. H. Hebb; Karl Ferdinand Herzfeld; J. H. Hibben; J. H. Hoge; D. R. Inglis; F. G. Keyes; F. C. Kracek; R. Myers; H. M. O’Bryan; E. Posnjak; A. E. Ruark; R. B. Scott; Francis B. Silsbee; C. Starr; Otto Stern; Edward Teller; Harold C. Urey; and B. D. van Evera.

Antonopoulos:1996:BNF

[Ant96] Constantin Antonopoulos. Bohr on nonlocality: the facts and the fiction. Philosophia Naturalis, 33(2):205–241, 1996. CODEN ???? ISSN 0031-8027.

Anthes:2015:NSU

[Ant15] Gary Anthes. News: Scientists update views of light. Communications of the ACM, 58(10):15–17, October 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://cacm.acm.org/

magazines/2015/10/192377/fulltext.

Amaldi:1996:ERC

[APB96] Edoardo Amaldi, Giovanni Paoloni, and Giovanni Battimelli, editors. Essays and recollections on 20th century physics: a selection of historical writings, volume 2 of Edoardo Amaldi Foundation series. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2369-2. xiii + 747 pp. LCCN Q127.I8 A53 1996.

Ariyama:1935:NBH

[Ari35] Kanetaka Ariyama. Niels Bohr on his fiftieth birthday. Journal of Jocular Physics, I:1–4, 1935.

Atmanspacher:1999:BSM

[ARK^{+}99] Harald Atmanspacher, Helmut Rechenberg, Horst Kant, Rolf Hempelmann, Reinhold Blümel, Gerhard Börner, Houjun Mo, Philippe Blanchard, Marius Grundmann, Dietrich Stauffer, and Christian Uebing. Bücher und Software: Meÿenn: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a., Band IV/Teil II: 1953– 1954/Mladjenovic: The Defining Years in Nuclear Physics. 1932–1960s/ Frenkel: Professor Friedrich Houtermans Arbeiten, Leben, Schicksal/ Kärger and Heitjans: Diffusion in Condensed Matter/Friedrich: Theoretical Atomic Physics/Peacock: Cosmological Physics/Grosse u Martin: Particle Physics and the Schrödinger Equation/Gaponenko: Optical Properties of Semiconductor Nanocrystals/Gershenfeld: The Nature of Mathematical Modeling/Origin 6.0: Vertrieb. Physikalische Blätter, 55 (9):79–83, September 1999. CODEN PHBLAG. ISSN 0031-9279 (print),

1521-3722 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1002/phbl.19990550919/abstract.

ArroyoCamejo:2006:BED

[Arr06a] Silvia Arroyo Camejo. Die Bohr–Einstein-Debatte. (German) [The Bohr–Einstein debate]. In Skurrile Quantenwelt. (German) [Crazy Quantum World] [Arr06b], pages 87–101. ISBN 3-540-29720-0. LCCN

QC174.12 .A77 2006. US$. URL http://www.springer.com/physics/
quantum+physics/book/978-3-540-29720-8.

ArroyoCamejo:2006:SQG

[Arr06b] Silvia Arroyo Camejo. Skurrile Quantenwelt. (German) [Crazy Quantum World]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-29720-0. xiii + 246 pp. LCCN QC174.12 .A77 2006. EUR 29.95, SFR 51.00. URL

http://www.springer.com/physics/quantum+physics/book/978-3-
540-29720-8.

Aylesworth:1926:DCA

[Ayl26] Evelyn Frances Aylesworth. The Dielectric Constant of Atomic Hydrogen from the Point of View of Bohr’s Quantum Theory. Ph.D dissertation, University of California, Berkeley, Berkeley, CA, USA, 1926. 8 pp.

Aylesworth:1927:DCA

[Ayl27] E. F. Aylesworth. The dielectric constant of atomic hydrogen from the point of view of Bohr’s quantum theory. Proceedings of the National Academy of Sciences of the United States of America, 13(6):438–445, June 1927. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Boggild:1955:SSH

[B^{+}55] J. K. Bøggild et al. Some studies on heavy meson events in stripped emulsions: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(3) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 28 pp. LCCN AS281.

Bohr:1990:GRS

[Bō90] Nīrusu Bōa [Niels Bohr]. Genshi riron to shizen kijutsu. Misuzu Shobō, Tokyo, Japan, 1990. ISBN ???? 486 pp. LCCN ???? Translated by Ken Inoue.

Bohr:2005:NBP

[BA05] Niels Bohr and Finn Aaserud. Niels Bohr: Politics, Popularization and People. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 0-444-51914-9. vi + 76 pp. LCCN QC16.B63 B64 2005.

URL http://www.elsevier.com/authored_subject_sections/P12/
worldyearofphysics/pdf/wyop_booklet.pdf.

Ben-Amotz:2006:RBS

[BA06] Dor Ben-Amotz. Revisiting Bohr’s semiclassical quantum theory. Journal of Physical Chemistry. B. Condensed matter, materials, surfaces, interfaces & biophysical, 110(40):19861–19866, October 12, 2006. CO-DEN JPCBFK. ISSN 1520-6106.

Babich:2002:HPS

[Bab02] Babette E. Babich. Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J., volume 225 of Boston Studies in the Philosophy and History of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346. xvii + 484 pp. LCCN Q174 .B67 vol. 225. URL https://link.springer.

com/book/10.1007/978-94-017-1767-0.

Bacciagaluppi:2015:DBU

[Bac15] Guido Bacciagaluppi. Did Bohr understand EPR? In Aaserud and Kragh [AK15], pages 377–396. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Badash:1971:IBE

[Bad71] Lawrence Badash. The importance of being Ernest Rutherford. Science, 173(4000):873, September 3, 1971. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://adsabs.harvard.

edu/abs/1971Sci...173..873B;   http://www.jstor.org/stable/
1731789.

Badash:1983:BRM

[Bad83] Lawrence Badash. Book review: Mathematical Sciences: Peter Robertson, The early years: the Niels Bohr Institute, 1921–1930. Copenhagen: Akademisk Forlag, Universitetsforlaget i København, 1979. Pp. 175. British Journal for the History of Science, 16(3):296–297, November 1983. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic). URL http://www.jstor.org/stable/4026166.

Baggott:2004:BMM

[Bag04] J. E. Baggott. Beyond measure: modern physics, philosophy, and the meaning of quantum theory. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-852927-9 (hardcover), 0-19-852536-2 (paperback). xvi + 379 pp. LCCN QC174.12 .B33

2004. URL http://catdir.loc.gov/catdir/enhancements/fy0615/
2003066207-d.html.

Baggott:2010:FWP

[Bag10] J. E. Baggott. The first war of physics: the secret history of the atom bomb, 1939–1949. Pegasus Books, New York, NY, USA, 2010. ISBN 1-60598-084-6. xiv + 576 + 16 pp. LCCN QC773 .B24 2010. US$35.00US$35.00.

Baggott:2011:QSH

[Bag11] Jim Baggott. The quantum story: a history in 40 moments. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-956684-4 (hardcover), 0-19-965597-9 (paperback). xix + 469 + 16 pp. LCCN QC173.98 .B34 2011.

Baily:2013:EAM

[Bai13] C. Baily. Early atomic models — from mechanical to quantum (1904– 1913). European Physical Journal H, 38:1–38, January 2013. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http:

//adsabs.harvard.edu/abs/2013EPJH...38....1B.

Ballantine:1939:BRB

[Bal39] Constance Ballantine. Book review: Encyclopedia and Unified Science, Otto Neurath, Niels Bohr, John Dewey, Bertrand Russell, Rudolf Carnap, Charles W. Morris, 1938, vii + 75 pp $1.00. American Mathematical Monthly, 46(3):162–163, March 1939. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.

org/stable/2302467.

Balibar:1985:BEE

[Bal85] Françoise Balibar. Bohr entre Einstein et Dirac. (French) [Bohr between Einstein and Dirac]. Revue d’Histoire des Sciences, 38 (3–4):293–307, 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0048-7996, 0151-4105, 1969-6582 (electronic). URL http:

//www.persee.fr/web/revues/home/prescript/article/rhs_0151-
4105_1985_num_38_3_4008.

Baldwin:2014:KRP

[Bal14] Melinda Baldwin. ‘Keeping in the race’: physics, publication speed and national publishing strategies in Nature, 1895–1939. British Journal for the History of Science, 47(2):257–279, June 2014. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:/

/journals.cambridge.org/abstract_S0007087413000381. Published

online, but not yet assigned to a journal volume.

Bandyopadhyay:2000:WWE

[Ban00] S. Bandyopadhyay. Welcher Weg experiments and the orthodox Bohr’s Complementarity Principle. Physics Letters A, 276(5–6):233–239, November 6, 2000. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-

2429 (electronic). URL http://arxiv.org/abs/quant-ph/0003073v2.

Barker:1923:RPR

[Bar23] E. F. Barker. Recent publications: Reviews: La Théorie de Bohr (Publications de la Société de Chimie-Physique, X), by E. Bauer. American Mathematical Monthly, 30(4):203, May/June 1923. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Barschall:1987:RED

[Bar87] Henry H. Barschall. Reminiscences of the early days of fission. Physics Today, 40(6):??, June 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bartelborth:1989:KRS

[Bar89] Thomas Bartelborth. Kann es rational sein, eine inkonsistente Theorie zu akzeptieren? (Eine Untersuchung zum frühen Bohrschen Atommodell). (German) [Can it be rational to accept an inconsistent theory? (An investigation to the early Bohr atomic model)]. Philosophia Naturalis, 26(1):91–120, 1989. CODEN ???? ISSN 0031-8027.

Bartlett:1999:JSY

[Bar99] Albert A. Bartlett. Just sixty years ago. The Physics Teacher, 37 (1):30–31, January 1999. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Barad:2007:MUH

[Bar07] Karen Michelle Barad, editor. Meeting the Universe Halfway: Quantum Physics and the Entanglement of Matter and Meaning. Duke University Press, Durham, NC, USA, 2007. ISBN 0-8223-3901-3 (hardcover), 0-8223-3917-X (paperback). xiii + 524 pp. LCCN QC6 .B328 2007.

Bastin:1971:QTB

[Bas71] Ted Bastin, editor. Quantum Theory and Beyond: Essays and Discussions Arising from a Colloquium. Cambridge University Press, Cambridge, UK, 1971. ISBN 0-521-07956-X. LCCN QC174.1

.Q37. URL http://www.loc.gov/catdir/enhancements/fy0907/
77127237-d.html;   http://www.loc.gov/catdir/enhancements/
fy0907/77127237-t.html.

Boisseau:2001:CBB

[BAV01] C. Boisseau, E. Audouard, and J. Vigue. Comment on “Breakdown of Bohr’s correspondence principle”. Physical Review Letters, 86(12):2694, March 19, 2001. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Bohr:1987:PWN

[BB87] Niels Bohr and Niels Bohr. The philosophical writings of Niels Bohr: Atomic theory and the description of nature, volume 1. Ox Bow Press, Woodbridge, CT, USA, 1987. ISBN 0-918024-51-X, 0-918024-50-1. 119 pp. LCCN QC5.58 .B64213 1987.

Bohr:1989:COW

[BB^{+}89] Niels Bohr, Niels Barfoed, et al., editors. The challenge of an open world: essays dedicated to Niels Bohr. Munksgaard, Copenhagen, Denmark, 1989. ISBN 87-16-10382-3. 235 pp. LCCN JX1974.7 .C45 1989.

Bokulich:2005:NBG

[BB05] Peter Bokulich and Alisa Bokulich. Niels Bohr’s generalization of classical mechanics. Foundations of Physics, 35(3):347–371, March 2005. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/s10701-004-1979-5.

Bohr:1940:VRR

[BBBL40] N. Bohr, J. K. Bøggild, K. J. Brostrøm, and T. Lauritsen. Velocityrange relation for fission fragments. Physical Review, 58(9):839–840, November 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://link.aps.org/doi/10.1103/PhysRev.58.
839;   http://prola.aps.org/abstract/PR/v58/i9/p839_1.

Bohr:1986:CNA

[BBCN86] Niels Bohr, Anders Boserup, L. (Leif) Christensen, and Ove Nathan, editors. The challenge of nuclear armaments: essays dedicated to Niels Bohr and his appeal for an open world. University of Copenhagen, Copenhagen, Denmark, 1986. ISBN 87-7245-142-4 (paperback). 346 + 16 pp. LCCN JX1974.7 .C46 1986.

Born:1969:AP

[BBSR69] Max Born, R. J. (Roger John) Blin-Stoyle, and J. M. Radcliffe. Atomic Physics. Dover books on physics and chemistry. Dover, New York, NY, USA, eighth edition, 1969. ISBN 0-486-65984-4. xiv + 495 + 11 pp. LCCN QC776 .B5713 1989. US$11.95. URL http://www.loc.

gov/catdir/description/dover031/89012033.html.

Bohr:1923:RPS

[BC23] Niels Bohr and Dirk Coster. Röntgenspektren und periodisches System der Elemente. (German) [X-ray spectra and the periodic system of elements]. Zeitschrift für Physik, 12(1):342–374, December 1923. CO-DEN ZEPYAA. ISSN 0044-3328. URL http://link.springer.com/

article/10.1007/BF01328104.

Bitbol:1992:ESP

[BD92] Michel Bitbol and Olivier Darrigol, editors. Erwin Schrödinger: philosophy and the birth of quantum mechanics = philosophie et naissance de la mécanique quantique. Editions Frontières, Gif-sur-Yvette, France, 1992. ISBN 2-86332-116-1. xiii + 460 pp. LCCN QC16.S265 E78 1992. 263.75F.

Blum:1986:WHG

[BDR86] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abteilung C, Band 4. Allgemeinverständliche Schriften. Biographisches und Kernphysik: Autobiographisches, Laudationes, Buchbesprechungen, Kernphysik, Münchener Festrede u.a. (German) [Werner Heisenberg: Collected Works: Series C, Volume 4: Philosophical and popular writings. Biographical and nuclear physics. Autobiographical, Laudations, Book Reviews, Nuclear Physics, Munich Festival Speeches, a.o.]. Piper, München, West Germany, 1986. ISBN 3-492-02928-0. x + 505 pp. LCCN ????

Becker:1957:PAF

[Bec57] Bill Becker. Pioneer of the atom: Forty-four years ago Niels Bohr unlocked the secret of atomic structure. this week he will receive the first Atoms for Peace Award. New York Times, ??(??):SM52–SM53, October 20, 1957. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/114089545/.

Becker:2018:WRU

[Bec18] Adam Becker. What is Real?: the Unfinished Quest for the Meaning of Quantum Physics. Basic Books, New York, NY, USA, first edition, 2018. ISBN 0-465-09605-0 (hardcover), 0-465-09606-9. ix + 370 pp. LCCN QC173.98 .B43 2018.

Behrens:1941:EDB

[Beh41] Carl Ernest Behrens. The early development of the Bohr atom, 1913– 1915. M.S. dissertation, The University of Chicago, Chicago, IL, USA, 1941. 61 pp.

Behrens:1943:AT

[Beh43a] Carl E. Behrens. Atomic theory from 1904 to 1913. American Journal of Physics, 11(2):60–66, April 1943. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/resource/

1/ajpias/v11/i2/p60_s1.

Behrens:1943:EDB

[Beh43b] Carl E. Behrens. The early development of the Bohr atom. American Journal of Physics, 11(3):135–147, June 1943. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v11/i3/p135_s1.

Behrens:1943:FDB

[Beh43c] Carl E. Behrens. Further developments of Bohr’s early atomic theory. American Journal of Physics, 11(5):272–281, October 1943. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:/

/ajp.aapt.org/resource/1/ajpias/v11/i5/p272_s1.

Beijerinck:2013:CB

[Bei13] Herman C. W. Beijerinck. Crossing borders. Europhysics News, 44 (4):21, July/August 2013. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic).

Bell:1964:EPR

[Bel64] John S. Bell. On the Einstein–Podolsky–Rosen paradox. Physics, 1(3):

195–200, 1964. ISSN ???? URL http://en.wikipedia.org/wiki/EPR_
paradox;   http://plato.stanford.edu/entries/bell-theorem/;
 http://scienceworld.wolfram.com/physics/Einstein-Podolsky-
RosenParadox.html;   http://www.ncsu.edu/felder-public/kenny/
papers/bell.html;   http://www.physics.princeton.edu/~mcdonald/
examples/QM/bell_physics_1_195_64.pdf. Reprinted in [Bel87, pp.

14–21].

Beller:1986:BRW

[Bel86] Mara Beller. Book review: Wie es eigentlich gewesen? [How was it actually?] The Creation of Quantum Mechanics and the Bohr–Pauli Dialogue. Isis, 77(1):107–109, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232508.

Bell:1987:SUQ

[Bel87] John S. Bell. Speakable and unspeakable in quantum mechanics: collected papers on quantum philosophy. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33495-0. xii + 212 pp. LCCN QC173.97 .B45 1987.

Beller:1992:BBC

[Bel92a] Mara Beller. The birth of Bohr’s complementarity: The context and the dialogues. Studies in History and Philosophy of Science Part A, 23(1):147–180, March 1992. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.com/

science/article/pii/0039368192900296.

Beller:1992:GBC

[Bel92b] Mara Beller. The genesis of Bohr’s complementarity principle and the Bohr–Heisenberg dialogue. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 18, pages 273–293. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346. LCCN Q174 .S354

1992. URL https://link.springer.com/chapter/10.1007/978-94-
011-2688-5_18.

Beller:1993:EBR

[Bel93] Mara Beller. Einstein and Bohr’s rhetoric of complementarity. Science in Context, 6(1):241–255, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Beller:1999:JCC

[Bel99a] Mara Beller. Jocular commemorations: The Copenhagen spirit. Osiris (Series 2), 14(??):252–273, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.jstor.org/

stable/301971.

Beller:1999:QDM

[Bel99b] Mara Beller. Quantum dialogue: the making of a revolution. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1999. ISBN 0-226-04181-6 (hardcover), 0-226-04182-4. xv + 365 + 8 pp. LCCN QC174.13 .B45 1999.

URL http://www.loc.gov/catdir/description/uchi052/99035499.
htm;   http://www.loc.gov/catdir/toc/uchi052/99035499.htm.

Bentzen:2000:LMN

[Ben00] Soren M. Bentzen. Lise Meitner and Niels Bohr: A historical note. Acta Oncologica (Stockholm), 39(8):1002, ???? 2000. ISSN 0284-186X.

Benedek:2013:LEA

[Ben13] Giorgio Benedek. Letter to the editor: About “Crossing borders”. Europhysics News, 44(6):32, November/December 2013. URL http:

//www.europhysicsnews.org/articles/epn/pdf/2013/06/epn2013-

44-6.pdf. Brief comment on [Bei13, Luc07].

Bergwitz:1915:AGP

[Ber15] Karl Bergwitz, editor. Arbeiten aus den Gebieten der Physik, Mathematik, Chemie: Festschrift Julius Elster und Hans Geitel zum 60. Geburtstag. (German) [Works from the fields of physics, mathematics, chemistry: Festschrift for Julius Elster and Hans Geitel’s 60th birthday]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1915. xi + 719 pp. LCCN QD455 .F47 1915. URL http://catalog.hathitrust.

org/api/volumes/oclc/47733990.html.

Berry:1989:HGN

[Ber89] R. Stephen Berry. How good is Niels Bohr’s atomic model? Contemporary Physics, 30(1):1–19, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bernstein:1991:KQ

[Ber91] J. Bernstein. King of the quantum: Niels Bohr’s Times, In Physics, Philosophy, and Polity by Abraham Pais. New York Review of Books, 38(15):61–63, September 26, 1991. ISSN 0028-7504 (print),

1944-7744 (electronic). URL http://www.nybooks.com/articles/
archives/1991/sep/26/king-of-the-quantum/.

Bernstein:1995:WDH

[Ber95] Jeremy Bernstein. What did Heisenberg tell Bohr about the bomb? Scientific American, 272(5):92–97 (Intl. ed. 72–??), May 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n5/pdf/scientificamerican0595-92.pdf.

Bernstein:1996:HUC

[Ber96] Jeremy Bernstein. Hitler’s uranium club: the secret recordings at Farm Hall. American Institute of Physics, Woodbury, NY, USA, 1996. ISBN 1-56396-258-6. xxx + 427 + 4 pp. LCCN QC773.3.G3 B47 1995. US$34.95. Introduction by David Cassidy.

Bernstein:2003:DWH

[Ber03] Jeremy Bernstein. The drawing, or why history is not mathematics. Physics in Perspective (PIP), 5(3):243–261, 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bethe:1937:NPB

[Bet37] H. A. Bethe. Nuclear physics B. Nuclear dynamics, theoretical. Reviews of Modern Physics, 9(2):69–244, April 1, 1937. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.9.69;
http://rmp.aps.org/abstract/RMP/v9/i2/p69_1.

Bethe:1985:NBH

[Bet85] Hans Bethe. Niels Bohr and his institute. In French and Kennedy [FK85], pages 232–234. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bethe:1991:RAP

[Bet91] Hans Albrecht Bethe. The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science. Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1991. ISBN 0-88318-707-8, 0-671-74012-1. xvii + 286 pp. LCCN U264 .B455 1991. US$24.95.

Beyer:1949:FNP

[Bey49] Robert T. (Robert Thomas) Beyer, editor. Foundations of nuclear physics: facsimiles of thirteen fundamental studies as they were originally reported in the scientific journals. Dover, New York, NY, USA, 1949. 272 pp. LCCN QC173 .B485. See original paper [HS39b] and later translation [Gra64].

Beller:1994:BRE

[BF94] M. Beller and A. Fine. Bohr’s response to EPR. In Faye and Folse [FF94], pages 1–32. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67

vol. 153 1994; QC16 1994. URL http://www.loc.gov/catdir/
enhancements/fy0823/93024825-d.html;   http://www.loc.gov/
catdir/enhancements/fy0823/93024825-t.html;   https://link.
springer.com/book/10.1007/978-94-015-8106-6.

Bohr:1998:CCS

[BFF98] Niels Bohr, Jan Faye, and Henry J. Folse. Causality and Complementarity: Supplementary Papers, volume 4 of The philosophical writings of Niels Bohr. Ox Bow Press, Woodbridge, CT, USA, 1998. ISBN 1-881987-14-0 (hardcover), 1-881987-13-2 (paperback). vii + 191 pp. LCCN QC6.4.C3 B64 1998.

Bothe:1924:WEN

[BG24] W. Bothe and H. Geiger. Ein Weg zur experimentellen Nachprüfung der Theorie von Bohr, Kramers, und Slater. (German) [One way of experimental verification of the theory of Bohr, Kramers, and Slater]. Zeitschrift für Physik, 26(1):44, December 1924. CODEN ZEPYAA. ISSN 0044-3328.

Bethe:1995:DBS

[BGS95] Hans A. Bethe, Kurt Gottfried, and Roald Z. Sagdeev. Did Bohr share nuclear secrets? Scientific American, 272(5):84–90, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n5/pdf/scientificamerican0595-84.pdf.

Born:1923:PBM

[BH23] Max Born and Werner Heisenberg. Über Phasenbeziehungen bei den Bohrschen Modellen von Atomen und Molekeln. (German) [On phase relations in the Bohr model of atoms and molecules]. Zeitschrift für Physik, 14(1):44–55, December 1923. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.com/

article/10.1007/BF01340032;   http://www.springerlink.com/
content/v71005762t157n89/.

Bohr:1964:BAG

[BH64] Niels Bohr and Armin Hermann. Das Bohrsche Atommodell. (German) [The Bohr atomic model], volume 5 of Dokumente der Naturwissenschaft. Abteilung Physik. E. Battenberg, Stuttgart, West Germany, 1964. 122 pp. LCCN QC173 .B537. Entwicklung der Atomtheorie bis Niels Bohr and Niels Bohr, 1885–1962. Abhandlungen über Atombau aus den Jahren 1913–1916. [Development of the atomic theory of Niels Bohr (1885–1962). Papers on atomic structure from the years 1913–1916.].

Brown:1982:BEP

[BH82] Laurie M. Brown and Lillian Hoddeson. The birth of elementaryparticle physics. Physics Today, 35(4):36–43, April 1982. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

www.physicstoday.org/resource/1/phtoad/v35/i4/p36_s1. This is

a much-abridged version of [BH83a].

Brown:1983:BEP

[BH83a] Laurie M. Brown and Lillian Hoddeson. The birth of elementary particle physics. In The Birth of Particle Physics [BH83b], pages 3–36. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Brown:1983:BPP

[BH83b] Laurie M. Brown and Lillian Hoddeson, editors. The Birth of Particle Physics. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Booss:2003:MW

[BH03] Bernhelm Booss and Jens Høyrup, editors. Mathematics and war. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2003. ISBN 3-7643-1634-9, 0-8176-1634-9. viii + 416 pp. LCCN QA10.8 .M38 2003.

Bianchi:2011:DVS

[BH11] Eugenio Bianchi and Hal M. Haggard. Discreteness of the volume of space from Bohr–Sommerfeld quantization. Physical Review Letters, 107 (1):011301, July 1, 2011. CODEN PRLTAO. ISSN 1079-7114.

Bohr:2008:NBC

[BHK^{+}08] Niels Bohr, Ulrich Hoyer, Jørgen Kalckar, J. Rud Nielsen, Rudolf Ernst Peierls, Léon Rosenfeld, Erik Rüdinger, Klaus Stolzenburg, Jens Thorsen, Finn Aaserud, and David Favrholdt, editors. Niels Bohr: collected works. Elsevier, Amsterdam, The Netherlands, 2008. ISBN 0-444-53286-2 (set), 0-444-53283-8 (vol. 1), 0-444-53282-X (vol. 2), 0-444-53287-0 (vol. 3), 0-444-53288-9 (vol. 4), 0-444-53279-X (vol. 5), 0-444-53289-7 (vol. 6), 0-444-53290-0 (vol. 7), 0-444-53284-6 (vol. 8), 0-444-53277-3 (vol. 9), 0-444-53278-1 (vol. 10), 0-444-53280-3 (vol. 11), 0-444-53281-1 (vol. 12), 0-444-53291-9 (vol. 13). various pp. LCCN QC3 .B584; QC771 .B63 2008. URL http://www.elsevier.

com/wps/find/bookdescription.cws_home/BS_NB/description;
http://www.elsevierdirect.com/brochures/nielsbohr/;   http:
//www.sciencedirect.com/science/bookseries/18760503.

Badash:1986:NFR

[BHT86] Lawrence Badash, Elizabeth Hodes, and Adolph Tiddens. Nuclear fission: Reaction to the discovery in 1939. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 130(2):196–231, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.org/stable/

987181.

Birks:1962:RM

[Bir62] J. B. Birks, editor. Rutherford at Manchester. Heywood and Company LTD, London, UK, 1962. x + 364 pp. LCCN ????

Birks:1963:RM

[Bir63] J. B. (John Betteley) Birks, editor. Rutherford at Manchester. Benjamin, New York, NY, USA, 1963. x + 364 pp. LCCN QC16.R8 B57 1962.

Bishop:2004:BRS

[Bis04] Robert C. Bishop. Book review: Steen Brock: Niels Bohr’s Philosophy of Quantum Physics in the Light of the Helmholtzian Tradition of Theoretical Physics. Isis, 95(2):334–335, June 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/10.1086/426269.

BastholmJensen:2003:DH

[BJ03] Mette Bastholm Jensen and Steven L. B. Jensen, editors. Denmark and the Holocaust, volume 3 of Danish genocide studies series. Institute for International Studies, Department for Holocaust and Genocide Studies, København, Danmark, 2003. ISBN 87-989305-1-6 (paperback). ISSN 1602-8031. 128 pp. LCCN DS135.D4 D465 2003. URL http://diis.

dk/sw12080.

Bergstrom:2001:PCS

[BJN01] L. Bergström, K. E. Johansson, and Ch. Nilsson. The physics of Copenhagen for students and the general public. Physics Education, 36(5): 388–393, September 2001. CODEN PHEDA7. ISSN 0031-9120 (print),

1361-6552 (electronic). URL http://stacks.iop.org/0031-9120/36/
i=5/a=303.

Bohr:1937:TANa

[BK37] Niels Bohr and F. Kalckar. On the transmutation of atomic nuclei by impact of material particles. I. General theoretical remarks. Mat.-Fys. Medd. Dan. Vidensk. Selsk., 14(10):1–40, ???? 1937.

Bitbol:2009:COT

[BKP09] Michel Bitbol, Pierre Kerszberg, and Jean Petitot, editors. Constituting Objectivity: Transcendental Perspectives on Modern Physics, volume 74 of The Western Ontario series in philosophy of science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-9509-0 (cased). viii + 544 pp. LCCN QC6 CON 2009.

URL http://link.springer.com/10.1007/978-1-4020-9510-8.

Bohr:1985:FQP

[BKR85] Niels Bohr, Jørgen Kalckar, and Erik Rüdinger. Foundations of quantum physics I (1926–1932). North-Holland Publishing Co., Amsterdam, The Netherlands, 1985. ISBN 0-444-86712-0, 0-08-087104-6 (e-book), 0-7204-1800-3 (set). xxvi + 495 pp. LCCN SCI; 98.B04026.

Bohr:1924:QTR

[BKS24a] Niels Bohr, Henrik A. Kramers, and John Clarke Slater. The quantum theory of radiation. Philosophical Magazine, 47(281):785–802, May 1924. CODEN PHMAA4. ISSN 0031-8086. Reprinted in [Boh67c], and in German translation in [BKS24b].

Bohr:1924:QSG

[BKS24b] Niels Bohr, Henrik A. Kramers, and John Clarke Slater. Über die Quantentheorie der Strahlung. (German) [On the quantum theory of radiation]. Zeitschrift für Physik, 24(1):69–87, December 1924. CO-DEN ZEPYAA. ISSN 0044-3328. URL http://link.springer.com/

article/10.1007/BF01327235.

Born:1918:KBA

[BL18a] Max Born and A. Landé. Kristallgitter und Bohrsches Atommodell. (German) [Crystal lattice and Bohr’s atomic model]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 20(21–24):202–209, December 30, 1918. CODEN VDPEAZ. ISSN 0372-5448.

Born:1918:ABK

[BL18b] Max Born and A. Landé. Über die absolute Berechnung der Kristalleigenschaften mit Hilfe Bohrscher Atommodelle. (German) [On the calculation of absolute crystal properties using Bohr atomic models]. S. B. preuss. Akad. Wiss. Berlin, ??(??):1048–1068, ???? 1918. CODEN ???? ISSN ????

Born:1919:ABH

[BL19] Max Born and A. Landé. Antwort auf die Bemerkungen des Herrn L. Vegard zu unseren Arbeiten über Kristallgitter und Bohrsches Atommodell. (German) [Response to the remarks of Mr. L. Vegard to our studies on the crystal lattice and Bohr atomic model]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 21(11–12):385–387, June 30, 1919. CODEN VDPEAZ. ISSN 0372-5448.

Bohr:1954:ECL

[BL54] Niels Bohr and Jens Lindhard. Electron capture and loss by heavy ions penetrating through matter. Matematisk-Fysiske Meddelelser Kongelige Danske Videnskabernes Selskab, 28(7):1–30, ???? 1954.

Born:1963:KBA

[BL63a] Max Born and A. Landé. Kristallgitter und Bohrsches Atommodell. (German) [Crystal lattice and Bohr’s atomic model]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63], chapter 16, pages 348–355. LCCN ???? Reprinted from Verh. Dtsch. Physik. Ges. 20, 202–209 (1918).

Born:1963:ABK

[BL63b] Max Born and A. Landé. Über die absolute Berechnung der Kristalleigenschaften mit Hilfe Bohrscher Atommodelle. (German) [On the absolute calculation of crystal properties using Bohrian atom models]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63], chapter 15, pages 327–347. LCCN ???? Reprinted from S. B. Preuß. Akad. Wiss. Berlin 1918, 1048–1068.

Bernstein:2003:HBE

[BL03] Jeremy Bernstein and Peter D. Lax. Heisenberg & the bomb: An exchange. New York Review of Books, 50(17):??, November 6, 2003. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http:

//www.nybooks.com/articles/archives/2003/nov/06/heisenberg-
the-bomb-an-exchange/. Reply by Thomas Powers.

Boya:2008:ARF

[BL08] Luis J. Boya and Amand A. Lucas. About “Revisiting Farm Hall”. Europhysics News, 39(2):13–15, March/April 2008. CODEN EUP-NAS. ISSN 0531-7479 (print), 1432-1092 (electronic). URL http:

//www.europhysicsnews.org/articles/epn/pdf/2008/02/epn2008-

39-2.pdf. See [Luc07].

Blay:1985:BC

[Bla85] Michel Blay. Bohr et la complémentarité. Revue d’Histoire des Sciences, 38(3–4):193–194, 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0048-7996, 0151-4105, 1969-6582 (electronic).

URL http://www.persee.fr/web/revues/home/prescript/article/
rhs_0151-4105_1985_num_38_3_4004.

Blaedel:1985:HOE

[Blæ85] Niels Blædel. Harmoni og enhed: Niels Bohr, en biografi. (Danish) [Harmony and unity: Niels Bohr, a biography]. Carlsbergfondet, Copenhagen, Danmark, 1985. ISBN 87-7245-006-1. 381 pp. LCCN QC16.B63 B571 1985; QC16.B63 B57 1985.

Blaedel:1988:HUL

[Blæ88] Niels Blædel. Harmony and unity: the life of Niels Bohr. Scientific revolutionaries. Science Tech Publishers, Madison, WI, USA, 1988. ISBN 0-910239-14-2, 3-540-19334-0. xi + 323 pp. LCCN QC16.B63 B5713 1988.

Bloch:1963:RNB

[Blo63] Felix Bloch. Reminiscences of Niels Bohr. Physics Today, 16(10):32–34, October 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bjorkman:2019:ICT

[BLW19] Maria Björkman, Patrik Lundell, and Sven Widmalm, editors. Intellectual collaboration with the Third Reich: treason or reason? Routledge studies in Second World War history. Routledge, Abingdon, Oxon, UK, 2019. ISBN 0-8153-9474-8 (hardcover), 1-351-18508-X (mobi), 1-351-18509-8 (e-pub), 1-351-18510-1 (Adobe), 1-351-18511-X (e-book). ???? pp. LCCN DD256.6 .H58 2019.

Bohr:1955:MIR

[BM55] Aage Bohr and Ben Mottelson. Moments of inertia of rotating nuclei: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(1) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 24 pp. LCCN AS281.

Baym:2009:TSD

[BO09] Gordon Baym and Tomoki Ozawa. Two-slit diffraction with highly charged particles: Niels Bohr’s consistency argument that the electromagnetic field must be quantized. Proceedings of the National Academy of Sciences of the United States of America, 106(9):3035–3040, ???? 2009. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Bodanis:2016:EGM

[Bod16] David Bodanis. Einstein’s greatest mistake: a biography. Houghton Mifflin Harcourt, Boston, MA, USA, 2016. ISBN 0-544-80856-8 (hardcover), 0-544-80858-4 (e-book). xiv + 280 pp. LCCN QC16.E5 B66 2016.

Bohr:1909:DSTa

[Boh09a] Niels Bohr. Determination of the surface-tension of water by the method of jet vibration. Philosophical Transactions of the Royal Society of London. Series A, Containing Papers of a Mathematical or Physical Character, 209(??):281–317, 1909. URL http://www.jstor.org/

stable/91039.

Bohr:1909:DSTb

[Boh09b] Niels Bohr. Determination of the surface-tension of water by the method of jet vibration. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 82(552):146, March 10, 1909. ISSN 0950-1207 (print), 2053-9150

(electronic). URL http://www.jstor.org/stable/93046.

Bohr:1910:DTR

[Boh10] Niels Bohr. On the determination of the tension of a recently formed water-surface. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 84(572):

395–403, December 15, 1910. URL http://www.jstor.org/stable/
93307.

Bohr:1911:SME

[Boh11] Niels Henrik David Bohr. Studier over Metallernes Elektronteori. (Danish) [Studies on the electron theory of metals]. Doktor disputats, Københavns Universitet, København, Danmark, 1911. 120 pp.

URL http://genealogy.math.ndsu.nodak.edu/id.php?id=105783;
 http://www.nba.nbi.dk/bohr1911a.jpg;   http://www.nba.nbi.

dk/bohr1911c.jpg. Afhandling for den filosofiske doktorgrad. [Thesis for the Doctor of Philosophy.].

Bohr:1912:NET

[Boh12] Niels Bohr. Note on the electron theory of thermoelectric phenomena. Philosophical Magazine, 23(6):984–986, ???? 1912. CODEN PHMAA4. ISSN 0031-8086.

Bohr:1913:CAMa

[Boh13a] Niels Bohr. On the constitution of atoms and molecules, Part I. The binding of electrons by positive nuclei. Philosophical Magazine, 26(??):1–25, July 1913. CODEN PHMAA4. ISSN 0031-8086.

URL http://home.tiscali.nl/physis/HistoricPaper/Bohr/Bohr%
20Model%20of%20the%20Atom.htm;   http://web.ihep.su/dbserv/
compas/src/bohr13/eng.pdf;   http://www.chemteam.info/Chem-
History/Bohr/Bohr-1913a.html;   http://www.sciencedirect.com/
science/article/pii/S187605030870046X;   http://www.sciencedirect.
com/science/article/pii/S1876050308700471.

Bohr:1913:CAMb

[Boh13b] Niels Bohr. On the constitution of atoms and molecules. Part II. Systems containing only a single nucleus. Philosophical Magazine, 26(??):476–502, September 1913. CODEN PHMAA4. ISSN 0031-

8086. URL http://www.sciencedirect.com/science/article/pii/
S187605030870046X;   http://www.sciencedirect.com/science/
article/pii/S1876050308700471.

Bohr:1913:CAMc

[Boh13c] Niels Bohr. On the constitution of atoms and molecules. Part III. Systems containing several nuclei. Philosophical Magazine, 26 (??):857–875, October 1913. CODEN PHMAA4. ISSN 0031-

8086. URL http://www.sciencedirect.com/science/article/pii/
S187605030870046X;   http://www.sciencedirect.com/science/
article/pii/S1876050308700471.

Bohr:1913:TDV

[Boh13d] Niels Bohr. On the theory of the decrease of velocity of moving electrified particles on passing through matter. Philosophical Magazine, 25(6):10–31, 1913. CODEN PHMAA4. ISSN 0031-

8086. URL http://www.sciencedirect.com/science/article/pii/
S1876050308700379;   http://www.sciencedirect.com/science/
article/pii/S1876050308700410.

Bohr:1913:SHH

[Boh13e] Niels Bohr. The spectra of helium and hydrogen. Nature, 92(2295): 231–232, October 23, 1913. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v92/n2295/pdf/092231d0.pdf;   http://www.sciencedirect.
com/science/article/pii/S1876050308700537.

Bohr:1914:AMX

[Boh14a] Niels Bohr. Atomic models and X-ray spectra. Nature, 92(2307): 553–554, January 15, 1914. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v92/n2307/pdf/092553d0.pdf;   http://www.sciencedirect.
com/science/article/pii/S1876050308700574.

Bohr:1914:EEM

[Boh14b] Niels Bohr. On the effect of electric and magnetic fields on spectral lines. Philosophical Magazine, 27(??):506–524, ???? 1914. CODEN

PHMAA4. ISSN 0031-8086. URL http://www.sciencedirect.com/
science/article/pii/S1876050308700628.

Bohr:1914:OBD

[Boh14c] Niels Bohr. XII. om brintspektret. (Danish) [XII. On the hydrogen spectrum]. Fysisk Tidsskrift, 12(??):97–114, ???? 1914. URL http:

//dieumsnh.qfb.umich.mx/archivoshistoricosMQ/bohr1914.htm;
 http://home.tiscali.nl/physis/HistoricPaper/Bohr%5CNiels%
20Bohr.htm;   http://www.sciencedirect.com/science/article/

pii/S1876050308700562. English translation in [Boh22g].

Bohr:1915:EEM

[Boh15a] Niels Bohr. On the effect of electric and magnetic fields on spectral lines. Philosophical Magazine, 30(??):394–415, ???? 1915. CODEN

PHMAA4. ISSN 0031-8086. URL http://www.sciencedirect.com/
science/article/pii/S1876050308700689.

Bohr:1915:QTR

[Boh15b] Niels Bohr. On the quantum theory of radiation and the structure of the atom. Philosophical Magazine, 30(??):394–415, ???? 1915. CODEN

PHMAA4. ISSN 0031-8086. URL http://www.sciencedirect.com/
science/article/pii/S1876050308700689.

Bohr:1915:SSH

[Boh15c] Niels Bohr. On the series spectrum of hydrogen and the structure of the atom. Philosophical Magazine, 30(??):332–335, ???? 1915. CODEN

PHMAA4. ISSN 0031-8086. URL http://www.sciencedirect.com/
science/article/pii/S1876050308700653.

Bohr:1915:SHH

[Boh15d] Niels Bohr. The spectra of hydrogen and helium. Nature, 95(2366): 6–7, March 4, 1915. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.nature.com/nature/journal/
v95/n2366/pdf/095006a0.pdf;   http://www.sciencedirect.com/
science/article/pii/S1876050308700677.

Bohr:1916:AQT

[Boh16] Niels Bohr. On the application of the quantum theory to periodic systems. Intended to appear in Philosophical Magazine, April 1916, but not published there. Printed in volume 2 of Bohr’s Collected Works, pp. 431–461. See [Bro15, pages 203–205] for the importance of this work for the concept of zero-point energy, for Planck’s correspondence principle of 1911, versus the better known of Bohr in 1918. See also [Kle66] for more on the role of Planck and Bohr in that import idea.,

1916. URL http://www.sciencedirect.com/science/article/pii/
S1876050308700720.

Bohr:1918:QTLc

[Boh18a] Niels Bohr. On the Quantum Theory of Line Spectra. A. F. Hølst, Copenhagen, Denmark, 1918. 118 pp. URL https://www.gutenberg.

org/ebooks/47167.

Bohr:1918:QTLa

[Boh18b] Niels Bohr. On the quantum theory of line spectra, part I: On the general theory. Det Kongelige Danske Videnskabernes Selskab, Matematiskfysiske Meddelser, 4(1):1–36, ???? 1918. URL http://dieumsnh.

qfb.umich.mx/archivoshistoricosMQ/ModernaHist/Bohr1918.pdf.

Reprinted in [vdW67, pp. 95–137].

Bohr:1918:QTLb

[Boh18c] Niels Bohr. On the quantum theory of line spectra, part II: On the hydrogen spectrum. Det Kongelige Danske Videnskabernes Selskab, Matematisk-fysiske Meddelser, 4(1):36–100, ???? 1918.

Bohr:1920:IBL

[Boh20a] Niels Bohr. On the interaction between light and matter. Translation of a lecture given on 13 Feb. 1920 before the Royal Danish Academy. Printed in volume 3 of Bohr’s Collected Works, pp. 227–240., February 13, 1920.

Bohr:1920:SCA

[Boh20b] Niels Bohr. Some considerations of atomic structure. Translation of a lecture given before the Physical Society of Copenhagen, 15 December 1920. Printed in volume 4 of Bohr’s Collected Works, pp. 43–69., December 15, 1920.

Bohr:1920:SEG

[Boh20c] Niels Bohr. Über die Serienspektra der Elemente. (German) [On the series spectra of the elements]. Zeitschrift für Physik, 2(5):423–469, October 1920. CODEN ZEPYAA. ISSN 0044-3328. URL http://

link.springer.com/article/10.1007/BF01329978. Translated by A.

D. Udden. Reprinted in [Boh22g].

Bohr:1921:AAJ

[Boh21a] Niels Bohr. Abhandlungen über Atombau aus den Jahren 1913–1918. (German). [Discussions on Atomic Structure from the Years 1913– 1918]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1921. xix + 155 pp. LCCN QC173 B676g. Authorized translation to German by Hugo Stintzing from the English of various articles published between 1913–1916 in Philosophical Magazine and in Nature, London, England. Contains a new foreword by Niels Bohr.

Bohr:1921:ASa

[Boh21b] Niels Bohr. Atomic structure. Nature, 107(2682):104–107, March 24, 1921. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://dieumsnh.qfb.umich.mx/archivoshistoricosMQ/
bohr1921.htm;   http://www.nature.com/nature/journal/v107/
n2682/pdf/107104a0.pdf.

Bohr:1921:ASb

[Boh21c] Niels Bohr. Atomic structure. Nature, 108(2711):208–209, October 13, 1921. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v108/n2711/
pdf/108208b0.pdf.

Bohr:1921:CA

[Boh21d] Niels Bohr. Constitution of atoms. Unpublished manuscript for the Solvay congress (1921). Printed in volume 4 of Bohr’s Collected Works, pp. 99–174., 1921.

Bohr:1921:FPS

[Boh21e] Niels Bohr. Zur Frage der Polarisation der Strahlung in der Quantentheorie. (German). [On the question of the polarisation of rays in the quantum theory]. Zeitschrift für Physik, 6(1):1–9, December 1921. CODEN ZEPYAA. ISSN 0044-3328. URL http://link.springer.

com/article/10.1007/BF01327957.

Bohr:1921:ABO

[Boh21f] Niels Henrik David Bohr. Atomernes Bygning og Stoffernes fysiske og kemiske Egenskaber. (Danish) [Atomic structure and physical and chemical properties of matter]. Gjellerup, København, Danmark, 1921. 70 pp. Foredrag holdt i Fysisk Forening den 18. Oktober 1921. [Lecture before the Physical Society, Copenhagen, 18 October 1921.] German translation in [Boh22b], and English translation, with modifications, in [Boh22g].

Bohr:1922:ANB

[Boh22a] Niels Bohr. Ausarbeitung of Niels Bohr’s lectures on atomic physics, 1922. ????, ????, 1922. 107 pp. LCCN ????

Bohr:1922:BAP

[Boh22b] Niels Bohr. Der Bau der Atome und die physikalischen und chemischen Eigenschaften der Elemente. (German) [The constitution of atoms and the physical and chemical properties of elements]. Zeitschrift für Physik, 9(1):1–67, December 1922. CODEN ZEPYAA. ISSN 0044-3328.

URL http://link.springer.com/article/10.1007/BF01326955.

Bohr:1922:DBS

[Boh22c] Niels Bohr. The difference between series spectra of isotopes. Nature, 109(2745):746, June 19, 1922. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v109/n2745/pdf/109746a0.pdf.

Bohr:1922:DAS

[Boh22d] Niels Bohr. Drei Aufsätze über Spektren und Atombau. (German) [Three lectures on spectra and atomic structure], volume 56 of Sammlung Vieweg. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1922. vi + 148 pp. LCCN ????

Bohr:1922:QTL

[Boh22e] Niels Bohr. On the quantum theory of line spectra, part III: On the spectra of elements of higher atomic number. Det Kongelige Danske Videnskabernes Selskab, Matematisk-fysiske Meddelser, 4(1):101–118, ???? 1922.

Bohr:1922:SLT

[Boh22f] Niels Bohr. Seven lectures on the theory of atomic structure. Given in Göttingen, Germany. Published in volume 4 of Bohr’s Collected Works, pp. 341–419., 1922.

Bohr:1922:TSA

[Boh22g] Niels Bohr. The theory of spectra and atomic constitution: three essays. Cambridge University Press, Cambridge, UK, 1922. x + 126 pp.

LCCN QC451 .B6. URL https://www.gutenberg.org/ebooks/47464.

Bohr:1923:LTQ

[Boh23a] Niels Bohr. L’application de la théorie des quanta aux problèmes atomiques. (French) [Application of quantum theory to atomic problems]. In Lorentz et al. [LRdB^{+}23], pages 364–380. LCCN QC1 .I6 1921. Proceedings of the Solvay III international conference.

Bohr:1923:SSA

[Boh23b] Niels Bohr. Les spectres et la structure de l’atome: Trois conférences. (French) [Spectra and Atomic Structure: Three lectures]. J. Hermann et cie, Paris, France, 1923. 152 pp. LCCN ???? Translation to French by A. Corvisy.

Bohr:1923:LAG

[Boh23c] Niels Bohr. Linienspektren und Atombau. (German) [Line spectra and atomic structure]. Annalen der Physik (1900), 376(9–12):228–288,

???? 1923. ISSN 1521-3889. URL http://onlinelibrary.wiley.com/
doi/10.1002/andp.19233760918/abstract.

Bohr:1923:OAB

[Boh23d] Niels Bohr. Om Atomernes Bygning. (Danish) [On the structure of atoms]. Fysisk Tidsskrift, 21(??):6–44, ???? 1923. English translation in [Boh23f], and German translation in [Boh22b].

Bohr:1923:EEM

[Boh23e] Niels Bohr. The seventh Guthrie Lecture, entitled “The effect of electric and magnetic fields on spectral lines”. Proceedings of the Physical Society, London, 35(1):275–302, ???? 1923. CODEN PPSOAU.

ISSN 0370-1328. URL http://iopscience.iop.org/1478-7814/35/1/
342;   http://stacks.iop.org/1478-7814/35/i=1/a=342. The Sev-

enth Guthrie Lecture.

Bohr:1923:SA

[Boh23f] Niels Bohr. The structure of the atom. Nature, 112(2801):29–44, July 7, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v112/n2801/
pdf/112029a0.pdf.

Bohr:1923:BAG

[Boh23g] Niels Bohr. Über den Bau der Atome. (German) [On the structure of atoms]. Die Naturwissenschaften, 11(27):606–624, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). Lecture held in Stockholm, 11 December 1922, on the occasion of the receipt of the Nobel Prize for physics for the year 1922. Translated to German by W. Pauli, Jr.

Bohr:1923:AQAa

[Boh23h] Niels Bohr. Über die Anwendung der Quantentheorie auf dem Atombau, I. Die Grundpostulate der Quantentheorie. (German) [On the application of quantum theory to atomic structure. I. The basic postulates of quantum theory]. Zeitschrift für Physik, 13(1):117–165, December 1923. CODEN ZEPYAA. ISSN 0044-3328. URL http://link.springer.

com/article/10.1007/BF01328209.

Bohr:1923:AQAb

[Boh23i] Niels Bohr. Über die Anwendung der Quantentheorie auf dem Atombau. II. Theorie der Serienspektren. (German) [On the application of quantum theory to atomic structure. II. Theory of line spectra]. Printed in volume 3 of Bohr’s Collected Works, pp. 502–531, 1923.

Bohr:1924:APD

[Boh24a] Niels Bohr. Atomteoretiske Problemer. (Danish) [Atomic theory problems]. Printed in volume 3 of Bohr’s Collected Works, pp. 569–574., 1924.

Bohr:1924:DAS

[Boh24b] Niels Bohr. Drei Aufsätze über Spektren und Atombau. (German) [Three lectures on spectra and atomic structure], volume 56 of Sammlung Vieweg. Friedrich Vieweg und Sohn, Braunschweig, Germany, second edition, 1924. 150 pp. LCCN ????

Bohr:1924:GMA

[Boh24c] Niels Bohr. Grundlaget for den moderne Atomforskning. (Danish) [The foundation of modern atomic research]. Lecture on the award of the Örsted medal. Printed in volume 5 of Bohr’s Collected Works, pp. ??–??., October 22, 1924.

Bohr:1924:AQT

[Boh24d] Niels Bohr. On the application of quantum theory to atomic structure. I. The fundamental postulates. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 22 (supplement)(??): 1–42, ???? 1924. CODEN PCPSA4. ISSN 0008-1981.

Bohr:1924:TSA

[Boh24e] Niels Bohr. The theory of spectra and atomic constitution: three essays. Cambridge University Press, Cambridge, UK, second edition, 1924. x + 138 pp. LCCN ????

Bohr:1924:BAG

[Boh24f] Niels Bohr. Über den Bau der Atome. (German) [On the structure of atoms]. Angewandte Chemie, 37(27):466, June 1924. CODEN AN-CEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Bohr:1924:PFG

[Boh24g] Niels Bohr. Zur Polarisation des Fluorescenzlichtes. (German) [The polarisation of fluorescent light]. Die Naturwissenschaften, 12(49):1115– 1117, December 1924. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bohr:1925:ATM

[Boh25a] Niels Bohr. Atomic theory and mechanics. Nature, 116(2927):845–852, December 5, 1925. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.nature.com/nature/journal/
v116/n2927/pdf/116845a0.pdf.

Bohr:1925:AOM

[Boh25b] Niels Bohr. Atomteori og Mekanik. (Danish) [Atomic theory and mechanics]. Mat. Tidsskrift B, 1925(??):103–107, ???? 1925.

Bohr:1925:BAG

[Boh25c] Niels Bohr. Über den Bau der Atome. (German) [On the structure of atoms]. J. Springer, Berlin, Germany, third edition, 1925. 60 pp. LCCN ????

Bohr:1925:WAS

[Boh25d] Niels Bohr. Über die Wirkung von Atomen bei Stössen. (German) [The effect of atoms on impacts]. Zeitschrift für Physik, 34(1):142–157, December 1925. CODEN ZEPYAA. ISSN 0044-3328. URL http://

link.springer.com/article/10.1007/BF01328464.

Bohr:1926:A

[Boh26a] Niels Bohr. Atom. In Encyclopedia Britannica, pages 262–?? Encyclopaedia Britannica, Chicago, IL, USA, 1926.

Bohr:1926:ATM

[Boh26b] Niels Bohr. Atomic theory and mechanics. In Matematiker kongressen i København: 31. August–4. September 1925: den sjette skandinaviske Matematikerkongres: beretning. (Danish) [Mathematicians Conference in Copenhagen: 31 August–4 September 1925: the sixth Scandinavian Mathematician Conference: report], pages 51–70. Det Hoffensbergske Etabl., København, Danmark, ???? 1926.

Bohr:1926:AMG

[Boh26c] Niels Bohr. Atomtheorie und Mechanik. (German) [Atomic theory and mechanics]. Die Naturwissenschaften, 14(1):1–10, January 1, 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bohr:1926:SER

[Boh26d] Niels Bohr. Sir Ernest Rutherford, O.M., P.R.S. Nature, 118 (2981S):51–52, December 18, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v118/n2981supp/pdf/118051a0.pdf.

Bohr:1926:SES

[Boh26e] Niels Bohr. Spinning electrons and the structure of spectra. Nature, 117(2938):265, February 20, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v117/n2938/pdf/117265a0.pdf.

Bohr:1928:QNE

[Boh28a] Niels Bohr. Das Quantenpostulat und die neuere Entwicklung der Atomistik. (German). [Quantum postulate and recent developments in atomism]. Die Naturwissenschaften, 16(15):245–257, April 13, 1928. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bohr:1928:PQNb

[Boh28b] Niels Bohr. Le postulat des quanta et le nouveau développement de l’atomistique. (French) [The postulate of quanta and new developments in atomic theory]. In Anonymous [Ano28], pages 215–247. LCCN ????

Bohr:1928:QPRa

[Boh28c] Niels Bohr. The quantum postulate and the recent development of atomic theory. In ????, editor, Atti del Congresso internazionale dei fisici, volume 2, pages 565–588. Zanichelli, Bologna, Italy, 1928. LCCN ???? Two volumes.

Bohr:1928:QPRb

[Boh28d] Niels Bohr. The quantum postulate and the recent development of atomic theory. Nature, 121(3050):580–590, April 14, 1928. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http:/

/www.nature.com/nature/journal/v121/n3050/pdf/121580a0.pdf.

Bohr:1928:SAG

[Boh28e] Niels Bohr. Sommerfeld und die Atomtheorie. (German) [Sommerfeld and atomic theory]. Die Naturwissenschaften, 16(49):1036, ???? 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bohr:1929:WNG

[Boh29a] Niels Bohr. Wirkungsquantum und Naturbeschreibung. (German) [The quantum of action and the description of nature]. Die Naturwissenschaften, 17(26):483–486, June 1929. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bohr:1929:AON

[Boh29b] Niels Henrik David Bohr. Atomteori og Naturbeskrivelse. (Danish) [Atomic theory and the description of nature]. In Niels Henrik David Bohr, editor, Festskrift udgivet af Københavns Universitet i Anledning af Universitetets Aarsfest November 1929. (Danish) [Festival volume issued by Copenhagen University in Connection with the University’s Yearly Festival, November 1929], pages 3–82. Bianco Lunos Bogtrykkeri, København, Danmark, 1929. ISSN 0109-2758. LCCN QC173 B54 1929.

Bohr:1929:KOA

[Boh29c] Niels Henrik David Bohr. Kvantenpostulatet og Atomteoriens seneste Udvikling. (Danish) [The quantum postulate and atomic theory’s latest developments]. In Niels Henrik David Bohr, editor, Atomteori og Naturbeskrivelse: 3 artiklar med en indledende oversigt. (Danish) [Atomic theory and description of nature: 3 articles with an introductory overview], Festskrift udgivet af Københavns universitet i anledning af universitetets årsfest, pages 40–68. Bianco Lunos Bogtrykkeri, København, Danmark, 1929. ISSN 0109-2758. LCCN QC173 B54 1929.

Bohr:1930:AOG

[Boh30a] Niels Bohr. Atomteorien og Grundprincipperne for Naturbeskrivelsen. (Danish) [Atomic Theory and the Description of Nature]. Særtryk af Fysisk Tidsskrift, København, Danmark, 1930. 12 pp. LCCN ???? Extended English translation in [Boh34].

Bohr:1930:APN

[Boh30b] Niels Bohr. Die Atomtheorie und die Prinzipien der Naturbeschreibung. (German) [Atomic theory and the principles of the description of nature]. Die Naturwissenschaften, 18(4):73–78, January 24, 1930. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bohr:1931:ANV

[Boh31a] Niels Bohr. Atomtheorie und Naturbeschreibung: Vier Aufsätze mit einer einleitenden Übersicht. (German) [Atomic Theory and the Description of Nature: Four lectures with an introductory overview]. J. Springer, Berlin, Germany, 1931. 77 pp. LCCN ????

Bohr:1931:AQP

[Boh31b] Niels Bohr. Die Andwendung der Quantentheorie auf periodische Systeme. (German) [The use of the quantum theory of periodic systems]. In Abhandlungen über Atombau aus den Jahren 1913–1916. (German) [Papers on atomic structure from the years 1913–1916], pages 123–151. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1931. LCCN ???? Translated by Hugo Stintzing.

Bohr:1931:MMT

[Boh31c] Niels Bohr. Maxwell and modern theoretical physics. Nature, 128 (3234):691–692, October 24, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v128/n3234/pdf/128691a0.pdf.

Bohr:1932:R

[Boh31d] Niels Bohr. On atomic stability. British Association for the Advancement of Science, page 333, 1931. URL http://biodiversitylibrary. org/page/30524583. Report of the Centenary Meeting, London, 23–30 September 1931.

Bohr:1931:NED

[Boh31e] Niels Bohr. On Neutronernes Egenskaber. (Danish) [On the neutron properties]. Oversigt over det Kongelige Danske Videnskabernes Selskabs Virksomhed, ??(??):52–??, ???? 1931. CODEN KVSVA4. ISSN 0368-7201.

Bohr:1932:ASC

[Boh32a] Niels Bohr. Atomic stability and conservation laws. In ????, editor, Atti del Convegno di Fisica Nucleare dell “Fondazion Alessandro Volta”, Ottobre 1931. (Italian) [Actes of the Congress on Nuclear physics of the “Fondazion Alexander Volta”, October 1931], pages 119–130. Reale Accademica d’Italia, Rome, Italy, 1932. LCCN ????

Bohr:1932:FLC

[Boh32b] Niels Bohr. Faraday Lecture: Chemistry and the quantum theory of atomic constitution. Journal of the Chemical Society, 131(??):349– 384, ???? 1932. CODEN JCSOA9. ISSN 0368-1769 (print), 2050-5574 (electronic). Faraday Lecture 1930, delivered before the Fellows of the Chemical Society at the Salters’ Hall on May 8th, 1930. Reprinted in [Boh65b].

Bohr:1932:TAD

[Boh32c] Niels Bohr. La théorie atomique et la description des phénomènes. Quatre articles précédés d’une introduction. (French) [Atomic Theory and the Description of Nature: Four lectures with an introductory overview]. Gauthier-Villars et cie, Paris, France, 1932. v + 111 pp. LCCN ???? Translation of [Boh30a] to French by A. Legros and L. Rosenfeld.

Bohr:1932:LL

[Boh32d] Niels Bohr. Light and life. In ????, editor, International Congress on Light Therapy, Copenhagen, Denmark, August 1932, page ?? ????, ????, 1932. LCCN ???? Reprinted in [Boh58b, pages 3–12] and [Boh85i].

Bohr:1933:LLa

[Boh33a] Niels Bohr. Light and life. Nature, 131(3308):421–423, March 25, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v131/n3308/
pdf/131421a0.pdf.

Bohr:1933:LLb

[Boh33b] Niels Bohr. Light and life. Nature, 131(3309):457–459, April 1, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v131/n3309/
pdf/131457a0.pdf.

Bohr:1933:CME

[Boh33c] Niels Bohr. On the correspondence method in electron theory. Address to seventh Solvay conference. Reprinted in [Pei08, pages 183–191]., October 1933.

Bohr:1934:ATD

[Boh34] Niels Bohr. Atomic Theory and the Description of Nature: Four essays, with an introductory survey. Cambridge University Press, Cambridge, UK, 1934. vi + 119 pp. LCCN QC173.18 .B64 1934.

Bohr:1935:CQM

[Boh35a] N. Bohr. Can quantum-mechanical description of physical reality be considered complete? Physical Review, 48(8):696–702, October 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://dieumsnh.qfb.umich.mx/archivoshistoricosMQ/ModernaHist/
Bohr1935.pdf;   http://link.aps.org/doi/10.1103/PhysRev.48.
696;   http://prola.aps.org/abstract/PR/v48/i8/p696_1.

Bohr:1935:FPS

[Boh35b] Niels Bohr. Friedrich Paschen zum siebzigsten Geburtstag. (German) [Friedrich Paschen’s sixtieth birthday]. Die Naturwissenschaften, 23(5): 73, February 1, 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bohr:1935:OAE

[Boh35c] Niels Bohr. Om Atomkernernes Egenskaber og Opbygning. (Danish) [On the properties and structure of atomic nuclei]. Oversigt over det Kongelige Danske Videnskabernes Selskabs Virksomhed, ??(??):39, ???? 1935. CODEN KVSVA4. ISSN 0368-7201.

Bohr:1935:QMP

[Boh35d] Niels Bohr. Quantum mechanics and physical reality. Nature, 136(3428):65, July 13, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v136/n3428/pdf/136065a0.pdf.

Bohr:1936:CLQ

[Boh36a] Niels Bohr. Conservation laws in quantum theory. Nature, 138(3479): 25–26, July 4, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.nature.com/nature/journal/
v138/n3479/pdf/138025b0.pdf.

Bohr:1936:KKG

[Boh36b] Niels Bohr. Kausalität und Komplementarität. (German) [Causality and complementarity]. Erkenntnis, 6(1):293–303, December 1936.

CODEN ???? ISSN 1876-2514. URL http://www.jstor.org/stable/
20011824.

Bohr:1936:NCN

[Boh36c] Niels Bohr. Neutron capture and nuclear constitution. Nature, 137(3461):344–348, February 29, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

www.nature.com/nature/journal/v137/n3461/pdf/137344a0.pdf;
 http://www.nature.com/nature/journal/v137/n3461/pdf/137351a0.
pdf;   http://www.nature.com/nature/journal/v137/n3461/pdf/

137351b0.pdf. See comments [Ano36a, Ano36b].

Bohr:1936:NBA

[Boh36d] Niels Bohr. Neutroneneinfang und Bau der Atomkerne. (German) [Neutron capture and structure of atomic nuclei]. Die Naturwissenschaften, 24(16):241–245, April 17, 1936. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bohr:1936:PCA

[Boh36e] Niels Bohr. Properties and constitution of atomic nuclei. Nature, 138(3494):695, October 17, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Bohr:1937:CC

[Boh37a] Niels Bohr. Causality and complementarity. Philosophy of Science, 4(3):289–298, July 1937. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184445.

Bohr:1937:MNF

[Boh37b] Niels Bohr. Mécanique nucléaire. (French). [Nuclear mechanics]. In ????, editor, Réunion internationale de physique–chemie–biologie, Congrès du Palais de la découverte, Paris, octobre 1937. (French) [International physics–chemistry–biology meeting, Congress of the Palace of Discovery, Paris, October 1937], pages 81–82. Hermann et cie, Paris, France, 1937. LCCN ????

Bohr:1937:ORH

[Boh37c] Niels Bohr. Obituary: The Right Hon. Lord Rutherford of Nelson, O.M., F.R.S. Nature, 140(3548):752–753, October 30, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http:/

/www.nature.com/nature/journal/v140/n3548/pdf/140752b0.pdf.

Bohr:1937:OAD

[Boh37d] Niels Bohr. Om Atomkernernereaktioner. (Danish) [On atomic nuclear reactions]. Oversigt over det Kongelige Danske Videnskabernes Selskabs Virksomhed, ??(??):32–??, ???? 1937. CODEN KVSVA4. ISSN 0368-7201.

Bohr:1937:OSA

[Boh37e] Niels Bohr. Om Spaltning af Atomkerner. (Danish) [On fission of atomic nuclei]. In ????, editor, 5. nordiske Elektroteknikermøde i Købehavn den 25.–27. August 1937. (Danish) [Fifth Electrotechnical Meeting in Copenhagen, 25–27 August 1937], pages 21–23. J. H. Schultz, København, Danmark, 1937. LCCN ????

Bohr:1937:TANb

[Boh37f] Niels Bohr. Transmutations of atomic nuclei. Science (New Series), 86 (2225):161–165, August 20, 1937. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/stable/

1664129.

Bohr:1938:BAP

[Boh38a] Niels Bohr. Biology and atomic physics. Il Nuovo Cimento (8), 15(7): 429–438, July 1938. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). Celebrazione del secondo centenario della nascita di Luigi Galvani Bologna 18–21 Ottobre 1937-XV.

Bohr:1938:BAD

[Boh38b] Niels Bohr. British Association discussions: Nuclear physics. Nature, 142(3594):520–521, September 17, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v142/n3594/pdf/142520a0.pdf.

Bohr:1938:NPE

[Boh38c] Niels Bohr. Nuclear photo-effects. Nature, 141(3564):326–327, February 19, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/v141/
n3564/pdf/141326a0.pdf.

Bohr:1938:RNP

[Boh38d] Niels Bohr. Resonance in nuclear photo-effects. Nature, 141(3581): 1096–1097, June 18, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v141/n3581/pdf/1411096b0.pdf.

Bohr:1938:WAG

[Boh38e] Niels Bohr. Wirkungsquantum und Atomkern. (German) [The quantum of action and the atomic nucleus]. Annalen der Physik (1900), 424 (1–2):5–19, ???? 1938. ISSN 1521-3889. URL http://onlinelibrary.

wiley.com/doi/10.1002/andp.19384240104/abstract.

Bohr:1939:RUT

[Boh39a] N. Bohr. Resonance in uranium and thorium disintegrations and the phenomenon of nuclear fission. Physical Review, 55(4):418–419, February 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v55/i4/p418_2.

Bohr:1939:CPA

[Boh39b] Niels Bohr. The causality problem in atomic physics. In ????, editor, The New Theories of Physics: Conference Organized in collaboration with The International Union of Physics and The Polish Intellectual Cooperation Committee, Warsaw, May 30th–June 3rd 1938, pages 11–45. International Institute of Intellectual Cooperation, Paris, France, 1939.

Bohr:1939:TFA

[Boh39c] Niels Bohr. Den teoretiske Forklaring af Atomkernernes Fission. (Danish) [The theoretical explanation of the fission of atomic nuclei]. Oversigt over det Kongelige Danske Videnskabernes Selskabs Virksomhed, ??(??): 28, ???? 1939. CODEN KVSVA4. ISSN 0368-7201.

Bohr:1939:DHN

[Boh39d] Niels Bohr. Disintegration of heavy nuclei. Nature, 143(3617):330, February 25, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.chemteam.info/Chem-History/
Bohr-Fission-1939.html;   http://www.nature.com/nature/journal/
v143/n3617/pdf/143330a0.pdf.

Bohr:1939:NPH

[Boh39e] Niels Bohr. Natural philosophy and human cultures. Nature, 143 (3616):268–272, February 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v143/n3616/pdf/143268a0.pdf.

Bohr:1939:QAN

[Boh39f] Niels Bohr. Quantum d’action et noyaux atomiques. (French) [Quantum of action and atomic nuclei]. Hermann et Cie., Paris, France, 1939. 28 pp. LCCN ????

Bohr:1939:RHN

[Boh39g] Niels Bohr. Reactions of heavy nuclei. Nature, 143(3614):215, February 4, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Bohr:1940:SSF

[Boh40a] N. Bohr. Scattering and stopping of fission fragments. Physical Review, 58(7):654–655, October 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/

10.1103/PhysRev.58.654;   http://prola.aps.org/abstract/PR/
v58/i7/p654_1.

Bohr:1940:STN

[Boh40b] N. Bohr. Successive transformations in nuclear fission. Physical Review, 58(10):864–866, November 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v58/i10/p864_1.

Bohr:1940:TAS

[Boh40c] Niels Bohr. Tunge Atomkerners Sønderdeling. (Danish) [The splitting of heavy atomic nuclei]. Oversigt over det Kongelige Danske Videnskabernes Selskabs Virksomhed, ??(??):38–??, 1940. CODEN KVSVA4. ISSN 0368-7201.

Bohr:1941:MDIa

[Boh41a] N. Bohr. Mechanism of deuteron-induced fission. Physical Review, 50 (12):1042, June 1941. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/PR/
v59/i12/p1042_1.

Bohr:1941:VRR

[Boh41b] N. Bohr. Velocity-range relation for fission fragments. Physical Review, 59(3):270–275, February 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v59/i3/p270_1.

Bohr:1941:DKN

[Boh41c] Niels Bohr. Dansk Kultur, Nogle indledende Betragtninger. (Danish) [Danish culture, some introductory remarks]. In Dahl [Dah43], page ?? LCCN DL131 .D2. Eight volumes.

Bohr:1941:MDIb

[Boh41d] Niels Bohr. Mechanism of deuteron-induced fission. Nature, 148 (3747):229, August 23, 1941. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v148/n3747/pdf/148229a0.pdf.

Bohr:1941:NUAa

[Boh41e] Niels Bohr. Nyere Underøgelser over Atomkernernes Omdannelser. (Danish) [Newer investigations of the transmutation of atomic nuclei]. Fysisk Tidsskrift, 39(??):3–32, ???? 1941.

Bohr:1941:NUAb

[Boh42a] Niels Bohr. Nyere Underøgelser over Atomkernernes Omdannelser. (Danish) [Newer investigations of the transmutation of atomic nuclei]. Fra Fysikkens Verden, 3(??):1–22, 81–96, ???? 1941–1942.

Bohr:1942:FDP

[Boh42b] Niels Bohr. Forord. (Danish) [preface]. In Mr. Tompkins i Drømmeland. (Danish) [Mr. Tompkins in Wonderland] [Gam42], pages 7–8. Forord af Niels Bohr.

Bohr:1944:MPR

[Boh44] Niels Bohr. Memorandum to President Roosevelt. Unpublished, July 3, 1944. Reprinted in [BHK^{+}08, volume 11, pages 101–108].

Bohr:1945:CC

[Boh45a] Niels Bohr. A challenge to civilization. Science (New Series), 102 (2650):363–364, October 12, 1945. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/stable/

1671741.

Bohr:1945:EAO

[Boh45b] Niels Bohr. Energy from the atom: an opportunity and a challenge. The Times [London, UK], ??(??):??, August 11, 1945. ISSN 0140-0460, 0956-1382. Reprinted in [Boh85h].

Bohr:1945:AOD

[Boh45c] Niels Bohr. On Atomkernernes Omdanndelser. (Danish) [On the transmutation of atomic nuclei]. Oversigt over det Kongelige Danske Videnskabernes Selskabs Virksomhed, ??(??):31, 1945. CODEN KVSVA4. ISSN 0368-7201.

Bohr:1945:SC

[Boh45d] Niels Bohr. Science and civilization. The Times [London, UK], ?? (??):??, August 11, 1945. ISSN 0140-0460, 0956-1382. Republished as [Boh45a].

Bohr:1946:FSC

[Boh46a] Niels Bohr. Foreword: Science and civilization. In Masters and Way [MW46a], page ?? LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Bohr:1946:F

[Boh46b] Niels Bohr. Forord. In Masters and Way [MW46b], page ?? LCCN ???? Danish translation of [MW46a] by Gudrun Frederiksen og Ebbe Rasmussen. Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Bohr:1946:NUA

[Boh46c] Niels Bohr. Nyere Underøgelser over Atomkernernes Omdannelser. (Danish) [Newer investigations of the transmutation of atomic nuclei]. Kosmos, 24(??):24–57, ???? 1946.

Bohr:1947:API

[Boh47] Niels H. D. Bohr. Atomic physics and international cooperation. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 91(2):137–138, April 1947. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://

www.jstor.org/stable/3143598.

Bohr:1948:NCC

[Boh48a] Niels Bohr. On the notions of causality and complementarity. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):312–319, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bohr:1948:PAP

[Boh48b] Niels Bohr. The penetration of atomic particles through matter, volume XVIII(8) of Kongelige Danske Videnskabernes Selskab. Mathematisk-fysiske meddelelser. E. Munksgaard, København, Danmark, 1948. 144 pp. LCCN ????

Bohr:1949:DEE

[Boh49] Niels Bohr. Discussion with Einstein on epistemological problems in atomic physics. In Schilpp [Sch49a], chapter 7, pages 199–241. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. URL

http://www.marxists.org/reference/subject/philosophy/works/

dk/bohr.htm. Reprinted 1951, 1969, and 1982.

Bohr:1950:ABT

[Boh50a] Niels Bohr. Åbent brev til de forenede nationer: 9. juni 1950. (Danish) [Open letter to the United Nations, 9 June 1950]. J. H. Schultz, København, Danmark, 1950. 14 pp. Danish translation of the English original [Boh50d].

Bohr:1950:OW

[Boh50b] Niels Bohr. For an open world. Bulletin of the Atomic Scientists, 6 (7):213–217, 219, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bohr:1950:NCC

[Boh50c] Niels Bohr. On the notions of causality and complementarity. Science (New Series), 111(2873):51–54, January 20, 1950. CODEN SCN-CAD. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/1677100.

Bohr:1950:OLU

[Boh50d] Niels Bohr. Open letter to the United Nations, 9 June 1950. Science (New Series), 112(2897):1–6, July 7, 1950. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/1678434.

Bohr:1950:UN

[Boh50e] Niels Bohr. To the United Nations. Impact of Science on Society, 1(2):68–??, ???? 1950. ISSN 0019-2872. URL http://neutrino.

aquaphoenix.com/un-esa/ws1999-letter-bohr.html.

Bohr:1951:CC

[Boh51] Niels Bohr. Causality and complementarity. Current Science, 20(5): 115–116, May 1951. CODEN CUSCAM. ISSN 0011-3891. URL https:

//www.jstor.org/stable/24212817.

Bohr:1952:MRN

[Boh52] Niels Bohr. Medical research and natural philosophy. Acta medica Scandinavica, 142(S266):967–972, January/December 1952. CODEN

AMSVAZ. ISSN 0001-6101. URL http://onlinelibrary.wiley.com/
doi/10.1111/j.0954-6820.1952.tb13446.x/abstract.

Bohr:1955:FGI

[Boh55a] Niels Bohr. Det fysiske grundlag for industriel udnyttelse af atomkerne-energien: foredrag på landsindustrimödet den 17. marts 1955. (Danish) [The physical basis for the industrial use of atomic nuclear energy: lecture at the industrial meeting, 17 March 1955]. Nielsen and Lydich, København, Danmark, 1955. 12 pp. LCCN ????

Bohr:1955:G

[Boh55b] Niels Bohr. On Geneva. Bulletin of the Atomic Scientists, 11(8):284, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bohr:1955:PSMa

[Boh55c] Niels Bohr. Physical science and man’s position. In ????, editor, Proceedings of the International Conference on the Peaceful Uses of Atomic Energy: Held in Geneva 8 August–20 August, 1955, vol. 16, pages 57–61. United Nations, New York, NY, USA, 1955.

Bohr:1955:PSMb

[Boh55d] Niels Bohr. Physical science and man’s position. Ingeniøren, 64(??): 810–814, ???? 1955. CODEN MNGEBU.

Bohr:1955:WIG

[Boh55e] Niels Bohr. Wer ist’s?. (German) [Who is it?]. Nachrichten aus Chemie und Technik, 3(23):236–237, December 7, 1955. CODEN NCH-

TAD. ISSN 0027-738X. URL http://onlinelibrary.wiley.com/doi/
10.1002/nadc.19550032303/abstract.

Bohr:1956:MNP

[Boh56] Niels Bohr. Mathematics and natural philosophy. The Scientific Monthly, 82(2):85–88, February 1956. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.jstor.org/

stable/22079.

Bohr:1958:APHa

[Boh58a] Niels Bohr. Atomic physics and human knowledge. Chapman and Hall, London, UK, 1958. 101 pp. LCCN ????

Bohr:1958:APHb

[Boh58b] Niels Bohr. Atomic Physics and Human Knowledge. Wiley, New York, NY, USA, 1958. viii + 101 pp. LCCN QC6 .B598. URL http://www.

questia.com/PM.qst?a=o&d=21083222.

Bohr:1958:AME

[Boh58c] Niels Bohr. Atomphysik und menschliche Erkenntnis. (German) [Atomic physics and human knowledge], volume 112 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1958. vii + 104 pp. LCCN ????

Bohr:1958:AON

[Boh58d] Niels Bohr. Atomteori og naturbeskrivelse. (Danish) [Atomic theory and the description of nature]. J. H. Schultz, København, Danmark, 1958. 96 pp. LCCN ????

Bohr:1958:AHK

[Boh58e] Niels Bohr. On atoms and human knowledge. Dædalus, 87(2):164– 175, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.org/stable/20026444.

Bohr:1958:AOM

[Boh58f] Niels Henrik David Bohr. Atomfysik og menneskelig erkendelse. (Danish) [Atomic physics and human knowledge]. J. H. Schultz, København, Danmark, fourth edition, 1958. 124 pp. LCCN ????

Bohr:1959:AOM

[Boh59a] Niels Bohr. Atomfysik och mänskligt vetande. (Swedish) [Atomic physics and human knowledge]. Aldus/Bonniers, Stockholm, Sweden, 1959. 119 pp. LCCN ????

Bohr:1959:EQGa

[Boh59b] Niels Bohr. Über Erkenntnisfragen der Quantenphysik. (German) [On Realization Questions of Quantum Physics]. VEB Deutscher Verlag der Wissenschaften, Berlin, West Germany, 1959. 7 pp.

Bohr:1959:EQGb

[Boh59c] Niels Bohr. Über Erkenntnisfragen der Quantenphysik. (German) [On realization questions of quantum physics]. In Max-Planck-Festschrift 1958 [KMP59], page ?? LCCN QC3 .M45.

Bohr:1960:API

[Boh60a] Niels Bohr. Atomic physics and international cooperation. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 91(2):137–138, April 5, 1960. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.

org/stable/3143598.

Bohr:1960:F

[Boh60b] Niels Bohr. Foreword. In Markus Fierz and Victor F. (Victor Frederick) Weisskopf, editors, Theoretical physics in the twentieth century: a memorial volume to Wolfgang Pauli, pages 1–4. Interscience Publishers, New York, NY, USA, 1960. LCCN QC3F54T 1960.

Bohr:1960:PAP

[Boh60c] Niels Bohr. The penetration of atomic particles through matter, volume 18(8). E. Munksgaard, København, Danmark, third edition, 1960. 144 pp. LCCN ????

Bohr:1961:ATD

[Boh61a] Niels Bohr. Atomic Theory and the Description of Nature: Four essays, with an introductory survey. Cambridge University Press, Cambridge, UK, 1961. vii + 119 pp. LCCN QC173.18 .B64 1961.

Bohr:1961:PAC

[Boh61b] Niels Bohr. Physique atomique et connaissance humaine. (French) [Atomic physics and human knowledge]. Gauthier-Villars et cie, Paris, France, 1961. 98 pp. LCCN ???? Translated from Danish to French by Edmond Bauer and Roland Omnes.

Bohr:1961:RML

[Boh61c] Niels Bohr. The Rutherford Memorial Lecture 1958: Reminiscences of the founder of nuclear science and of some developments based on his work. Proceedings of the Physical Society, London, 78(6):1083–1115, December 1, 1961. CODEN PPSOAU. ISSN 0370-1328. URL http://

stacks.iop.org/0370-1328/78/i=6/a=301n.

Bohr:1962:GSD

[Boh62a] Niels Bohr. The general significance of the discovery of the atomic nucleus. In Birks [Bir62], pages 43–44. LCCN ????

Bohr:1962:IKP

[Boh62b] Niels Bohr. Interview with T. S. Kuhn, A. Petersen, and E. Rudinger. In ????, editor, Archive for the History of Quantum Physics, page ?? University of California Press, Berkeley, CA, USA, 1962. LCCN ????

Bohr:1962:CAM

[Boh62c] Niels Bohr. On the constitution of atoms and molecules (1913). In Birks [Bir62], pages 228–256. LCCN ????

Bohr:1962:QTR

[Boh62d] Niels Bohr. On the quantum theory of radiation and the structure of the atom (1915). In Birks [Bir62], pages 283–307. LCCN ????

Bohr:1962:RFN

[Boh62e] Niels Bohr. Reminiscences of the founder of nuclear science and of some developments based on his work. In Birks [Bir62], pages 114–167. LCCN ????

Bohr:1963:RAB

[Boh63a] Aage Bohr. Remarks of Aage Bohr at Niels Bohr Memorial Session. Physics Today, 16(10):33, October 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bohr:1963:EAP

[Boh63b] Niels Bohr. Essays, 1958–1962, on Atomic Physics and Human Knowledge. Interscience Publishers, New York, NY, USA, 1963. x + 100 pp. LCCN QC6 .B599 1963.

Bohr:1963:FAW

[Boh63c] Niels Bohr. Fizyka atomowa a wiedza ludzka. (Polish) [Atomic physics and human knowledge]. Państwowe Wydawnictwo Naukowe, Warsawa, Poland, 1963. 151 + 1 pp. LCCN ???? Translated by Wacław Staszewski, Stanisław Szpikowski (1926–), and Armin Teske (1910–1967).

Bohr:1963:GSD

[Boh63d] Niels Bohr. The general significance of the discovery of the atomic nucleus. In Birks [Bir63], page ?? LCCN QC16.R8 B57 1962.

Bohr:1963:LLN

[Boh63e] Niels Bohr. Licht und Leben-noch einmal. (German) [Light and life: again]. Die Naturwissenschaften, 50(24):725–727, December 2, 1963. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bohr:1963:CAM

[Boh63f] Niels Bohr. On the Constitution of Atoms and Molecules. W. A. Benjamin, New York, NY, USA, 1963. lii + 77 pp. LCCN QC173 B677. Papers of 1913 reprinted from the Philosophical Magazine, and draft of an unpublished section on magnetism, with an introduction by Léon Rosenfeld.

Bohr:1964:KAOb

[Boh64a] Aage Bohr. Krigens år og atomvåbnenes perspektiver. (Danish) [The war years and the perspectives for atomic weapons]. In Bohr and Rozental [BR64], pages 184–206. LCCN QC16.B63 N5. English translation in [Roz67].

Bohr:1964:KAOa

[Boh64b] Aage Bohr. Krigens år og atomvåbnenes perspektiver. (Danish) [The war years and the prospects of atomic weapons]. In Niels Bohr: Hans liv or virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: His life and works told by a group of friends and co-workers] [RRK^{+}64], pages 191–214. LCCN ???? URL http://www.nba.nbi. dk/papers/rozental.htm. Contains comments on the Bohr–Heisenberg meeting in the summer of 1941.

Bohr:1964:OFD

[Boh64c] Hans Bohr. Om far. (Danish) [About Father]. In Bohr and Rozental [BR64], pages 315–327. LCCN QC16.B63 N5. English translation in [Roz67].

Bohr:1964:ABT

[Boh64d] Niels Bohr. Åbent brev til De Forenede Nationer. (Danish) [Open letter to the United Nations]. In Bohr and Rozental [BR64], pages 328– 340. LCCN QC16.B63 N5. English translation in [Roz67].

Bohr:1964:AOM

[Boh64e] Niels Bohr. Atomfysik og menneskelig erkendelse II: Artikler fra årene 1958–1962. (Danish) [Atomic physics and human knowledge II: Articles from the years 1958–1962]. J. H. Schultz, København, Danmark, second edition, 1964. 124 pp. LCCN ????

Bohr:1964:YZL

[Boh64f] Niels Bohr. Yuan zi lun he zi ran de miao shu. Shang wu yin shu guan, Beijing, China, 1964. ix + 85 pp. LCCN QC173.18 .B6412 1964.

Bohr:1965:NLS

[Boh65a] Niels Bohr. Nobel lecture: The structure of the atom. In Anonymous, editor, Nobel lectures: including presentation speeches and laureates’ biographies: physics 1922–1941, pages 7–43. Elsevier, Amsterdam, The Netherlands, 1965. LCCN QC71 .P455 1965. URL http://nobelprize.

org/nobel_prizes/physics/laureates/1922/bohr-lecture.html;
 http://nobelprize.org/nobel_prizes/physics/laureates/1922/

bohr-lecture.pdf. Foreword by Arne Tiselius.

Bohr:1965:QTA

[Boh65b] Niels Bohr. Quantum theory of atomic constitution [Faraday lecture]. In Nuclear Forces [Bri65], pages 138–143. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965. Reprint of [Boh32b].

Bohr:1966:AME

[Boh66a] Niels Bohr. Atomphysik und menschliche Erkenntnis II: Aufsätze und Vorträge aus den Jahren 1958–1962. (German). [Atomic physics and human knowledge II. Essays and lectures from the years 1958–1962], volume 123 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1966. viii + 104 pp. LCCN ????

Bohr:1966:EAP

[Boh66b] Niels Bohr. Essays (1958-1962) on atomic physics and human knowledge. Random House, New York, NY, USA, 1966. x + 100 pp. LCCN ????

Bohr:1967:QTL

[Boh67a] Niels Bohr. On the quantum theory of line-spectra. In van der Waerden [vdW67], pages 95–137. LCCN QC174.12 S655.

Bohr:1967:OLU

[Boh67b] Niels Bohr. Open letter to the United Nations. In Rozental [Roz67], pages 340–352. LCCN QC16.B63 N53. Printed in Danish and English in 1950, and reprinted in 1967.

Bohr:1967:QTR

[Boh67c] Niels Bohr. The quantum theory of radiation. In van der Waerden [vdW67], pages 159–176. LCCN QC174.12 S655.

Bohm:1971:BVC

[Boh71a] David Bohm. On Bohr’s views concerning the quantum theory. In Bastin [Bas71], page ?? ISBN 0-521-07956-X. LCCN QC174.1

.Q37. URL http://www.loc.gov/catdir/enhancements/fy0907/
77127237-d.html;   http://www.loc.gov/catdir/enhancements/
fy0907/77127237-t.html.

Bohr:1970:INT

[Boh71b] Niels Bohr. Izbrannye naucnye trudy. (Russian) [Selected scientific works]. Akademija nauk Sojuza SSR, Moskva, USSR, 1970–1971. ISBN ???? ???? pp. LCCN ???? Two volumes. Volume 1: Stat’i 1909–1925 (1970). Volume 2: Stat’i 1925–1961 (1971).

Bohr:1972:PAC

[Boh72] Niels Bohr. Physique atomique et connaissance humaine. (French) [Atomic physics and human knowledge]. Gauthier-Villars et cie, Paris, France, 1972. 196 pp. LCCN ????

Bohr:1975:QPN

[Boh75a] Niels Bohr. 7. the quantum postulate and the new development of atomic theory. In Mehra [Meh75], chapter 6, pages 151–179. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Bohr:1975:ATQ

[Boh75b] Niels Bohr. 8. application of the theory of quanta to atomic problems. In Mehra [Meh75], chapter 4, pages 109–111. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Bohr:1977:AAC

[Boh77] N. Bohr. Actions of atoms at collisions. Uspekhi Fizicheskikh Nauk, 122 (4):571–574, ???? 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Bohr:1978:ATD

[Boh78] Niels Bohr. Atomic Theory and the Description of Nature. Four Essays, volume 1. AMS Press, New York, NY, USA, 1978. ISBN 0-404-14737-2. 119 pp. LCCN ????

Bohr:1981:CWV

[Boh81] Niels Bohr. Collected works. Vol. 2: Work on Atomic Physics (1912– 1917). North-Holland Publishing Co., Amsterdam, The Netherlands, 1981. ISBN 0-7204-1802-X. xv + 647 + 1 pp. URL http://

www.sciencedirect.com/science/bookseries/18760503/2. Work on

atomic physics (1912–1917), edited and with introductory material by L. Rosenfeld and Ulrich Hoyer.

Bohr:1984:AEM

[Boh84] Niels Bohr. Atomfizika és emberi megismerés: második kiadás. Gondolat, Budapest, Hungary, 1984. ISBN ???? 192 pp. LCCN ???? Translated from German to Hungarian.

Bohm:1985:BVC

[Boh85a] David Bohm. On Bohr’s views concerning the quantum theory. In French and Kennedy [FK85], pages 153–159. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohr:1985:CPA

[Boh85b] N. Bohr. The causality problem in atomic physics. Uspekhi Fizicheskikh Nauk, 147(2):343–355, ???? 1985. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). Reprinted from New Theories in Physics, Pg 11–30, 1939.

Bohr:1985:IOW

[Boh85c] N. Bohr. The ideal of an open world. Impact of Science on Society, 35(1):27–??, ???? 1985. ISSN 0019-2872. Reprinted from Impact of Science on Society, Vol 1, Pg 68, 1950.

Bohr:1985:OLUa

[Boh85d] N. Bohr. Open letter to the United Nations. Uspekhi Fizicheskikh Nauk, 147(2):357–366, ???? 1985. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). Reprinted from Science, Vol 112, Pg 1–6, 1950.

Bohr:1985:AME

[Boh85e] Niels Bohr. Atomphysik und menschliche Erkenntnis: Aufsätze und Vorträge aus den Jahren 1930 bis 1961. (German). [Atomic physics and human knowledge II. Essays and lectures from the years 1930–1961], volume 20 of Facetten der Physik. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1985. ISBN 3-528-08910-5. 160 pp. LCCN ???? With a new foreword by Karl von Meyenn.

Bohr:1985:BED

[Boh85f] Niels Bohr. The Bohr–Einstein dialogue. In French and Kennedy [FK85], pages 121–140. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohr:1985:CL

[Boh85g] Niels Bohr. The Como lecture. In French and Kennedy [FK85], pages 191–194. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohr:1985:EAO

[Boh85h] Niels Bohr. Energy from the atom: an opportunity and a challenge. In French and Kennedy [FK85], pages 261–265. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohr:1985:LL

[Boh85i] Niels Bohr. Light and life. In French and Kennedy [FK85], pages 311– 319. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohr:1985:NPL

[Boh85j] Niels Bohr. Nobel Prize lecture: the structure of the atom. In French and Kennedy [FK85], pages 91–97. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohr:1985:OLUb

[Boh85k] Niels Bohr. Open letter to the United Nations. In French and Kennedy [FK85], pages 288–296. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohr:1985:SRR

[Boh85l] Niels Bohr, editor. Sixteen research reports, volume 509; vol. 41:1– 10; 25 of v. 1: Contributions to geology; v. 1: Matematisk-fysiske Meddelelser; v. 2: Biologiske skrifter. Det Kongelige Danske videnskabernes selskab, Munksgaard, Denmark, 1985. ISBN 87-7304-156-4 (paperback: part 1). ISSN 0023-3323. ???? pp. LCCN AS281 .D215 Bd. 41 1985.

Bohr:1985:TAN

[Boh85m] Niels Bohr. Transmutations of atomic nuclei. In French and Kennedy [FK85], pages 235–239. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohr:1985:T

[Boh85n] Niels Bohr. The trilogy. In French and Kennedy [FK85], pages 76–90. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bohr:1987:QPR

[Boh87a] N. Bohr. The quantum postulate and the recent development of atomic theory [Nature 121 (1928), Suppl., 580–590]. In Pekka (Pekka Johannes) Lahti and Peter Mittelstaedt, editors, Symposium on the Foundations of Modern Physics, 1987: the Copenhagen interpretation 60 years after the Como lecture, Joensuu, Finland, 6–8 August, 1987, pages 1–18. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1987. ISBN 9971-5-0382-4, 9971-5-0460-X. LCCN QC173.96 .S9551 1987; QC173.96 .F68 1987.

Bohr:1987:ATD

[Boh87b] Niels Bohr. Atomic theory and the description of nature, volume 1 of Philosophical writings of Niels Bohr. Ox Bow Press, Woodbridge, CT, USA, 1987. ISBN 0-918024-51-X, 0-918024-50-1 (paperback). 119 pp. LCCN QC5.58 .B64213 1987 v.; QC173 .B57213 1987; QC173 .B63a 1987.

Bohr:1987:EAPa

[Boh87c] Niels Bohr. Essays 1932–1957 on atomic physics and human knowledge: Philosophical writings, volume 2 of The Philosophical writings of Niels Bohr. Ox Bow Press, Woodbridge, CT, USA, 1987. ISBN 0-918024-53-6, 0-918024-52-8 (paperback). viii + 101 pp. LCCN QC5.58 .B64213 1987 vol. 2; QC173 1987.

Bohr:1987:EAPb

[Boh87d] Niels Bohr. Essays 1958–1962 on atomic physics and human knowledge, volume 3 of The Philosophical writings of Niels Bohr. Ox Bow Press, Woodbridge, CT, USA, 1987. ISBN 0-918024-55-2, 0-918024-54-4 (paperback). x + 100 pp. LCCN QC5.58 .B64213 1987 vol. 3; QC174.12 1987.

Bohr:1988:TAD

[Boh88] Niels Bohr. La teoría atómica y la descripción de la naturaleza: cuatro ensayos precedidos de una introducción. (Spanish) [Atomic Theory and the Description of Nature: Four essays, with an introductory survey], volume 525 of Alianza Universidad. Ciencias. Alianza Editorial, Madrid, Spain, 1988. ISBN 84-206-2525-6. 156 pp. LCCN QC173.18 .B6418 1988.

Bohr:1991:CC

[Boh91] Niels Bohr. Causality and complementarity. In Ferris and Fadiman [FF91], pages 801–807. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Bohr:1993:TAD

[Boh93] Niels Bohr. La théorie atomique et la description des phénomènes. (French) [Atomic theory and the description of phenomena]. Les Grands Classiques Gauthier-Villars. [Gauthier-Villars Great Classics]. Éditions Jacques Gabay, Sceaux, France, French edition, 1993. ISBN 2-87647-153-1. vi + 115 pp. Translated from the Danish and German by Andrée Legros and Léon Rosenfeld.

Bohr:2001:A

[Boh01] Niels Henrik David Bohr. Atom. In Hoiberg [Hoi01], pages 97–106. ISBN 81-88237-00-0. LCCN AS911.N9 A15 2001.

Bohr:2005:NBC

[Boh05] Niels Bohr. Niels Bohr — collected works. Vol. 11. The political arena (1934–1961). Elsevier/North-Holland, Amsterdam, 2005. ISBN 0-444-51336-1. xxiv + 754 pp. Edited by Finn Aaserud.

Bohr:2007:FSC

[Boh07a] Niels Bohr. Foreword: Science and civilization. In Masters and Way [MW07], page ?? ISBN 1-59558-227-4 (hardcover). LCCN UG1282.A8

O54 2007. URL http://thenewpress.com/index.php?option=com_
title&task=view_title&metaproductid=1703;   http://www.loc.
gov/catdir/toc/ecip0718/2007020838.htm. Foreword by Niels Bohr.

Introduction by Arthur H. Compton. Reprint of [MW46a].

Bohr:2007:NBC

[Boh07b] Niels Bohr. Niels Bohr — collected works. Vol. 12. Popularization and people (1911–1962). Elsevier/North-Holland, Amsterdam, The Netherlands, 2007. ISBN 0-444-52946-2. xxiv + 585 pp. LCCN QC3 .B584; QC3 .B677; QC782.5.P66 B64 2007. URL http://www.sciencedirect.

com/science/book/9780444529466. Edited by Finn Aaserud.

Bohr:2008:GRS

[Boh08] Niels Henrik David Bohr. Genshi riron to shizen kijutsu. Misuzu Shobō, Tōkyō, Japan, 2008. ISBN 4-622-07357-9. 486 pp. LCCN ???? Translation to Japanese by Takeshi Inoue of [Boh34].

Bohr:2010:APH

[Boh10] Niels Bohr. Atomic Physics and Human Knowledge. Dover books on physics. Dover, New York, NY, USA, 2010. ISBN 0-486-47928-5. viii + 101 pp. LCCN QC6 .B598 2010. URL http://catdir.loc.gov/

catdir/enhancements/fy1010/2010022935-d.html.

Bohr:2011:VDV

[Boh11a] Hanne Bohr. Vil du vide noget om Niels Bohr? (Danish) [Would you like to know something about Niels Bohr?]. Web document., February

2, 2011. URL http://www.theofilus.dk/nielsbohr/.

Bohr:2011:ATD

[Boh11b] Niels Bohr. Atomic Theory and the Description of Nature: Four Essays with an Introductory Survey. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-62805-9 (paperback). 119 pp. LCCN ????

Bokulich:2008:PDE

[Bok08a] Alisa Bokulich. Paul Dirac and the Einstein–Bohr debate. Perspectives on Science, 16(1):103–114, Spring 2008. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). URL http://muse.jhu.edu/

journals/perspectives_on_science/v016/16.1bokulich.html;
http://muse.jhu.edu/journals/perspectives_on_science/v016/
16.1bokulich.pdf;   http://www.mitpressjournals.org/doi/abs/
10.1162/posc.2008.16.1.103.

Bokulich:2008:RQC

[Bok08b] Alisa Bokulich. Reexamining the Quantum–Classical Relation: Beyond Reductionism and Pluralism. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-85720-1. x + 195 pp. LCCN QA805 .B685

2008. URL http://www.loc.gov/catdir/enhancements/fy0834/
2008018902-b.html;   http://www.loc.gov/catdir/enhancements/
fy0834/2008018902-d.html;   http://www.loc.gov/catdir/enhancements/
fy0834/2008018902-t.html.

Bolles:2004:EDB

[Bol04] Edmund Blair Bolles. Einstein defiant (Bohr unyielding): genius versus genius in the quantum revolution. Joseph Henry Press, Washington, DC, USA, 2004. ISBN 0-309-08998-0 (hardcover). viii + 348 pp. LCCN

QC173.98 .B65 2004. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http://www.loc.gov/catdir/toc/ecip0410/
2003023735.html.

Bonnet:2010:CPD

[Bon10] L. Bonnet. Classical photodissociation dynamics with Bohr quantization. Journal of Chemical Physics, 133(17):174108, November 7, 2010. CODEN JCPSA6. ISSN 1089-7690.

Born:1923:QSGa

[Bor23] Max Born. Quantentheorie und Störungsrechnung. (German) [Quantum theory and perturbation calculation ]. Die Naturwissenschaften, 11 (27):537–542, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Borgers:1958:BRB

[Bor58] Alfons Borgers. Book review: The Causality Problem in Atomic Physics, by Niels Bohr; Language and Reality in Modern Physics, by Werner Heisenberg; Die Stellung der Logik im Gebaude der Heutigen Wissenschaft, by Evert Willem Beth. Journal of Symbolic Logic, 23(1): 66, March 1958. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886

(electronic). URL http://www.jstor.org/stable/2964516.

Born:1963:AAGa

[Bor63] Max Born. Ausgewählte Abhandlungen. (German) [Selected works], volume 1. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1963. xxiv + 718 pp. LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Borresen:2012:FNP

[Bør12] Hans Christofer Børresen. Flawed nuclear physics and atomic intelligence in the campaign to deny Norwegian heavy water to Germany, 1942–1944. Physics in Perspective (PIP), 14(4):471–497, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00016-
012-0094-9.

Borrelli:2013:BRN

[Bor13] Arianna Borrelli. Book review: Niels Bohr and the Quantum Atom: The Bohr Model of Atomic Structure 1913–1925 [Helge Kragh, Oxford, Oxford University Press, 2012. vi + 410 pp. ISBN 978-0-19-965498-7, £35.50, US$62.99 (hardcover)]. International Studies in the Philosophy of Science, 27(2):222–224, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/02698595.2013.813250.

Borrelli:2016:BRF

[Bor16] Arianna Borrelli. Book review: Finn Aaserud; Helge Kragh, eds. One Hundred Years of the Bohr Atom: Proceedings from a Conference. Isis, 107(3):671–672, September 2016. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bouvier:1949:AOBa

[Bou49a] Robert Bouvier. Analyses d’ouvrages: L’Évolution de la notion de phénomène physique, des Primitifs à Bohr et Louis de Broglie. Leçons sur l’histoire de la pensée scientifique, professées à l’Université libre de Bruxelles par Jean Pelseneer. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):283–286, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23904777.

Bouvier:1949:AOL

[Bou49b] Robert Bouvier. Analyses d’ouvrages: L’Évolution de la notion de phénomène physique, des Primitifs à Bohr et Louis de Broglie. Leçons sur l’histoire de la pensée scientifique, professées à l’Université libre de Bruxelles par Jean Pelseneer. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):283–286, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23904777.

Bohr:1939:NRC

[BPP39] Niels Bohr, Rudolf Peierls, and George Placzek. Nuclear reactions in the continuous energy region. Nature, 144(3639):200–201, July 29, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v144/n3639/
pdf/144200a0.pdf;   https://ui.adsabs.harvard.edu/#abs/1939Natur.
144..200B.

Bohr:1997:NRC

[BPP97] N. Bohr, R. Peierls, and G. Placzek. Nuclear reactions in the continuous energy region. In Dalitz and Peierls [DP97], pages 260–263. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (ebook). LCCN QC21.2 .P42 1997. URL https://ui.adsabs.harvard.

edu/#abs/1997sspr.book..260B.

Bohr:1986:NP

[BPR86] Niels Bohr, Rudolf E. (Rudolf Ernst) Peierls, and Erik Rüdinger. Nuclear Physics (1929–1952). North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISBN 0-7204-1800-3 (set), 0-444-86929-8. xviii + 693 pp. LCCN QC776; SCI; M88.D00077.

Bohr:1933:FME

[BR33] N. Bohr and L. Rosenfeld. Zur Frage der Meßbarkeit der elektromagnetischen Feldgrößen. (German) [On the question of measurability of electromagnetic field quantities]. Kongelige Danske Videnskabernes Selskab. Mathematisk-fysiske meddelelser, 12(8):1–65, 1933. Reprinted in [BR79b].

Bohr:1950:FCM

[BR50] Niels Bohr and Léon Rosenfeld. Field and charge measurements in quantum electrodynamics. Physical Review, 78(6):794–798, June 1950. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://dieumsnh.qfb.umich.mx/archivoshistoricosMQ/ModernaHist/
ModernaHist/Bohr%20rosenfeld.pdf;   http://prola.aps.org/abstract/

PR/v78/i6/p794_1. Reprinted in [BR79a].

Bohr:1955:AE

[BR55] Niels Bohr and I. I. Rabi. Albert Einstein: 1879–1955. Scientific American, 192(6):31–32, June 1955. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v192/n6/pdf/scientificamerican0655-
31.pdf.

Bohr:1964:NBH

[BR64] Niels Bohr and S. (Stefan) Rozental, editors. Niels Bohr: hans liv og virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: his life and works told by a circle of friends and colleagues]. J. H. Schultz, København, Danmark, 1964. 341 pp. LCCN QC16.B63 N5. English translation in [Roz67].

Bohr:1979:FCM

[BR79a] Niels Bohr and Léon Rosenfeld. Field and charge measurements in quantum electrodynamics. In Cohen and Stachel [CS79a], chapter II.2, pages 401–412. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21

QC7. URL http://link.springer.com/chapter/10.1007/978-94-

009-9349-5_27. Reprint of [BR50].

Bohr:1979:QME

[BR79b] Niels Bohr and Léon Rosenfeld. On the question of the measurability of electromagnetic field quantities. In Cohen and Stachel [CS79a], chapter II.1, pages 357–400. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21

QC7. URL http://link.springer.com/chapter/10.1007/978-94-

009-9349-5_27. Reprint of [BR33].

Bohr:1985:NBO

[BR^{+}85] Niels Bohr, Erik Rüdinger, et al., editors. Niels Bohr og den moderne atomfysik: fem offentlige foredrag i Videnskabernes selskab holdt i 100året for Niels Bohrs fødsel. (Danish) [Niels Bohr and modern atomic physics: five public lecture in the Academy of Sciences held on the 100th anniversary of Niels Bohr’s birth]. Det Kongelige Danske videnskabernes selskab, København, Danmark, 1985. ISBN 87-7245-095-9. 116 pp. LCCN QC7.5 .N54 1985.

Bragg:1961:RML

[Bra61] Sir Lawrence Bragg, F.R.S. The Rutherford Memorial Lecture, 1960. The development of X-ray analysis. Proceedings of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 262(1309):145–158, 1961. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (elec-

tronic). URL http://rspa.royalsocietypublishing.org/content/

262/1309/145. Delivered at the University of Canterbury, Christchurch, New Zealand, on 21 September 1960.

Brandt:2009:HCD

[Bra09] Siegmund Brandt. The harvest of a century: discoveries of modern physics in 100 episodes. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954469-7 (hardcover). xiv + 500 pp. LCCN QC7 .B64 2009.

Bhattacharjee:2012:BSQ

[BRB12] Shayak Bhattacharjee, D. S. Ray, and J. K. Bhattacharjee. Bohr– Sommerfeld quantisation and molecular potentials. Journal of Mathematical Chemistry, 50(4):819–832, April 2012. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (electronic). URL http://link.

springer.com/article/10.1007/s10910-011-9926-0.

Brennan:1997:HPS

[Bre97] Richard P. Brennan. Heisenberg probably slept here: the lives, times, and ideas of the great physicists of the 20th Century. Wiley, New York, NY, USA, 1997. ISBN 0-471-15709-0 (cloth). xi + 274 pp. LCCN QC15 .B74 1997; 97.E02533.

Bohr:1972:CW

[BRH^{+}99] Niels Bohr, L. (Léon) Rosenfeld, Ulrich Hoyer, Erik Rüdinger, Rudolf Peierls, Jørgen Kalckar, Jens Thorsen, Finn Aaserud, J. Rud Nielsen, Klaus Stolzenburg, and David Favrholdt. Collected works. North-Holland Publishing Co., Amsterdam, The Netherlands, 1972– 1999. ISBN 0-7204-1800-3 (US: set). various pp. LCCN QC3 .B584.

Bridgman:1958:RNB

[Bri58] P. W. Bridgman. Remarks on Niels Bohr’s talk. Dædalus, 87(2):175– 177, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.org/stable/20026445.

Brink:1965:NF

[Bri65] David Maurice Brink. Nuclear Forces, volume 354 of Commonwealth and international library: Selected readings in physics. Pergamon Press, New York, NY, USA, 1965. ISBN 0-08-011034-7. viii + 232 pp. LCCN QC173 .B8513 1965.

Brown:1936:RBS

[Bro36] G. Burniston Brown. Review: Where Is Science Going? by Max Planck, James Murphy; Atomic Theory and the Description of Nature by Niels Bohr; Science and the Human Temperament by Erwin Schrödinger, James Murphy. Philosophy, 11(43):366–367, July 1936. ISSN 0031-8191 (print), 1469-817X (electronic). URL http://www.jstor.org/stable/

3746205.

Bromberg:1971:INB

[Bro71] Joan Bromberg. The impact of the neutron: Bohr and Heisenberg. Historical Studies in the Physical Sciences, 3(??):307–341, ???? 1971. CO-

DEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.org/stable/
27757321.

Bronowski:1973:AM

[Bro73a] Jacob Bronowski. The Ascent of Man. British Broadcasting Corporation, London, UK, 1973. ISBN 0-563-10498-8. 448 pp. LCCN Q175 .B7918 1973; CB151.

Bronowski:1973:SAR

[Bro73b] Jacob Bronowski. Structure in the atom: Rutherford and Niels Bohr. In The Ascent of Man [Bro73a], pages 333–341. ISBN 0-563-10498-8. LCCN Q175 .B7918 1973; CB151.

Brown:1986:HJC

[Bro86] Harvey R. Brown. Book review: Hendry, John [1984]: The Creation of Quantum Mechanics and the Bohr–Pauli Dialogue. D. Reidel Publishing Company. xi + 177 pp. (ISBN 9O-277-1648-X). British Journal for the Philosophy of Science, 37(4):497–506, December 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic).

Brown:1995:BRK

[Bro95] Laurie M. Brown. Book review: Karl von Meyenn, Wolfgang Pauli: Scientific Correspondence with Bohr, Einstein, Heisenberg a. o.; Vol. III: 1940–1949. Physics Today, 48(4):86, April 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v48/i4/p86/s1.

Brown:1999:BRN

[Bro99] Laurie M. Brown. Book review: Niels Bohr and Jørgen Kalckar: Niels Bohr: Collected Works. Volume 7: Foundations of Quantum Physics II (1933–1958). Isis, 90(1):143–144, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/237526.

Brown:2000:BRBc

[Bro00a] Laurie M. Brown. Book review: Niels Bohr: Collected Works. Volume 10: Complementarity beyond Physics (1928–1962), by Niels Bohr and David Favrholdt. Isis, 91(3):619–620, September 2000. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237970.

Brown:2000:BRBb

[Bro00b] Neil Brown. Book review: Quantum Generations. A history of physics in the twentieth century, by Helge Kragh. Physics Education, 35(2): 145, 2000. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (elec-

tronic). URL http://iopscience.iop.org/0031-9120/35/2/703.

Brock:2003:NBP

[Bro03] Steen Brock. Niels Bohr’s philosophy of quantum physics in the light of the Helmholtzian tradition of theoretical physics. Logos, Berlin, Germany, 2003. ISBN 3-8325-0200-9 (paperback). 301 pp. LCCN QC174.13 .B763 2003.

Brock:2009:OWE

[Bro09] Steen Brock. Old wine enriched in new bottles: Kantian flavors in Bohr’s viewpoint of complementarity. In Bitbol et al. [BKP09], pages 301–316. ISBN 1-4020-9509-0 (cased). LCCN QC6 CON 2009. URL

http://link.springer.com/10.1007/978-1-4020-9510-8.

Brown:2015:PDV

[Bro15] Brandon R. Brown. Planck: driven by vision, broken by war. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-021947-5 (hardcover), 0-19-021948-3 (e-book). xviii + 258 pp. LCCN QC16.P6 B76 2015.

Bohr:1938:PNF

[BS38] Niels Bohr and P. Scherrer. Physique nucléaire. (French) [Nuclear physics]. Hermann et Cie., Paris, France, 1938. 56 pp. LCCN ????

Bohr:1936:CNB

[BS50] Niels Bohr and Leo Szilard. [correspondence: Niels Bohr and Leo Szilard]. Five letters between the authors held in the Leo Szilard Pa-

pers archive., 1936–1950. URL http://library.ucsd.edu/dc/object/
bb25612766.

Bohr:1987:NBE

[BS87] Niels Bohr and John T. Sanders. Niels Bohr: essays and papers: typescript. ????, ????, 1987. ???? pp. LCCN ???? 2 volumes.

Bricmont:1999:SDC

[BSMB99] Jean Bricmont, Alan Sokal, N. David Mermin, and Mara Beller. Sokalratic debate continues, fueled by Latour and Copenhagen interpretations. Physics Today, 52(8):15–17, 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v52/i8/p15/s1.

Bub:1973:USB

[Bub73] Jeffrey Bub. Under the spell of Bohr. British Journal for the Philosophy of Science, 24(1):78–90, March 1973. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://bjps.oxfordjournals.

org/content/24/1/78.full.pdf+html.

Bub:1989:BRE

[Bub89] Jeffrey Bub. On Bohr’s response to EPR: A quantum logical analysis. Foundations of Physics, 19(7):793–805, July 1989. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF01889301.

Bub:1990:BRB

[Bub90a] Jeffrey Bub. Book review: Niels Bohr’s Philosophy of Physics by Dugald Murdoch. Philosophy of Science, 57(2):344–347, June 1990. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187848.

Bub:1990:BRE

[Bub90b] Jeffrey Bub. On Bohr’s response to EPR: II. Foundations of Physics, 20(8):929–941, August 1990. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/
10.1007/BF00738373.

Buhrke:1998:LIE

[Büh98a] Thomas Bührke. “Im ersten Augenblick eine ungeheuerliche und für das Vorstellungsvermögen fast unerträgliche Zumutung.” Niels Bohr (1885–1962). (German) [“At first, an outrageous and almost unbearable impossibility for imagination.” Niels Bohr (1885–1962)]. In Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner] [Büh98b], pages 162–183. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). LCCN ????

Buhrke:1998:NAS

[Büh98b] Thomas Bührke. Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner], volume 1202 of Beck’sche Reihe. Verlag C. H. Beck, München, Germany, 1998. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). 258 pp. LCCN ????

Bunge:1979:EBD

[Bun79] Mario Bunge. The Einstein–Bohr debate over quantum mechanics: who was right about what? In H. Nelkowski, A. Hermann, H. Poser, R. Schrader, and R. Seiler, editors, Einstein Symposion, Berlin (1979), volume 100 of Lecture Notes in Physics, pages 204–219. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979.

URL https://ui.adsabs.harvard.edu/#abs/1979LNP...100..204B.

Bunge:1985:PNB

[Bun85] Mario Bunge, editor. The Philosophy of Niels Bohr, volume 85/03 of Cahiers d’épistémologie. Université du Québec à Montréal, Montréal, QC, Canada, 1985. ISBN 2-920250-23-X. ISSN 0228-7080. LCCN QC16.B63 B86 1985.

Bunge:1988:NBP

[Bun88] Mario Bunge. Niels Bohr’s philosophy. Philosophia Naturalis, 25(3–4): 399–415, 1988. CODEN ???? ISSN 0031-8027.

Bunge:1992:PNB

[Bun92] Mario Bunge. La philosophie de Niels Bohr. Horizons philosophiques, 2(2):27–50, 1992. ISSN 1181-9227 (print), 1920-2954 (electronic). URL

https://id.erudit.org/iderudit/800894ar.

Burgers:1918:AVR

[Bur18] Johannes Martinus Burgers. Het Atoommodel van Rutherford–Bohr. (Dutch) [The Atomic Model of Rutherford–Bohr]. PhD thesis, Rijksuniversiteit te Leiden, Leiden, The Netherlands, 1918. xix + 265 pp. Uit Archives du Musee Teyler, series 3, vol. 4.

Bussey:2012:QRB

[Bus12] Peter J. Bussey. The Quest for Reality: Bohr and Wittgenstein — Two Complementary Views, by S. Stenholm, Scope: general interest. Level: scientist. Contemporary Physics, 53(3):254–255, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2020:BRN

[Bus20] Peter J. Bussey. Book review: Niels Bohr: a very short introduction, by J. L. Heilbron, Oxford, Oxford University Press, 2020, 160 pp., $11.95 (paperback), ISBN 978-0-19-881926-4. Scope: general interest, biography. Level: scientist. Contemporary Physics, 61(1):67–68, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bensaude-Vincent:1985:LCT

[BV85] Bernadette Bensaude-Vincent. L’évolution de la complémentarité dans les textes de Bohr (1927–1939). (French) [The evolution of complementarity in Bohr’s papers (1927–1939)]. Revue d’Histoire des Sciences, 38(3–4):231–250, 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0048-7996, 0151-4105, 1969-6582 (electronic).

URL http://www.persee.fr/web/revues/home/prescript/article/
rhs_0151-4105_1985_num_38_3_4006.

Bacciagaluppi:2009:QTC

[BV09] Guido Bacciagaluppi and Antony Valentini, editors. Quantum theory at the crossroads: reconsidering the 1927 Solvay conference. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-81421-9 (hardcover). LCCN QC173.96 .B33 2009. URL http://www.loc.gov/

catdir/toc/ecip0817/2008019585.html.

Bohr:1939:FP

[BW39a] Niels Bohr and John A. Wheeler. The fission of protactinium. Physical Review, 56(10):1065–1066, November 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v56/i10/p1065_2.

Bohr:1939:MNF

[BW39b] Niels Bohr and John Archibald Wheeler. The mechanism of nuclear fission. Physical Review (2), 56(5):426–450, September 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//prola.aps.org/abstract/PR/v56/i5/p426_1.

Bohr:1985:MNF

[BW85] Niels Bohr and John A. Wheeler. The mechanism of nuclear fission. In French and Kennedy [FK85], pages 240–243. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Born:2000:BRQ

[BW00a] Gustav Born and Alvin M. Weinberg. Copenhagen — roles in quantum mechanics. Physics Today, 53(7):74–75, 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v53/i7/p74/s1.

Born:2000:CRQ

[BW00b] Gustav Born and Alvin M. Weinberg. Copenhagen — roles in quantum mechanics. Physics Today, 53(7):74–75, 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v53/i7/p74/s1.

Byrne:2010:MWH

[Byr10] Peter Byrne. The many worlds of Hugh Everett III: multiple universes, mutual assured destruction, and the meltdown of a nuclear family. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-955227-4 (hardcover), 0-19-965924-9 (paperback). xiii + 436 pp. LCCN QC774.E94 .B97 2012. URL

https://global.oup.com/academic/product/the-many-worlds-of-
hugh-everett-iii-9780199552276.

Belyaev:1985:NBP

[BZ85] S. T. Belyaev and V. G. Zelevinski. Niels Bohr and the physics of the atomic nucleus. Soviet Physics. Uspekhi, 28(10):854–??, ???? 1985. CODEN SOPUAP. ISSN 0038-5670. URL http://stacks.iop.org/

0038-5670/28/i=10/a=A02.

Campbell:1999:RSS

[Cam99] John Campbell. Rutherford: scientist supreme. AAS Publications, Christchurch, New Zealand, 1999. ISBN 0-473-05700-X (hardcover). xvi + 516 + 32 + 16 pp. LCCN ????

Camilleri:2007:BHD

[Cam07] Kristian Camilleri. Bohr, Heisenberg and the divergent views of complementarity. Studies in History and Philosophy of Modern Physics, 38(3):514–528, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S135521980600092X.

Camejo:2008:DTB

[Cam08] Silvia Arroyo Camejo. Il dibattito tra Bohr e Einstein. (Italian). [The debate between Bohr and Einstein]. In Il bizzarro mondo dei quanti. (Italian) [The strange world of quanta], I blu, pagine di scienza, pages 87–101. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 88-470-0644-9. LCCN QC174.12 .C36 2008.

URL http://www.ellibs.com/book/9788847006447.

Camilleri:2009:HIQ

[Cam09] Kristian Camilleri. Heisenberg and the interpretation of quantum mechanics: the physicist as philosopher. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-88484-5 (hardcover). xii + 199 pp.

LCCN QC16.H395 .C34 2009. URL http://www.loc.gov/catdir/toc/
ecip0825/2008034113.html.

Canaday:2000:NML

[Can00] John Canaday. The Nuclear Muse: Literature, Physics, and the First Atomic Bombs. University of Wisconsin Press, Madison, WI, USA, 2000. ISBN 0-299-16850-6 (hardcover), 0-299-16854-9 (paperback). xviii + 310 pp. LCCN QC791.96 .C36 2000. URL http://www.e-streams.com/

es0409/es0409_1493.htm.

Cao:1990:BRB

[Cao90] Tian Yu Cao. Book review: The Description of Nature: Niels Bohr and the Philosophy of Quantum Physics by John Honner. Isis, 81(1):151– 152, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234153.

Carrier:1983:LBP

[Car83] Martin Carrier. Lakatos und Bohrs Programm: Entgegnung auf eine Kritik von Hans Radder. (German) [Lakatos and Bohr’s program: a response to a criticism of Hans Radder]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2): 368–371, ???? 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170654. See

[Rad82].

Carson:1997:BRBb

[Car97] Cathryn Carson. Book review: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg, u.a. Volume 4, Part 1: 1950–1952. [Wolfgang Pauli: Scientific Correspondence with Bohr, Einstein, Heisenberg, et al. Volume 4, Part 1: 1950–1952.] by Wolfgang Pauli; Karl von Meyenn. Isis, 88(4):726–727, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237871.

Carruthers:2004:BED

[Car04] Rebecca Carruthers. The Bohr–Einstein dialogue: a rhetorical and genre analysis. M.A. dissertation, Simon Fraser University, Burnaby, BC, Canada, 2004. 233 pp.

Casimir:1935:WBB

[Cas35] Hendrik B. G. Casimir. Über eine weniger bekannt Bohr’sche Theorie und ihre experimentelle Bestätigung. (German) [A less well-known Bohr theory and its experimental verification]. Journal of Jocular Physics, I: 24–25, 1935.

Casimir:1955:BE

[Cas55] Hendrik B. G. Casimir. Broken English. Journal of Jocular Physics, III:14–18, October 7, 1955. Reprinted in [Cas83, 122–125].

Casimir:1964:EFA

[Cas64] Hendrik G. B. Casimir. Erindringer fra årene 1929–1931. (Danish) [Memories from the years 1929–1931]. In Bohr and Rozental [BR64], pages 104–108. LCCN QC16.B63 N5. English translation in [Roz67].

Casimir:1983:HRH

[Cas83] H. B. G. (Hendrik Brugt Gerhard) Casimir. Haphazard reality: half a century of science. The Alfred P. Sloan Foundation series. Harper & Row, New York, NY, USA, 1983. ISBN 0-06-015028-9, 0-06-337031-X. xii + 356 pp. LCCN QC16.C36 A33 1983.

Casimir:1985:NBP

[Cas85] H. B. G. Casimir. Niels Bohr and the physics of simple phenomena. In French and Kennedy [FK85], pages 175–180. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Cassidy:2000:HPC

[Cas00] David C. Cassidy. A historical perspective on Copenhagen. Physics Today, 53(7):28–32, July 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/link/

phtoad/v53/i7/p28/s1.

Cassidy:2002:NLC

[Cas02] David C. Cassidy. New light on “Copenhagen” and the German Nuclear Project. Physics in Perspective (PIP), 4(4):447–455, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/s000160200004.

Cassidy:2009:BUH

[Cas09] David C. Cassidy. Beyond uncertainty: Heisenberg, quantum physics, and the bomb. Bellevue Literary Press, New York, NY, USA, 2009. ISBN 1-934137-13-8. 480 pp. LCCN QC16.W518 C37 2008.

Castellari:2014:FMF

[Cas14] Marco Castellari, editor. Formula e metafora: figure di scienziati nelle letterature e culture contemporanee. (Italian) [Formula and metaphor: figures of scientists in the literature and contemporary cultures], volume 8 of Di/segni. Ledizioni, Milano, Italy, 2014. ISBN 88-6705-207-1. LCCN ????

Cavecchi:2014:BHD

[Cav14] Mariacristina Cavecchi. Bohr e Heisenberg, o dell’indeterminazione. (Italian) [Bohr and Heisenberg, of the indetermination]. In Castellari [Cas14], pages 363–376. ISBN 88-6705-207-1. LCCN ????

Cockcroft:1934:SPN

[CCJ^{+}34] J. Cockcroft, J. Chadwick, F. Joliot, J. Joliot, N. Bohr, G. Gamov, P. A. M. Dirac, and W. Heisenberg, editors. Structure et propriétés des noyaux atomiques. Rapports et discussions du septième conseil de physique tenu à Bruxelles du 22 au 29 octobre 1933 sous les auspices de l’institut international de physique Solvay. (French) [Structure and properties of atomic nuclei. Reports and discussions of the Seventh Meeting on Physics held in Brussels from 22 to 29 October 1933 under the auspices of the Solvay International Institute of Physics]. Gauthier-Villars et cie, Paris, France, 1934. LCCN ???? Publiés par la commission administrative de l’institut.

Chen:2006:MAB

[CDH^{+}06] Goong Chen, Zhonghai Ding, Sze-Bi Hsu, Moochan Kim, and Jianxin Zhou. Mathematical analysis of a Bohr atom model. Journal of Mathematical Physics, 47(2):022107, February 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i2/p022107_s1.

Craig:1998:REP

[CF98] Edward Craig and Luciano Floridi, editors. Routledge Encyclopedia of Philosophy Online. Routledge, London, UK, version 2.0 edition, 1998. ISBN 0-415-16916-X (CD-ROM), 0-415-19608-6 (user guide), 0-415-16917-8 (10 vol set + CD-ROM). LCCN B51. URL http://www.

rep.routledge.com/.

Crease:2014:QMH

[CG14] Robert P. Crease and Alfred S. Goldhaber. The quantum moment: how Planck, Bohr, Einstein, and Heisenberg taught us to love uncertainty. W. W. Norton & Co., New York, NY, USA, 2014. ISBN 0-393-06792-0 (hardcover). vii + 332 pp. LCCN QC174.123 .C74 2014.

Chang:2009:ABP

[Cha09] Kenneth Chang. Aage Bohr, physicist’s son who won Nobel, dies at 87. New York Times, September 11, 2009. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/docview/1030645710.

Chevalley:1985:BRJ

[Che85a] Catherine Chevalley. Book review: John Hendry, The Creation of Quantum Mechanics and the Bohr–Pauli Dialogue. Dordrecht/Boston/ Lancaster: D. Reidel Publishing Company, 1984. Pp. 177. ISBN 90-277-1648-X. DF1 95. £23.50. British Journal for the History of Science, 18(3):362–364, November 1985. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/stable/

4026401.

Chevalley:1985:CLI

[Che85b] Catherine Chevalley. Complémentarité et langage dans l’interprétation de Copenhague. (French) [complementarity and language in the Copenhagen interpretation]. Revue d’Histoire des Sciences, 38(3–4):251–292, 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0048-7996, 0151-4105, 1969-6582 (electronic). URL http://www.persee.fr/web/

revues/home/prescript/article/rhs_0151-4105_1985_num_38_3_
4007.

Chevalley:1992:CEB

[Che92] C. Chevalley. Le conflit de 1926 entre Bohr et Schrödinger: un exemple de sous-détermination des théories. (French) [The conflict of 1926 between Bohr and Schrödinger: an example of underdetermination of theories]. In Bitbol and Darrigol [BD92], pages 81–94. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

Chevalley:1993:AOB

[Che93] Catherine Chevalley. Analyses d’ouvrages: Niels Bohr. His Heritage and Legacy: An Anti-realist View of Quantum Mechanics, ^{≪}Science and Philosophy ^{≫}, vol. 6 par Jan Faye; Niels Bohr’s Philosophical Background, ^{≪}Mat. Fys. Medd. Dan. Vid. Selsk. ^{≫}, 63 par David Favrholdt. Revue d’Histoire des Sciences, 46(1):111–113, January 1993. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23633753.

Chevalley:1994:NBC

[Che94a] Catherine Chevalley. Niels Bohr and contemporary philosophy. In Faye and Folse [FF94], pages 33–55. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67

vol. 153 1994; QC16 1994. URL http://www.loc.gov/catdir/
enhancements/fy0823/93024825-d.html;   http://www.loc.gov/
catdir/enhancements/fy0823/93024825-t.html;   https://link.
springer.com/book/10.1007/978-94-015-8106-6.

Chevalley:1994:NBW

[Che94b] Catherine Chevalley. Niels Bohr’s words and the Atlantis of Kantianism. In Faye and Folse [FF94], chapter 2, pages 33–55. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https://link.springer.

com/chapter/10.1007/978-94-015-8106-6_2.

Chevalley:1997:MPC

[Che97] Catherine Chevalley. Mythe et philosophie. La construction de “Niels Bohr” dans la doxographie. (French) [Myth and philosophy. the construction of “Niels Bohr” in the doxography]. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 34(3):569–603 (1998), 1997. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Chevalley:1999:DBB

[Che99a] Catherine Chevalley. 5. On the difference between Bohr’s epistemology and the so-called Copenhagen interpretation of quantum theory: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN QC174.13 .E65 1999.

Chevalley:1999:WDW

[Che99b] Catherine Chevalley. Why do we find Bohr obscure? In Greenberger et al. [GRZ99], pages 59–73. ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN QC174.13 .E65 1999.

Chiu:1960:BEH

[Chi60] Ying-Nan Chiu. Bohr Effect in Heme Compounds and the Imidazole Theory of Hemoglobin. Ph.D dissertation, Yale University, New Haven, CT, USA, 1960.

Cho:2002:LAP

[Cho02] Adrian Cho. Letters aver physicist supported Nazi bomb. Science, 295(5558):1211, February 15, 2002. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/295/5558/1211.1.summary.

Cohen:1996:RAR

[CHR96] Robert S. Cohen, Risto Hilpinen, and Qiu Renzong, editors. Realism and Anti-Realism in the Philosophy of Science: Beijing International Conference, 1992, volume 169 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346. LCCN ???? URL https://link.springer.

com/book/10.1007/978-94-015-8638-2.

Cohen:1983:PPP

[CL83] R. S. Cohen and L. Laudan, editors. Physics, Philosophy and Psychoanalysis: Essays in Honour of Adolf Grünbaum, volume 76 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346. xviii + 339 pp. LCCN ????

URL https://link.springer.com/book/10.1007/978-94-009-7055-

Clark:1988:BPC

[Cla88] S. J. Clark. Bohr’s principle of complementarity. Annals of Internal Medicine, 109(12):994–995, December 15, 1988. CODEN AIMEAS. ISSN 0003-4819 (print), 1539-3704 (electronic).

Clarke:2013:BRH

[Cla13] Imogen Clarke. Book review: Helge Kragh, Niels Bohr and the Quantum Atom. The Bohr Model of Atomic Structure 1913–1925. Oxford: Oxford University Press, 2012. vi + 410 pp., ISBN 978-0-19-965498-7. Nuncius, 28(2):511–512, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-02802020.

Cline:1965:QPQ

[Cli65] Barbara Lovett Cline. The questioners: physicists and the quantum theory. Crowell, New York, 1965. vii + 274 pp. LCCN QC15 .C4. URL

http://catalog.hathitrust.org/api/volumes/oclc/372589.html.

Cline:1987:MWM

[Cli87] Barbara Lovett Cline. Men who made a new physics: physicists and the quantum theory. University of Chicago Press, Chicago, IL, USA and London, UK, 1987. ISBN 0-226-11027-3 (paperback). xii + 274 pp. LCCN QC15 .C4 1987. US$11.95. URL http://www.loc.

gov/catdir/description/uchi051/87010786.htm;   http://www.
loc.gov/catdir/enhancements/fy0608/87010786-t.htm;   http:/
/www.loc.gov/catdir/enhancements/fy0609/87010786-b.htm.

Christmas-Mller:1985:NBO

[CM85] Wilhelm Christmas-Møller. Niels Bohr og atomvåbnet. (Danish) [Niels Bohr and the atomic weapon]. Vindrose, Copenhagen, Denmark, 1985. ISBN 87-7456-219-3. 246 pp. LCCN QC776 .C55 1985.

Carroll:2007:IMA

[CO07] Bradley W. Carroll and Dale A. Ostlie. An Introduction to Modern Astrophysics. Pearson Addison-Wesley, San Francisco, CA, USA, second edition, 2007. ISBN 0-8053-0402-9, 0-321-44284-9 (paperback). ???? pp. LCCN QB461 .C35 2007. URL http://

catdir.loc.gov/catdir/toc/ecip0613/2006015391.html;   http:
//www.gbv.de/dms/ilmenau/toc/512485305.PDF.

Cockcroft:1963:NHD

[Coc63] John Douglas Cockcroft. Niels Henrik David Bohr, 1885–1962. Biographical Memoirs of Fellows of the Royal Society, 9:37–53, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/10.
1098/rsbm.1963.0002.

Cohen:1942:BRBb

[Coh42] I. Bernard Cohen. Book review: International Encyclopedia of Unified Science by Otto Neurath; Rudolf Carnap; Charles W. Morris; Niels Bohr; John Dewey; Bertrand Russell; Leonard Bloomfield; Victor F. Lenzen; Ernest Nagel; J. H. Woodger. Isis, 33(6):721–723, June 1942. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330723.

Cohen:1988:MDE

[Coh88] Montague Cohen. My Dear Eve…: The letters of Ernest Rutherford to Arthur Eve, 1907–1908. Fontanus: from the collections of McGill University, 1:3–37, 1988. ISSN 0838-2026. URL http://fontanus.

mcgill.ca/article/download/1/1. See comments [dR92].

Cohen:1991:MDE

[Coh91] Montague Cohen. My Dear Eve…: The letters of Ernest Rutherford to Arthur Eve. Part II, 1909–1911. Fontanus: from the collections of McGill University, 4:69–108, 1991. ISSN 0838-2026. URL http://fontanus.

mcgill.ca/article/download/48/51. See comments [dR92].

Cohen:1992:MDE

[Coh92] Montague Cohen. My Dear Eve…: The letters of Ernest Rutherford to Arthur Eve. Part IV, 1915–1919. Fontanus: from the collections of McGill University, 5:123–159, 1992. URL http://fontanus.mcgill.

ca/article/view/64/72.

Compton:1956:AQP

[Com56] Arthur Holly Compton. Atomic quest, a personal narrative. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1956. xix + 370 pp. LCCN QC773.A1 C65.

Condon:1949:SBM

[Con49] E. U. Condon. Superconductivity and the Bohr magneton. Proceedings of the National Academy of Sciences of the United States of America, 35 (8):488–490, August 1949. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Condon:1962:YQP

[Con62] Edward U. Condon. 60 years of quantum physics. Physics Today, 15(10):37–49, October 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.org/ resource/1/PHTOAD/v15/i10. Delayed 1951 Presidential address at the 1500th regular meeting of the American Philosophical Society of Washington, 2 December 1962, at the Natural History Museum Auditorium of the Smithsonian Institution, on the 60th anniversary of Planck’s constant, h. Reprinted in [WP85, pages 310–318].

Coster:1922:RAV

[Cos22] Dirk Coster. Rontgenspectra en de Atoomtheorie van Bohr. (Dutch) [Xray spectra and Bohr’s atomic theory]. Ph.D dissertation, Universiteit Leiden, Leiden, The Netherlands, 1922.

Courant:1964:HAV

[Cou64] Richard Courant. Halvtreds års venskab. (Danish) [Fifty years of friendship]. In Bohr and Rozental [BR64], pages 292–296. LCCN QC16.B63 N5. English translation in [Roz67].

Crawford:2002:HSN

[Cra02] Elisabeth T. Crawford, editor. Historical Studies in the Nobel Archives: the Prizes in Science and Medicine, volume 31 of Uppsala studies in the history of science. Universal Academy Press, Tokyo, Japan, 2002. ISBN 4-946443-69-X. ISSN 0282-1036. vii + 161 pp. LCCN AS911.N9 .H57 2002.

Cropper:2001:GPL

[Cro01] William H. Cropper. Great physicists: the life and times of leading physicists from Galileo to Hawking. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-513748-5. xii + 500 pp. LCCN QC15 .C76 2001. URL http://www.loc.gov/

catdir/enhancements/fy0637/2001021611-d.html;   http://www.
loc.gov/catdir/enhancements/fy0724/2001021611-b.html.

Cohen:1979:SPL

[CS79a] R. S. (Robert Sonné) Cohen and John J. Stachel, editors. Selected Papers of Léon Rosenfeld, volume 21 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. xxxiv + 929 pp.

LCCN Q174 .B67 vol. 21 QC7. URL https://link.springer.com/
book/10.1007/978-94-009-9349-5.

Cohen:1979:ECB

[CS79b] Robert S. Cohen and John J. Stachel. The epistemological conflict between Einstein and Bohr (dedicated to Max Born on his 80th birthday) [1963c]. In Cohen and Stachel [CS79a], chapter 35, pages 517– 521. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (ebook). ISSN 0068-0346. LCCN Q174 .B67 vol. 21 QC7. URL https:/

/link.springer.com/chapter/10.1007/978-94-009-9349-5_35.

Cohen:1979:FCM

[CS79c] Robert S. Cohen and John J. Stachel. Field and charge measurements in quantum electrodynamics [with N. Bohr] [1950a]. In Cohen and Stachel [CS79a], chapter 27, pages 401–412. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174

.B67 vol. 21 QC7. URL https://link.springer.com/chapter/10.
1007/978-94-009-9349-5_27.

Cohen:1979:NBE

[CS79d] Robert S. Cohen and John J. Stachel. Niels Bohr: An essay dedicated to him on the occasion of his sixtieth birthday, October 7, 1945 [second edition corrected] 1961 [1945e]. In Cohen and Stachel [CS79a], chapter 22, pages 313–326. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol.

21 QC7. URL https://link.springer.com/chapter/10.1007/978-
94-009-9349-5_22.

Cohen:1979:NBC

[CS79e] Robert S. Cohen and John J. Stachel. Niels Bohr’s contribution to epistemology [1963h]. In Cohen and Stachel [CS79a], chapter 36, pages 522–535. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21 QC7. URL https:

//link.springer.com/chapter/10.1007/978-94-009-9349-5_36.

Cohen:1979:QEA

[CS79f] Robert S. Cohen and John J. Stachel. On quantum electrodynamics [among essays dedicated to Niels Bohr on the occasion of his 70th birthday] [1955b]. In Cohen and Stachel [CS79a], chapter 28, pages 413–441. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21 QC7. URL https:

//link.springer.com/chapter/10.1007/978-94-009-9349-5_28.

Cohen:1979:QME

[CS79g] Robert S. Cohen and John J. Stachel. On the question of the measurability of electromagnetic field quantities [with Niels Bohr] [1933b]. In Cohen and Stachel [CS79a], chapter 26, pages 357–400. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21 QC7. URL https://link.springer.

com/chapter/10.1007/978-94-009-9349-5_26.

Camilleri:2015:NBP

[CS15] Kristian Camilleri and Maximilian Schlosshauer. Niels Bohr as philosopher of experiment: Does decoherence theory challenge Bohr’s doctrine of classical concepts? Studies in History and Philosophy of Modern Physics, 49(??):73–83, February 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1355219815000064.

Crawford:1996:NTW

[CSW96] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. A Nobel tale of wartime injustice. Nature, 382(6590):393–395, August 1, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v382/n6590/
abs/382393a0.html.

Crawford:1997:KIP

[CSW97a] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. Die Kernspaltung und ihr Preis. Warum nur Otto Hahn den Nobelpreis erhielt, Otto Frisch, Lise Meitner und Fritz Straßmann dagegen nicht berücksichtigt werden. (German) [Fission and its price. Why only Otto Hahn received the Nobel Prize: Otto Frisch, Lise Meitner and Fritz Strassmann are not taken into consideration]. Kultur & Technik. Zeitschrift des Deutschen Museums München, 21(2):30–35, ???? 1997.

ISSN 0344-5690. URL http://www.deutsches-museum.de/fileadmin/
Content/data/Insel/Information/KT/heftarchiv/1997/21-2-30.

pdf. Translation from English by Dieter Beisel of [CSW96].

Crawford:1997:NTP

[CSW97b] Elisabeth Crawford, Ruth Lewin Sime, and Mark Walker. A Nobel tale of postwar injustice: Recently released Swedish documents reveal why Lise Meitner, codiscoverer of nuclear fission, did not receive the 1946 Physics Prize for her theoretical interpretation of the process. Physics Today, 50(9):26–32, September 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Conn:1965:ENA

[CT65] G. K. T. (George Keith Thurburn) Conn and H. D. (Henry Dicken) Turner. The Evolution of the Nuclear Atom. Iliffe Books, London, UK, 1965. 266 pp. LCCN QC173 .C613 1966.

Cuffaro:2024:RSP

[Cuf24] Michael E. Cuffaro. Review of Slobodan Perović’s From Data to Quanta — Niels Bohr’s Vision of Physics. Chicago, IL: The University of Chicago Press (2021), 280 pp., $45 (cloth). Philosophy of Science, 91 (2):525–529, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/review-of-slobodan-
perovics-from-data-to-quanta-niels-bohrs-vision-of-physics-
slobodan-perovi-c-from-data-to-quanta-niels-bohrs-vision-of-
physics-chicago-il-the-university-of-chicago-press-2021-280-
pp-45-cloth/F70FD92FEA6B1ECC0A3F80FFE82A8124.

Cushing:1994:BRB

[Cus94a] James T. Cushing. Book review: Niels Bohr: His Heritage and Legacy by Jan Faye. Philosophy of Science, 61(1):149–150, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http:

//www.jstor.org/stable/188298.

Cushing:1994:QMH

[Cus94b] James T. Cushing. Quantum mechanics: historical contingency and the Copenhagen hegemony. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1994. ISBN 0-226-13202-1, 0-226-13204-8 (paperback). xvi + 317 pp. LCCN QC173.98 .C87 1994. URL http://www.loc.gov/

catdir/description/uchi051/94008427.html;   http://www.loc.
gov/catdir/toc/uchi051/94008427.html.

Cohen:1985:PTF

[CW85] Robert S. Cohen and Marx W. Wartofsky, editors. A Portrait of Twenty-five Years: Boston Colloquium for the Philosophy of Science 1960–1985, volume 89 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1985. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN

0068-0346. LCCN ???? URL https://link.springer.com/book/10.
1007/978-94-009-5345-1.

Cavalcanti:2012:BNS

[CW12] Eric G. Cavalcanti and Howard M. Wiseman. Bell nonlocality, signal locality and unpredictability (or what Bohr could have told Einstein at Solvay had He known about Bell experiments). Foundations of Physics, 42(10):1329–1338, October 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-012-9669-1.

DAgostino:2000:HIT

[D’A00] Salvo D’Agostino. A History of the Ideas of Theoretical Physics: Essays on the Nineteenth and Twentieth Century Physics, volume 213 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6094-X, 94-010-9034-3. ISSN 0068-0346. xviii + 381 pp. LCCN

Q174 .B67 no. 213. URL https://link.springer.com/book/10.1007/
978-94-010-9034-6.

DAgostino:2001:CCN

[D’A01a] Salvo D’Agostino. Correspondence and complementarity in Niels Bohr’s papers: 1925–1927. In A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics [D’A01b], chapter 12, pages 253–271. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346. LCCN QC5.53; QC6. URL https://link.springer.

com/chapter/10.1007/978-94-010-9034-6_12.

DAgostino:2001:HIT

[D’A01b] Salvo D’Agostino. A history of the ideas of theoretical physics: essays on the nineteenth and twentieth century physics, volume 213 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0244-0, 94-010-9034-3 (e-book). ISSN 0068-0346. xviii + 381 pp. LCCN QC5.53; QC6.

Dahl:1941:DKV

[Dah43] Svend Dahl, editor. Danmarks Kultur ved Aar 1940. (Danish) [Denmark’s culture in the year 1940], volume 1. Det Danske Selskab, København, Danmark, 1941–1943. ???? pp. LCCN DL131 .D2. Eight volumes.

Dahl:2002:BHCa

[Dah02a] Jens Peder Dahl. The Bohr–Heisenberg Correspondence Principle viewed from phase space. Fortschritte der Physik = Progress of Physics, 50(5–7):630–635, May 2002. CODEN FPYKA6. ISSN 0015-8208 (print), 1521-3978 (electronic).

Dahl:2002:BHCb

[Dah02b] Jens Peder Dahl. The Bohr-Heisenberg Correspondence Principle viewed from phase space. In Papenfuss et al. [PLS02], pages 201–206. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.loc.gov/

catdir/bios/wiley046/2003268171.html;   http://www.loc.gov/
catdir/description/wiley0310/2003268171.html;   http://www.
loc.gov/catdir/toc/wiley031/2003268171.html.

Dahl:2014:SZG

[Dah14] Izabela A. Dahl. Schweden als Zufluchtsland 1933–1945. (German) [Sweden as a refuge 1933–1945]. In Skandinavien als Zuflucht für jüdische Intellektuelle 1933–1945. (German) [Scandinavia as a Refuge for Jewish Intellectuals 1933–1945] [DF14], pages 14–31. ISBN 3-86331-194-9. LCCN DS135.S32 S53 2014.

Dainian:1996:NBR

[Dai96] Fan Dainian. Niels Bohr and realism. In Cohen et al. [CHR96], pages 279–287. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346.

LCCN ???? URL https://link.springer.com/chapter/10.1007/
978-94-015-8638-2_19.

Dam:1985:NBA

[Dam85] Poul Dam. Niels Bohr (1885–1962): atomic theorist, inspirator, rallying point. Royal Danish Ministry of Foreign Affairs, Press and Cultural Relations Dept., Copenhagen, Denmark, 1985. ISBN 87-87646-53-6 (paperback). 102 + 1 pp. LCCN QC16.B63 D35 1985. kr35.00. Translation from Danish to English by Gitte and Norman Shine.

Daniel:1989:BER

[Dan89] Wojciech Daniel. Bohr, Einstein and realism. Dialectica: International Review of Philosophy of Knowledge, 43(3):249–261, September 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Darwin:1927:FMW

[Dar27] Charles G. Darwin. Free motion in the wave mechanics. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 117(766):258–293, December 1, 1927. ISSN 0950-1207 (print), 2053-9150 (electronic).

URL http://rspa.royalsocietypublishing.org/content/117/776/

258.full.pdf+html. According to [De 14, page 76, column 2], this is the paper “where […] Bohr’s idea of complementary aspects of atomic theory is published for the first time.”.

Darwin:1955:DAN

[Dar55] Charles G. Darwin. The discovery of atomic number. In Pauli et al. [PRW55], pages 1–11. LCCN QC71 .P3 1955.

Darrigol:1985:CCA

[Dar85] Olivier Darrigol. La complémentarité comme argument d’autorité (1927–1934). (French) [Complementarity as an argument of authority (1927–1934)]. Revue d’Histoire des Sciences, 38(3–4):309–323, 1985. CO-DEN RHSAAM. ISBN 2-13-039137-0. ISSN 0048-7996, 0151-4105, 1969-

6582 (electronic). URL http://www.persee.fr/web/revues/home/
prescript/article/rhs_0151-4105_1985_num_38_3_4009.

Darrigol:1987:AOB

[Dar87] Olivier Darrigol. Analyses d’ouvrages: Niels Bohr, 1885–1962; Leben und Werk eines Atomphysikers. ^{≪}Grosse Naturforscher ^{≫}, Bd. 47 par Ulrich Röseberg. Revue d’Histoire des Sciences, 40(1):139–140, January 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632769.

Darrigol:1988:QEA

[Dar88] Olivier Darrigol. The quantum electrodynamical analogy in early nuclear theory or the roots of Yukawa’s theory. Revue d’Histoire des Sciences, 41(3–4):225–297, juillet–décembre 1988. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23633104.

Darrigol:1990:AOBb

[Dar90] Olivier Darrigol. Analyses d’ouvrages: Atomi metafore paradossi: Niels Bohr e la costruzione di una nuova fisica par Sandro Petruccioli. Revue d’Histoire des Sciences, 43(4):490–492, October 1990. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http:

//www.jstor.org/stable/23632942.

Darrigol:1991:AOBb

[Dar91a] Olivier Darrigol. Analyses d’ouvrages: Niels Bohr’s philosophy of physics, 2nd ed. par Dugald Murdoch. Revue d’Histoire des Sciences, 44 (3–4):503, juillet–décembre 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632893.

Darrigol:1991:CCM

[Dar91b] Olivier Darrigol. Cohérence et complétude de la mécanique quantique : l’exemple de ^{≪}Bohr–Rosenfeld ^{≫}. Revue d’Histoire des Sciences, 44(2):137–179, avril–juin 1991. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633147.

Darrigol:1992:AOBd

[Dar92a] Olivier Darrigol. Analyses d’ouvrages: Niels Bohr’s times, in physics, philosophy, and polity par Abraham Pais. Revue d’Histoire des Sciences, 45(2–3):382–383, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23633016.

Darrigol:1992:AOBe

[Dar92b] Olivier Darrigol. Analyses d’ouvrages: Physique atomique et connaissance humaine, ^{≪}Folio/essais ^{≫}, 157 par Niels Bohr; Edmond Bauer; Roland Omnes; Catherine Chevalley. Revue d’Histoire des Sciences, 45 (2–3):383, avril–septembre 1992. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23633017.

Darrigol:1992:BA

[Dar92c] Olivier Darrigol. The Bohr atom (1913–1916). In From c-numbers to q-numbers: the Classical Analogy in the History of Quantum Theory [Dar92d], chapter V, pages 85–101. ISBN 0-520-07822-5. LCCN QC173.98 .D37 1992.

Darrigol:1992:NNC

[Dar92d] Olivier Darrigol. From c-numbers to q-numbers: the Classical Analogy in the History of Quantum Theory. California studies in the history of science. University of California Press, Berkeley, CA, USA, 1992. ISBN 0-520-07822-5. 388 pp. LCCN QC173.98 .D37 1992.

Darrigol:1997:CCB

[Dar97] Olivier Darrigol. Classical concepts in Bohr’s atomic theory (1913– 1925). Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 34(3):545–567 (1998), 1997. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic). Reprinted in [GGK02].

Dass:2015:SPQ

[Das15] N. D. Hari Dass. The superposition principle in quantum mechanics — did the rock enter the foundation surreptitiously? In Aaserud and Kragh [AK15], pages 435–449. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Davisson:1916:DHH

[Dav16] C. Davisson. The dispersion of hydrogen and helium on Bohr’s theory. Physical Review (2), 8(1):20–27, July 1916. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL https://journals.

aps.org/pr/abstract/10.1103/PhysRev.8.20.

Davis:1986:ITD

[Dav86] H. Davis. If they didn’t listen to Einstein or Bohr, they are not going to listen to you. Medicine and War, 2(3):207–208, ???? 1986. CODEN MEWAE4. ISSN 0748-8009 (print), 2378-9980 (electronic).

deBroglie:1924:TB

[dB24] Louis de Broglie. Sur un théorème de Bohr. (French) [On a theorem of Bohr]. Comptes rendus de l’Académie des sciences, Paris, 179(??): 676–??, ???? 1924.

deBroglie:1925:RTQa

[dB25] Louis de Broglie. Recherches sur la théorie des quanta. (French) [Research on quantum theory]. Doctoral thesis, La Sorbonne, Paris, France,

1925. ???? pp. URL https://www.annphys.org/articles/anphys/
abs/1925/03/anphys19251003p22/anphys19251003p22.html. Pre-

sented at the Sorbonne on 25 November 1924. German translation in [?]. English translation in [?].

deBroglie:1963:VOU

[dB63] Louis de Broglie. La vie et l’œuvre de Niels Bohr. (French) [The life and works of Niels Bohr]. Annals of Physics, 13(8):117–120, 1963. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035X (electronic).

deBeauregard:1986:AOBc

[dB86a] O. Costa de Beauregard. Analyses d’ouvrages: The creation of Quantum Mechanics and the Bohr–Pauli Dialogue, (^{≪}Studies in the History of Modern Science ^{≫}, 14) par John Hendry. Revue d’Histoire des Sciences, 39(4):367–368, October 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/stable/

23632444.

deBeauregard:1986:BDF

[dB86b] O. Costa de Beauregard. Bohr’s discussion of the fourth uncertainty relation revisited. Foundations of Physics, 16(9):937–939, September 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL http://link.springer.com/article/10.1007/BF00765340.

deBerg:2013:BRN

[dB13] K. C. de Berg. Book review: Niaz, M. & Marcano, C. (2012): Reconstruction of Wave–Particle Duality and Its Implications for General Chemistry Textbooks. Science & Education (Springer), 22(8):2031–2033, August 2013. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

deBroglie:1921:MAB

[dBdB21] Louis de Broglie and Maurice de Broglie. Sur le modèle d’atome de Bohr et les spectres corpusculaires. (French) [On the Bohr atomic model and particle spectra]. Comptes rendus de l’Académie des sciences, Paris, 172(??):746–??, ???? 1921.

Dainian:1996:CSH

[DC96] Fan Dainian and Robert S. Cohen. Chinese Studies in the History and Philosophy of Science and Technology, volume 179 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346. xiii + 483 pp. LCCN Q174 .B67

vol. 180; U390 .N19 1996. URL https://link.springer.com/book/
10.1007/978-94-015-8717-4.

DeBoer:1986:LQT

[DDU86] Jorrit De Boer, Erik Dal, and Ole Ulfbeck, editors. The lesson of quantum theory: Niels Bohr Centenary Symposium, October 3–7, 1985. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISBN 0-444-87012-1 (Elsevier). LCCN QC16.B63 L47 1986.

DeGregorio:2012:BWD

[De 12] Alberto De Gregorio. Bohr’s way to defining complementarity. arXiv.org, ??(??):1–25, December 28, 2012. URL https://arxiv.org/

abs/1212.6481.

DeGregorio:2014:BWD

[De 14] Alberto De Gregorio. Bohr’s way to defining complementarity. Studies in History and Philosophy of Modern Physics, 45(??):72–82, February 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1355219813000865.

Dear:2006:INH

[Dea06] Peter Dear. The intelligibility of nature: how science makes sense of the world. Science.culture. University of Chicago Press, Chicago, IL, USA and London, UK, 2006. ISBN 0-226-13948-4. xii + 242 pp. LCCN Q175.32.R45 D43 2006. URL http://www.loc.

gov/catdir/enhancements/fy0702/2005026032-b.html;   http://
www.loc.gov/catdir/enhancements/fy0702/2005026032-d.html;
http://www.loc.gov/catdir/toc/ecip0518/2005026032.html.

Dean:2013:NYT

[Dea13] Cornelia Dean, editor. The New York Times book of physics and astronomy: more than 100 years of covering the expanding universe. Sterling, New York, NY, USA, 2013. ISBN 1-4027-9320-0 (hardcover). xviii + 557 pp. LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Anton:1989:IQP

[Deg89] Degen, Peter Anton. Interpretations of quantum physics, the mystical and the paranormal: Einstein, Schrödinger, Bohr, Pauli and Jordan. Ph.D. dissertation, Drew University, Madison, NJ, USA, December 1989. 319 pp.

delRegato:1981:NB

[del81] J. A. del Regato. Niels Bohr. International Journal of Radiation Oncology, Biology, Physics, 7(4):509–530, April 1981. CODEN IOBPD3. ISSN 0360-3016 (print), 1879-355x (electronic).

Demortier:2015:QEQ

[Dem15] Guy Demortier. Qui est qui dans ^{≪}the letter to Roosevelt ^{≫}?. (French) [Who is who in “the letter to Roosevelt”?]. Revue des Questions Scientifiques, 186(1):125–152, ???? 2015. CODEN RQSCAN. ISSN 0035-

2160. URL https://www.unamur.be/sciences/philosoc/revueqs/
textes-en-ligne/rqs_186_1_demortier.

Denekamp:1999:ETM

[Den99] Juliana Denekamp. Entering the third millennium after a century of ionizing radiation in science and oncology. Acta Oncologica (Stockholm), 38(7):819–821, 1999. CODEN ACTOEL. ISSN 0284-186X (print), 1651-226X (electronic). See comments and historical correction, and brief response by Denekamp [Ben00].

Destouches:1948:QAT

[Des48] Jean-Louis Destouches. Quelques aspects théoriques de la notion de complémentarité I. la complémentarité de Bohr. (French) [Some theoretical aspects of the notion of complementarity. I. The complementarity of Bohr]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):351–382, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dahl:2014:SZJ

[DF14] Izabela A. Dahl and Jorunn Sem Fure. Skandinavien als Zuflucht für jüdische Intellektuelle 1933–1945. (German) [Scandinavia as a Refuge for Jewish Intellectuals 1933–1945]. Metropol-Verlag, Berlin, Germany, 2014. ISBN 3-86331-194-9. 326 pp. LCCN DS135.S32 S53 2014.

Donovan:2011:PVD

[DG11] Anne L. Donovan and Michael A. Gropper. Prediction is very difficult, especially about the future — Niels Bohr. Critical Care Medicine, 39(8): 2005–2007, August 2011. ISSN 1530-0293 (print), 1530-0293 (electronic).

Dirac:1924:NDP

[Dir24] P. A. M. Dirac. Note on the Doppler principle and Bohr’s frequency condition. Mathematical proceedings of the Cambridge Philosophical Society, 22(3):432–433, ???? 1924. CODEN PCPSA4. ISSN 0008-1981.

Dirac:1928:DFR

[Dir28] P. A. M. Dirac. Discussion following reports of Bohr and Heisenberg. In ????, editor, Electrons and photons (5th Conseil de physique de l’Institut International de Physique, Solvay, 24–29 October 1927, Bruxelles), pages 258–263. Gauthier-Villars et cie, Paris, France, 1928. LCCN ????

Dirac:1964:NBA

[Dir64] Paul A. M. Dirac. Niels Bohrs alsidighed. (Danish) [Niels Bohr’s versatility]. In Bohr and Rozental [BR64], pages 297–300. LCCN QC16.B63 N5. English translation in [Roz67].

Dirac:1967:VNB

[Dir67] P. A. M. Dirac. The versatility of Niels Bohr. In Rozental [Roz67], pages 306–309. LCCN QC16.B63 N53.

Duncan:2015:SEB

[DJ15] Anthony Duncan and Michel Janssen. The Stark effect in the Bohr– Sommerfeld theory and in Schrödinger’s wave mechanics. In Aaserud and Kragh [AK15], pages 217–271. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Djerassi:2007:BRW

[Dje07] Carl Djerassi. Book reviews: When acting speaks louder than words: Science on stage: From ‘Doctor Faustus’ to ‘Copenhagen’, by Kirsten Shepherd-Barr. Google’s PageRank and Beyond: The Science of Search Engine Rankings, by Amy N. Langville and Carl D. Meyer. Broken Genius The Rise and Fall of William Shockley, Creator of the Electronic Age, by Joel N. Shurkin. Physics Today, 60(2):63–64, 2007. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//link.aip.org/link/phtoad/v60/i2/p63/s1.

Durran:2008:DCB

[DNT08] Richard Durran, Andrew Neate, and Aubrey Truman. The divine clockwork: Bohr’s correspondence principle and Nelson’s stochastic mechanics for the atomic elliptic state. Journal of Mathematical Physics, 49(3):032102, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i3/p032102_s1.

Dolling:1995:RUP

[Dol95] Lisa M. Dolling. The role of understanding in the philosophical writings of Niels Bohr: a place for hermeneutics in the natural sciences. Ph.D. dissertation, City University of New York, New York, NY, USA, 1995. 242 pp.

Domondon:2006:BPB

[Dom06] Andrew T. Domondon. Bringing physics to bear on the phenomenon of life: the divergent positions of Bohr, Delbrück, and Schödinger. Studies in History and Philosophy of Science Part C: Studies in History and Philosophy of Biological and Biomedical Sciences, 37(3):433– 458, September 2006. CODEN ???? ISSN 1369-8486 (print), 1879-

2499 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848606000434.

Dorling:1979:BRN

[Dor79] Jon Dorling. Book review: Nineteenth and Twentieth Centuries: Niels Bohr Collected Works. Volume III: The Correspondence Principle (1918– 1923). Edited by J. Rud Nielsen. Amsterdam: North-Holland, 1976. Pp. xii + 702. $105.95. British Journal for the History of Science, 12(1):100– 102, March 1979. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4025674.

Dorries:2005:MFC

[Dör05] Matthias Dörries, editor. Michael Frayn’s Copenhagen in debate: historical essays and documents on the 1941 meeting between Niels Bohr and Werner Heisenberg, volume 20, 33 of Berkeley papers in history of science; Uppsala studies in history of science. Office for History of Science and Technology, University of California, Berkeley, Berkeley, CA, USA, 2005. ISBN 0-9672617-2-4 (paperback). viii + 195 pp. LCCN PR6056.R3 C6436 2005.

Dotson:2008:RBQ

[Dot08] Allen C. Dotson. Refocusing Bohr’s quantum postulate. Studies in History and Philosophy of Modern Physics, 39(3):610–619, September 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1355219808000208.

Dow:2009:HBU

[Dow09] Tsung-I Dow. Harmonious balance as the ultimate reality in artistic and philosophical interpretation of the “taiji diagram”. In Anna-Teresa Tymieniecka, editor, Existence, Historical Fabulation, Destiny, volume 99 of Analecta Husserliana, pages 247–257. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-9802-2 (e-book), 1-4020-9801-4 (hardcover). LCCN B3279.H94 E95 2009.

Dalitz:1997:SSP

[DP97] Richard Henry Dalitz and Sir Rudolf Peierls, editors. Selected scientific papers of Sir Rudolf Peierls: with commentary, volume 19 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). xxiii + 805 pp. LCCN QC21.2 .P42 1997. URL https:

//ui.adsabs.harvard.edu/#abs/1997sspr.book.....D;   https:
//ui.adsabs.harvard.edu/#abs/1997ssps.book.....P;   https:/
/ui.adsabs.harvard.edu/#abs/1997WSSP...19.....D;   https://
www.worldscientific.com/worldscibooks/10.1142/3128.

delRegato:1985:RP

[dR85] Juan A. del Regato. Radiological physicists. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-469-9. 188 pp. LCCN QC774.A2 D44 1985.

delRegato:1992:CMD

[dR92] Juan A. del Regato. Comments on “My dear Eve… ”. Letters of Ernest Rutherford to Arthur Stewart Eve [FONTANUS I, 3–37 (1988), continued in II, 111–138 (1989) and lastly in FONTANUS IV, 69–108 (1991)]. Medical Physics, 19(2):261–262, March 1992. CODEN MPHYA6. ISSN 0094-2405 (print), 1522-8541 (electronic). See [Coh88].

DeGregorio:2009:CEB

[DS09] Alberto De Gregorio and Fabio Sebastiani. La complementarità: la esposizione di Bohr a Como nel 1927, tra storiogra a e documenti di archivio. (Italian) [Complementarity: the exposure of Bohr in Como in 1927, between historiography and archival documents]. Quaderni di Storia della fisica, 15(??):3–33, ???? 2009. ISSN 1594-9974 (print), 1827-

6164 (electronic). URL http://www.sif.it/riviste/qsf/econtents/
2009/015/01/article/2.

Dale-Trotter:2018:BRQ

[DT18] Iain Dale-Trotter. Book review: The quantum heretics: What is Real: the Unfinished Quest for the Meaning of Quantum Physics by Adam Becker. Physics World, 31(12):??, December 2018. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL https:/

/physicsworld.com/a/the-quantum-heretics/.

Duffus:1946:ABV

[Duf46] R. L. Duffus. Atom bomb versus ‘human race’: The scientists who fashioned it remind us of its terrible threat: Review of One World or None. A Report to the Public on the Full Meaning of the Atomic Bomb. Edited by Dexer Masters and Kathlerine Way. Foreword by Niels Bohr. Introduction by Arthur H. Compton. x + 79 pp. New York: Whittlesey House. $1. New York Times, ??(??):BR1–BR2, March 17, 1946. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:/

/search.proquest.com/docview/107455627.

Durrani:2000:HSH

[Dur00] Matin Durrani. History of science: Heisenberg hoax puts author in a spin. Physics World, 13(7):7, July 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.iop.

org/2058-7058/13/i=7/a=7.

Eckert:2013:ASA

[Eck13a] Michael Eckert. Arnold Sommerfeld: Atomphysiker und Kulturbote, 1868–1951: eine Biografie. (German) [Arnold Sommerfeld: atomic physicist and cultural messenger 1868–1951: a biography], volume 29 of Abhandlungen und Berichte (Deutsches Museum (Germany)). Wallstein Verlag, Göttingen, Germany, 2013. ISBN 3-8353-1206-5 (hardcover). 604 pp. LCCN QC16.S76 E25 2013.

Eckert:2013:ASS

[Eck13b] Michael Eckert. Arnold Sommerfeld: science, life and turbulent times 1868—1951. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-7460-4 (paperback), 1-4614-7461-2 (e-book). xiv + 471 pp. LCCN QC16.S76 E2513 2013. Translation by Tom Artin of [Eck13a].

Eckert:2014:HSE

[Eck14] Michael Eckert. How Sommerfeld extended Bohr’s model of the atom (1913–1916). European Physical Journal H, 39(2):141–156, April 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/epjh/
e2013-40052-4.

Eckert:2015:EBS

[Eck15] Michael Eckert. Extending Bohr: Sommerfeld’s early atomic theory, 1913–1916. In Aaserud and Kragh [AK15], pages 161–174. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63 2015. URL http://

www.ijqf.org/wps/wp-content/uploads/2015/06/Bohr-Book.pdf.

Eltschka:2001:CBB

[EFM01] C. Eltschka, H. Friedrich, and M. J. Moritz. Comment on “Breakdown of Bohr’s correspondence principle”. Physical Review Letters, 86(12): 2693, March 19, 2001. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Ehrenfest:1923:ATQ

[Ehr23] Paul Ehrenfest. Adiabatische Transformationen in der Quantentheorie und ihre Behandlung durch Niels Bohr. (German) [Adiabatic transformations in quantum theory and their treatment by Niels Bohr]. Die Naturwissenschaften, 11(27):543–550, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Eisner:2000:RBP

[Eis00] Werner Eisner. Rezension: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Herausgegeben von Karl von Meyenn (Band IV, Teil I: 1950–1952; Teil II: 1953–1954). Berlin/Heidelberg/New York: Springer 1996 und 1999; 968 und 1100 Seiten. (German) [Wolfgang Pauli: Scientific Correspondence with Bohr, Einstein, Heisenberg, and others. Edited by Karl of Meyenn (Volume IV, part I: 1950–1952, Part II: 1953–1954). Berlin / Heidelberg / New York: Springer 1996 and 1999, 968 and 1100 pages]. Berichte zur Wissenschaftsgeschichte, 23(1):59–62, ???? 2000. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Elyashevich:1985:NBD

[El’85] M. A. El’yashevich. Niels Bohr’s development of the quantum theory of the atom and the correspondence principle (his 1912–1923 work in atomic physics and its significance). Soviet Physics. Uspekhi, 28(10):879–??, ???? 1985. CODEN SOPUAP. ISSN 0038-5670.

URL http://iopscience.iop.org/0038-5670/28/10/A03;   http:/
/stacks.iop.org/0038-5670/28/i=10/a=A03.

Engler:2003:SHP

[Eng03] F. O. Engler. Structure and heuristic: In praise of structural reallism in the case of Niels Bohr. Vienna Circle Institute Yearbook, 10(A):297–309, ???? 2003. ISSN 0929-6328 (print), 2215-1818 (electronic).

Enz:2002:NTB

[Enz02a] Charles P. (Charles Paul) Enz. No time to be brief: a scientific biography of Wolfgang Pauli. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-856479-1. viii + 573 pp. LCCN QC16.P37 E59 2002. URL http://www.loc.gov/

catdir/enhancements/fy0614/2002726902-d.html;   http://www.
loc.gov/catdir/enhancements/fy0614/2002726902-t.html.

Enz:2002:USN

[Enz02b] Charles P. (Charles Paul) Enz. Under the spell of Niels Bohr. In No time to be brief: a scientific biography of Wolfgang Pauli [Enz02a], pages 84–90. ISBN 0-19-856479-1. LCCN QC16.P37 E59 2002. URL http://

www.loc.gov/catdir/enhancements/fy0614/2002726902-d.html;
http://www.loc.gov/catdir/enhancements/fy0614/2002726902-t.
html.

Einstein:1935:CQM

[EPR35] Albert Einstein, Boris Podolsky, and Nathan Rosen. Can quantum mechanical description of physical reality be considered complete? Physical Review (2), 47(10):777–780, May 15, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v47/i10/p777_1.

Epstein:1924:SJT

[Eps24] P. S. Epstein. On the simultaneous jumping of two electrons in Bohr’s model. Proceedings of the National Academy of Sciences of the United States of America, 10(8):337–342, August 1924. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Erlichson:1972:BEP

[Erl72] Herman Erlichson. Bohr and the Einstein–Podolsky–Rosen paradox. American Journal of Physics, 40(4):634–636, 1972. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:/

/scitation.aip.org/content/aapt/journal/ajp/40/4/10.1119/1.
1988078.

Eaton:1974:SDN

[ES74] T. W. Eaton and R. Smith. A simple determination of the nuclear mass of deuterium. Physics Education, 9(7):485–486, 1974. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http:

//stacks.iop.org/0031-9120/9/i=7/a=011.

Evans:2007:QMC

[ET07] James Evans and A. S. (Alan S.) Thorndike, editors. Quantum mechanics at the crossroads: new perspectives from history, philosophy and physics. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-32663-4 (hardcover), 3-540-32665-0 (e-book). ISSN 1612-3018. x + 249 pp. LCCN QC174.12 .Q346 2007. URL http://www.loc.

gov/catdir/enhancements/fy0824/2006934045-b.html;   http://
www.loc.gov/catdir/enhancements/fy0824/2006934045-d.html;
http://www.loc.gov/catdir/toc/fy0708/2006934045.html.

Evans:1923:BLT

[Eva23] G. C. Evans. A Bohr–Langmuir transformation. Proceedings of the National Academy of Sciences of the United States of America, 9(7): 230–236, July 1923. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Evans:1996:EHR

[Eva96] C. H. Evans, editor. Episodes from the History of the Rare Earth Elements, volume 15 of Chemists and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-4101-5. xxv + 240 pp. LCCN QD172.R2

E65 1996. URL http://www.loc.gov/catdir/enhancements/fy1006/
96210275-d.html;   http://www.loc.gov/catdir/enhancements/
fy1006/96210275-t.html.

Everett:1957:RSF

[Eve57] Hugh Everett III. “relative state” formulation of quantum mechanics. Reviews of Modern Physics, 29(3):454–462, July 1957. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.29.454;
 http://rmp.aps.org/abstract/RMP/v29/i3/p454_1. See assessment

[Whe57].

Enz:1994:WPW

[EvM94] Charles P. (Charles Paul) Enz and K. von Meyenn. Wolfgang Pauli: Writings on physics and philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). vi + 289 pp. LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/fy0812/
94015098-d.htm.

Fafarman:1958:RBO

[Faf58] Alfred Fafarman. Radiative Bohr Orbit Transitions of Mesonic Atoms Formed by Capture of Cosmic Ray Negative Muons in Lead, Carbon, and Oxygen. Ph.D dissertation, New York University, New York, NY, USA, March 1958. 134 pp.

Falkenburg:1998:BPU

[Fal98] Brigitte Falkenburg. Bohr’s principles of unifying quantum disunities. Philosophia Naturalis, 35(1):95–120, 1998. CODEN ???? ISSN 0031-8027. Models, theories and disunity in physics (Berlin, 1996).

Fara:2001:GPI

[Far01] Patricia Fara. Group portraits IV — the Seventh Solvay Conference [Bruxelles, 22–29 October 1933]. Endeavour, 25(4):137–138, December 2001. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Farmelo:2013:BRB

[Far13a] Graham Farmelo. Book review: Love, Literature and the Quantum Atom: Niels Bohr’s 1913 Trilogy Revisited, by Finn Aaserud and John L. Heilbron. Times Higher Education, ??(??):??, August 1,

2013. URL http://www.timeshighereducation.co.uk/books/love-
literature-and-the-quantum-atom-niels-bohrs-1913-trilogy-revisited-
by-finn-aaserud-and-john-l-heilbron/2006035.article.

Farmelo:2013:CBHa

[Far13b] Graham Farmelo. Churchill’s bomb: a hidden history of science, war and politics. Faber and Faber, London, UK, 2013. ISBN 0-571-24978-7 (hardcover). x + 554 + 12 pp. LCCN ????

Farmelo:2013:CBHb

[Far13c] Graham Farmelo. Churchill’s bomb: how the United States overtook Britain in the first nuclear arms race. Basic Books, New York, NY, USA, 2013. ISBN 0-465-02195-6 (hardcover), 0-465-06989-4 (e-book). vi + 554 pp. LCCN UA647 .F28 2013.

Farmelo:2013:NSBa

[Far13d] Graham Farmelo. November 1938 to September 1939 Bohr thinks the Bomb is ‘inconceivable’. In Churchill’s bomb: a hidden history of science, war and politics [Far13b], page ?? ISBN 0-571-24978-7 (hardcover). LCCN ????

Farmelo:2013:NSBb

[Far13e] Graham Farmelo. November 1938 to September 1939 Bohr thinks the Bomb is ‘inconceivable’. In Churchill’s bomb: how the United States overtook Britain in the first nuclear arms race [Far13c], page ?? ISBN 0-465-02195-6 (hardcover), 0-465-06989-4 (e-book). LCCN UA647 .F28 2013.

Farmelo:2013:SMBa

[Far13f] Graham Farmelo. September 1943 to May 1944 Bohr takes a political initiative. In Churchill’s bomb: a hidden history of science, war and politics [Far13b], page ?? ISBN 0-571-24978-7 (hardcover). LCCN ????

Farmelo:2013:SMBb

[Far13g] Graham Farmelo. September 1943 to May 1944 Bohr takes a political initiative. In Churchill’s bomb: how the United States overtook Britain in the first nuclear arms race [Far13c], page ?? ISBN 0-465-02195-6 (hardcover), 0-465-06989-4 (e-book). LCCN UA647 .F28 2013.

Favrholdt:1991:RBH

[Fav91] David Favrholdt. Remarks on the Bohr–Høffding relationship. Studies in History and Philosophy of Science Part A, 22(3):399–414, September 1991. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/003936819190002A.

Favrholdt:1992:NBP

[Fav92] David Favrholdt. Niels Bohr’s philosophical background, volume 63 of Historisk-filosofiske meddelelser. Royal Danish Academy of Sciences and Letters, Copenhagen, Denmark, 1992. ISBN 87-7304-228-5. 147 pp. LCCN QC16.B63 F3 1992; AS281 .C79h; AS281 .D201.

Favrholdt:1994:NBR

[Fav94] D. Favrholdt. Niels Bohr and realism. In Faye and Folse [FF94], pages 77–96. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (ebook). ISSN 0068-0346. LCCN Q174 .B67 vol. 153 1994; QC16

1994. URL http://www.loc.gov/catdir/enhancements/fy0823/
93024825-d.html;   http://www.loc.gov/catdir/enhancements/
fy0823/93024825-t.html;   https://link.springer.com/book/10.
1007/978-94-015-8106-6.

Favrholdt:1999:CBP

[Fav99] David Favrholdt. Complementarity Beyond Physics: 1928–1962, volume 10 of Collected works / Niels Bohr. Elsevier, Amsterdam, The Netherlands, 1999. ISBN 0-08-087108-9 (e-book), 0-7204-1800-3 (series), 0-444-89972-3 (v. 10). xlix + 613 pp. LCCN QC174.17.C63 B64

1999. URL http://store.elsevier.com/Complementarity-Beyond-
Physics-1928-1962/isbn-9780080871080/.

Favrholdt:2005:SNB

[Fav05] David Favrholdt. Spaltningen: Niels Bohr og Werner Heisenberg i videnskab og politik. (Danish) [Fission: Niels Bohr and Werner Heisenberg in science and politics]. Lindhardt og Ringhof, Copenhagen, Denmark, 2005. ISBN 87-595-2450-2. 282 pp. LCCN QC16.B63 F387 2005.

Favrholdt:2009:FNB

[Fav09] David Favrholdt. Filosoffen Niels Bohr. (Danish) [The philosopher Niels Bohr]. Informations Forlag, København, Danmark, 2009. ISBN 87-7514-216-3. 459 pp. LCCN QC16.B63 F378 2009. DKR 399,00.

Faye:1988:BHR

[Fay88] Jan Faye. The Bohr–Høffding relationship reconsidered. Studies in History and Philosophy of Science Part A, 19(3):321–346, September 1988. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0039368188900039.

Faye:1990:BRB

[Fay90] Jan Faye. Book review: Niels Bohr’s Philosophy of Physics by Dugald Murdoch. Isis, 81(2):378–379, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/233778.

Faye:1991:NBH

[Fay91] Jan Faye. Niels Bohr, His Heritage and Legacy, an Anti-realist View of Quantum Mechanics, volume 6 of Science and philosophy. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1294-5. xxii + 263 pp. LCCN QC173.98 .F38 1991.

URL http://elib.uni-stuttgart.de/opus/volltexte/2010/5370/.

Faye:2007:BRG

[Fay07] Jan Faye. Book review: Gerald Holton: Victory and Vexation in Science: Einstein, Bohr, Heisenberg, and Others. Isis, 98(3):660–661, September 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/524275.

Frayn:2000:CSI

[FB00] Michael Frayn and David Burke. Celia’s secret: an investigation. Faber, London, UK, 2000. ISBN 0-571-20530-5. viii + 110 pp. LCCN PR6056.R3 C45 2000.

Frayn:2001:CPI

[FB01] Michael Frayn and David Burke. The Copenhagen papers: an intrigue. Metropolitan Books, New York, NY, USA, 2001. ISBN 0-312-42124-9, 0-8050-6752-3. 129 pp. LCCN PR6056.R3 C45 2001. URL http://www.

loc.gov/catdir/enhancements/fy0667/00051965-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0667/00051965-d.html.

Fermi:1968:III

[Fer68] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA and London, UK, 1968. xi + 440 pp. LCCN E184.A1 F47.

Fermi:1971:III

[Fer71] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA and London, UK, second edition, 1971. ISBN 0-226-24376-1, 0-226-24378-8 (paperback). xi + 431 pp. LCCN E184.A1 F47 1971.

Ferreira:2007:BRD

[Fer07] Pedro G. Ferreira. Book review: David Lindley, Uncertainty: Einstein, Heisenberg, Bohr, and the Struggle for The Soul of Science. Physics Today, 60(8):57–??, August 2007. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/link/

phtoad/v60/i8/p57/s1.

Feynman:1933:CPM

[Fey33] Richard P. Feynman. The calculus for the practical man. High-school notebook., 1933. Copy in CalTech archive. Original in Niels Bohr Library, American Institute of Physics.

Froman:1978:PIQ

[FF78a] N. Froman and P. O. Froman. On phase-integral quantization conditions for bound states in one-dimensional smooth single-well potentials. Journal of Mathematical Physics, 19(9):1830–1837, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Froman:1978:AGQ

[FF78b] N. Froman and P. O. Froman. On the application of the generalized quantal Bohr–Sommerfeld quantization condition to single-well potentials with very steep walls. Journal of Mathematical Physics, 19(9):1823– 1829, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Faye:1994:NBC

[FF94] Jan Faye and Henry J. Folse, editors. Niels Bohr and Contemporary Philosophy, volume 153 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. xxvii + 377 pp. LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL http://www.

loc.gov/catdir/enhancements/fy0823/93024825-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0823/93024825-t.html;
https://link.springer.com/book/10.1007/978-94-015-8106-6.

Faye:2017:NBP

[FF17a] Jan Faye and Henry J. Folse, editors. Niels Bohr and the Philosophy of Physics: Twenty-first-century Perspectives. Bloomsbury Academic, London, UK, 2017. ISBN 1-350-03511-4 (hardcover), 1-350-03512-2 (ebook), 1-350-03513-0. vi + 384 pp. LCCN QC16.B63 N532 2017.

Frederick-Frost:2017:LMH

[FF17b] K. M. Frederick-Frost. For the love of a mother — Henry Moseley’s rare earth research. Historical Studies in the Natural Sciences, 47(4):529–567, September 2017. ISSN 1939-1811 (print), 1939-182X (electronic).

Farouki:2021:EBD

[FG21] Nayla Farouki and Philippe Grangier. The Einstein–Bohr debate: Finding a common ground of understanding? Foundations of Science, 26 (1):97–101, March 2021. CODEN FOSCFI. ISSN 1233-1821 (print),

1572-8471 (electronic). URL https://link.springer.com/article/
10.1007/s10699-020-09716-7.

Figurovskij:1960:SBG

[FH60] Nikolaj Aleksandrovic Figurovskij and Gerhard Harig. Sowjetische Beiträge zur Geschichte der Naturwissenschaft. (German) [Soviet contributions to the history of natural science]. Dt. Verl. der Wiss., Berlin, Germany, 1960. viii + 242 pp. LCCN Q125 1960. DM-Ost 17.50.

Fine:1986:SGE

[Fin86] Arthur Fine. The shaky game: Einstein, realism, and the quantum theory. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1986. ISBN 0-226-24946-8. xi + 186 pp. LCCN QC6 .F54 1986.

Fischer:2010:HQE

[Fis10a] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2010:NB

[Fis10b] Ernst Peter Fischer. Niels Bohr (1885–1962). In Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger] [Fis10a], pages 106–132. ISBN 3-7766-2643-7. LCCN ????

Fischer:2012:HQE

[Fis12a] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Fischer:2012:NB

[Fis12b] Ernst Peter Fischer. Niels Bohr (1885–1962). In Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger] [Fis12a], pages 106–132. ISBN 3-596-19406-7. LCCN ????

Faye:2021:BBQ

[FJ21a] Jan Faye and Rasmus Jaksland. Barad, Bohr, and quantum mechanics. Synthese, 199(3–4):8231–8255, December 2021. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL https://link.

springer.com/article/10.1007/s11229-021-03160-1.

Faye:2021:WBW

[FJ21b] Jan Faye and Rasmus Jaksland. What Bohr wanted Carnap to learn from quantum mechanics. Studies in History and Philosophy of Science Part A, 88(??):110–119, August 2021. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0039368121000741.

French:1985:NBC

[FK85] A. P. (Anthony Philip) French and P. J. Kennedy, editors. Niels Bohr: a centenary volume. Harvard University Press, Cambridge, MA, USA, 1985. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). xiv + 403 pp. LCCN QC16.B63 N49 1985. US$27.50.

Freire:2010:DMM

[FL10] Olival Freire and Christoph Lehner. ‘Dialectical materialism and modern physics’, an unpublished text by Max Born. Notes and Records of the Royal Society of London, 64(2):155–162, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Featonby:2012:WHN

[FM12] David Featonby and Rick Marshall. What happens next?: Mirrors produce a real image. Insights and conundrums: Bohr’s theory explores the mighty atom. Physics Education, 47(2):249, 2012. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http:

//iopscience.iop.org/0031-9120/47/2/M03.

Feshbach:1988:NBP

[FMO88] Herman Feshbach, Tetsuo Matsui, and Alexandra Oleson, editors. Niels Bohr, physics and the world: proceedings of the Niels Bohr Centennial Symposium, Boston, MA, USA, November 12–14, 1985, American Academy of Arts and Sciences, Cambridge, Massachusetts. Harwood Academic Publishers, Chur, Switzerland and New York, NY, USA, 1988. ISBN 3-7186-0484-1. LCCN QC16.B63 N54 1985; QC16.B63 N67 1985.

Fock:1957:IQM

[Foc57] V. A. Fock. On the interpretation of quantum mechanics. Československy çasopis pro fyziku, 7(6):643–656, November 1957. CODEN CKCFAH. ISSN 0009-0700 (print), 1804-8536 (electronic). Russian original in Uspekhi Fisichkich Nauk 62, 461–?? (August 1957).

Folse:1985:PNB

[Fol85] Henry J. Folse. The philosophy of Niels Bohr: the framework of complementarity. North-Holland personal library. North-Holland Publishing Co., Amsterdam, The Netherlands, 1985. ISBN 0-444-86914-X, 0-444-86938-7 (paperback). x + 281 pp. LCCN QC174.17.C3 F65 1985.

Folse:1994:BFC

[Fol94] Henry J. Folse. Bohr’s framework of complementarity and the realism debate. In Faye and Folse [FF94], chapter 6, pages 119–139. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https://link.

springer.com/chapter/10.1007/978-94-015-8106-6_6.

Folse:1995:NBC

[Fol95] Henry J. Folse. Niels Bohr and the construction of a new philosophy. Studies in History and Philosophy of Science Part B: Studies In History and Philosophy of Modern Physics, 26(1):107–116, April 1995. CODEN ???? ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/1355219895967438.

Folse:1996:BED

[Fol96] Henry J. Folse. The Bohr–Einstein debate and the philosophers’ debate over realism versus anti-realism. In Cohen et al. [CHR96], pages 289–298. ISBN 90-481-4493-0, 94-015-8638-1 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/978-94-
015-8638-2_20.

Foppl:1915:SBA

[Föp15] Ludwig Föppl. Über die Stabilität des Bohrschen Atommodelles. (German) [On the stability of Bohr atomic models]. Physikalische Zeitschrift, 15(15):707–712, July 15, 1915. CODEN PHZTAO.

ISSN 0369-982X. URL http://babel.hathitrust.org/cgi/pt?id=
uc1.c2831834%3Bview=1up%3Bseq=753.

Forman:1991:BRB

[For91] Paul Forman. Book review: Redirecting Science: Niels Bohr, Philanthropy and the Rise of Nuclear Physics, by Finn Aaserud. Physics Today, 44(11):??, November 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Forman:2009:BRD

[For09] Paul Forman. Book review: David Lindley: Uncertainty: Einstein, Heisenberg, Bohr, and the Struggle for the Soul of Science. Isis, 100(1):180–181, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

10.1086/599679.

Fox:1980:HIS

[Fox80] Timothy R. Fox. Hydrogen Ion-Solid Interactions Near the Bohr Velocity. Ph.D dissertation, The University of Chicago, Chicago, IL, USA, 1980.

Frisch:1940:MPR

[FP40a] Otto Robert Frisch and Rudolf Peierls. Memorandum on the properties of a radioactive “super-bomb”. Technical report, University of Birmingham, Birmingham, UK, March 1940. URL http:

//en.wikipedia.org/wiki/Frisch%E2%80%93Peierls_memorandum;
 http://www.atomicarchive.com/Bios/Frisch.shtml;   http://
www.atomicarchive.com/Bios/Peierls.shtml;   http://www.atomicarchive.
com/Docs/Begin/FrischPeierls.shtml;   http://www.atomicarchive.
com/Docs/Begin/FrischPeierls2.shtml.

Frisch:1940:PRS

[FP40b] Otto Robert Frisch and Rudolf Peierls. The properties of a radioactive “super-bomb”. Memorandum, University of Birmingham, Birming-

ham, UK, March 1940. URL http://en.wikipedia.org/wiki/Frisch%
E2%80%93Peierls_memorandum;   http://www.atomicarchive.com/
Bios/Frisch.shtml;   http://www.atomicarchive.com/Bios/Peierls.
shtml;   http://www.atomicarchive.com/Docs/Begin/FrischPeierls.
shtml;   http://www.atomicarchive.com/Docs/Begin/FrischPeierls2.

shtml. Reprinted in [?].

Flynn:1998:BRM

[FR98] Michael Flynn and Linda Rothstein. Bulletin: The real mystery science theater. Bulletin of the Atomic Scientists, 54(6):9–10, November/ December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fernandez:2013:PTY

[FR13a] Bernard Fernandez and Georges Ripka. A proliferation of theories: Yukawa, Breit and Wigner, Bohr. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13b], page 331. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:UMA

[FR13b] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). xviii + 522 pp. LCCN QC773 .F47 2013.

Frank:1955:HCR

[Fra55] V. Frank. Hall Coefficient and resistivity of alpha- and beta-brass from 20–600^{∘} C: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(4) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 12 pp. LCCN AS281.

Franck:1963:NBP

[Fra63] James Franck. Niels Bohrs Persönlichkeit. (German) [Niels Bohr’s personality]. Die Naturwissenschaften, 50(9):341–343, May 1, 1963. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Franck:1985:PM

[Fra85] James Franck. A personal memoir. In French and Kennedy [FK85], pages 16–18. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Frayn:1998:C

[Fra98] Michael Frayn. Copenhagen. Methuen drama. Methuen Drama, London, UK, 1998. ISBN 0-413-72490-5. 116 pp. LCCN PR6056.R3 C64 1998.

Frayn:2000:C

[Fra00] Michael Frayn. Copenhagen. Anchor Books, New York, NY, USA, 2000. ISBN 0-385-72079-3 (paperback). 132 pp. LCCN PR6056.R3 C64

2000. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/description/random044/00055814.
html;   http://www.pbs.org/hollywoodpresents/copenhagen/.

Franklin:2004:TRN

[Fra04] Allan Franklin. Are there really neutrinos?: an evidential history. Perseus Publishers, Cambridge, MA, USA, 2004. ISBN 0-7382-0265-7, 0-8133-4128-0 (paperback). ix + 371 pp. LCCN QC793.5.N42

F73 2004. URL http://www.loc.gov/catdir/enhancements/fy0837/
2008530710-d.html;   http://www.loc.gov/catdir/toc/fy0805/
2008530710.html.

French:1985:SCR

[Fre85] A. P. French. Some closing reflections. In French and Kennedy [FK85], pages 351–353. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

French:1986:NBH

[Fre86] A. P. French. Niels Bohr at 100: his life and work. Physics Education, 21(4):220–226, July 1986. CODEN PHEDA7. ISSN 0031-9120 (print),

1361-6552 (electronic). URL http://stacks.iop.org/0031-9120/21/
i=4/a=005.

French:1989:BRB

[Fre89] Anthony P. French. Book review: Harmony and Unity: The Life of Niels Bohr, by Nils Blædel. Physics Today, 42(9):??, September 1989.

CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [Blæ88].

Frenkel:1990:BRB

[Fre90] Victor Ya. Frenkel. Book review: Niels Bohr: Physics and the World, by Herman Feshbach, Tetsuo Matsui, and Alexandra Oleson. Physics Today, 43(6):??, June 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [FMO88].

Frenkel:1996:YIF

[Fre96] Viktor Iakovlevich Frenkel. Yakov Ilich Frenkel: his work, life, and letters. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 3-7643-2741-3 (Basel), 0-8176-2741-3 (Boston). viii + 323 pp. LCCN QC16.F715 A3 1996.

FreireJunior:2015:QDR

[Fre15] Olival Freire Junior. The Quantum Dissidents: Rebuilding the Foundations of Quantum Mechanics (1950–1990). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-662-44661-8, 3-662-44662-6 (e-book). xvi + 356 pp. LCCN QC173.98 .F74 2015.

Frisch:1939:PED

[Fri39] Otto Robert Frisch. Physical evidence for the division of heavy nuclei under neutron bombardment. Nature, 143(3616):276, February 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.chemteam.info/Chem-History/Frisch-
Fission-1939.html;   http://www.nature.com/nature/journal/
v143/n3616/pdf/143276a0.pdf.

Frisch:1954:AEH

[Fri54] Professor O. R. Frisch, O.B.E., F.R.S. Atomic energy — how it all began. British Journal of Applied Physics, 5(3):81–84, March 1954. CODEN BJAPAJ. ISSN 0508-3443 (print), 2057-7656 (elec-

tronic). URL http://iopscience.iop.org/0508-3443/5/3/301;
http://stacks.iop.org/0508-3443/5/i=3/a=301. A lecture delivered

in London to the Education Group of The Institute of Physics on 20 October, 1953.

Frisch:1964:ISS

[Fri64] Otto Robert Frisch. Interessen samler sig omkring atomkernen. (Danish) [Interest collects around the atomic nucleus]. In Bohr and Rozental [BR64], pages 132–144. LCCN QC16.B63 N5. English translation in [Roz67].

Frisch:1967:BNB

[Fri67] O. R. Frisch. Books: Niels Bohr, his life and work, edited by S. Rozental. Niels Bohr, by Ruth Moore. Scientific American, 216(6):145–153, June 1967. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v216/n6/pdf/scientificamerican0667-145.pdf.

Fischer:2006:EPF

[FS06] Ernst Peter Fischer and Klaus Sander. Ernst Peter Fischer erzählt, Paarläufe der Wissenschaft. (German) [Ernst Peter Fischer tells the pair runs of science]. Suppoš, Köln, Germany, 2006. ISBN 3-932513-68-

1. LCCN ???? URL http://deposit.ddb.de/cgi-bin/dokserv?id=
2758637\%26prov=M\%26dok\_var=1\%26dok\_ext=htm. 4 CDs.

Fano:2019:BRN

[FT19] Vincenzo Fano and Gino Tarozzi. Book review: Niels Bohr and the Philosophy of Physics: Twenty-First Century Perspectives — By J. Faye and H. J. Folse. Dialectica: International Review of Philosophy of Knowledge, 73(1-2):267–273, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fuchs:1985:RCN

[Fuc85] Klaus Fuchs. Reflections concerning Niels Bohr’s refutation of the Einstein–Podolsky–Rosen-Paradox. Annalen der Physik (1900), 497(1): 73–81, ???? 1985. ISSN 1521-3889. URL http://onlinelibrary.

wiley.com/doi/10.1002/andp.19854970112/abstract.

Fierz:1960:TPT

[FW60] Markus Fierz and Victor Frederick Weisskopf. Theoretical physics in the Twentieth Century: a memorial volume to Wolfgang Pauli. Interscience Publishers, New York, NY, USA, 1960. x + 328 pp. LCCN QC3 .F52. See [?] for comments on some inaccuracies in this book, and on the history of the discovery of the spin of the electron and the nucleus.

Frisch:1967:DFH

[FW67] Otto Robert Frisch and John A. Wheeler. The discovery of fission: How it all began and mechanism of fission. Physics Today, 20(11):43– 52, November 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://www.physicstoday.org/resource/1/

phtoad/v20/i11/p43_s1. Reprinted in [WP85, pages 272–281].

Frisch:1985:DF

[FW85] Otto R. Frisch and John A. Wheeler. The discovery of fission. In Weart and Phillips [WP85], pages 272–281. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [FW67].

Graetzer:1971:DNF

[GA71] Hans G. Graetzer and David L. Anderson. The discovery of nuclear fission: a documentary history, volume 20 of Van Nostrand Reinhold momentum books. Van Nostrand Reinhold, New York, NY, USA, 1971. viii + 120 pp. LCCN QC790 .G68.

Gamow:1929:DSA

[Gam29] George Gamow. Discussion on the structure of atomic nuclei. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 123(792):386–387, April 6, 1929. ISSN 0950-1207 (print), 2053-9150 (electronic). URL http://

www.jstor.org/stable/pdfplus/95202.pdf.

Gamow:1939:MTW

[Gam39] George Gamow. Mr. Tompkins in Wonderland: or, Stories of c, G, and h. Cambridge University Press, Cambridge, UK, 1939. x + 91 pp. LCCN ???? Illustrated by John Hookham.

Gamow:1942:MTD

[Gam42] George Gamow. Mr. Tompkins i Drømmeland. (Danish) [Mr. Tompkins in Wonderland]. Gyldendalske Boghandel Nordisk Forlag, København, Danmark, 1942. 95 pp. Forord af Niels Bohr.

Gamow:1960:JNB

[Gam60] George Gamow. Der junge Niels Bohr Zum 75. Geburtstag, nach Erinnerungen. (German) [The young Niels Bohr: 75th birthday memoirs]. Physikalische Blätter, 16(10):525–527, October 1960. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19600161008/abstract.

Gamow:1961:BPa

[Gam61a] George Gamow. Biography of Physics, volume TB567 of Harper torchbooks. The Cloister library. Harper, New York, NY, USA, 1961. 338 pp. LCCN QC7 .G263 1964.

Gamow:1961:BP

[Gam61b] George Gamow. Biography of physics, volume TB567 of Harper torchbooks. The Cloister library. Harper, New York, 1961. 338 pp. LCCN QC7 .G263 1964.

Gamow:1962:BPb

[Gam62a] George Gamow. Biography of Physics. Harper torchbooks. Science library. Hutchinson Science Library, London, UK, 1962. x + 338 pp. LCCN ????

Gamow:1962:BP

[Gam62b] George Gamow. Biography of physics. Harper torchbooks. Science library. Hutchinson, ????, 1962. 339 pp. LCCN Hist.

Gamow:1963:NBM

[Gam63] George Gamow. Niels Bohr, the man who explained the atom. Science Digest, ??(??):??, May 1963. CODEN ???? ISSN ????

Gamow:1966:RBB

[Gam66] George Gamow. Review of Niels Bohr: The Man. His Science, and the World They Changed, by Ruth Moore. New York Times Book Review, October 23, 1966. CODEN ???? ISSN 0028-7806.

Gamow:1985:TYS

[Gam85] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory. Dover, New York, NY, USA, 1985. ISBN 0-486-24895-X (paperback). xiv + 224 + 9 pp. LCCN QC174.12 .G35 1985.

US$4.95. URL http://www.loc.gov/catdir/description/dover032/
85006797.html.

Gamow:1988:GPG

[Gam88] George Gamow. The great physicists from Galileo to Einstein. Dover, New York, NY, USA, 1988. ISBN 0-486-25767-3. vi + 338 pp. LCCN QC7 .G27 1988; QC7 .G14b 1988.

Ganesh:2017:SAH

[Gan17] A. S. Ganesh. From the structure of the atom to hafnium. The Hindu, ??(??):??, December 11, 2017. URL http://www.thehindu.com/

children/from-the-structure-of-the-atom-to-hafnium/article21389576.
ece.

Gao:2015:HDE

[Gao15] Shan Gao. How do electrons move in atoms? From the Bohr model to quantum mechanics. In Aaserud and Kragh [AK15], pages 450– 464. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/2015/
06/Bohr-Book.pdf.

Garrett:1963:BMA

[Gar63a] Alfred Benjamin Garrett. The Bohr model of the atom: Niels Bohr. In The flash of genius [Gar63c], pages 116–121. LCCN Q125 .G38.

Garrett:1963:DFP

[Gar63b] Alfred Benjamin Garrett. The discovery of the fission process: Lise Meitner and O. R. Frisch. In The flash of genius [Gar63c], pages 227– 232. LCCN Q125 .G38.

Garrett:1963:FG

[Gar63c] Alfred Benjamin Garrett. The flash of genius. Van Nostrand, Princeton, NJ, USA, 1963. x + 249 pp. LCCN Q125 .G38.

Gardner:1972:TDL

[Gar72] Michael R. Gardner. Two deviant logics for quantum theory: Bohr and Reichenbach. British Journal for the Philosophy of Science, 23(2):89– 109, May 1972. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537

(electronic). URL http://bjps.oxfordjournals.org/content/23/2/
89.full.pdf+html;   http://www.jstor.org/stable/686433.

Gauderis:2014:ENP

[Gau14] Tjerk Gauderis. To envision a new particle or change an existing law? Hypothesis formation and anomaly resolution for the curious case of the β decay spectrum. Studies in History and Philosophy of Modern Physics, 45(??):27–45, February 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1355219813000920.

Gearhart:2014:FHE

[Gea14] Clayton A. Gearhart. The Franck–Hertz experiments, 1911–1914 experimentalists in search of a theory. Physics in Perspective (PIP), 16(3): 293–343, September 2014. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-014-0139-3.

Galison:2002:QMS

[GGK02] Peter Galison, Michael Gordin, and David Kaiser, editors. Quantum Mechanics: Science and Society. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2002. ISBN 1-136-70972-X. 433 (est.) pp. LCCN ????

Ghirardi:2005:SLG

[Ghi05] G. C. Ghirardi. Sneaking a Look at God’s Cards: Unraveling the Mysteries of Quantum Mechanics. Princeton University Press, Princeton, NJ, USA, revised edition, 2005. ISBN 0-691-04934-3, 0-691-12139-7, 0-691-13037-X. xix + 488 pp. LCCN Q173 .G4813 2004.

Greenberger:2009:CQP

[GHW09] Daniel M. Greenberger, Klaus Hentschel, and Friedel Weinert, editors. Compendium of Quantum Physics: Concepts, Experiments, History and Philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-70626-7, 3-540-70622-4. xvi + 901 pp. LCCN C174.12 .C66 2009.

Gibson:2019:SIH

[Gib19] Susannah Gibson. The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2019. ISBN 0-19-883337-7 (hardcover). xxi + 377 pp. LCCN Q41.C194 G537 2019.

Ginzburg:2001:PLR

[Gin01] V. L. (Vitaliĭ Lazarevich) Ginzburg. The Physics of a Lifetime: Reflections on the Problems and Personalities of 20th Century Physics. Physics and astronomy online library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-67534-5, 3-642-08699-3. xiii + 513 pp. LCCN QC7 .G59 2001. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0812/00066112-d.html;
 http://catdir.loc.gov/catdir/enhancements/fy0812/00066112-
t.html.

Garwin:2003:CNT

[GL03] Laura Garwin and Tim Lincoln, editors. A century of nature: twentyone discoveries that changed science and the world. University of Chicago Press, Chicago, IL, USA and London, UK, 2003. ISBN 0-226-28413-1 (hardcover), 0-226-28415-8 (paperback). xviii + 360 pp. LCCN Q125

.C384 2003. URL http://catdir.loc.gov/catdir/bios/uchi051/
2003014899.html;   http://catdir.loc.gov/catdir/description/

uchi052/2003014899.html;   http://catdir.loc.gov/catdir/toc/
ecip046/2003014899.html.

Golub:2023:HPF

[GL23] Robert Golub and Steve Keith Lamoreaux. The Historical and Physical Foundations of Quantum Mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2023. ISBN 0-19-186123-5, 0-19-255536-7, 0-19-882218-9 (hardcover), 0-19-882219-7 (paperback). xiii + 747 pp. LCCN QC174.12 .G65 2023.

Glass:1960:BRB

[Gla60] Bentley Glass. Book review: Atomic Physics and Human Knowledge, by Niels Bohr. Quarterly Review of Biology, 35(3):223–224, September 1960. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (elec-

tronic). URL http://www.jstor.org/stable/2816035.

Glanz:2000:PFA

[Gla00] James Glanz. Of physics, friendship and the atomic bomb: 60 years later, a mystery abides of physics, friendship and Nazi Germany’s atomic bomb efforts mystery of a 1941 meeting becomes drama on the stage. New York Times, ??(??):F1–F2, March 21, 2000. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.

proquest.com/hnpnewyorktimes/docview/91737705/.

Guerra:2014:WEC

[GLR14] Francesco Guerra, Matteo Leone, and Nadia Robotti. When energy conservation seems to fail: The prediction of the neutrino. Science & Education (Springer), 23(6):1339–1359, June 2014. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Gottstein:2001:HMB

[GLSC01] Klaus Gottstein, Harry J. Lipkin, Donald C. Sachs, and David C. Cassidy. Heisenberg’s message to Bohr: Who knows? Physics Today, 54(4):14, 92–93, April 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/link/

phtoad/v54/i4/p14/s1.

Grene:1986:SS

[GN86] Marjorie Grene and Debra Nails. Spinoza and the Sciences, volume 91 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346.

xix + 339 pp. LCCN ???? URL https://link.springer.com/book/
10.1007/978-94-009-4514-2.

Goldberger:1987:RDJ

[Gol87] Leo Goldberger, editor. The Rescue of the Danish Jews: moral courage under stress. New York University Press, New York, NY, USA, 1987. ISBN 0-8147-3010-8, 0-8147-3011-6 (paperback). xxvii + 222 pp. LCCN DS135.D4 R47 1987. US$40.00; US$20.00. URL http://www.

loc.gov/catdir/enhancements/fy0808/87011253-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0808/87011253-d.html.

Goldberg:1994:LWA

[Gol94] Stanley Goldberg. Letter: Wrong about Bohr, too. Bulletin of the Atomic Scientists, 50(5):3, 59–60, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomatam:2006:NBI

[Gom06] Ravi Gomatam. Niels Bohr’s interpretation and the Copenhagen interpretation — are the two incompatible? Philosophy of Science, 74(5):736–748, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/525618.

Goodstein:1992:BRB

[Goo92] Judith R. Goodstein. Book review: Atomi metafori paradossi: Niels Bohr e la costruzione di una nuova fisica by Sandro Petruccioli. Isis, 83 (1):158, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234048.

Gordin:2017:BRT

[Gor17] Michael D. Gordin. Book review: A Tale of Seven Scientists and a New Philosophy of Science, Scerri, Eric. 262 pp. Oxford U.P., New York, 2016. Price: $29.95 (hardcover). ISBN 978-0-19-023299-3. American Journal of Physics, 85(10):801–802, October 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Gottstein:2002:HBA

[Got02] Klaus Gottstein. Heisenberg and Bohr — another view. Science, 295 (5563):2211, March 22, 2002. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/content/
295/5563/2211.2.full. Rebuttal to [Cho02].

Goudsmit:1971:DES

[Gou71] Samuel A. Goudsmit. The discovery of the electron spin. Web

document, April 1971. URL http://www.lorentz.leidenuniv.nl/

history/spin/goudsmit.html. English translation by J. H. van der Waals of Goudsmit’s lecture in Dutch for the golden jubilee of the Dutch Physical Society in April 1971.

Gowing:1985:NBN

[Gow85] Margaret Gowing. Niels Bohr and nuclear weapons. In French and Kennedy [FK85], pages 266–277. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Guerra:2008:EMF

[GR08] Francesco Guerra and Nadia Robotti. Ettore Majorana’s forgotten publication on the Thomas–Fermi model. Physics in Perspective (PIP), 10 (1):56–76, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-007-0340-8;   http://www.springerlink.com/content/
53l26g8353l2421g/.

Graetzer:1964:DNF

[Gra64] Hans G. Graetzer. Discovery of nuclear fission. American Journal of Physics, 32(9):9–15, January 1964. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See [HS39b, Bey49].

Graham:1985:CML

[Gra85] Loren Graham. Complementarity and Marxism–Leninism. In French and Kennedy [FK85], pages 332–341. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Grandin:2015:SAF

[Gra15] Karl Grandin. “I shall always follow your progress with warm interest”: Niels Bohr as seen from a Swedish perspective until 1930. In Aaserud and Kragh [AK15], pages 522–?? ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Greenberger:2000:BIB

[Gre00] Daniel Greenberger. Bohr the innovator? Or Bohr the intimidator? Science (New Series), 287(5461):2166–2167, March 24, 2000. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Gornitz:2001:BSM

[GRE^{+}01] Thomas Görnitz, Helmut Rechenberg, Berthold-Georg Englert, Thomas Walcher, Katja Bammel, Thomas W. Beneke, Wolfgang W. Schwippert, and Filip Floecel. Bücher und Software: Meyenn: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg U. A./Kraus: Von der Uranspaltung zur Göttinger Erklärung. Otto Hahn, Werner Heisenberg und Carl Friedrich von Weizsäcker und die Verantwortung des Wissenschaftlers/Dürr: Bohmsche Mechanik als Grundlage der Quantenmechanik/Thomas/Weise: The Structure of the Nucleon/Sube: Langenscheidts Fachwörterbuch Physik. Deutsch-Englisch/ ConceptDraw 1.6 Zeichnen mit intelligenten Objekten /Rath: Quantenphysik für Windows Version 1.5. Physikalische Blätter, 57(11):82–86, November 2001. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.20010571123/abstract.

Grujic:2014:BMM

[Gru14] Petar Grujic. Bohr’s molecular model and the melding of classical and quantum mechanics. Physics Today, 67(8):8–10, August 2014. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [SSH14a].

Greenberger:1999:EEP

[GRZ99] Daniel M. Greenberger, Wolfgang L. Reiter, and Anton Zeilinger, editors. Epistemological and experimental perspectives on quantum physics, volume 7 of Vienna Circle Institute yearbook. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). x + 377 pp. LCCN QC174.13 .E65 1999.

Garg:2004:BSQ

[GS04] Anupam Garg and Michael Stone. Bohr–Sommerfeld quantization of spin Hamiltonians. Physical Review Letters, 92(1):010401, January 9, 2004. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Gavroglu:1995:PPS

[GSW95] Kōstas Gavroglu, John J. Stachel, and Marx W. Wartofsky, editors. Physics, Philosophy, and the Scientific Community: Essays in the Philosophy and History of the Natural Sciences and Mathematics in Honor of Robert S. Cohen, volume 163 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 0-7923-2991-0 (set), 0-7923-2988-0. xxvii + 383 pp. LCCN Q174 .B67 vol. 163. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0823/94022250-d.html;
 http://catdir.loc.gov/catdir/enhancements/fy0823/94022250-
t.html.

Gustafson:1955:PCF

[Gus55] Torsten Gustafson. On the potential collective flow of a rotating nucleus with non-ellipsoidal boundary: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(5) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 15 pp. LCCN AS281.

Haberer:1969:PCS

[Hab69] Joseph Haberer. Politics and the Community of Science. Van Nostrand Reinhold, New York, NY, USA, 1969. vi + 337 pp. LCCN Q125 .H23 1969.

Hall:2009:GSP

[Hal09] Derek Hall, editor. Great scientists. Physical sciences. Facts at your fingertips. Brown Bear Books, Redding, CT, USA, 2009. ISBN 1-933834-46-3. 64 pp. LCCN QC15 .P468 2009.

Halpern:2012:QHP

[Hal12] Paul Halpern. Quantum humor: The playful side of physics at Bohr’s Institute for Theoretical Physics. Physics in Perspective (PIP), 14(3): 279–299, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-011-0071-8.

Halpern:2017:QLH

[Hal17] Paul Halpern. The Quantum Labyrinth: How Richard Feynman and John Wheeler Revolutionized Time and Reality. Basic Books, New York, NY, USA, 2017. ISBN 0-465-09758-8 (hardcover), 0-465-09759-6 (ebook). ix + 311 pp. LCCN QC174.12 .H347 2017.

Hamilton:2005:SBS

[Ham05] Mark Hamilton. Singular Bohr–Sommerfeld leaves and geometric quantization. Ph.D. dissertation, University of Toronto, Toronto, ON, Canada, 2005. 89 pp.

Hartree:1923:SAN

[Har23] D. R. Hartree. Some approximate numerical applications of Bohr’s theory of spectra. Proceedings of the Cambridge Philosophical Society.

Mathematical and physical sciences, 21(??):625–641, ???? 1923. CODEN PCPSA4. ISSN 0008-1981.

Hartree:1926:SQA

[Har26] D. R. Hartree. Some quantitative applications of Bohr’s theory of spectra. Ph.D. thesis, Faculty of Physics and Chemistry, University of Cambridge, Cambridge, UK, October 9, 1926.

???? pp. URL https://idiscover.lib.cam.ac.uk/;   https:
//www.proquest.com/dissertations-theses/some-quantitative-
applications-bohr-s-theory/docview/301187054.

Hartree:1928:WMA

[Har28] D. R. Hartree. The wave mechanics of an atom with a non-coulomb central field. Part I. Theory and methods. Mathematical proceedings of the Cambridge Philosophical Society, 24(1):89–110, ???? 1928. CODEN PCPSA4. ISSN 0008-1981. Page 91 of this paper introduces atomic units for length (a_{H} = h^{2}/(4π^{2}me^{2}), the radius of first Bohr orbit of hydrogen), charge (e, the electron charge), and mass (m, electron mass). From these are derived the unit of action (h/(2π) ), the unit of energy (e^{2}/a, twice the ionization energy of hydrogen), and the unit of time (1/(4πcR)). The value R is the Rydberg constant.

Harcourt:1983:BOT

[Har83] R. D. Harcourt. Bohr orbit theory revisited. I. Ground-state energies for the helium isoelectronic sequence. Journal of Physics B: Atomic and Molecular Physics, 16(15):2647–2657, August 14, 1983. CODEN JPAMA4. ISSN 0953-4075 (print), 1361-6455 (electronic). URL http://

iopscience.iop.org/0022-3700/16/15/009/. See also Part II [Har87].

Harcourt:1987:BOT

[Har87] Richard D. Harcourt. Bohr orbit theory revisited. II. Energies for 1S, 2P, 3D, and 4F states of helium. International Journal of Quantum Chemistry, 31(3):445–453, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See also Part I [Har83].

Hartree:19xx:HDRc

[Harxx] D. R. Hartree. Hartree, Douglas Rayner, (1897–1958), mathematical physicist: correspondence with Niels Bohr (1928–1947). Niels Bohr Archive, Blegdamsvej 17, Copenhagen, Denmark, DK-2100, 19xx. URL

https://www.nbarchive.dk/. Reference NRA 33014 NAHC.

Harper:2001:AGG

[Har01] Eamon Harper. In appreciation: George Gamow: Scientific amateur and polymath. Physics in Perspective (PIP), 3(3):335–372, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic).  URL http://adsabs.harvard.edu/abs/
2001PhP.....3..335H;   http://link.springer.com/article/10.
1007/PL00000536;   http://www.springerlink.com/content/m0htmt9ww9b9jjv1/
;   http://www.springerlink.com/openurl.asp?genre=journal&
issn=1422-6944.

Hastrup:2015:P

[Has15] Kirsten Hastrup. Prelude. In Aaserud and Kragh [AK15], pages 51–60. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/2015/
06/Bohr-Book.pdf.

Hawkings:1994:LOB

[Haw94] David Hawkings. Letter: Oppenheimer and Bohr. Bulletin of the Atomic Scientists, 50(5):60, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Heisenberg:1963:KDQ

[HB63] Werner Heisenberg and Niels Bohr, editors. Die Kopenhagener Deutung der Quantentheorie. (German) [The Copenhagen Interpretation of Quantum Theory]. E. Battenberg, Stuttgart, Germany, 1963. 67 pp. LCCN QC174.1 .K65.

Heisenberg:1996:FCS

[HBS96] Werner Heisenberg, Niels Bohr, and Erwin Schrödinger. Física cuántica: (Spanish) [Quantum physics]. Círculo de Lectores, Barcelona, Spain, 1996. ISBN ???? ???? pp. LCCN ????

Haensel:1980:BFK

[HEB^{+}80] R. Haensel, Karl Ehrlich, K. Bethge, H. Soffel, W. von Witsch, and H. Rechenberg. Bücher Filme: Kunz: Synchrotron Radiation Techniques and Applications/Lehmann: Interaction of Radiation with Solids and Elementary Defect Production/Olson u. Schumacher: Collective Ion Acceleration/Brosche u Sündermann: Tidal Friction and the Earth’s Rotation/Haas u. Koch: Physik-Lehrbuch für Pharmazeuten und Mediziner/Hermann, Meyenn u. Weisskopf: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a., Bd. I: 1919–1929. Physikalische Blätter, 36(10):318–320, October 1980. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19800361013/abstract.

Heelan:1995:QMS

[Hee95] Patrick A. Heelan. Quantum mechanics and the social sciences: After hermeneutics. Science & Education (Springer), 4(2):127–136, April 1995. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Heisenberg:1928:DSC

[Hei28] Werner Heisenberg. Discussione sulla communicazione Bohr. (Italian) [Discussions of Bohr’s communication]. In ????, editor, Atti del Congresso Internazionale dei Fisici, 11–20 Settembre 1927-V, Como– Pavia–Roma, pages 593–594, 597. Nicola Zanichelli, Bologna, Italy, 1928. LCCN ????

Heisenberg:1935:NBF

[Hei35] Werner Heisenberg. Niels Bohr zum fünfzigsten Geburtstage am 7. Oktober 1935. (German) [Niels Bohr’s fiftieth birthday, 7 October 1935]. Die Naturwissenschaften, 23(40):679, October 4, 1935. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1948:LBV

[Hei48] Werner Heisenberg. [Letter to B. L. van der Waerden]. Cassidy [Cas09, page 314] writes of this letter “Heisenberg himself seems to suggest another purpose for his visit [to Niels Bohr in Copenhagen] … Heisenberg implied that he was attempting to stave off an Allied crash program on nuclear fission research.”, April 28, 1948.

Heisenberg:1955:DIQ

[Hei55] Werner Heisenberg. The development of the interpretation of the quantum theory. In Pauli et al. [PRW55], pages 12–29. LCCN QC71

.P3 1955. URL http://labs.adsabs.harvard.edu/ui/abs/1955nbdp.
book...12H.

Heisenberg:1962:BIQ

[Hei62] Werner Heisenberg. Bohrs Interpretation der Quantentheorie und die Physik der Elementarteilchen. (German) [Bohr’s interpretation of quantum theory and the physics of elementary particles]. Fysisk Tidsskrift, 60(??):47–53, ???? 1962.

Heisenberg:1963:NB

[Hei63a] Werner Heisenberg. Niels Bohr 7.10.1885–18.11.1962. In ????, editor, Bayerische Akademie der Wissenschaften. Jahrbuch 1963. (German), pages 204–207. Bayerische Akademie der Wissenschaften, Munich, West Germany, 1963.

Heisenberg:1963:PST

[Hei63b] Werner Heisenberg. The present situation in the theory of elementary particles. In ????, editor, Commemoration of the Fiftieth Anniversary of Niels Bohr’s First Papers on Atomic Constitution Held in Copenhagen on 8–15 July 1963. Part 2: Session on Elementary Particles, pages 1–21. Institute for Theoretical Physics, Copenhagen, DK, 1963. URL ????.

Heilbron:1964:HAM

[Hei64a] John L. Heilbron. A History of Atomic Models from the Discovery of the Electron to the Beginnings of Quantum Mechanics. Dissertation, University of California, Berkeley, Berkeley, CA, USA, ???? 1964. ???? pp.

Heisenberg:1964:KOD

[Hei64b] Werner Heisenberg. Kvanteteorien og dens fortolkning. (Danish) [Quantum theory and its interpretation]. In Bohr and Rozental [BR64], pages 90–103. LCCN QC16.B63 N5. English translation in [Roz67].

Heisenberg:1970:PHL

[Hei70] Werner Heisenberg. Professor Heisenberg’s lecture (atoms and peace). In ????, editor, Dansk Ingeniørforening. Award of the Niels Bohr Gold Medal to Werner Karl Heisenberg. 7th October 1970, pages 20–33. Dansk Ingeniørforening, Copenhagen, DK, 1970. LCCN ????

Heilbron:1974:HGJ

[Hei74] J. L. Heilbron. H. G. J. Moseley: The life and letters of an English physicist, 1887–1915. University of California Press, Berkeley, CA, USA, 1974. ISBN 0-520-02375-7. xiii + 312 pp. LCCN QC16.M68 H44.

Heisenberg:1975:BRU

[Hei75] Werner Heisenberg. Book review: U. Hoyer, Die Geschichte der Bohrschen Atomtheorie. Physikalische Blätter, 31(1):45, January 1975. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19750310111/abstract.

Heilbron:1977:JJT

[Hei77] John L. Heilbron. J. J. Thomson and the Bohr atom. Physics Today, 30(4):23–30, April 1977. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.org/

resource/1/phtoad/v30/i4/p23_s1.

Heilbron:1981:RBA

[Hei81] John L. Heilbron. Rutherford–Bohr atom. American Journal of Physics, 49(3):223–231, March 1981. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/resource/

1/ajpias/v49/i3/p223_s1.

Heisenberg:1984:IER

[Hei84] Elisabeth Heisenberg. Inner exile: recollections of a life with Werner Heisenberg. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1984. ISBN 0-8176-3146-1. xvii + 170 + 20 pp. LCCN QC16.H35 H4413 1984.

Heilbron:1985:EMC

[Hei85a] J. L. Heilbron. The earliest missionaries of the Copenhagen spirit. Revue d’Histoire des Sciences, 38(3–4):195–230, 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0048-7996, 0151-4105, 1969-6582 (electronic).

URL http://www.persee.fr/web/revues/home/prescript/article/
rhs_0151-4105_1985_num_38_3_4005.

Heilbron:1985:BFTa

[Hei85b] John L. Heilbron. Bohr’s first theories of the atom. Physics Today, 38 (10):28–36, October 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Heilbron:1985:BFTb

[Hei85c] John L. Heilbron. Bohr’s first theories of the atom. In French and Kennedy [FK85], pages 33–49. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Heilbron:1985:JJT

[Hei85d] John L. Heilbron. J. J. Thomson and the Bohr atom. In Weart and Phillips [WP85], pages 303–309. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [Hei77].

Heisenberg:1985:R

[Hei85e] Werner Heisenberg. Reminiscences from 1926 and 1927. In French and Kennedy [FK85], pages 163–171. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Heims:1986:BRB

[Hei86a] Steve J. Heims. Book review: Niels Bohr: A Centenary Volume by A. P. French; P. J. Kennedy. Isis, 77(2):388–389, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232722.

Heisenberg:1986:GPN

[Hei86b] Werner Heisenberg. Glückwunschschreiben an Professor Niels Bohr (1961). (German) [Congratulations to Professor Niels Bohr (1961)]. In Blum et al. [BDR86], pages 129–130. ISBN 3-492-02928-0. LCCN ????

Heilbron:1996:BRBb

[Hei96] J. L. Heilbron. Book review: A Guide to the Archival Collections in the Niels Bohr Library at the American Institute of Physics. Isis, 87(2): 384–385, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236131.

Heisenberg:2010:SNB

[Hei10a] Werner Heisenberg. In der Spur von Niels Bohrs Physik und Philosophie. (German) [On the track of Niels Bohr’ physics and philosophy]. In Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die “fröhliche Wissenschaft” (Jugend bis Nobelpreis). (German) [Werner Heisenberg — the language of atoms: Life and Work — a scientific biography: the “Happy Science” (youth to Nobel Prize)] [Rec10], pages 223–297. ISBN 3-540-69221-5 (hardcover), 3-540-69222-3 (e-book). LCCN QC16.H35 R434 2010.

URL http://ebooks.ciando.com/book/index.cfm/bok_id/46829;
 http://www.ciando.com/img/books/width167/3540692223_k.jpg;
 http://www.ciando.com/pictures/bib/3540692223bib_t_1.jpg.

Heisenberg:2010:QWA

[Hei10b] Werner Heisenberg. Quantenmechanik, Wellenmechanik und Anschauung. (German) [Quantum mechanics, wave mechanics and opinion]. In Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die “fröhliche Wissenschaft” (Jugend bis Nobelpreis). (German) [Werner Heisenberg — the language of atoms: Life and Work — a scientific biography: the “Happy Science”

(youth to Nobel Prize)] [Rec10], pages 449–525. ISBN 3-540-69221-5 (hardcover), 3-540-69222-3 (e-book). LCCN QC16.H35 R434 2010.

URL http://ebooks.ciando.com/book/index.cfm/bok_id/46829;
 http://www.ciando.com/img/books/width167/3540692223_k.jpg;
 http://www.ciando.com/pictures/bib/3540692223bib_t_1.jpg.

Heilbron:2013:BCH

[Hei13] John L. Heilbron. Bohr’s creation of his quantum atom. In APS April Meeting Abstracts. American Physical Society, Ridge, NY 11961, USA,

April 2013. URL http://adsabs.harvard.edu/abs/2013APS..APR.
P1001H.

Heilbron:2015:MCA

[Hei15] J. L. Heilbron. “My courage is ablaze so wildly”: Niels Bohr en route to his quantum atom. In Aaserud and Kragh [AK15], pages 27–50. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63 2015. URL http:/

/www.ijqf.org/wps/wp-content/uploads/2015/06/Bohr-Book.pdf.

Heilbron:2020:NBV

[Hei20] J. L. Heilbron. Niels Bohr: a very short introduction, volume 627 of Very short introductions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2020. ISBN 0-19-881926-9 (paperback), 0-19-185986-9 (e-book). xxii + 132 pp. LCCN QC16.B63 H45 2020.

Hellemans:1985:BRB

[Hel85] Alexander Hellemans. Book review: Niels Bohr: The Man, His Science and the World They Changed, by Ruth Moore. Physics Today, 38(10):??, October 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Held:1998:BED

[Hel98a] Carsten Held. Die Bohr–Einstein-Debatte: Quantenmechanik und physikalische Wirklichkeit. (German) [The Bohr–Einstein Debate: Quantum Mechanics and Physical Reality]. Ferdinand Schöningh, Paderborn, Germany, 1998. ISBN 3-506-73823-2 (paperback). 292 pp. LCCN 0.2hel a0160 a0165 a0170 a0365.

Hellwig:1998:CCC

[Hel98b] Karl-Eberhard Hellwig. Correspondence and complementarity. Comment on Brigitte Falkenburg: “Bohr’s principles of unifying quantum disunities” [Philos. Natur. 35 (1998), no. 1, 95–120]. Philosophia Naturalis, 35(1):121–125, 1998. CODEN ???? ISSN 0031-8027. Models, theories and disunity in physics (Berlin, 1996).

Hendry:1981:BKS

[Hen81] John Hendry. Bohr–Kramers–Slater: a virtual theory of virtual oscillators and its role in the history of quantum mechanics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(2):189–221, July 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hendry:1984:CQM

[Hen84] John Hendry. The creation of quantum mechanics and the Bohr–Pauli dialogue, volume 14 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 90-277-1648-X. xi + 177 pp. LCCN QC173.98 .H46 1984.

Hendry:1986:BRJ

[Hen86a] John Hendry. Book review: Jagdish Mehra & Helmut Rechenberg. The Historical Development of Quantum Theory. Volume 1, The Quantum Theory of Planck, Einstein, Bohr and Sommerfeld: Its Foundations and the Rise of its Difficulties 1900–1925. Volume 2, The Discovery of the Quantum Mechanics 1925. Volume 3, The Formulation of Matrix Mechanics and its Modifications 1925–1926. Volume 4, The Fundamental Equations of Quantum Mechanics 1925–1926 and The Reception of the New Quantum Mechanics 1925–1926. New York, Heidelberg and Berlin: Springer-Verlag, 1982. Volume 1 in two parts, pp. xlvii + 372, vi + 506. ISBN 3-540-90642-8, 3-540-90667-3. DM 75, DM 85. Volume 2, pp. vii + 355. ISBN 3-540-90674-6. DM 65. Volume 3, pp. vii + 334. ISBN 3-540-90675-4. DM 75. Volume 4, pp. viii + 322. ISBN 3-540-90680-0. DM 75. Andrew Pickering. Constructing Quarks. A Sociological History of Particle Physics. Edinburgh: Edinburgh University Press, 1984. Pp. xi + 468. ISBN 0-85224-458-4 £20. British Journal for the History of Science, 19(2):206–208, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/stable/

4026596.

Hendry:1986:BRW

[Hen86b] John Hendry. Book review: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Scientific Correspondence with Bohr, Einstein, Heisenberg a.o. Berlin, Heidelberg and New York: Springer, 1985. Pp. xxix + 783. ISBN 3-540-13609-6. DM 298.00. British Journal for the History of Science, 19(3):348, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4026519.

Hendry:1992:BRA

[Hen92] John Hendry. Book review: Abraham Pais. Niels Bohr’s Times, in Physics, Philosophy and Polity. Oxford: Oxford University Press, 1991. Pp. xvii + 565. British Journal for the History of Science, 25(4):490–491, December 1992. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4027078.

Hendry:1993:BRB

[Hen93] John Hendry. Book review: Niels Bohr: His Heritage and Legacy: An Anti-Realist View of Quantum Mechanics by Jan Faye. Isis, 84(1): 169, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235605.

Hentschel:2009:QTC

[Hen09a] Klaus Hentschel. Quantum theory, crisis period 1923–early 1925. In Greenberger et al. [GHW09], pages 613–617. ISBN 3-540-70626-7, 3-540-70622-4. LCCN C174.12 .C66 2009.

Hentschel:2009:SR

[Hen09b] Klaus Hentschel. Selection rules. In Greenberger et al. [GHW09], pages 690–692. ISBN 3-540-70626-7, 3-540-70622-4. LCCN C174.12 .C66 2009.

Hentschel:2009:S

[Hen09c] Klaus Hentschel. Spin. In Greenberger et al. [GHW09], pages 726–731. ISBN 3-540-70626-7, 3-540-70622-4. LCCN C174.12 .C66 2009.

Hertz:1923:BTE

[Her23] G. Hertz. Bohrsche Theorie und Elektronenstoß. (German) [Bohr theory and electron collision]. Die Naturwissenschaften, 11(27):564–567, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Hermansen:1955:PMT

[Her55] Alfred Hermansen. A polarimetric method for thickness control in the production of interference filters: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(6) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 41 pp. LCCN AS281.

Hermann:1971:GQT

[Her71] Armin Hermann. The Genesis of Quantum Theory (1899–1913). MIT Press, Cambridge, MA, USA, 1971. ISBN 0-262-08047-8. ix + 165 pp. LCCN QC174 .H4613 1971.

Herzfeld:1972:BAR

[Her72] Karl F. Herzfeld. Bohr atom: a remark on the early history. Science (New Series), 175(4028):1393–1394, March 24, 1972. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Hettema:1995:BTA

[Het95] Hinne Hettema. Bohr’s theory of the atom 1913–1923: a case study in the progress of scientific research programmes. Studies in History and Philosophy of Science Part B: Studies In History and Philosophy of Modern Physics, 26(3):307–323, December 1995. CODEN ???? ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.sciencedirect.

com/science/article/pii/1355219895000186.

Hartz:2015:UAN

[HF15] Thiago Hartz and Olival Freire, Jr. Uses and appropriations of Niels Bohr’s ideas about quantum field measurement, 1930–1965. In Aaserud and Kragh [AK15], pages 397–418. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Hon:2015:CMB

[HG15] Giora Hon and Bernard R. Goldstein. Constitution and model: Bohr’s quantum theory and imagining the atom. In Aaserud and Kragh [AK15], pages 347–359. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172

.O63 2015. URL http://www.ijqf.org/wps/wp-content/uploads/
2015/06/Bohr-Book.pdf.

Hansen:1955:FTN

[HH55] H. M. Hansen and K. G. Hansen, editors. Festskrift til Niels Bohr på hans halvfjerdsårsdagen den 7 oktober 1955. (Danish) [Festival volume for Niels Bohr’s 70th birthday, 7 October 1955]. Ejnar Munksgaard, København, Danmark, 1955. 302 pp. LCCN QC3 .F43.

Harper:1976:FPT

[HH76] William L. Harper and C. A. (Clifford Alan) Hooker, editors. Foundations of Probability Theory, Statistical Inference, and Statistical Theories of Science: Volume III Foundations and Philosophy of Statistical Theories in the Physical Sciences, volume 6c of The University of Western Ontario Series in Philosophy of Science, A Series of Books on Philosophy of Science, Methodology, and Epistemology Published in Connection with The University of Western Ontario Philosophy of Science Programme. Springer Netherlands, Dordrecht, 1976. ISBN 90-277-0621-2, 94-010-1438-8 (e-book). ISSN 1566-659X. 253 pp. LCCN B67. URL

http://www.springerlink.com/content/978-94-010-1438-0.

Hoyningen-Huene:1994:NBA

[HH94] Paul Hoyningen-Huene. Niels Bohr’s argument for the irreducibility of biology to physics. In Faye and Folse [FF94], chapter 10, pages 231– 255. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https:/

/link.springer.com/chapter/10.1007/978-94-015-8106-6_10.

Hargittai:2004:CSI

[HH04a] Magdolna Hargittai and István Hargittai, editors. Candid science IV: conversations with famous physicists. Imperial College Press, London, UK, 2004. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). xvi + 711

pp. LCCN QC15 .H295 2004. URL http://www.worldscibooks.com/
physics/p304.html.

Hargittai:2004:EPW

[HH04b] Magdolna Hargittai and István Hargittai. Eugene P. Wigner. In Candid science IV: conversations with famous physicists [HH04a], pages 1– 19. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). LCCN QC15 .H295

2004. URL http://www.worldscibooks.com/physics/p304.html.

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Holst:1922:BAA

[HK22] Helge Holst and Hendrich Anthony Kramers. Bohrs atomteori: almenfatteligt fremstillet. (Danish) [Bohr’s atomic theory: easy presentation]. Gyldendal, Kjøbenhavn, Danmark, 1922. 134 pp. LCCN ????

Holst:1926:BAA

[HK26] Helge Holst and Hendrik Anthony Kramers. Das Atom und die Bohrsche Theorie seines Baues. (German) [The Atom and the Bohr Theory: Its Construction]. Julius Springer, Berlin, Germany, 1926. ???? pp. LCCN ????

Heilbron:1969:GBA

[HK69] John L. Heilbron and Thomas S. Kuhn. The genesis of the Bohr atom. Historical Studies in the Physical Sciences, 1(??):vi, 211–290, ???? 1969. CODEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.org/

stable/27757291.

Holst:1929:BAA

[HKK29] Helge Holst, Hendrik Anthony Kramers, and Otto Klein. Bohrs atomteori: almenfatteligt fremstillet. (Danish) [Bohr’s atomic theory: easy presentation]. Gyldendal, Kjøbenhavn, Danmark, second edition, 1929. 168 pp. LCCN ????

Holbrow:2010:SBA

[HLA^{+}10] Charles H. Holbrow, James N. Lloyd, Joseph C. Amato, Enrique Galvez, and M. Elizabeth Parks. Spectra and the Bohr atom. In Modern Introductory Physics, chapter 17, pages 517–551. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-387-79080-2. LCCN QC21.3 .M63 2010.

Heunen:2012:BOA

[HLS12] Chris Heunen, Nicolaas P. Landsman, and Bas Spitters. Bohrification of operator algebras and quantum logic. Synthese, 186(3):719–752, June 2012. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11229-
011-9918-4;   http://link.springer.com/content/pdf/10.1007/
s11229-011-9918-4.pdf.

Hirosige:1964:FBT

[HN64] T. Hirosige and S. Niso. Formation of Bohr’s theory of atomic constitution. Jap. Studies Hist. Sci., 3(??):6–28, ???? 1964.

Hirosige:1970:GBA

[HN70] Tetu Hirosige and Sigeko Nisio. The genesis of the Bohr atom model and Planck’s theory of radiation. Japanese Studies in the History of Science, 9(??):35–47, ???? 1970. CODEN JSHIAE. ISSN 0090-0176.

Hoyer:1974:GBA

[HnB74] Ulrich Hoyer and Niels Associated name Bohr. Die Geschichte der Bohrschen Atomtheorie. (German) [The History of the Bohr Atomic Theory]. Physik-Verlag, Weinheim, West Germany, 1974. ISBN 3-87664-022-9. 267 pp. LCCN QC173.

Hoch:1983:BDY

[Hoc83] Paul Hoch. Bohr’s decisive years. Contemporary Physics, 24(2):203– 204, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hoffmann:1947:SSQ

[Hof47] Banesh Hoffmann. The Strange Story of the Quantum, an Account for the General Reader of the Growth of the Ideas Underlying Our Present Atomic Knowledge. Harper and Brothers, New York, 1947. xi + 1 + 239 pp. LCCN QC174.1 .H6.

Hoffmann:1959:SSQ

[Hof59] Banesh Hoffmann. The Strange Story of the Quantum; an Account for the General Reader of the Growth of the Ideas Underlying Our Present Atomic Knowledge. Dover, New York, NY, USA, second edition, 1959. 285 pp. LCCN QC174.1 .H6 1959.

Hofmann:1994:BRBb

[Hof94] James R. Hofmann. Book review: The Physics of Chance: From Blaise Pascal to Niels Bohr by Charles Ruhla; G. Barton. Isis, 85(4):680–681, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235297.

Hoffmann:2014:SSQ

[Hof14] Banesh Hoffmann. The Strange Story of the Quantum: an Account for the General Reader of the Growth of the Ideas Underlying Our Present Atomic Knowledge. Dover Books on Science. Dover, New York, NY, USA, second edition, 2014. ISBN 0-486-20518-5. xi + 285 + 5 pp. LCCN ????

Hoiberg:2001:YNL

[Hoi01] Dale Hoiberg, editor. 100 years with Nobel laureates. Encyclopaedia Britannica (India) and I.K. International, New Delhi, India, 2001. ISBN 81-88237-00-0. xiii + 1098 pp. LCCN AS911.N9 A15 2001.

Holton:1970:RC

[Hol70] Gerald Holton. The roots of complementarity. Dædalus, 99(4):1015– 1055, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.org/stable/20023980.

Holzner:1980:BEB

[Hol80] Michael Holzner. Bohr effect in blood oxygenation in the presence of heavy metal ions: a computer simulation. M.S. dissertation, California State University, Long Beach, Long Beach, CA, USA, 1980. 105 pp.

Holcomb:1986:LVC

[Hol86] Harmon R. Holcomb, III. Latency versus complementarity: Margenau and Bohr on quantum mechanics. British Journal for the Philosophy of Science, 37(2):193–206, June 1986. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.jstor.org/stable/

686977.

Holton:1988:NBI

[Hol88a] Gerald James Holton. Niels Bohr and the integrity of science. In Thematic origins of scientific thought: Kepler to Einstein [Hol88b], page ?? ISBN 0-674-87747-0, 0-674-87748-9 (paperback). LCCN Q125 .H722 1988.

Holton:1988:TOS

[Hol88b] Gerald James Holton. Thematic origins of scientific thought: Kepler to Einstein. Harvard University Press, Cambridge, MA, USA, revised edition, 1988. ISBN 0-674-87747-0, 0-674-87748-9 (paperback). 499 pp. LCCN Q125 .H722 1988.

Holton:2000:BRW

[Hol00] Gerald Holton. Book review: What is Copenhagen trying to tell us? Los Angeles Times, ??(??):??, December 31, 2000. URL https://www.

latimes.com/archives/la-xpm-2000-dec-31-bk-6553-story.html.

Holton:2005:VVS

[Hol05] Gerald James Holton. Victory and vexation in science: Einstein, Bohr, Heisenberg, and others. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01519-3. xi + 229 pp. LCCN Q180.A3 H65 2005.

Honner:1981:TAI

[Hon81] John Roderick Honner. Transcendental arguments in the interpretation of quantum theory and the foundations of theology: Niels Bohr, Karl Rahner, and the philosophy of science. Thesis (Ph.D.), Oxford University, Oxford, UK, 1981. ???? pp. 4 microfiches.

Honner:1982:NBM

[Hon82a] John Honner. Niels Bohr and the mysticism of nature. Zygon Journal of Religion and Science, 17(3):243–253, September 1982. ISSN 0591-2385 (print), 1467-9744 (electronic). URL http://onlinelibrary.wiley.

com/doi/10.1111/j.1467-9744.1982.tb00481.x/abstract.

Honner:1982:TPN

[Hon82b] John Honner. The transcendental philosophy of Niels Bohr. Studies in History and Philosophy of Science Part A, 13(1):1–29, March 1982. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0039368182900024.

Honner:1988:DNN

[Hon88] John Honner. The description of nature: Niels Bohr and the philosophy of quantum physics. Clarendon Press, Oxford, UK, 1988. ISBN 0-19-824976-4. 256 pp. LCCN QC171.2.

Hooker:1991:PPI

[Hoo91] C. A. Hooker. Projection, physical intelligibility, objectivity and completeness: the divergent ideals of Bohr and Einstein. British Journal for the Philosophy of Science, 42(4):491–511, December 1991. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:/

/bjps.oxfordjournals.org/content/42/4/491.full.pdf+html.

Hooker:1994:BCE

[Hoo94] Clifford A. Hooker. Bohr and the crisis of empirical intelligibility: An essay on the depth of Bohr’s thought and our philosophical ignorance. In Faye and Folse [FF94], chapter 8, pages 155–199. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https://link.springer.com/

chapter/10.1007/978-94-015-8106-6_8.

Houtermans:1930:NAB

[Hou30] Friedrich G. Houtermans. Neuere Arbeiten über Quanthentheorie des Atomkerns. (German) [Recent work on quantum theory of the atomic nucleus]. Ergebnisse der exakten Naturwissenschaften, 9:123–221, 1930. CODEN EENAA3. ISBN 3-662-37481-1 (print), 3-662-38246-6 (e-

book). ISSN 0367-0325. URL https://link.springer.com/chapter/
10.1007/978-3-662-38246-2_4.

Howard:1979:CON

[How79] Don Armin Howard. Complementarity and Ontology: Niels Bohr and the Problem of Scientific Realism in Quantum Physics. Ph.D dissertation, Boston University Graduate School, Boston, MA, USA, June 1979. 438 pp.

Howard:1986:BRB

[How86] Don Howard. Book review: The Philosophy of Niels Bohr: The Framework of Complementarity by Henry J. Folse. Isis, 77(1):117–118, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232518.

Howard:1999:BBN

[How99] Don Howard. 4. A brief on behalf of Niels Bohr: 4. Sept. 1998. In Greenberger et al. [GRZ99], page ?? ISBN 0-7923-6063-X. ISSN 0929-6328 (print), 2215-1818 (electronic). LCCN QC174.13 .E65 1999.

Howard:2004:WIC

[How04] Don Howard. Who invented the “Copenhagen Interpretation”? A study in mythology. Philosophy of Science, 71(5):669–682, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/425941.

Howson:2018:PCS

[How18] Colin Howson. The primacy of the classical? Saul Kripke meets Niels Bohr. International Studies in the Philosophy of Science, 32(3– 4):141–153, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
02698595.2020.1767890.

Hoyer:1973:RSE

[Hoy73] Ulrich Hoyer. Über die Rolle der Stabilitätsbetrachtungen in der Entwicklung der Bohrschen Atomtheorie. (German) [On the role of stability conditions in the development of Bohr’s atomic theory]. Archive for History of Exact Sciences, 10(3–5):177–206, January 1973. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0003-9519&volume=10&issue=3&spage=177.

Hoffmann:1981:LHQ

[HP81] Banesh Hoffmann and Michel Paty. L’étrange histoire des quanta. (French) [The strange history of quanta], volume 26 of Points. Sciences.

Éditions du Seuil, Paris, France, 1981. ISBN 2-02-005417-5. 282 pp. LCCN ????

Hermann:1985:WPW

[HP85] Armin Hermann and Wolfgang Pauli, editors. Wolfgang Pauli, wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Bd. 2, 1930–1939. (German) [Wolfgang Pauli, scientific correspondence with Bohr, Einstein, Heisenberg and others, volume 2, 1930–1939], volume Vol. 6 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-13609-6 (Berlin), 0-387-13609-6 (New York). xxxix + 783 pp. LCCN ????

Hess:2001:BTP

[HP01] Karl Hess and Walter Philipp. Bell’s theorem and the problem of decidability between the views of Einstein and Bohr. Proceedings of the National Academy of Sciences of the United States of America, 98(25): 14228–14233, ???? 2001. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Hund:1974:GQHa

[HR74] Friedrich Hund and Gordon Reece. The History of Quantum Theory. Harrap, London, UK, 1974. ISBN 0-245-50902-X. 260 pp. LCCN ???? Translation by Gordon Reece of Geschichte der Quantentheorie [Hun67].

Hahn:1939:NEA

[HS39a] Otto Hahn and Fritz Strassmann. Nachweis der Entstehung aktiver Bariumisotope aus Uran und Thorium durch Neutronenbestrahlung; Nachweis weiterer aktiver Bruchstücke bei der Uranspaltung. (German) [Evidence of formation of active barium isotopes of uranium and thorium by neutron irradiation: further evidence of active fragments from the fission of uranium]. Die Naturwissenschaften, 27(6):89–95, February 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

URL http://www.springerlink.com/content/w406757166152l83/.

Hahn:1939:NVB

[HS39b] Otto Hahn and Fritz Strassmann. Über den Nachweis und das Verhalten der bei der Bestrahlung des Urans mittels Neutronen entstehenden Erdalkalimetalle. (German) [Concerning the existence of alkaline earth metals resulting from the neutron irradiation of uranium]. Die Naturwissenschaften, 27(1):11–15, January 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). A facsimile is also available in [Bey49, pages 87–91] and in [Gra64]. Abridged English translation in [GA71, pages 44–47].

Hubisz:2014:BRP

[Hub14] John L. Hubisz. Book review: The Physics of Chance: From Pascal to Niels Bohr, by Charles Ruhla and translated from the French by G. Barton (1992) with a Foreword by Alain Aspect and published by Oxford University Press, New York, NY (1989), $21.95, pp. xi + 222, paperback, ISBN: 0-19-853977-0. The Physics Teacher, 52(8):510, November 2014. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Hughes:1989:SIQ

[Hug89] R. I. G. Hughes. The Structure and Interpretation of Quantum Mechanics. Harvard University Press, Cambridge, MA, USA, 1989. ISBN 0-674-84391-6 (hardcover), 0-674-84392-4 (paperback). xiii + 369 pp. LCCN QC174.12 .H82 1989.

Hughes:1990:BAM

[Hug90] R. I. G. Hughes. The Bohr atom, models, and realism. Philosophical Topics, 18(2):71–84, 1990. ISSN 0276-2080 (print), 2154-154X (elec-

tronic). URL http://www.jstor.org/stable/43154077.

Hund:1967:GQ

[Hun67] Friedrich Hund. Geschichte der Quantentheorie. (German) [History of quantum theory]. B. I.-Hochschultaschenbücher, 200/200a. Bibliographisches Institut, Mannheim, West Germany, 1967. 239 pp. LCCN QC174.1 .H77.

Hund:1974:HQTb

[Hun74] Friedrich Hund. The History of Quantum Theory. Barnes and Noble, New York, NY, USA, 1974. ISBN 0-06-493060-2. 260 pp. LCCN QC173.98 .H8613 1974b. Translation by Gordon Reece of Geschichte der Quantentheorie [Hun67].

Hund:1985:BGQ

[Hun85] Friedrich Hund. Bohr, Göttingen, and quantum mechanics. In French and Kennedy [FK85], pages 71–75. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Hurter:2022:TBS

[Hür22] Tobias Hürter. Too Big for a Single Mind: How the Greatest Generation of Physicists Uncovered the Quantum World. The Experiment, New York, NY, USA, 2022. ISBN 1-61519-920-9 (hardcover), 1-61519-921-7 (e-book). 357 pp. LCCN QC7 .H88813 2022. Translation to English by David Shaw of Das Zeitalter der Unschärfe, published by Klett-Cotta (2021).

Hermann:1979:WPW

[HvMW79] Armin Hermann, Karl von Meyenn, and Victor F. Weisskopf, editors. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band I: 1919–1929 Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume I: 1919–1929, volume 2 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-08962-4, 3-540-08962-4. xlvii + 577 pp. LCCN

QC16.P37 W64. URL http://www.springer.com/physics/book/978-
3-540-08962-9.

Hey:1996:EM

[HW96] Anthony J. G. Hey and Patrick Walters. Einstein’s mirror. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-43504-8 (hardcover), 0-521-43532-3 (paperback). xii + 291 pp. LCCN QC173.55 .H49

1996. URL http://www.zentralblattmath.org/zmath/en/search/?
an=0968.83002.

Home:2007:ESQ

[HW07a] Dipankar Home and Andrew Whitaker. Einstein’s Struggles with Quantum Theory: a Reappraisal. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-71520-7. 300 pp. LCCN QC16.E5 2007.

Home:2007:PBE

[HW07b] Dipankar Home and Andrew Whitaker. The philosophical background: Einstein and Mach. In Einstein’s Struggles with Quantum Theory: a Reappraisal [HW07a], page 300. ISBN 0-387-71520-7. LCCN QC16.E5 2007.

Iskov:2007:SDD

[IH07] Brian Iskov and Christoffer Hegnsvad, editors. Store danskere. (Danish) [Great Danes]. DR, København, Danmark, 2007. ISBN 87-7680-357-0. 173 pp. LCCN ???? DKR 299,00.

Inamura:2016:NAM

[Ina16] Takashi T. Inamura. Nagaoka’s atomic model and hyperfine interactions. Proceedings of the Japan Academy, Series B, 92(4):121–134, 2016.

Infeld:1933:NDN

[Inf33] Leopold Infeld. Nowe drogi nauki: kwanty i materja (Polish) [New pathways of science: quanta and matter], volume 2 of Z Dziedziny Nauki i Techniki. nakł. Mathesis Polskiej, Warszawa, Poland, 1933. x + 284 pp. LCCN ????

Infeld:1934:WMS

[Inf34] Leopold Infeld. The world of modern science: matter and quanta. G. P. Putnam’s Sons, New York, NY, USA, 1934. 287 pp. LCCN ???? Translation by Louis Infield of [Inf33]. Introduction by Albert Einstein.

Infeld:1978:WLC

[Inf78] Leopold Infeld. Why I left Canada: reflections on science and politics. McGill-Queen’s University Press, Montréal, Québec, Canada, 1978. ISBN 0-7735-0272-6. xii + 212 pp. LCCN QC16.I6 A3213. Translations from Polish to English by Helen Infeld. Edited with introduction and notes by Lewis Pyenson. Foreword by Alfred Schild (1921–1977).

Ishiwara:1915:UBW

[Ish15] Jun Ishiwara. Universelle Bedeutung des Wirkungsquantums. (German) [The universal meaning of the quantum of action]. Proceedings of Tokyo Mathematico-Physical Society, 8(??):106–116, ???? 1915. See also English translation [Ish17] and commentary [PB17].

Ishiwara:2017:UMQ

[Ish17] Jun Ishiwara. The universal meaning of the quantum of action. European Physical Journal H, 42(4–5):523–536, December 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https:

//link.springer.com/article/10.1140/epjh/e2017-80041-1. See

analysis [PB17].

Jacob:1985:TNB

[Jac85] M. Jacob. A tribute to Niels Bohr (on the hundredth anniversary of his birth). Report 85-17, CERN, Geneva, Switzerland, 1985. ix + 13 pp. Special colloquium held at CERN on 6 May 1985.

Jacobsen:2012:LRP

[Jac12] Anja Skaar Jacobsen. Léon Rosenfeld: physics, philosophy, and politics in the twentieth century. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2012. ISBN 981-4307-81-5 (hardcover), 981-4307-82-3 (e-book), 1-280-66920-9. xii + 354 pp.

LCCN QC16.R493 J326 2012. URL http://www.worldscientific.
com/worldscibooks/10.1142/7776.

Jaffe:1972:MNE

[Jaf72] Bernard Jaffe. Moseley and the Numbering of the Elements, volume 40 of Science study series. Heinemann, London, UK, 1972. ISBN 0-435-55073-X. x + 178 pp. LCCN QC16.M68 J3 1972.

Jahnert:2015:PCP

[Jäh15] Martin Jähnert. Practising the correspondence principle in the old quantum theory: Franck, Hund and the Ramsauer effect. In Aaserud and Kragh [AK15], pages 200–216. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Jakobsen:2013:DBB

[Jak13] Knud Jakobsen. Danskeren bag bomben: Om tømrerlærlingen fra Holstebro, der hjalp atombomben til verden og bestemte, hvor den skulle kastes. (Danish) [The Dane behind the bomb: about the carpenter’s apprentice from Holstebro who helped develop the atomic bomb and decided where it should be dropped]. Berlingske Media Forlag, København, Danmark, 2013. ISBN 87-7108-983-7. 279 pp. LCCN C774.L39 .J35 2013.

Jammer:1966:CDQ

[Jam66] Max Jammer. The conceptual development of quantum mechanics. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1966. xii + 399 pp. LCCN QC174.1 .J26.

Jammer:1967:BRN

[Jam67] Max Jammer. Book review: Niels Bohr. His Life and Work as Seen by His Friends and Colleagues. S. Rozental, Ed. North-Holland, Amsterdam; Interscience (Wiley), New York, 1967. 355 pp., illus. $9. Science (New Series), 158(3803):899, November 17, 1967. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-9203 (electronic).

Jammer:1974:BED

[Jam74a] Max Jammer. The Bohr–Einstein debate. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74d], chapter 5, pages 108–158. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:EVC

[Jam74b] Max Jammer. Early versions of the complementarity interpretation. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74d], chapter 4, pages 85– 107. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:IOL

[Jam74c] Max Jammer. The incompleteness objection and later versions of the complementarity interpretation. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74d], chapter 6, pages 159–252. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:PQM

[Jam74d] Max Jammer. The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective. Wiley, New York, NY, USA, 1974. ISBN 0-471-43958-4. xi + 536 pp. LCCN QC173.98 .J35.

Jammer:1989:CDQ

[Jam89] Max Jammer. The Conceptual Development of Quantum Mechanics, volume 12 of The History of modern physics, 1800-1950. Tomash Publishers, Los Angeles, CA, USA, second edition, 1989. ISBN 0-88318-617-9. xvii + 436 pp. LCCN QC174.12 .J36 1989.

Janssen:2017:BRH

[Jan17] Michel Janssen. Book review: Helge Kragh. Niels Bohr and the Quantum Atom: The Bohr Model of Atomic Structures, 1913–1925. Isis, 108 (1):219–220, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jeans:1919:QTN

[Jea19] James Hopwood Jeans. The quantum theory and new theories of atomic structure. Journal of the Chemical Society, 115:865–871, 1919. CODEN JCSOA9. ISSN 0368-1769 (print), 2050-5574 (electronic).

Jensen:2000:CCN

[Jen00] Carsten Jensen. Controversy and Consensus: Nuclear Beta Decay 1911– 1934, volume 24 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2000. ISBN 3-0348-9569-0 (paperback), 3-7643-5313-9 (hardcover), 3-0348-8444-3 (ebook). xv + 217 pp. LCCN QC793.5.B425 J46 2000. URL http://

www.springerlink.com/content/978-3-0348-8444-0. Carsten Jensen

died of cancer a few months after presenting his doctoral dissertation in 1990 at the University of Copenhagen. Finn Aaserud, Helge Kragh, Erik Rüdinger, and Roger H. Stuewer produced this book as a slightly edited version of that work, supplying additional figures, but leaving the prose largely untouched.

Jensen:2003:AHK

[Jen03] William B. Jensen. Ask the historian: The KLM shell labels. Journal of Chemical Education, 80(??):996–997, 2003. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Joaquim:2015:QEB

[JFEH15] Leyla Joaquim, Olival Freire, Jr., and Charbel N. El-Hani. Quantum explorers: Bohr, Jordan, and Delbrück venturing into biology. Physics in Perspective (PIP), 17(3):236–250, September 2015. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-015-0167-7.

Johnson:2013:CMG

[Joh13] Jeffrey Allan Johnson. The case of the missing German quantum chemists: On molecular models, mobilization, and the paradoxes of modernizing chemistry in Nazi Germany. Historical Studies in the Natural Sciences, 43(4):391–452, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.jstor.org/stable/

10.1525/hsns.2013.43.4.391.

Jones:1978:WWB

[Jon78] R. V. (Reginald Victor) Jones. The Wizard War: British Scientific Intelligence, 1939–1945. Coward, McCann and Geoghegan, New York, NY, USA, 1978. ISBN 0-698-10896-5. xx + 556 + 16 pp. LCCN D810.C88

J66 1978. URL https://en.wikipedia.org/wiki/Reginald_Victor_
Jones.

Jones:1985:CWL

[Jon85a] R. V. Jones. Complementarity as a way of life. In French and Kennedy [FK85], pages 320–324. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Jones:1985:MWA

[Jon85b] R. V. Jones. Meetings in wartime and after. In French and Kennedy [FK85], pages 278–287. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Jonas:1986:PCP

[Jon86] Hans Jonas. Parallelism and complementarity: The psycho-physical problem in Spinoza and in the succession of Niels Bohr. In Spinoza and the Sciences [GN86], chapter 10, pages 237–247. ISBN 94-010-8511-0, 94-009-4514-0 (e-book). ISSN 0068-0346. LCCN ???? URL https://

link.springer.com/chapter/10.1007/978-94-009-4514-2_10.

Jones:2008:QTS

[Jon08] Sheilla Jones. The Quantum Ten: a Story of Passion, Tragedy, Ambition and Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-536909-2. xii + 323 + 8 pp. LCCN QC174.12 .J66 2008.

Jones:2014:GBA

[Jon14] Derry W. Jones. The genesis of the Bohr atom, Scope: review. Level: general readership, undergraduate, postgraduate, advanced undergraduate. Contemporary Physics, 55(3):222–225, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2015:MJO

[Jon15] Alexander Jones. A Mathematician’s Journeys: Otto Neugebauer and Modern Transformations of Ancient Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-25863-X (hardcover), 3-319-25865-6 (e-book). xi + 342 pp. LCCN Q125 .M38 2016. URL http://www.loc.gov/

catdir/enhancements/fy1608/2015959787-d.html;   http://www.
loc.gov/catdir/enhancements/fy1608/2015959787-t.html.

Jahanpanah:2022:RRV

[JVK22] Jafar Jahanpanah, A. Vahedi, and H. Khosrojerdi. Relativistic rovibrational feature of electron in Bohr’s orbits of hydrogen-like atoms in Heisenberg picture. International Journal of Quantum Chemistry, 122 (14):e26911:1–e26911:??, July 15, 2022. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kamefuchi:1984:PIS

[K^{+}84] S. (Sususmu) Kamefuchi et al., editors. Proceedings of the International Symposium Foundations of Quantum Mechanics in the Light of New Technology: Central Research Laboratory, Hitachi, Ltd., Kokubunji, Tokyo, Japan, August 29–31, 1983. Physical Society of Japan, Tokyo, Japan, 1984. ISBN 4-89027-001-9. LCCN QC173.96 .I57 1983.

Khristov:1986:NBF

[K^{+}86] Khristo Iĭankov Khristov et al., editors. Nils Bor i fizikata na XX vek. Izd-vo na Bulgarskata akademiĭa na naukite, Sofía, Bulgaria, 1986. ISBN ???? 191 pp. LCCN QC16.B63 N57 1986. 3.11 lv.

Kaempffert:1939:WSW

[Kae39] Waldemar Kaempffert. This week in science: When uranium splits: Doubtful source of power; cancer and x-rays; neutron possibilities. New York Times, ??(??):D9, March 5, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/102937542.

Kaempffert:1945:SSB

[Kae45] Waldemar Kaempffert. Story of scientists’ ‘battle’ for atom bomb secret revealed in Smyth Report. New York Times, ??(??):8, August 16, 1945. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/107074385/.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/108348269/.

Kaiser:1994:BHA

[Kai94a] David Kaiser. Bringing the human actors back on stage: the personal context of the Einstein–Bohr debate. British Journal for the History of Science, 27(2):129–152, June 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/stable/

4027432.

Kaiser:1994:NBC

[Kai94b] David Kaiser. Niels Bohr’s conceptual legacy in contemporary particle physics. In Faye and Folse [FF94], chapter 11, pages 257–268. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https://

link.springer.com/chapter/10.1007/978-94-015-8106-6_11.

Kalckar:1964:FTY

[Kal64] Jørgen Kalckar. Forholdet til de yngste disciple. (Danish) [Relationship to the youngest disciples]. In Bohr and Rozental [BR64], pages 219–233. LCCN QC16.B63 N5. English translation in [Roz67].

Kalckar:1985:FQP

[Kal85] Jørgen Kalckar, editor. Foundations of Quantum Physics I (1926– 1932), volume 6 of Niels Bohr — Collected Works. Elsevier, Amsterdam, The Netherlands, 1985. ISBN 0-08-087104-6, 0-444-86712-0. 523

pp. LCCN QC3.B584. URL http://public.eblib.com/EBLPublic/
PublicView.do?ptiID=349337.

Kalckar:1996:FQP

[Kal96] Jørgen Kalckar, editor. Foundations of Quantum Physics II (1933– 1958), volume 7 of Niels Bohr — Collected Works. Elsevier, Amsterdam, The Netherlands, 1996. ISBN 0-444-89892-1, 0-08-087105-4 (e-book), 0-7204-1800-3 (set). xix + 537 pp. LCCN QC174.12 .B64 1996. URL http:

//public.eblib.com/EBLPublic/PublicView.do?ptiID=349338.

Kaloyerou:2016:CQO

[Kal16] P. N. Kaloyerou. Critique of quantum optical experimental refutations of Bohr’s principle of complementarity, of the Wootters–Zurek principle of complementarity, and of the particle–wave duality relation. Foundations of Physics, 46(2):138–175, February 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-015-9959-5.

Pihl:1964:NBOa

[Kam64] Viggo Kampmann. Niels Bohr og Risø. (Danish) [Niels Bohr and Risø [Danish Atomic Energy Research Establishment]]. In Bohr and Rozental [BR64], pages 272–280. LCCN QC16.B63 N5. English translation in [Roz67].

Kangnian:1996:NBC

[Kan96] Yan Kangnian. Niels Bohr in China. In Chinese Studies in the History and Philosophy of Science and Technology [DC96], chapter 33, pages 433–437. ISBN 90-481-4546-5, 94-015-8717-5 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 180; U390 .N19 1996. URL https://link.

springer.com/chapter/10.1007/978-94-015-8717-4_33.

Karlsch:2005:GTK

[Kar05a] Rainer Karlsch. Ein geheimnisvolles Treffen in Kopenhagen. (German) [A mysterious meeting in Copenhagen]. In Hitlers Bombe: die geheime Geschichte der deutschen Kernwaffenversuche. (German) [Hitler’s bomb: The Secret History of German Nuclear Tests] [Kar05b], pages 78–81. ISBN 3-421-05809-1. LCCN QC773.3.G3 K37 2005. URL

http://www.loc.gov/catdir/toc/fy054/2005433177.html.

Karlsch:2005:HBG

[Kar05b] Rainer Karlsch. Hitlers Bombe: die geheime Geschichte der deutschen Kernwaffenversuche. (German) [Hitler’s bomb: The Secret History of German Nuclear Tests]. Deutsche Verlags-Anstalt, München, Germany, 2005. ISBN 3-421-05809-1. 415 pp. LCCN QC773.3.G3 K37 2005. URL

http://www.loc.gov/catdir/toc/fy054/2005433177.html.

Karlsch:2006:ABH

[Kar06] Rainer Karlsch. Atomowa bomba Hitlera: historia tajnych niemieckich prób z bronia̧ ja̧drowa̧. Wojny — konflikty. Wydawn. Dolnośla̧skie, Wrocław, Poland, 2006. ISBN 83-7384-512-7. 366 pp. LCCN ???? Polish translation by Jerzy Pasieka of [Kar05b].

Karlsch:2007:BHH

[Kar07] Rainer Karlsch. La bombe de Hitler: histoire secrète des tentatives allemandes pour obtenir l’arme nucléaire. (French) [Hitler’s bomb: the secret history of German attempts to get the nuclear weapon]. Le Grand livre du mois, Paris, France, 2007. ISBN 2-286-03945-3 (hardcover).

522 pp. LCCN ???? URL http://catalogue.bnf.fr/ark:/12148/
cb41169004k.

Katz:1986:NBP

[Kat86] Ellen Laura Katz. Niels Bohr: Philosopher-Physicist (Complementarity, Quantum Mechanics). Ph.D dissertation, New York University, New York, NY, USA, October 1986. 189 pp.

Katsumori:1998:BEC

[Kat98] Makoto Katsumori. Bohr’s early complementarity argument. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 8(1):1–19, August 1998. CODEN HISCDU. ISSN 0285-4821.

Katsumori:2011:BDQ

[Kat11a] Makoto Katsumori. Bohr and the Development of Quantum Theory: a Brief Review, chapter 1, pages 1–10. Volume 286 of Boston Studies in the Philosophy of Science [Kat11b], 2011. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). LCCN QC174.17.C63 K38 2011. URL http://www.

springerlink.com/content/m25j025004858544/.

Katsumori:2011:NBC

[Kat11b] Makoto Katsumori. Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction, volume 286 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). xiv + 176 pp. LCCN QC174.17.C63 K38 2011.

Katsumori:2011:OBC

[Kat11c] Makoto Katsumori. An overview of Bohr’s complementarity. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11b], chapter 2, pages 11–37. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). LCCN QC174.17.C63 K38

2011. URL https://link.springer.com/chapter/10.1007/978-94-
007-1748-0_2.

Katsumori:2011:PHA

[Kat11d] Makoto Katsumori. A philosophical–historical analysis of complementarity. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11b], chapter 4, pages 61–88. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). LCCN

QC174.17.C63 K38 2011. URL https://link.springer.com/chapter/
10.1007/978-94-007-1748-0_4.

Katsumori:2011:PIC

[Kat11e] Makoto Katsumori. Prior interpretations of complementarity. In Niels Bohr’s Complementarity: Its Structure, History, and Intersections with Hermeneutics and Deconstruction [Kat11b], chapter 3, pages 39– 60. ISBN 94-007-1747-4, 94-007-1748-2 (e-book). LCCN QC174.17.C63

K38 2011. URL https://link.springer.com/chapter/10.1007/978-
94-007-1748-0_3.

Katzir:2015:MWB

[Kat15] Shaul Katzir. Manchester at war: Bohr and Rutherford on problems of science, war and international communication. In Aaserud and Kragh [AK15], pages 495–510. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Kaveh:2014:ICS

[Kav14] Shahin Kaveh. The incongruent correspondence: Seven non-classical years of old quantum theory. Studies in History and Philosophy of Modern Physics, ??(??):??, ???? 2014. ISSN 1355-2198 (print), 1879-

2502 (electronic). URL http://www.sciencedirect.com/science/

article/pii/S1355219814000227. Online, but not yet assigned volume and pages on 10 March 2014.

Kaye:1914:XRI

[Kay14] George William Clarkson Kaye. X-Rays: an Introduction to the Study of Röntgen Rays. ????, London, UK, 1914. ???? pp. LCCN ????

Keller:1958:CBS

[Kel58] Joseph B. Keller. Corrected Bohr–Sommerfeld quantum conditions for nonseparable systems. Annals of Physics, 4:180–188, 1958. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035X (electronic). URL http:

//www.maths.ed.ac.uk/~aar/papers/keller1.pdf;   http://www.
sciencedirect.com/science/article/pii/0003491658900320.

Keller:1983:IAP

[Kel83a] Alex Keller. The Infancy of Atomic Physics: Hercules in His Cradle. Clarendon Press, Oxford, UK, 1983. ISBN 0-19-853904-5. 230 pp. LCCN QC773 .K4 1983. US$19.50.

Keller:1983:NB

[Kel83b] Alex Keller. Niels Bohr. In The Infancy of Atomic Physics: Hercules in His Cradle [Kel83a], chapter 9, pages 147–155. ISBN 0-19-853904-5. LCCN QC773 .K4 1983. US$19.50.

Keller:1985:SM

[Kel85] Joseph B. Keller. Semiclassical mechanics. SIAM Review, 27(4):485– 504, December 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelly:2007:MPB

[Kel07] Cynthia C. Kelly, editor. The Manhattan Project: the birth of the atomic bomb by its creators, eyewitnesses, and historians. Black Dog and Leventhal Publishers, New York, NY, USA, 2007. ISBN 1-57912-747-9, 1-57912-808-4 (paperback), 1-60376-206-X (e-book). xiv + 495 pp. LCCN QC773.3.U5 M27 2007. URL http://www.loc.gov/catdir/

toc/ecip0720/2007023984.html.

Kelly:2009:MPB

[Kel09] Cynthia C. Kelly, editor. The Manhattan Project: the birth of the atomic bomb in the words of its creators, eyewitnesses, and historians. Black Dog and Leventhal, New York, NY, USA, 2009. ISBN 1-57912-808-4 (paperback), 1-57912-747-9, 1-60376-206-X (e-book). xiv + 495 pp. LCCN QC773.3.U6 M27 2009.

Kennedy:1985:DCE

[Ken85a] P. J. Kennedy. Delayed-choice experiments. In French and Kennedy [FK85], pages 148–152. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Kennedy:1985:SB

[Ken85b] P. J. Kennedy. A short biography. In French and Kennedy [FK85], pages 3–15. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Kramers:1923:ABT

[KH23] Hendrik Anthony Kramers and Helge Holst. The Atom and the Bohr Theory of its Structure: an Elementary Presentation. Gyldendal, Copenhagen, Denmark, 1923. xii + 210 pp. URL https://www.gutenberg. org/ebooks/70708. With a foreword by Sir Ernest Rutherford. Translated by Robert Bruce Lindsay and Rachel T. Lindsay from the Danish original, Bohrs Atomteori, almenfatteligt fremstillet [HK22]. Copublished in 1923 by Alfred A. Knopf, New York, NY, USA, and by Morrison and Gibb Ltd., Edinburgh, UK.

Kramers:1925:ABT

[KH25a] Hendrik Anthony Kramers and Helge Holst. Das Atom und die Bohrsche Theorie seines Baues. (German) [The Atom and Bohr’s theory of its construction]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1925. vii + 192 + 1 pp. LCCN ????

Kramers:1925:ASE

[KH25b] Hendrik Anthony Kramers and Helge Holst. El átomo y su estructura: Según la teoría de N. Bohr. (Spanish) [The atom and its structure: According to the theory of N. Bohr]. Revista de Occidente, Madrid, España, 1925. 260 pp. Translation to Spanish by Tomás Rodríguez Bachiller.

Khalatnikov:1989:LPM

[Kha89] I. M. (Isaak Markovich) Khalatnikov, editor. Landau, the physicist and the man: recollections of L. D. Landau. Pergamon Press, New York, NY, USA, 1989. ISBN 0-08-036383-0. viii + 323 + 24 pp. LCCN QC16.L25 L38 1989.

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN

QC174.1 .S66. URL http://www.amphilsoc.org/guides/ahqp/;
http://www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Kieler:2007:RFP

[Kie07] Jørgen Kieler, editor. Resistance Fighter: a Personal History of the Danish Resistance Movement, 1940–1945. Gefen Publishing House, Jerusalem, Israel, 2007. ISBN 965-229-397-0 (paperback), 965-229-991-X (e-book). xiv + 354 pp. LCCN D802.D4 K48 2007.

Konotop:2003:BSQ

[KK03] V. V. Konotop and P. G. Kevrekidis. Bohr–Sommerfeld quantization condition for the Gross–Pitaevskii equation. Physical Review Letters, 91(23):230402, December 5, 2003. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Klein:1935:PQ

[Kle35] Oscar Klein. On political quantization. Unpublished manuscript submitted to the Journal of Jocular Physics in Copenhagen, but refused for fear of political consequences, particularly in Nazi Germany. See [Hal12]., 1935.

Klein:1964:GAN

[Kle64] Oskar Klein. Glimt af Niels Bohr som forsker og tænker. (Danish) [A glimpse of Niels Bohr as researcher and philosopher]. In Bohr and Rozental [BR64], pages 71–89. LCCN QC16.B63 N5. English translation in [Roz67].

Klein:1966:TQP

[Kle66] Martin J. Klein. Thermodynamics and quanta in Planck’s work. Physics Today, 19(11):23–32, November 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.

aip.org/content/aip/magazine/physicstoday/article/19/11/10.

1063/1.3047812. Reprinted in [WP85, 294–302].

Klein:1970:FPB

[Kle70] Martin J. Klein. The first phase of the Bohr–Einstein dialogue. Historical Studies in the Physical Sciences, 2(??):iv + 1–39, ???? 1970. CO-

DEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.org/stable/
27757302.

Klein:1978:WAB

[Kle78] A. Klein. WKB approximation for bound states by Heisenberg matrix mechanics. Journal of Mathematical Physics, 19(1):292–297, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klein:1993:BRB

[Kle93] Martin J. Klein. Book review: Niels Bohr’s Times: In Physics, Philosophy, and Polity by Abraham Pais. Isis, 84(3):606–607, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235708.

Klein:2010:PEN

[Kle10] Martin J. Klein. Paul Ehrenfest, Niels Bohr, and Albert Einstein: Colleagues and friends. Physics in Perspective (PIP), 12(3):307–337, September 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.1007/
s00016-010-0025-6;   http://www.springerlink.com/content/g08710644434v387/
fulltext.pdf.

Katzir:2013:TTH

[KLR13] Shaul Katzir, Christoph Lehner, and Jürgen Renn, editors. Traditions and transformations in the history of quantum physics: HQ-3, Third International Conference on the History of Quantum Physics, Berlin, June 28–July 2, 2010, volume 5 of Max Planck research library for the history and development of knowledge. Proceedings. Edition Open Access, Berlin, Germany, 2013. ISBN 3-8442-5134-0. LCCN QC173.98. URL

http://www.edition-open-access.de/proceedings/5/.

Kockel:1959:MPF

[KMP59] B. (Bernhard) Kockel, Wilhelm Macke, and Achilles Papapetrou. Max-Planck-Festschrift 1958. Deutscher Verlag der Wissenschaften, Berlin, West Germany, 1959. 412 pp. LCCN QC3 .M45.

Kragh:2011:SGB

[KN11] Helge Kragh and Kristian Hvidtfelt Nielsen. Spreading the gospel: The Bohr atom popularised. ArXiv e-prints, December 2011. URL http://

adsabs.harvard.edu/abs/2011arXiv1112.2499K.

Knudsen:2012:PCC

[KN12] Henrik Knudsen and Henry Neilsen. Pursuing common cultural ideals: Niels Bohr, neutrality, and international scientific collaboration during the interwar period. In Lettevall et al. [LSW12], chapter 6, pages 115–

139. ISBN 0-203-11679-8, 0-415-89377-1. ISSN 1404-7586. LCCN D723

.N49 2012. URL http://libanswers.liverpool.ac.uk/faq/182315.

Kragh:2013:SGP

[KN13] Helge Kragh and Kristian Hvidtfelt Nielsen. Spreading the gospel: a popular book on the Bohr atom in its historical context. Annals of Science, 70(2):257–283, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:2001:NBJ

[KNA01] Helge Kragh, Jesper Degn Nielsen, and Finn Aaserud. Niels Bohr (1922): “Jeg ved, hvor lidt jeg har fortjent dette”. Aage Bohr og Ben Mottelson (1975): et kernekollektiv. (Danish) [Niels Bohr (1922): “I know how little I deserve this”. Aage Bohr and Ben Mottelson (1975): a nuclear collective]. In Nielsen and Nielsen [NN01], page ?? ISBN 87-7288-900-4. LCCN AS911.N9 N33 2001. DKR 398,00. URL http:/

/da.unipress.dk/udgivelser/n/nabo-til-nobel/.

Koch:1964:TAS

[Koc64] Hans Henrik Koch. Træk af et samarbejde. (Danish) [View of a collaboration]. In Bohr and Rozental [BR64], pages 301–305. LCCN QC16.B63 N5. English translation in [Roz67].

Komar:1970:QEC

[Kom70] Arthur Komar. The quantitative epistemological content of Bohr’s Correspondence Principle. Synthese, 21(1):83–92, March 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:

//link.springer.com/article/10.1007/BF00414189.

Konno:1983:SEG

[Kon83] Hiroyuki Konno. Slater’s evidence for the genesis of the Bohr–Kramers– Slater theory. Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):39–52, September 1983. CO-DEN HISCDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/

110009837441.

Konno:2000:BSQ

[Kon00] Hiroyuki Konno. Bohr’s search for the quantum theory of dispersion: the number of dispersion electrons, absorption and emission of light and the oscillator model. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(2):163– 176, November 2000. CODEN HISCDU. ISSN 0285-4821. Discovery of energy quanta and development of early quantum theory.

Konno:2002:RDL

[Kon02] Hiroyuki Konno. Ritz’s discovery of the Lyman series before 1913 and Lyman’s indifference to the Bohr theory. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 44(1–2):127– 139, July 2002. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Konitzer:2015:BSE

[Kon15] Franziska Konitzer. Bohr schlägt Einstein im Doppelspalt: mit Sauerstoff-Molekülen haben Forscher ein Gedankenexperiment von Albert Einstein realisiert und die spukhaften Ergebnisse der Quantenwelt erneut bestätigt. (German) [Bohr proposes Einstein in the double slit: with oxygen molecules, researchers have realized a thought experiment by Albert Einstein and reconfirmed the spooky results of the quantum world]. Bild der Wissenschaft, 52:40–41, 2015. ISSN 0006-2375.

Kossel:1923:BBA

[Kos23] W. Kossel. Die Beziehungen der Bohrschen Atomtheorie zur Deutung chemischer Vorgänge. (German) [The relations of the Bohr atomic theory to the interpretation of chemical procedures]. Die Naturwissenschaften, 11(27):598–604, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Kothari:1985:CPE

[Kot85] D. S. Kothari. The complementarity principle and eastern philosophy. In French and Kennedy [FK85], pages 325–331. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Kowarski:1953:HAN

[Kow53] Lew Kowarski. Hitting the atomic nucleus. UNESCO Courier, 6(12): 3–4, December 1953. ISSN 0041-5278. URL http://unesdoc.unesco.

org/images/0007/000708/070862eo.pdf.

Kallen:1955:MSD

[KP55] Gunnar Källén and Wolfgang Pauli. On the mathematical structure of T. D. Lee’s model of a renormalizable field theory: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(7) of Det Kongelige Danske videnskabernes selskab. Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 23 pp. LCCN AS281 .D215 bd. 30, no. 7.

Kramers:1923:KSA

[Kra23a] H. A. Kramers. Das Korrespondenzprinzip und der Schalenbau des Atoms. (German) [The Correspondence Principle and the scaling of atomic structure]. Die Naturwissenschaften, 11(27):550–559, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Kratzer:1923:BMG

[Kra23b] A. Kratzer. Bandenspektren und molekülmodelle. (German) [Band spectra and molecular models]. Die Naturwissenschaften, 11(27):577– 584, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Kramers:1924:LDB

[Kra24] H. A. Kramers. The law of dispersion and Bohr’s theory of spectra. Nature, 113(2845):673–674, May 10, 1924. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/articles/113673a0.

Kramers:1925:OVM

[Kra25] H. A. Kramers. Om Vekselvirkningen mellem Lys og Stof. (Danish) [On the interaction between light and matter]. Fysisk Tidsskrift, 23(??): 26–40, ???? 1925.

Kragh:1976:END

[Kra76] Helge Kragh. Elements no. 70, 71 and 72: Discoveries and controversies. In Evans [Eva96], chapter 5, pages 67–89. ISBN 0-7923-4101-5. LCCN QD172.R2 E65 1996. URL http://www.

loc.gov/catdir/enhancements/fy1006/96210275-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1006/96210275-t.html.

Kragh:1979:NBS

[Kra79] Helge Kragh. Niels Bohr’s second atomic theory. Historical Studies in the Physical Sciences, 10(??):123–186, ???? 1979. CODEN HSPSAS.

ISSN 0073-2672. URL http://www.jstor.org/stable/27757389.

Kragh:1985:TPS

[Kra85] Helge Kragh. The theory of the periodic system. In French and Kennedy [FK85], pages 50–67. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Kragh:1996:BRBa

[Kra96] Helge Kragh. Book review: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Volume 3: 1940–1949 by Wolfgang Pauli; Karl von Mayenn. Isis, 87(1):196, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/235799.

Kragh:1997:JJT

[Kra97] Helge Kragh. J. J. Thomson, the electron, and atomic architecture. The Physics Teacher, 35(6):328–332, September 1997. CODEN PH-TEAH. ISSN 0031-921X (print), 1943-4928 (electronic). URL http://

tpt.aapt.org/resource/1/phteah/v35/i6/p328_s1.

Kragh:2006:FSE

[Kra06] Helge Kragh. The first subatomic explanations of the periodic system. Foundations of Chemistry, 3(2):129–143, 2006. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic).

Kragh:2007:BRM

[Kra07] Helge Kragh. Book review: Matthias Dörries (ed.), Michael Frayn’s Copenhagen in Debate: Historical Essays and Documents on the 1941 Meeting between Niels Bohr and Werner Heisenberg. Berkeley Papers in History of Science Vol. 20. Berkeley, CA: Office for History of Science and Technology, 2005. Pp. viii + 195. ISBN 0-9672617-2-4. $12.00 (paperback). British Journal for the History of Science, 40(1):115–116, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4500686.

Kragh:2009:BKS

[Kra09] Helge Kragh. Bohr–Kramers–Slater theory. In Greenberger et al. [GHW09], pages 62–64. ISBN 3-540-70626-7, 3-540-70622-4. LCCN C174.12 .C66 2009.

Kragh:2010:BBT

[Kra10a] Helge Kragh. Before Bohr: Theories of atomic structure 1850–1913. Report RePOSS #10, Centre for Science Studies, University of Aarhus, Denmark. Research group: History and philosophy of science, Århus,

Denmark, October 2010. URL http://css.au.dk/fileadmin/reposs/
reposs-010.pdf.

Kragh:2010:ERB

[Kra10b] Helge Kragh. The early reception of Bohr’s atomic theory (1913– 1915). A preliminary investigation. Report RePOSS #9, Centre for Science Studies, University of Aarhus, Denmark. Research group: History and philosophy of science, Århus, Denmark, July 2010. URL http://

css.au.dk/fileadmin/reposs/reposs-009.pdf.

Kragh:2011:COB

[Kra11a] Helge Kragh. Conceptual objections to the Bohr atomic theory — do electrons have a “free will”? European Physical Journal H, 36(3): 327–352, November 2011. CODEN EPJHAD. ISSN 2102-6459 (print),

2102-6467 (electronic). URL http://link.springer.com/article/
10.1140/epjh/e2011-20031-x.

Kragh:2011:QSV

[Kra11b] Helge Kragh. Quantenspringerei: Schrödinger vs. Bohr. Report Re-POSS #14, Centre for Science Studies, University of Aarhus, Denmark. Research group: History and philosophy of science, Århus, Denmark,

February 2011. URL http://css.au.dk/fileadmin/reposs/reposs-
014.pdf.

Kragh:2011:RBA

[Kra11c] Helge Kragh. Resisting the Bohr atom: The early British opposition. Physics in Perspective (PIP), 13(1):4–35, March 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http:/

/link.springer.com/article/10.1007/s00016-010-0048-z.

Kragh:2012:NBQ

[Kra12a] Helge Kragh. Niels Bohr and the Quantum Atom: The Bohr Model of Atomic Structure 1913–1925. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-965498-0 (hardcover), 0-19-163045-4 (e-book), 0-19-174169-8 (e-book). vi + 410 pp. LCCN QC774.A2.

Kragh:2012:RRA

[Kra12b] Helge Kragh. Rutherford, radioactivity, and the atomic nucleus. ArXiv e-prints, February 2012. URL http://adsabs.harvard.edu/

abs/2012arXiv1202.0954K.

Kragh:2013:BAF

[Kra13a] Helge Kragh. Bohrs atomteori fylder 100 år. (Danish) [Bohr’s atomic theory is 100 years old]. Aktuel Naturvidenskab, 1(??):10–13, ???? 2013.

URL http://www.royalacademy.dk/Files/Billeder/Niels%20Bohr%
202013/2013%20AN%20Bohr_100_%C3%A5r.pdf.

Kragh:2013:NBB

[Kra13b] Helge Kragh. Niels Bohr between physics and chemistry. Physics Today, 66(5):36–41, May 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Kragh:2013:QDB

[Kra13c] Helge Kragh. A quantum discontinuity: the Bohr–Schrödinger dialogue. In Reiter and Yngvason [RY13], pages 135–152. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Krause:2014:BAM

[Kra14a] Michael Krause. Bohr’s atomic model. In CERN: how we found the Higgs boson [Kra14b], chapter 5, page ?? ISBN 981-4623-55-5 (hardcover), 981-4623-46-6 (paperback), 981-4623-48-2 (e-book). LCCN QC793 .K73 2014.

Krause:2014:CHW

[Kra14b] Michael Krause. CERN: how we found the Higgs boson. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2014. ISBN 981-4623-55-5 (hardcover), 981-4623-46-6 (paperback), 981-4623-48-2 (e-book). xiii + 243 pp. LCCN QC793 .K73 2014.

Krause:2014:DTR

[Kra14c] Michael Krause. Dalton, Thomson, Rutherford, Bohr. In CERN: how we found the Higgs boson [Kra14b], chapter 5, page ?? ISBN 981-4623-55-5 (hardcover), 981-4623-46-6 (paperback), 981-4623-48-2 (ebook). LCCN QC793 .K73 2014.

Kragh:2015:RBS

[Kra15a] Helge Kragh. From Ørsted to Bohr:the sciences and the Danish university system, 1800–1920. In Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes [SDG15], chapter 3, pages 31–47. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346. LCCN Q183.4.E85 2015. URL https://link.springer.

com/chapter/10.1007/978-94-017-9636-1_3.

Kragh:2015:I

[Kra15b] Helge Kragh. Introduction. In Aaserud and Kragh [AK15], pages 13–26. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/2015/
06/Bohr-Book.pdf.

Kragh:2015:MFB

[Kra15c] Helge Kragh. The many faces of the Bohr atom. In Aaserud and Kragh [AK15], pages 95–110. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Kragh:2016:BQP

[Kra16a] Helge Kragh. Bohr’s quantum philosophy. Science & Education (Springer), 25(7–8):937–938, October 2016. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Kragh:2016:BRF

[Kra16b] Helge Kragh. Book review: Finn Aaserud and John Heilbron, Love, Literature, and the Quantum Atom: Bohr’s 1913 Trilogy Revisited. Oxford: Oxford University Press, 2013, x + 284 pages, $61 (hardcover). Physics in Perspective (PIP), 18(3):351–354, September 2016. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http:/

/link.springer.com/article/10.1007/s00016-016-0188-x.

Kragh:2017:LSS

[Kra17] Helge Kragh. ‘Let the stars shine in peace!’ Niels Bohr and stellar energy, 1929–1934. Annals of Science, 74(2):126–148, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kuhn:1962:INB

[KRBR62] Thomas S. Kuhn, Léon Rosenfeld, Aage Bohr, and Erik Rudinger. [interview with Niels Bohr [17 November 1962]]. Web docu-

ment, 1962. URL https://www.aip.org/history-programs/oral-
histories/search?search_api_views_fulltext=Niels+Bohr. Recorded

one day before the death of Niels Bohr.

Krige:1991:BRF

[Kri91] John Krige. Book review: Finn Aaserud. Redirecting Science: Niels Bohr, Philanthropy and the Rise of Nuclear Physics. Cambridge: Cambridge University Press, 1990. Pp. xiii + 356. ISBN 0-521-35366-1. £35.00, $47.50. British Journal for the History of Science, 24(4):475– 476, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-

001X (electronic). URL http://www.jstor.org/stable/4027136.

Krips:1994:CBL

[Kri94] Henry Krips. A critique of Bohr’s local realism. In Faye and Folse [FF94], chapter 12, pages 269–277. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153

1994; QC16 1994. URL https://link.springer.com/chapter/10.
1007/978-94-015-8106-6_12.

Kleint:2005:WHB

[KRW05] Christian Kleint, Helmut Rechenberg, and Gerald Wiemers, editors. Werner Heisenberg, 1901–1976: Beiträge, Berichte, Briefe: Festschrift zu seinem 100. Geburtstag. (German) [Werner Heisenberg, 1901–1976: Contributions, Reports, and Letters: Celebration of his 100th Birthday], volume 62 of Abhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Naturwissenschaftliche Klasse; Aufsatzsammlung. Verlag der Sächsischen Akademie der Wissenschaften zu Leipzig, Leipzig, Germany, 2005. ISBN 3-7776-1402-5. ISSN 0365-6470. 424 pp. LCCN QC16.H35 W474 2005. URL http://www.loc.

gov/catdir/toc/fy0613/2006402966.html.

Kubbinga:2008:TNB

[Kub08] Henk Kubbinga. A tribute to Niels Bohr. Europhysics News, 39(5): 32–34, September/October 2008. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic). URL http://www.europhysicsnews.

org/articles/epn/abs/2008/05/epn080506/epn080506.html.

Kubbinga:2009:ERB

[Kub09] Henk Kubbinga. Essay review: Niels Bohr’s Collected Works, Volumes 11 and 12, edited by Finn Aaserud. Isis, 100(1):119–126, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/597565.

Kuhn:1967:RTR

[Kuh67] Thomas S. Kuhn. Review: The turn to recent science: The Questioners: Physicists and the Quantum Theory by Barbara Lovett Cline Thirty Years That Shook Physics: The Story of Quantum Theory by George Gamow; The Conceptual Development of Quantum Mechanics by Max Jammer; Korrespondenz, Individualitat, und Komplementaritat: Eine Studie zur Geistesgeschichte der Quantentheorie in den Beitragen Niels Bohrs by Klaus Michael Meyer-Abich; Niels Bohr: The Man, His Science, and the World They Changed by Ruth Moore; Sources of Quantum Mechanics by B. L. Van der Waerden. Isis, 58(3):409–419, ???? 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228002;   http://www.jstor.
org/stable/pdfplus/228002.pdf.

Kumar:2010:QEB

[Kum10] Manjit Kumar. Quantum: Einstein, Bohr and the great debate about the nature of reality. W. W. Norton & Co., New York, NY, USA, 2010. ISBN 0-393-07829-9 (hardcover). xvi + 448 + 16 pp. LCCN QC173.98 .K86 2010; QC173.98.K86.

Kumar:2011:GRP

[Kum11] Manjit Kumar. Le grand roman de la physique quantique: Einstein, Bohr et le dêbat sur la nature de la réalité. (French) [Quantum: Einstein, Bohr and the great debate about the nature of reality]. J. C. Lattés, Paris, France, 2011. ISBN 2-7096-2465-6. 524 + 16 pp. LCCN ????

Kalckar:1985:IBA

[KvG85] Jørgen Kalckar and Johann Wolfgang von Goethe. Det inkommensurable: brudstykker af et tonedigt in d-Moll: basso ostinato, Goethetemaer fra vekselsange med Niels Bohr. (Danish) [The incommensurable: pieces of a composition in D-flat: basso ostinato, Goethe themes from antiphonies with Niels Bohr]. Rhodos, København, Danmark, 1985. ISBN 87-7245-090-8. 158 pp. LCCN PT2177 .K35 1985.

Krafft:1988:KLA

[KZN^{+}88] F. Krafft, H.-D. Zeh, H. P. Nilles, J. A. Maruhn, H. Rechenberg, F. Pobell, and A. Hofzumahaus. Kerner: Lise, Atomphysikerin. Die Lebensgeschichte der Lise Meitner/Hawking: Eine kurze Geschichte der Zeit: Die Suche nach der Urkraft des Universums/Davies u. Brown: Superstrings: A Theory of Everything?/Hasse u. Myers: Geometrical Relationships of Macroscopic Nuclear Physics/Thorsen: Niels Bohr Collected Works, Vol. 8/Richardson u. Smith: Experimental Techniques in Condensed Matter Physics at Low Temperatures/Jaenicke: Ergebnisse aus dem gleichnamigen Sonderforschungsbereich. (German) [Kerner: Lise, nuclear physicist. The Life Story of Lise Meitner / Hawking: A Brief History of Time: The Quest for the Primal Power of the Universe/ Davies & Brown: Superstrings: A Theory of Everything?. Myers: Geometrical Relationships of Macroscopic Nuclear Physics/Thorsen: Niels Bohr Collected Works, Vol. 8/Richardson & Smith: Experimental Techniques in Condensed Matter Physics at Low Temperatures/Jaenicke: Results from the eponymous special research area]. Physik Journal, 44(12): 466–468, December 1988. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Lakhtakia:1996:MMH

[Lak96] A. (Akhlesh) Lakhtakia, editor. Models and Modelers of Hydrogen: Thales, Thomson, Rutherford, Bohr, Sommerfeld, Goudsmit, Heisenberg, Schrödinger, Dirac, Sallhofer. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2302-1. xv + 424 pp. LCCN QC171.2 .M63 1996. URL http:/

/www.worldscientific.com/worldscibooks/10.1142/2788.

Landau:1955:QTF

[Lan55] Lev D. Landau. On the quantum theory of fields. In Pauli et al. [PRW55], pages 52–69. LCCN QC71 .P3 1955.

Lande:1959:BRB

[Lan59] Alfred Landé. Book review: Atomic Physics and Human Knowledge by Niels Bohr. Philosophy of Science, 26(2):150–153, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http:

//www.jstor.org/stable/185263.

Landea:1972:EQG

[Lan72] Alfred Landea. Einheit in der Quantenwelt (Gegen den Bohr– Heisenberg’schen Positivismus). (German) [Unity in the quantum world (through Bohr–Heisenberg positivism)]. Dialectica: International Review of Philosophy of Knowledge, 26(2):115–130, June 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lanouette:1994:AS

[Lan94] William Lanouette. Atomic spies. Debate with the authors of Special Tasks by Soviet spymaster Pavel Sudoplatov [SSSS95], which falsely alleged that Szilard, Niels Bohr, J. Robert Oppenheimer and Enrico Fermi were Soviet agents within the Manhattan Project. MacNeil/Lehrer NewsHour, April 26, 1994.

Landsman:2006:WCM

[Lan06] N. P. Landsman. When champions meet: Rethinking the Bohr–Einstein debate. Studies in History and Philosophy of Modern Physics, 37(1): 212–242, March 2006. CODEN ???? ISSN 1355-2198 (print), 1879-

2502 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000869.

Landsman:2008:BRM

[Lan08] N. P. Landsman. Book review: Matthias Dörries (Ed.), Michael Frayn’s Copenhagen in Debate: Historical Essays and Documents on the 1941 Meeting Between Niels Bohr and Werner Heisenberg. Office for History of Science and Technology, University of California, Berkeley, ISBN 0-9672617-2-4, 2005 (viii + 195 pp., $12.00 pbk). Studies in History and Philosophy of Modern Physics, 39(2):462–464, May 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219807001025. See

[Dör05].

Lassen:1955:TCE

[Las55] N. O. Lassen. Total charges and electron capture cross-sections of fission fragments in gases: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(8) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 12 pp. LCCN AS281.

Latimer:1983:TBT

[Lat83] Colin J. Latimer. Teaching Bohr theory. Physics Education, 18(2): 86–90, 1983. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552

(electronic). URL http://stacks.iop.org/0031-9120/18/i=2/a=312.

Laurence:1933:JHM

[Lau33] William L. Laurence. Jekyll–Hyde mind attributed to man: ‘complementarity,’ new theory of knowledge, is presented by Prof. Niels Bohr. All things dual in aspect we can know only one at any one time — scientists hail theory as revolutionary. New York Times, ??(??):1, 13, June 23, 1933. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/100684498/.

Laurence:1951:DHR

[Lau51] William L. Laurence. Day of Hiroshima recalls atom race: Two leaders in atomic work at Columbia University. New York Times, ??(??):3, August 6, 1951. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X,

1553-8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/112188459/fulltextPDF.

Laubichler:2007:BRM

[Lau07] Manfred D. Laubichler. Book review: Matthias Dörries: Michael Frayn’s Copenhagen in Debate: Historical Essays and Documents on the 1941 Meeting between Niels Bohr and Werner Heisenberg. Isis, 98(2): 401–402, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/10.1086/521471.

Laurence:2013:FMS

[Lau13] William L. Laurence. Fermi measures speed of neutron / Bohr and Einstein at odds. In Dean [Dea13], page ?? ISBN 1-4027-9320-0 (hardcover). LCCN QC7 .D43 2013. Foreword by Neil deGrasse Tyson.

Loeffler:1999:BNK

[LDN99] Martha Loeffler, Knud Dyby, and John Mark Nielsen. Boats in the night: Knud Dyby’s involvement in the rescue of the Danish Jews and the Danish resistance. Lur Publications, Danish Immigrant Archive, Dana College, Blair, NE, USA, 1999. ISBN 0-930697-06-5. xviii + 140 pp. LCCN D804.66.D93 L64 1999.

Lee:2006:BVB

[Lee06] Jeongmin Lee. Bohr vs. Bohm: Interpreting quantum theory through the philosophical tradition. Ph.D. dissertation, Indiana University, Bloomington, IN, USA, 2006. 248 pp.

Lee:2007:BPC

[Lee07] Sabine Lee. The Bethe–Peierls correspondence. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007. ISBN 981-277-135-2. xi + 506 pp. LCCN QC16.B46 A4 2007eb. URL

http://site.ebrary.com/lib/yale/Doc?id=10255467;   http://
www.loc.gov/catdir/toc/fy0805/2008274437.html;   http://www.
worldscientific.com/worldscibooks/10.1142/6595.

Lee:2007:SRP

[Lee09] Sabine Lee, editor. Sir Rudolf Peierls: selected private and scientific correspondence. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007–2009. ISBN 981-256-503-5 (vol. 1), 981-279-706-8 (vol. 2). xiv + 866 (volume 1), xii + 1108 (volume 2) pp. LCCN QC16.P375 A4 2007. URL http://

wspc-prod.literatumonline.com/worldscibooks/10.1142/6790;
http://www.worldscientific.com/worldscibooks/10.1142/5941.

Lemaitre:1931:BWP

[Lem31] G. Lemaître. The beginning of the world from the point of view of quantum theory. Nature, 127(3210):706, May 9, 1931. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Lenzen:1948:BRB

[Len48a] V. F. Lenzen. Book review: L’évolution de la notion de phénomène physique, des primitifs à Bohr et Louis de Broglie by Jean Pelseneer. Isis, 39(3):194–196, August 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226331.

Lenzen:1948:BRL

[Len48b] V. F. Lenzen. Book review: L’évolution de la notion de phénomène physique, des primitifs à Bohr et Louis de Broglie by Jean Pelseneer. Isis, 39(3):194–196, August 1948. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/226331.

Leskov:1994:SFU

[Les94] Sergei Leskov. The Sudoplatov file: An unreliable witness. Bulletin of the Atomic Scientists, 50(4):33–36, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levidow:1986:SP

[Lev86] Les Levidow, editor. Science as politics, volume 20 of Radical science series, 0305-0963. Free Association Books, London, UK, 1986. ISBN 0-946960-49-6. 180 pp. LCCN Q175.5 .S349 1986.

Levitt:1995:BRJ

[Lev95] Norman Levitt. Book review: James T. Cushing, Quantum Mechanics: Historical Contingency and the Copenhagen Hegemony. Physics Today, 48(11):84–85, 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/v48/i11/
p84/s1.

Levine:2000:DDD

[Lev00] Ellen Levine. Darkness over Denmark: the Danish Resistance and the Rescue of the Jews. Holiday House, New York, NY, USA, 2000. ISBN 0-8234-1447-7. x + 164 pp. LCCN DS135.D4 L4 2000.

Levi:1955:QBT

[LH55] Hilde Levi and Anne S. Hogben. Quantitative beta track autoradiography with nuclear track emulsions: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(9) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 23 pp. LCCN AS281.

Liboff:1975:BCP

[Lib75] Richard L. Liboff. Bohr correspondence principle for large quantum numbers. Foundations of Physics, 5(2):271–293, June 1975. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http:/

/link.springer.com/article/10.1007/BF00717443.

Lightman:2005:DGBb

[Lig05] Alan P. Lightman. The discoveries: great breakthroughs in twentiethcentury science. Pantheon, New York, NY, USA, 2005. ISBN 0-676-97789-8, 0-375-42168-8 (hardcover). xviii + 553 + 16 pp. LCCN

Q180.55.D57 .L53. URL http://www.loc.gov/catdir/enhancements/
fy0623/2005040854-b.html;   http://www.loc.gov/catdir/enhancements/
fy0623/2005040854-d.html;   http://www.loc.gov/catdir/enhancements/
fy0623/2005040854-t.html.

Lindsay:1924:QNB

[Lin24] R. Bruce Lindsay. The quantum numbers of the Bohr orbits in the alkali atoms. Science (New Series), 60(1560):475–476, November 21, 1924. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Lindblad:1955:AMC

[Lin55] J. Lindblad. To all members of CERN: Standardization of papers. Journal of Jocular Physics, III:10–13, October 7, 1955.

Lindsay:1964:BRB

[Lin64] R. B. Lindsay. Book review: Essays 1958/1962 on Atomic Physics and Human Knowledge, by Niels Bohr. Physics Today, 17(12):55, December 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/phtoad/
v17/i12/p55_s1.

Lindsay:1981:BRB

[Lin81] R. Bruce Lindsay. Book review: The Early Years: The Niels Bohr Institute 1921–1930, by Peter Robertson. Physics Today, 34(6):??, June 1981. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [Rob79].

Lindley:2007:UEH

[Lin07] David Lindley. Uncertainty: Einstein, Heisenberg, Bohr, and the struggle for the soul of science. Doubleday, Garden City, NY, USA, 2007. ISBN 0-385-51506-5. vii + 257 pp. LCCN QC174.17.H4 L56 2007. URL

http://www.loc.gov/catdir/toc/ecip0614/2006017029.html.

Lindley:2008:UEH

[Lin08] David Lindley. Uncertainty: Einstein, Heisenberg, Bohr, and the struggle for the soul of science. Anchor Books, New York, NY, USA, 2008. ISBN 1-4000-7996-9. viii + 257 pp. LCCN QC174.17.H4 L56 2007. URL

http://www.loc.gov/catdir/toc/ecip0614/2006017029.html.

Lahti:1985:SFM

[LM85] Pekka (Pekka Johannes) Lahti and Peter Mittelstaedt, editors. Symposium on the Foundations of Modern Physics: 50 Years of the Einstein– Podolsky–Rosen Gedankenexperiment, Joensuu, Finland, 16–20 June 1985. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1985. ISBN 9971-5-0004-3. LCCN QC173.96 .S96 1985.

Loder:1996:BBD

[LN96] James E. Loder and W. Jim Neidhardt. Barth, Bohr, and dialectic. In Richardson and Wildman [RW96], pages 271–290. ISBN 0-415-91666-6 (hardcover), 0-415-91667-4 (paperback). LCCN BL240.2 .R43

1996. URL http://catdir.loc.gov/catdir/enhancements/fy0651/
95047045-d.html.

Logan:2000:RSN

[Log00] Jonothan Logan. Review: “a strange new quantum ethics” [Copenhagen. Michael Frayn. 132 pp. Methuen Publishing Ltd., 2000]. American Scientist, 88(4):356–359, July/August 2000. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL https://www.

jstor.org/stable/27858060.

Losee:1961:CMP

[Los61] John Price Losee, Jr. A Comparison of Methodological Principles Basic to the Quantum Mechanics of Bohr and Heisenberg, the Metaphysics of Emmet, and the Theology of Tillich. Ph.D dissertation, Drew University, Madison, NJ, USA, November 1961. 253 pp.

Losee:2011:TCA

[Los11] John Losee. Theories of causality: from antiquity to the present. Transaction Publishers, New Brunswick, NJ, USA, 2011. ISBN 1-4128-1832-X. viii + 210 pp. LCCN BD531 .L67 2011.

Ladenburg:1923:AZD

[LR23] R. Ladenburg and F. Reiche. Absorption, Zerstreuung und Dispersion in der Bohrschen Atomtheorie. (German) [Absorption, scattering and dispersion in the Bohr atomic theory]. Die Naturwissenschaften, 11(27): 584–598, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Lorentz:1923:AER

[LRdB^{+}23] H. A. Lorentz, E. Rutherford, M. de Broglie, R. A. Millikan, H. Kamerlingh Onnes, P. Weiss, L. Brillouin, W. H. Bragg, W. J. de Haas, N. Bohr, and P. Ehrenfest, editors. Atomes et Électrons: Rapports et discussions du conseil de physique tenu à Bruxelles du 1er au 6 avril 1921. (French) [Atoms and electrons: reports and discussions of the physics meeting held in Brussels from 1st to 6th April, 1921]. Gauthier-Villars et cie, Paris, France, 1923. vii + 271 + i pp. LCCN QC1 .I6 1921. Proceedings of the Solvay III international conference.

Lettevall:2012:NTC

[LSW12] Rebecka Lettevall, Geert Somsen, and Sven Widmalm, editors. Neutrality in twentieth-century Europe: intersections of science, culture, and politics after the First World War, volume 18 of Routledge studies in cultural history. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2012. ISBN 0-203-11679-8, 0-415-89377-1. ISSN 1404-7586. ix +

351 pp. LCCN D723 .N49 2012. URL http://libanswers.liverpool.
ac.uk/faq/182315.

Lucas:2007:RFH

[Luc07] Amand A. Lucas. Revisiting Farm Hall. Europhysics News, 38(4):25–29, July/August 2007. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-

1092 (electronic). URL http://www.europhysicsnews.org/articles/
epn/abs/2007/04/epn07402/epn07402.html. See comment and reply

[BL08].

Lucas:2011:BS

[Luc11] Amand Lucas. The Bomb and the Swastika. CreateSpace Independent Publishing, ????, 2011. 68 pp. URL http://www.amazon.com/The-

Bomb-Swastikahistorys-scientists/dp/1466426675. A play in four

acts.

Meyer-Abich:1965:KIK

[MA65] Klaus Michael Meyer-Abich. Korrespondenz, Individualität und Komplementarität: eine Studie zur Geistesgeschichte der Quantentheorie in den Beiträgen Niels Bohrs, volume 5 of Boethius. Steiner, Wiesbaden, West Germany, 1965. xi + 209 pp. LCCN ???? Based on the author’s Ph.D. dissertation, Hamburg (1964).

Maleeh:2013:PBC

[MA13] Reza Maleeh and Parisa Amani. Pragmatism, Bohr, and the Copenhagen interpretation of quantum mechanics. International Studies in the Philosophy of Science, 27(4):353–367, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/02698595.2013.868182.

MacKinnon:1985:BFQ

[Mac85] Edward MacKinnon. Bohr on the foundations of quantum theory. In French and Kennedy [FK85], pages 101–120. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

MacKinnon:1986:BRB

[Mac86] Edward MacKinnon. Book review: The Philosophy of Niels Bohr: The Framework of Complementarity by Henry J. Folse. Philosophy of Science, 53(3):458–459, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.

org/stable/187685.

Mackinnon:1994:BRD

[Mac94] Edward Mackinnon. Bohr and the realism debates. In Faye and Folse [FF94], chapter 13, pages 279–302. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153

1994; QC16 1994. URL https://link.springer.com/chapter/10.
1007/978-94-015-8106-6_13.

MacKinnon:1995:BRB

[Mac95] Edward MacKinnon. Book review: Atoms, Metaphors, and Paradoxes: Niels Bohr and the Construction of a New Physics by Sandro Petruccioli. Isis, 86(3):515, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235084.

MacKinnon:1996:C

[Mac96] Edward MacKinnon. Complementarity. In Richardson and Wildman [RW96], pages 255–270. ISBN 0-415-91666-6 (hardcover), 0-415-91667-4 (paperback). LCCN BL240.2 .R43 1996. URL http://catdir.loc.

gov/catdir/enhancements/fy0651/95047045-d.html.

Mackintosh:1997:CE

[Mac97] A. R. Mackintosh. The crocodile and the elephant. Notes and Records of the Royal Society of London, 51(2):309–316, July 22, 1997. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Maiocchi:1993:BRD

[Mai93] Roberto Maiocchi. Book review: David Favrholdt, Niels Bohr’s philosophical background, Copenhagen, the Royal Danish Academy of Sciences and Letters, 1992, 147 pp. Nuncius, 8(1):367–370, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00352.

Malhotra:1970:BRB

[Mal70] M. K. Malhotra. Book review: Atomphysik und menschliche Erkenntnis II. Aufsätze und Vorträge aus den Jahren 1958–1962, by Niels Bohr.

Zeitschrift für philosophische Forschung, 24(1):157–158, January/March

1970. URL http://www.jstor.org/stable/20481844.

Maleeh:2015:BPL

[Mal15] Reza Maleeh. Bohr’s philosophy in the light of Peircean pragmatism. Journal for General Philosophy of Science = Zeitschrift für allgemeine Wissenschaftstheorie, 46(1):3–21, April 2015. CO-DEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL

http://link.springer.com/article/10.1007/s10838-014-9274-4;
 https://www.math.utah.edu/pub/bibnet/authors/b/bohr-niels.
bib.

Margenau:1954:ADV

[Mar54] Henry Margenau. Advantages and disadvantages of various interpretations of the quantum theory. Physics Today, 7(10):6–13, October 1954. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Margenau:1958:PPC

[Mar58] Henry Margenau. Philosophical problems concerning the meaning of measurement in physics. Philosophy of Science, 25(1):23–33, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185334.

Margenau:1963:MQSa

[Mar63a] Henry Margenau. Measurements and quantum states: Part I. Philosophy of Science, 30(1):1–16, January 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.

org/stable/186615.

Margenau:1963:MQSb

[Mar63b] Henry Margenau. Measurements and quantum states: Part II. Philosophy of Science, 30(2):138–157, April 1963. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.

org/stable/186421.

Marshak:1988:PHB

[Mar88] Robert E. Marshak. The pragmatic humanism of Bohr, Einstein and Sakharov. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 132(3):268–275, September 1988. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL http://www.jstor.org/stable/3143854.

Massimi:2004:WDI

[Mas04] M. Massimi. What demonstrative induction can do against the threat of underdetermination: Bohr, Heisenberg, and Pauli on spectroscopic anomalies (1921–24). Synthese, 140(3):243–277, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3ASYNT.
0000031319.64615.49.pdf.

Millikan:1924:SCS

[MB24] R. A. Millikan and I. S. Bowen. Some conspicuous successes of the Bohr atom and a serious difficulty. Physical Review, 24(3):223–228, September 1924. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.24.223.

Montwill:2011:NBI

[MB11a] Alex Montwill and Ann Breslin. Niels Bohr introduces the quantum into atomic physics. In The quantum adventure: does God play dice? [MB11b], chapter 7, pages 95–108. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (e-book). LCCN QC174.12 .M66 2012.

Montwill:2011:QAD

[MB11b] Alex Montwill and Ann Breslin. The quantum adventure: does God play dice? World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (e-book). ix + 248 pp. LCCN QC174.12 .M66 2012.

McCrea:1934:BRB

[McC34] W. H. McCrea. Book review: Atomic Theory and the Description of Nature, by N. Bohr. Mathematical Gazette, 18(230):279–280, October 1934. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL http://www.jstor.org/stable/3605384.

McEvoy:1972:PNB

[McE72] Paul Poynter McEvoy. The Philosophy of Niels Bohr. Ph.D dissertation, Massachusetts Institute of Technology, Cambridge, MA, USA, 1972.

McKaughan:2005:INB

[McK05] Daniel J. McKaughan. The influence of Niels Bohr on Max Delbrück: revisiting the hopes inspired by “light and life”. Isis, 96(4):507–529, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McMillan:1994:SFF

[McM94] Priscilla Johnson McMillan. The Sudoplatov file: Flimsy memories. Bulletin of the Atomic Scientists, 50(4):30–33, July/August 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McWeeny:1973:NUA

[McW73] Roy McWeeny. Natural units in atomic and molecular physics. Nature, 243(5404):196–198, May 25, 1973. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

McDayter:1967:GBB

[MD67] Walt McDayter and Norman Drew. The giants: The bomb builders. Denver Post, ??(??):??, February 3, 1967. URL http://library.ucsd. edu/dc/object/bb0103915g. This is a reasonably accurate 83-frame comic strip on the history of the building of the atomic bomb, with Leo Szilard as the central figure of the story.

Mehra:1975:SCP

[Meh75] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Mehra:1987:NBD

[Meh87] Jagdish Mehra. Niels Bohr’s discussions with Albert Einstein, Werner Heisenberg, and Erwin Schrödinger: the origins of the Principles of Uncertainty and Complementarity. Foundations of Physics, 17(5):461–506, May 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Meitner:1939:NPF

[Mei39] Lise Meitner. New products of the fission of the thorium nucleus. Nature, 143(3624):637, April 15, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v143/n3624/pdf/143637a0.pdf.

Meitner:1962:RWR

[Mei62] Lise Meitner. Right and wrong roads to the discovery of nuclear energy. International Atomic Energy Agency Bulletin, 4(0):4–6, ???? 1962. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690

(electronic). URL http://www.iaea.org/Publications/Magazines/
Bulletin/Bull040su/04004790608su.pdf.

Meitner:1964:LMLa

[Mei64a] Lise Meitner. Lise Meitner looks back. Advancement of Science, 21 (??):39–46, ???? 1964. CODEN ADSCAH. ISSN 0001-866X.

Meitner:1964:LMLb

[Mei64b] Lise Meitner. Lise Meitner looks back. International Atomic Energy Agency Bulletin, 6(1):4–12, ???? 1964. CODEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). URL http://www.iaea.org/

Publications/Magazines/Bulletin/Bull061/06101400412.pdf.

Melzer:2000:BHS

[Mel00] Richard Melzer. Breakdown: how the secret of the atomic bomb was stolen during World War II. Sunstone Press, Santa Fe, NM, 2000. ISBN 0-86534-304-7. 160 pp. LCCN QC789.2.U62 M45 2000.

Mermin:1985:BBP

[Mer85] N. David Mermin. A bolt from the blue: the E-P-R paradox. In French and Kennedy [FK85], pages 141–147. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Mermin:1989:BRB

[Mer89] David Mermin. Book review: The Philosophical Writings of Niels Bohr. Physics Today, 42(2):??, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Mermin:2004:CCH

[Mer04] N. D. Mermin. Copenhagen computation: How I learned to stop worrying and love Bohr. IBM Journal of Research and Development, 48(1): 53–??, ???? 2004. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-

8556 (electronic). URL http://www.research.ibm.com/journal/rd/
481/mermin.pdf.

Meitner:1939:DUN

[MF39a] Lise Meitner and Otto Robert Frisch. Disintegration of uranium by neutrons: a new type of nuclear reaction. Nature, 143(3615):239–240, February 11, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.nature.com/nature/journal/

v143/n3615/pdf/143239a0.pdf. This paper, and [MF39c], both submitted 16 January 1939 (see [Mei62]), provided the first published explanation of nuclear disintegration, called ‘nuclear fission’ by Frisch, that was first observed experimentally by Hahn and Strassmann in December 1939 [HS39b]. Frisch’s paper [Fri39] describes the first experimental confirmation. It was these results that Niels Bohr intended to hold confidential until their journal publication during his January 1939 trip to the USA, but his traveling companion Léon Rosenfeld [Ros72] wasn’t informed of that intent, and the news escaped and spread quickly.

Meitner:1939:PFUa

[MF39b] Lise Meitner and Otto Robert Frisch. On the products of the fission of uranium and thorium under neutron bombardment. Math.-fys. Meddr, 17(5):1–13, 1939.

Meitner:1939:PFUb

[MF39c] Lise Meitner and Otto Robert Frisch. Products of the fission of the uranium nucleus. Nature, 143(3620):471–472, March 18, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v143/n3620/

pdf/143471a0.pdf. See note in [MF39a].

Martienssen:1995:BAF

[MGG^{+}95] W. Martienssen, S. Großmann, P. Grassberger, D. Meissner, W. Sandhas, H. Nicolai, D. Hoffmann, H. Kant, P. Richter, H. Atmanspacher, M. Eckert, H. Rechenberg, H. Maier-Leibnitz, and D. Haarer. Buchbesprechungen: Argyris/Faust: Die Erforschung des Chaos/Plaschko/Brod: Nichtlineare Dynamik, Bifurkation und Chaotische Systeme/Bunde/Havlin: Fractals in Science/Goetzberger/Voß/Knobloch: Sonnenenergie: Photovoltaik/Lindner: Grundkurs Theoretische Physik/Vilenkin/Shellard: Cosmic Strings and other Topological Defects/Weiss: Großforschung in Deutschland Geschichte des Hahn-Meitner-Instituts/Fölsing: Wilhelm Conrad Röntgen Aufbruch ins Innere der Materie/Brandmüller: Galilei und die Kirche Ein Fall und seine Lösung/Meÿenn: Wolfgang Pauli Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a. Band 3: 1940 49/Atmanspacher/Primas: Der Pauli-Jung-Dialog und seine Bedeutung für die moderne Wissenschaft/Cassidy: Werner Heisenberg Leben und Werk/Weinberg: The First Nuclear Era The Life and Times of a Technological Fixer/Wenske: Wörterbuch Chemie/Dictionary of Chemistry. Physikalische Blätter, 51(11):1101–1106, November 1995. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19950511117/abstract.

Maeda:2009:NBW

[MGG09] H. Maeda, J. H. Gurian, and T. F. Gallagher. Nondispersing Bohr wave packets. Physical Review Letters, 102(10):103001, March 13, 2009. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Migdal:1985:NBQ

[Mig85] A. B. Migdal. Niels Bohr and quantum physics. Soviet Physics. Uspekhi, 28(10):910–934, October 1985. CODEN SOPUAP. ISSN 0038-5670.

URL http://stacks.iop.org/0038-5670/28/i=10/a=A04.

Milne:1924:BRB

[Mil24] E. A. Milne. Book review: Report on Radiation and the Quantum Theory, by J. H. Jeans; On the Application of the Quantum Theory to Atomic Structure. Part I.: The Fundamental Postulates, by Niels Bohr (translated by L. F. Curtiss). Mathematical Gazette, 12(172):217–220, October 1924. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL http://www.jstor.org/stable/3604572.

Mladjenovic:1998:DYN

[Mla98] Milorad Mladjenović. The Defining Years in Nuclear Physics, 1932– 1960s. IOP Publishing, Bristol, UK, 1998. ISBN 0-7503-0472-3 (hardcover). xx + 441 pp. LCCN QC773 .M54 1998.

Mlodinow:2014:BMM

[Mlo14] Leonard Mlodinow. Bohr’s molecular model and the melding of classical and quantum mechanics. Physics Today, 67(8):8, August 2014. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [SSH14a].

Minev:2019:CRQ

[MMS^{+}19] Z. K. Minev, S. O. Mundhada, S. Shankar, P. Reinhold, R. Gutiérrez-Jáuregui, R. J. Schoelkopf, M. Mirrahimi, H. J. Carmichael, and M. H. Devoret. To catch and reverse a quantum jump mid-flight. Nature, 570(7760):200–204, June 13, 2019. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/articles/s41586-019-1287-z.

Moffett:2006:WBQ

[Mof06] Pamela Valerie Moffett. Wittgenstein, Bohr and the quest for objectivity in education research: a new framework for educational and psychological attributes. Ph.D. dissertation, Queen’s University Belfast, Belfast, Northern Ireland, 2006. 304 pp.

Moller:1955:OPG

[Møl55] C. Møller. Old problems in the General Theory of Relativity viewed from a new angle: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(10) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 28 pp. LCCN AS281.

Moore:1966:NBM

[Moo66] Ruth E. Moore. Niels Bohr: the man, his science, and the world they changed. Alfred A. Knopf, New York, NY, USA, 1966. xi + 436 + vii + 8 pp. LCCN QC16.B55 M6.

Moore:1967:NBM

[Moo67] R. Moore. Niels Bohr: the man and the scientist. Hodder and Stoughton, London, UK, 1967. xi + 436 + vii pp. LCCN QC16.B63 M6 1967.

Moore:1969:NBC

[Moo69] Ruth E. Moore. Nil’s Bor — chelovek i ucheny. (Russian) [Niels Bohr — man and scientist]. Mir, Moskva, USSR, 1969. 468 pp. LCCN ????

Moore:1985:NBP

[Moo85a] Ruth Moore. Niels Bohr as a political figure. In French and Kennedy [FK85], pages 253–260. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Moore:1985:NBM

[Moo85b] Ruth Moore. Niels Bohr: the man, his science, and the world they changed. Biography. Science. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-63101-6. xvi + 436 + vii + 16 pp. LCCN ????

Morrison:1956:BRB

[Mor56] P. Morrison. Book review: Niels Bohr and the Development of Physics: Essays Dedicated to Niels Bohr on the Occasion of his Seventieth Birthday, edited by W. Pauli, L. Rosenfeld, and V. Weisskopf. Physics Today, 9(8):32–??, August 1956. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [PRW55].

Morrison:1985:GOS

[Mor85a] Philip Morrison. A glimpse on the other side. In French and Kennedy [FK85], pages 345–350. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Mortensen:1985:KOH

[Mor85b] Viggo Mortensen, editor. Kontrast og harmoni: Niels Bohr som fysiker og tænker. (Danish) [Contrast and harmony: Niels Bohr as physicist and philosopher], volume 1 of Complementa. Anis, Århus, Danmark, 1985. ISBN 87-7457-038-2. 110 pp. LCCN QC16.B63 K66 1985. kr98.00.

Morrison:1995:NTW

[Mor95] Philip Morrison. Nothing is too wonderful to be true, volume 11 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN 1-56396-363-9 (hardcover). xi + 446 pp. LCCN Q173.

Morrison:2007:SAW

[Mor07] Margaret Morrison. Spin: All is not what it seems. Studies in History and Philosophy of Modern Physics, 38(3):529–557, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1355219806000931.

Morgenweck:2011:RLS

[Mor11] Vera Morgenweck. Ruth Lewin Simes Darstellung des Verhältnisses Lise Meitner–Otto Hahn. (German) [Ruth Lewin Simes’s representation of the relationship Lise Meitner–Otto Hahn]. Wissenschaftliche Hausarbeit für das Lehramt an Haupt- und Realschulen (L2) eingereicht dem Wissenschaftlichen Prüfungsamt für das Lehramt an Grund-Schulen, Haupt- und Realschulen in Frankfurt am Main, August 21,

2011. URL http://www.psychologie.uni-frankfurt.de/61744137/
Antragsunterlagen_fuer_die_Wissenschaftliche_Hausarbeit.pdf.

Abgabetermin: 21. September 1998, Überarbeitet (2011).

Mott:1985:NBI

[Mot85] Nevill Mott. At the Niels Bohr Institute in 1929. In French and Kennedy [FK85], pages 172–174. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Mottelson:2008:NBD

[Mot08] Ben R. Mottelson. Niels Bohr and the development of concepts in nuclear physics. Lecture notes in physics, 746:115–136, 2008. CODEN LNPHA4. ISSN 0075-8450 (print), 1616-6361 (electronic).

URL http://www.springerlink.com/content/n443430v62773h03/.

Nishina Memorial Lectures.

Moyer:1981:EDE

[Moy81a] Donald Franklin Moyer. Evaluations of Dirac’s electron, 1928–1932. American Journal of Physics, 49(11):1055–1062, November 1981. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://ajp.aapt.org/resource/1/ajpias/v49/i11/p1055_s1. Part

2 of 3; see also part 1 [Moy81b] and part 3 [Moy81c]. The paper reprints two long letters between Niels Bohr and Paul Dirac from November– December 1929 on the problem of negative energies in Dirac’s quantum theory.

Moyer:1981:ODE

[Moy81b] Donald Franklin Moyer. Origins of Dirac’s electron, 1925–1928. American Journal of Physics, 49(10):944–949, October 1981. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:

//ajp.aapt.org/resource/1/ajpias/v49/i10/p944_s1. Part 1 of 3;

see also part 2 [Moy81a] and part 3 [Moy81c].

Moyer:1981:VDE

[Moy81c] Donald Franklin Moyer. Vindications of Dirac’s electron, 1932–1934. American Journal of Physics, 49(12):1120–1125, December 1981. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://ajp.aapt.org/resource/1/ajpias/v49/i12/p1120_s1. Part

3 of 3; see also part 1 [Moy81b] and part 2 [Moy81a].

Merleau-Ponty:1997:AOBa

[MP97] Jacques Merleau-Ponty. Analyses d’ouvrages: Niels Bohr and contemporary philosophy, ^{≪}Boston Studies in the Philos. of Sc. ^{≫}, vol. 153 par Jan Faye; Henry J. Folse. Revue d’Histoire des Sciences, 50(1–2):221– 223, January 1997. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-

6582 (electronic). URL http://www.jstor.org/stable/23633283.

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Mehra:1982:DQM

[MR82a] Jagdish Mehra and Helmut Rechenberg. The discovery of quantum mechanics, 1925, volume 2 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90674-6. vi + 355 pp. LCCN QC173.98 .M44 vol. 2. See other volumes in this series [MR82c, MR82b, MR82e, MR82f, MR87, MR00].

Mehra:1982:FMM

[MR82b] Jagdish Mehra and Helmut Rechenberg. The Formulation of Matrix Mechanics and Its Modifications, 1925–1926, volume 3 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90675-4, 3-540-90675-4. viii + 334 pp. LCCN QC173.98 .M44

vol. 3. URL http://www.loc.gov/catdir/enhancements/fy1005/
82003253-d.html;   http://www.loc.gov/catdir/enhancements/

fy1005/82003253-t.html. See other volumes in this series [MR82a, MR82c, MR82e, MR82f, MR87, MR00].

Mehra:1982:FEQ

[MR82c] Jagdish Mehra and Helmut Rechenberg. The fundamental equations of quantum mechanics, 1925–1926. The reception of the new quantum mechanics, 1925–1926, volume 4 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90680-0. viii + 322 pp. LCCN QC173.98 .M44 vol. 4. See other volumes in this series [MR82a, MR82b, MR82e, MR82f, MR87, MR00].

Mehra:1982:HDQ

[MR82d] Jagdish Mehra and Helmut Rechenberg. The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-90675-4 (Berlin), 0-387-90675-4 (New York). vi + 334 pp. LCCN QC173.98 .M44 vol. 3.

Mehra:1982:QTPa

[MR82e] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 1) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-90642-8 (Berlin), 0-387-90642-8 (New York). 372 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82f, MR87, MR00].

Mehra:1982:QTPb

[MR82f] Jagdish Mehra and Helmut Rechenberg. The quantum theory of Planck, Einstein, Bohr and Sommerfeld: its foundation and the rise of its difficulties 1900–1925, volume 1 (part 2) of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90667-3 (New York), 3-540-90667-3 (Berlin). 375–878 pp. LCCN QC 174.12. See other volumes in this series [MR82a, MR82c, MR82b, MR82e, MR87, MR00].

Mehra:1987:ESR

[MR87] Jagdish Mehra and Helmut Rechenberg. Erwin Schrödinger and the rise of wave mechanics. Part 1, Schrödinger in Vienna and Zurich 1887–1925, volume 5, part 1 of The Historical Development of Quantum Theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96284-0 (vol. 1, New York), 3-540-96284-0 (vol. 1, Berlin), 0-387-96377-4 (vol. 2, New York), 3-387-96377-4 (vol. 2, Berlin). xx + 366 pp. LCCN QC173.98 .M44 vol. 5. US$48.00, US$49.50. See other volumes in this series [MR82a, MR82c, MR82b, MR82e, MR82f, MR00].

Mehra:2000:CQM

[MR00] Jagdish Mehra and Helmut Rechenberg. The Completion of Quantum Mechanics, 1926–1941, volume 6 of The historical development of quantum theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-95086-9 (part 1), 0-387-95182-2 (part 2). xxxvi + 1612 pp. LCCN QC173.98 .M44 vol. 6. See other volumes in this series [MR82a, MR82c, MR82b, MR82e, MR82f, MR87].

Mataix:2002:FCR

[MR02] Carmen Mataix and Andrés Rivadulla Rodríguez, editors. Fisica cuántica y realidad = Quantum physics and reality, volume 18 of Philosophica Complutensia. Facultad de Filosofía, Universidad Complutense, Madrid, Spain, 2002. ISBN 84-7491-640-2. 369 pp. LCCN QC174.12 .F57 2002.

Meyer:1937:FTL

[MSB^{+}37] Stefan Meyer, A. Norman Shaw, Niels Bohr, George Hevesy, le Duc de Broglie, Johannes Stark, Otto Hahn, Enrico Fermi, L. Wertenstein, and Peter Kapitza. Further tributes to the late Lord Rutherford. Energia elettrica. Tribute to Lord Rutherford. Nature, 140(3555): 1047–1054, December 18, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/nature/

journal/v140/n3555/index.html.

Murdoch:1987:NBP

[Mur87] Dugald Murdoch. Niels Bohr’s philosophy of physics. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-33320-2. x + 294 pp. LCCN QC16.B63 M87 1987. URL http:

//catdir.loc.gov/catdir/description/cam032/87011717.html;
http://catdir.loc.gov/catdir/toc/cam032/87011717.html.

Murdoch:1989:NBP

[Mur89] Dugald Murdoch. Niels Bohr’s philosophy of physics. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-37927-X, 0-521-33320-2. x + 294 pp. LCCN QC16.B63 M87 1989.

Murdoch:1990:BRB

[Mur90a] Dugald Murdoch. Book review: Einstein versus Bohr: The Continuing Controversies in Physics by Mendel Sachs. Isis, 81(3):597, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233482.

Murdoch:1990:NBP

[Mur90b] Dugald Murdoch. Niels Bohr’s Philosophy of Physics. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-33320-2 (hardcover), 0-521-37927-X (paperback). x + 294 pp. LCCN QC16.B6 M87 1987. US$16.95.

Murdoch:1994:BED

[Mur94] Dugald Murdoch. The Bohr–Einstein dispute. In Faye and Folse [FF94], chapter 14, pages 303–324. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153

1994; QC16 1994. URL https://link.springer.com/chapter/10.
1007/978-94-015-8106-6_14.

Masters:1946:OWN

[MW46a] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. Whittlesey House, McGraw-Hill Book Co., Inc., New York, NY, USA, 1946. x + 79 pp. LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Masters:1946:VEI

[MW46b] Dexter Masters and Katherine Way, editors. Een Verden eller ingen. (Danish) [One world or none]. ????, København, Danmark, 1946. 220 pp. LCCN ???? Danish translation of [MW46a] by Gudrun Frederiksen og Ebbe Rasmussen. Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Miller:1984:DCE

[MW84] W. A. Miller and J. A. Wheeler. Delayed choice experiments and Bohr’s elementary phenomenon. In Kamefuchi et al. [K^{+}84], pages 140– 152. ISBN 4-89027-001-9. LCCN QC173.96 .I57 1983.

Masters:2007:OWN

[MW07] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. New Press, New York, NY, USA, 2007. ISBN 1-59558-227-4 (hardcover). xx + 220 pp. LCCN UG1282.A8 O54 2007.

URL http://thenewpress.com/index.php?option=com_title&task=
view_title&metaproductid=1703;   http://www.loc.gov/catdir/

toc/ecip0718/2007020838.htm. Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46a].

Mestayer:2008:RLB

[MWL^{+}08] J. J. Mestayer, B. Wyker, J. C. Lancaster, F. B. Dunning, C. O. Reinhold, S. Yoshida, and J. Burgdorfer. Realization of localized Bohrlike wave packets. Physical Review Letters, 100(24):243004, June 20, 2008. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

N:1938:BRV

[N.38] E. N. Book review: Volume I, Number 1: Encyclopedia and Unified Science, Otto Neurath, Niels Bohr, John Dewey, Bertrand Russell, Rudolf Carnap, Charles W. Morris, 1938, vii + 75 pp $1.00. Journal of Philosophy, 35(25):689–693, December 8, 1938. URL http://www.jstor.

org/stable/2017994.

Nakane:2015:OAA

[Nak15] Michiyo Nakane. The origins of action-angle variables and Bohr’s introduction of them in a 1918 paper. In Aaserud and Kragh [AK15], pages 290–309. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/2015/
06/Bohr-Book.pdf.

Nathan:1999:BRB

[Nat99] Ove Nathan. Book review: Niels Bohr: Memoirs of a Working Relationship by Stefan Rozental. Isis, 90(4):836–837, December 1999. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237724.

Nauenberg:2013:ESB

[Nau13] Michael Nauenberg. Edmund Stoner and the Bohr atom. Physics Today, 66(4):10, April 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Nauenberg:2015:WHB

[Nau15] Michael Nauenberg. What happened to the Bohr–Sommerfeld elliptic orbits in Schrödinger’s wave mechanics? In Aaserud and Kragh [AK15], pages 465–480. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172

.O63 2015. URL http://www.ijqf.org/wps/wp-content/uploads/
2015/06/Bohr-Book.pdf.

Navarro:2006:EAD

[Nav06] Jaume Navarro. Early attempts to detect the neutrino at the Cavendish Laboratory. Physics in Perspective (PIP), 8(1):64–82, March 2006. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/s00016-005-0249-z.

Navarro:2015:PPB

[Nav15] Jaume Navarro. Plum puddings and Bohr’s atom. In Aaserud and Kragh [AK15], pages 75–94. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Neurath:1938:IEU

[NBD^{+}38] Otto Neurath, Niels Bohr, John Dewey, Bertrand Russell, Rudolf Carnap, and Charles W. Morris, editors. International Encyclopedia of Unified Science. University of Chicago Press, Chicago, IL, USA and London, UK, 1938. 2 pp. LCCN Q121 .I5 vol. 1, no. 1 1938.

Nier:1940:NFS

[NBDG40] Alfred O. Nier, E. T. Booth, J. R. Dunning, and A. V. Grosse. Nuclear fission of separated uranium isotopes. Physical Review, 57(6): 546, March 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/PhysRev.

57.546. This paper reports experiments that confirm the Bohr–Wheeler prediction [BW39b] that U-235 is much more fissile than U-238. That information was critical for the production of both a nuclear reactor, and a nuclear bomb.

Neurath:1938:EUS

[NCM38] Otto Neurath, Rudolf Carnap, and Charles W. (Charles William) Morris, editors. Encyclopedia and unified science, volume I(1) of International encyclopedia of unified science. University of Chicago Press, Chicago, IL, USA and London, UK, 1938. viii + 75 pp. LCCN Q121 .I5 v.1, no.1.

Neurath:1955:IEU

[NCM55] Otto Neurath, Rudolf Carnap, and Charles W. (Charles William) Morris, editors. International encyclopedia of unified science. University of Chicago Press, Chicago, IL, USA and London, UK, 1955. ???? pp. LCCN Q175 I58.

Neurath:1969:FUS

[NCM70] Otto Neurath, Rudolf Carnap, and Charles W. (Charles William) Morris, editors. Foundations of the unity of science: toward an international encyclopedia of unified science. Foundations of the unity of science: toward an International encyclopedia of unified science. University of Chicago Press, Chicago, IL, USA and London, UK, 1969–1970. ISBN 0-226-57586-1 (vol. 1), 0-226-57588-8 (vol. 2). ???? pp. LCCN Q175 .F686 1969.

Neurath:1944:FSS

[Neu44] Otto Neurath, editor. Foundations of the social sciences, volume II(1) of International encyclopedia of unified science. University of Chicago Press, Chicago, IL, USA and London, UK, 1944. 51 pp. LCCN Q121 .I5 v.2, no.1.

Newton:2009:HPC

[New09] Roger G. Newton. How Physics Confronts Reality: Einstein Was Correct, But Bohr Won the Game. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2009. ISBN 981-4277-02-9 (hardcover), 981-4277-03-7 (paperback). ix + 147 pp. LCCN QC173.98 .N49 2009.

Niaz:1998:CRA

[Nia98] Mansoor Niaz. From cathode rays to alpha particles to quantum of action: A rational reconstruction of structure of the atom and its implications for chemistry textbooks. Science Education, 82(5):527–552, September 1998. CODEN SEDUAV. ISSN 0036-8326 (print), 1098-237X

(electronic). URL http://www.umich.edu/~chemstu/content_weeks/
F_06_Week4/Thompson_Rutherford_Bohr.pdf.

Nielsen:1955:MOG

[Nie55] O. B. Nielsen. Multipole order of the gamma-rays from 81 Tl 208: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(11) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 15 pp. LCCN ????

Nielson:1963:MNB

[Nie63] J. Rud Nielson. Memories of Niels Bohr. Physics Today, 16(10):22–30, October 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Nielsen:1965:BRO

[Nie65] J. Rud Nielsen. Book review: The occasional Bohr: The unity of knowledge: Niels Bohr: Essays 1958/1962 on Atomic Physics and Human Knowledge. Isis, 56(2):214–216, Summer 1965. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/227923.

Nielsen:1968:BRB

[Nie68] J. Rud Nielsen. Book review: Niels Bohr. His Life and Work as Seen by His Friends and Colleagues by S. Rozental. Isis, 59(1):121–123, Spring 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/227877.

NBL:1997:SGA

[Nie97] Niels Bohr Library. Supplement to the Guide to the archival collections in the Niels Bohr Library at the American Institute of Physics, 1993– 1996, volume 8 of International catalog of sources for history of physics and allied sciences. American Institute of Physics, College Park, MD, USA, 1997. ISBN 1-56396-787-1 (paperback). x + 95 pp. LCCN QC7 .N48 1994; QC7 .N48 1994 Suppl.

Nielsen:2015:BAB

[Nie15] Kristian H. Nielsen. The Bohr atom bound in cloth: Textual exposition of quantum theory in popular science books, 1918–1924. In Aaserud and Kragh [AK15], pages 141–160. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Nikolic:2008:WBB

[Nik08] Hrvoje Nikolić. Would Bohr be born if Bohm were born before Born? American Journal of Physics, 76(2):143–??, February 2008. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:/

/ajp.aapt.org/resource/1/ajpias/v76/i2/p143_s1.

Nikolic:2012:EBE

[Nik12] Hrvoje Nikolić. EPR before EPR: a 1930 Einstein–Bohr thought experiment revisited. European Journal of Physics, 33(5):1089, 2012. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http:/

/stacks.iop.org/0143-0807/33/i=5/a=1089.

Nishina:1937:PNB

[Nis37] Yoshio Nishina. Professor Niels Bohr’s visit to Japan. Kagaku, 7(??): 207–210, ???? 1937.

Nishio:1967:RCC

[Nis67] Sigeko Nishio. The role of the chemical considerations in the development of Bohr atom model. Japanese Studies in the History of Science, 6(??):26–40, ???? 1967. CODEN JSHIAE. ISSN 0090-0176.

Niaz:2012:RWP

[NM12] Mansoor Niaz and Cecilia Marcano. Reconstruction of Wave-particle Duality and Its Implications for General Chemistry Textbooks. Springer briefs in education. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-4395-5 (paperback). viii + 46 pp. LCCN QC476.W38 N53 2012. URL http://www.springerlink.

com/content/978-94-007-4395-3.

Nielsen:2001:NTN

[NN01] Keld Nielsen and Henry Nielsen, editors. Nabo til Nobel: historien om tretten danske Nobelpriser. (Danish) [Neighbor of Nobel: the story of thirteen Danish Nobel Prize winners]. Aarhus Universitetsforlag, Århus, Danmark, 2001. ISBN 87-7288-900-4. 570 pp. LCCN AS911.N9 N33

2001. DKR 398,00. URL http://da.unipress.dk/udgivelser/n/
nabo-til-nobel/.

Oppenheimer:19xx:JRO

[OBxx] J. Robert Oppenheimer and Niels Bohr. J. Robert Oppenheimer papers (1799–1980) (bulk 1947–1967) and Niels Henrik David Bohr papers. ????, ????, 19xx. ???? pp. LCCN 0456G; Vault 0202A; Microfilm

16,646-1P. URL http://hdl.loc.gov/loc.mss/eadmss.ms998007.

Rudinger:1964:GDB

[oER64] Léon Rosenfeld og Erik Rüdinger. Gennembrudsårene. (Danish) [The breakthrough years]. In Bohr and Rozental [BR64], pages 36–70. LCCN QC16.B63 N5. English translation in [Roz67].

Ozawa:2012:RBR

[OK12] Masanao Ozawa and Yuichiro Kitajima. Reconstructing Bohr’s reply to EPR in algebraic quantum theory. Foundations of Physics, 42 (4):475–487, April 2012. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/
10.1007/s10701-011-9615-7.

Okun:2001:RBE

[Oku01] L. B. Okun. Relation between energy and mass in Bohr’s essay on his debate with Einstein. Physics of Atomic Nuclei, 64(3):536– 539, ???? 2001. CODEN PANUEO. ISSN 1063-7788 (print), 1562-

692X (electronic). URL http://www.springerlink.com/content/

y8117632147335h3/. Translated from Russian journal Yadernaya Fizika, Vol. 64, No. 3, 2001, pp. 590–593. 90th Anniversary of A. B. Migdal’s Birthday Elementary Particles and Fields.

Oliphant:1985:BR

[Oli85] Mark Oliphant. Bohr and Rutherford. In French and Kennedy [FK85], pages 68–70. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Pihl:1964:NBI

[oMP64] Christian Møller og Mogens Pihl. Niels Bohrs indsats i fysikken. (Danish) [Niels Bohr contributions to physics]. In Bohr and Rozental [BR64], pages 234–253. LCCN QC16.B63 N5. English translation in [Roz67].

Ottaviani:2004:SLN

[OP04] Jim Ottaviani and Leland Purvis, editors. Suspended in Language: Niels Bohr’s Life, Discoveries, and the Century He Shaped. G. T. Labs, Ann Arbor, MI, USA, 2004. ISBN 0-9660106-5-5. 318 pp. LCCN QC773 2004. Additional art by Jay Hosler, Roger Langridge, Steve Leialoha, Linda Medley, and Jeff Parker.

Ottaviani:2009:SLN

[OPH^{+}09] Jim Ottaviani, Leland Purvis, Jay (Jay S.) Hosler, et al. Suspended in language: Niels Bohr’s life, discoveries, and the century he shaped. Gt Labs, Ann Arbor, MI, USA, second edition, 2009. ISBN 0-9788037-2-8. 318 pp. LCCN QC16.B63 O77 2009.

Oppenheimer:1950:AS

[Opp50] J. Robert Oppenheimer. The age of science: 1900–1950. Scientific American, 183(3):20–23, September 1950. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Oppenheimer:1963:NBH

[Opp63a] J. R. Oppenheimer. Niels Bohr and his times. Pegram Lectures, Brookhaven., 1963.

Oppenheimer:1963:NBM

[Opp63b] J. R. Oppenheimer. Niels Bohr memoir. In Year Book, American Philosophical Society, page ?? American Philosophical Society, Philadelphia, PA, USA, 1963.

Oppenheimer:1963:NB

[Opp63c] J. Robert Oppenheimer. Niels Bohr. Bulletin of the Atomic Scientists, 19(7):9, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1964:NBN

[Opp64] J. Robert Oppenheimer. Niels Bohr and nuclear weapons. New York Review of Books, 3(9):6–8, December 17, 1964. CODEN ???? ISSN 0028-7504 (print), 1944-7744 (electronic). URL http:

//www.nybooks.com/articles/archives/1964/dec/17/niels-bohr-
and-atomic-weapons/.

Oechsner:1986:BKP

[OSWR86] H. Oechsner, A. Schlachetzki, H. Walther, and H. Rechenberg. Buchbesprechungen: Kirschner: Polarized Electrons at Surfaces/Paul, Teubner: Optoelektronische Halbleiterbauelemente/Stitch, Bass: Laser Handbook, Vol. 4/Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a.: Teil II: 1930–1939. Physikalische Blätter, 42 (7):252–253, July 1986. CODEN PHBLAG. ISSN 0031-9279 (print),

1521-3722 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1002/phbl.19860420726/abstract.

Otte:2003:CSN

[Ott03] M. Otte. Complementarity, sets and numbers. Educational Studies in Mathematics, 53(3):203–228, September 2003. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1026001332585.

Pors:2001:PTH

[PA01] Felicity Pors and Finn Aaserud. The physical tourist: Historical sites of physical science in Copenhagen. Physics in Perspective (PIP), 3(2): 230–248, 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pais:1948:DTE

[Pai48] Abraham Pais. Developments in the theory of the electron: Containing also an abstract of field and charge measurements in quantum theory: a paper by N. Bohr and L. Rosenfeld. Princeton University Press, Princeton, NJ, USA, 1948. 45 pp. LCCN ????

Pais:1964:MFE

[Pai64] Abraham Pais. Minder fra efterkrigstiden. (Danish) [Memories from after the war]. In Bohr and Rozental [BR64], pages 207–218. LCCN QC16.B63 N5. English translation in [Roz67].

Pais:1979:EQT

[Pai79] A. Pais. Einstein and the quantum theory. Reviews of Modern Physics, 51(4):863–914, October 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/abstract/RMP/v51/p863;   http://prola.aps.org/
abstract/RMP/v51/i4/p863_1;   http://prola.aps.org/pdf/RMP/
v51/i4/p863_1.

Pais:1985:NBD

[Pai85a] A. Pais. Niels Bohr and the development of physics. Report, CERN, Geneva, Switzerland, 1985. ix + 13 pp. A tribute to Niels Bohr, special colloquium held at CERN on 6 May 1985.

Pais:1985:RPY

[Pai85b] Abraham Pais. Reminiscences from the postwar years. In French and Kennedy [FK85], pages 244–250. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Pais:1986:IBM

[Pai86] Abraham Pais. Inward bound: of matter and forces in the physical world. Clarendon Press, Oxford, UK, 1986. ISBN 0-19-851971-0 (hardcover), 0-19-851997-4 (paperback). xiv + 666 pp. LCCN QC7 .P27 1986.

URL http://www.loc.gov/catdir/enhancements/fy0638/85021587-
d.html.

Pais:1991:NBT

[Pai91] Abraham Pais. Niels Bohr’s times: in physics, philosophy, and polity. Clarendon Press, Oxford, UK, 1991. ISBN 0-19-852048-4 (paperback), 0-19-852049-2. xvii + 565 pp. LCCN QC773 .P35 1991. US$35.00.

URL http://www.loc.gov/catdir/enhancements/fy0635/90027248-
d.html.

Pais:1994:ELH

[Pai94] Abraham Pais. Einstein Lived Here: Essays for the Layman. Clarendon Press, Oxford, UK, 1994. ISBN 0-19-853994-0, 0-19-851770-X (T-shirt). xvi + 282 pp. LCCN QC16.E5 P25 1994. US$22.42.

Pais:2000:GSP

[Pai00] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp. LCCN Q141 .P29 2000. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http://www.
loc.gov/catdir/toc/fy02/99046603.html.

Palevsky:2000:AFD

[Pal00] Mary Palevsky. Atomic Fragments: a Daughter’s Questions. University of California Press, Berkeley, CA, USA, 2000. ISBN 0-520-22055-2 (hardcover), 0-250-22055-2 (hardcover). xiv + 289 pp.

LCCN QC773.3.U5 P35 2000. URL http://catdir.loc.gov/catdir/
bios/ucal052/99087422.html;   http://catdir.loc.gov/catdir/
description/ucal042/99087422.html.

Pallo:2015:BME

[Pal15] Gábor Palló. The Bohr model’s early reception in Hungary: Hevesy and Bohr. In Aaserud and Kragh [AK15], pages 511–521. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63 2015. URL http://www.ijqf.

org/wps/wp-content/uploads/2015/06/Bohr-Book.pdf.

Park:2006:BRG

[Par06] David Park. Book review: Gerald Holton, Victory and Vexation in Science: Einstein, Bohr, Heisenberg, and Others. Cambridge, Mass. and London: Harvard University Press, 2005. xiv + 229 pages. $35.00 (cloth). Physics in Perspective (PIP), 8(4):481–483, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pasachoff:2003:NBP

[Pas03] Naomi E. Pasachoff. Niels Bohr: physicist and humanitarian. Great minds of science. Enslow Publishers, Berkeley Heights, NJ, 2003. ISBN 0-7660-1997-7. 128 pp. LCCN QC16.B63 P37 2003. URL http://www.

loc.gov/catdir/toc/fy035/2002003887.html.

Paty:1985:ECA

[Pat85] Michel Paty. Einstein et la complémentarité au sens de Bohr: du retrait dans le tumulte aux arguments d’incomplétude. (French) [Einstein and complementarity according to Bohr: withdrawal in the tumult of the incompleteness arguments]. Revue d’Histoire des Sciences, 38(3–4):325–351, juillet–décembre 1985. CODEN RHSAAM. ISBN 2-13-039137-0. ISSN 0151-4105 (print), 1969-6582 (elec-

tronic). URL http://www.jstor.org/stable/23632502;   http:
//www.persee.fr/web/revues/home/prescript/article/rhs_0151-
4105_1985_num_38_3_4010.

Pauli:1933:APW

[Pau33] Wolfgang Pauli. Die allgemeinen Prinzipien der Wellenmechanik. (German) [The general principles of wave mechanics]. In Handbuch der Physik, volume XXIV(1), pages 83–272. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1933. According to [De 14, page 76, column 1], this may be the publication where “wave– particle duality” was first defined.

Pauli:1945:NBH

[Pau45] W. Pauli. Niels Bohr on his 60th birthday. Reviews of Modern Physics, 17(2–3):97–101, April 1, 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.17.97;   http://rmp.aps.org/
abstract/RMP/v17/i2-3/p97_1.

Pauli:1955:EPL

[Pau55] W. Pauli. Exclusion principle, Lorentz group and reflection of spacetime and charge. In Niels Bohr and the development of physics, pages 30–51. McGraw-Hill, New York, NY, USA, 1955.

Pauli:1985:JNG

[Pau85a] Wolfgang Pauli. Das Jahr 1930 Die Neutrinohypothese. (German) [The year 1930: The neutrino hypothesis]. In Sources in the History of Mathematics and Physical Sciences [Pau85b], pages 1–48. ISBN 3-540-78801-8.

Pauli:1985:SHM

[Pau85b] Wolfgang Pauli, editor. Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 3-540-78801-8. ???? pp.

Pauli:1994:NBH

[Pau94] Wolfgang Pauli. Niels Bohr on his 60th birthday. In Wolfgang Pauli: Writings on physics and philosophy [EvM94], chapter 4, pages 49–58. ISBN 3-540-56859-X (Berlin), 0-387-56859-X (New York). LCCN QC6.2

.P38 1994. URL http://www.loc.gov/catdir/enhancements/fy0812/
94015098-d.htm.

Pauli:2001:JBR

[Pau01] Wolfgang Pauli. Das Jahr 1955 Der Berner Relativitätskongreß und der Beitrag zur Bohr-Festschrift. (German) [The year 1955: The Bern Relativity Conference and the contribution to the Bohr Anniversary Publication]. In von Meyenn [vM01], pages 1–461. ISBN 3-540-67591-4, 3-540-78805-0. ISSN 0172-6315. LCCN QC16.P37 A34.

Pawlak:2002:BFB

[Paw02] Alexander Pawlak. Bohrende Fragen: Bohr-Dokumente zum Treffen mit Heisenberg 1941 in Kopenhagen freigegeben. (German) [Nagging questions: Bohr documents on the meeting with Heisenberg in 1941 released in Copenhagen]. Physik Journal, 1(3):7–8, ???? 2002. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL http:/

/www.pro-physik.de/details/articlePdf/1108683/issue.html.

Pauli:1958:NBR

[PB58] V. Pauli and Niels Bohr, editors. Nil’s Bor i razvitie fiziki: sbornik, posvjascennyj Nil’su Boru v svjazi s ego semidesjatiletiem. (Russian) [Niels Bohr and the development of physics]. Izdatelstvo Inostrannoj Literatury, Moskva, USSR, 1958. 258 pp. LCCN ????

Pelogia:2017:AJI

[PB17] Karla Pelogia and Carlos Alexandre Brasil. Analysis of the Jun Ishiwara’s “The universal meaning of the quantum of action”. European Physical Journal H, 42(4–5):507–521, December 2017. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL https:

//link.springer.com/article/10.1140/epjh/e2017-80034-x. See

[Ish17].

Pais:2006:JRO

[PC06] Abraham Pais and Robert P. Crease. J. Robert Oppenheimer: a life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-516673-6. xxii + 353 + 16 pp. LCCN QC16.O62 P35 2006. URL

http://www.loc.gov/catdir/enhancements/fy0636/2005002173-d.
html;   http://www.loc.gov/catdir/enhancements/fy0723/2005002173-
b.html;   http://www.loc.gov/catdir/toc/ecip056/2005002173.
html.

Peierls:1997:SSP

[PD97] Sir Rudolf Ernst Peierls and Richard Henry Dalitz, editors. Selected scientific papers of Sir Rudolf Peierls: with commentary, volume 19 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2692-6. xxiii + 805 pp. LCCN QC21.2 .P42 1997.

Peacock:2008:QRH

[Pea08] Kent A. Peacock. The Quantum Revolution: a Historical Perspective. Greenwood guides to great ideas in science. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2008. ISBN 0-313-33448-X. ISSN 1559-5374. xviii + 220 pp. LCCN QC173.98 .P43 2008. URL

http://www.loc.gov/catdir/toc/ecip081/2007039786.html.

Pedersen:1964:NBO

[Ped64] Johannes Pedersen. Niels Bohr og Det Kongelige Danske Videnskabernes Selskab. (Danish) [Niels Bohr and the Royal Danish Academy of Sciences and Letters]. In Bohr and Rozental [BR64], pages 259–271. LCCN QC16.B63 N5. English translation in [Roz67].

Peierls:1939:CCN

[Pei39] Rudolf Peierls. Critical conditions in neutron multiplication. Mathematical proceedings of the Cambridge Philosophical Society, 35(4):610–615, November 1939. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic). This is the first paper with an estimate of the critical mass of uranium needed to sustain a chain reaction. It contains a note added in proof to the fission model of Bohr and Wheeler [BW39b] that appeared after this paper was submitted.

Peierls:1940:BTN

[Pei40] Rudolf Peierls. The Bohr theory of nuclear reactions. Reports on Progress in Physics, 7(??):87–106, ???? 1940. CODEN RPPHAG. ISSN 0034-4885 (print), 1361-6633 (electronic). URL http://iopscience.

iop.org/0034-4885/7/1/306.

Peierls:1955:ABB

[Pei55] Rudolf E. Peierls. The atom that Bohr built. Journal of Jocular Physics, III:47–48, October 7, 1955.

Peierls:1963:ANB

[Pei63] R. E. Peierls. An appreciation of Niels Bohr. Proceedings of the Physical Society, London, 81(5):793–??, May 1, 1963. CODEN PPSOAU. ISSN

0370-1328. URL http://stacks.iop.org/0370-1328/81/i=5/a=301.

Peierls:1985:SRB

[Pei85] Rudolf Peierls. Some recollections of Bohr. In French and Kennedy [FK85], pages 227–231. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Peierls:1986:TC

[Pei86a] R. Peierls. Truth and clarity. In Jorrit de Boer, Erik Dal, and Ole Ulfbeck, editors, The Lesson of Quantum Theory, Niels Bohr Centenary Symposium held 3–7 October, 1985 in Copenhagen, Denmark, pages 379–?? North-Holland Publishing Co., Amsterdam, The Netherlands,

January 1986. URL https://ui.adsabs.harvard.edu/#abs/1986lqt.
.conf..379P.

Peierls:1986:F

[Pei86b] Rudolf Peierls. Foreword. In Niels Bohr: collected works, volume 9, pages 1–83. Elsevier, Amsterdam, The Netherlands, 1986.

Peierls:1986:I

[Pei86c] Rudolf Peierls. Introduction. In Niels Bohr: collected works, volume 9, pages 1–83. Elsevier, Amsterdam, The Netherlands, 1986.

Peierls:1988:RB

[Pei88] Rudolf Peierls. Rutherford and Bohr. Notes and Records of the Royal Society of London, 42(2):229–241, July 1, 1988. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Peierls:1991:MST

[Pei91] Rudolf Peierls. More Surprises in Theoretical Physics. Princeton Series in Physics. Princeton University Press, Princeton, NJ, USA, 1991. ISBN 0-691-08576-5 (hardcover), 0-691-02522-3 (paperback). viii + 106 pp. LCCN QC20 .P345 1991. URL http://www.loc.

gov/catdir/description/prin021/90023189.html;   https://ui.
adsabs.harvard.edu/#abs/1991mstp.book.....P.

Peierls:1997:BTN

[Pei97a] R. Peierls. The Bohr theory of nuclear reactions. In Dalitz and Peierls [DP97], pages 283–302. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL

https://ui.adsabs.harvard.edu/#abs/1997sspr.book..283P.

Peierls:1997:RB

[Pei97b] Rudolf Peierls. Rutherford and Bohr. Current Science, 73(8): 707–712, October 25, 1997. CODEN CUSCAM. ISSN 0011-3891.

URL http://www.currentscience.ac.in/Downloads/download_pdf.
php?titleid=id_073_08_0707_0712_0. Text of Rutherford Memo-

rial Lecture, given at National Physical Laboratory, New Delhi, India, November 10, 1987.

Peierls:1997:RBT

[Pei97c] Rudolf Peierls. Rutherford and Bohr. Current Science (Bangalore), 73 (8):707–712, ???? 1997. CODEN CUSCAM. ISSN 0011-3891. Text of Rutherford Memorial Lecture, given at National Physical Laboratory, New Delhi, India, November 10, 1987.

Peierls:1997:AH

[Pei97d] Sir Rudolf Ernst Peierls. Atomic Histories, volume 18 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1997. ISBN 1-56396-243-8 (hardcover). xvii + 378 pp. LCCN QC71 .P38 1997.

Peierls:2008:NBC

[Pei08] Sir Rudolf Peierls, editor. Niels Bohr collected works. Volume 9, Nuclear physics (1929–1952). Elsevier, Amsterdam, The Netherlands, 2008. ISBN 0-444-53277-3. xviii + 693 pp. LCCN ????

Peierls:2010:RB

[Pei10] Rudolf Peierls. Rutherford and Bohr. Resonance, 15(5):476–487, May 2010. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). Text of the Rutherford Memorial Lecture delivered by Sir Rudolf Peierls, FRS, on 10 November 1987 at the National Physical Laboratory, New Delhi. Reproduced from Current Science, Vol.73, pp.707–712, 25 October, 1997.

Pelseneer:1947:LNP

[Pel47a] Jean Pelseneer. L’évolution de la notion de phénomène physique des primitifs à Bohr et Louis de Broglie. Leçons sur l’histoire de la pensée scientifique professées à l’Université libre de Bruxelles. Office international de librairie, Bruxelles, Belgique, 1947. 176 pp. LCCN ????

Pelseneer:1947:LND

[Pel47b] Jean Pelseneer. L’évolution de la notion du phénomène physique des primitifs à Bohr et Louis de Broglie. (French) [Evolution of the concept of the physical phenomenon of primitives in Bohr and Louis de Broglie]. Office des cours du Cercle des Sciences de l’Université Libre de Bruxelles, Bruxelles, Belgium, 1947.

Peres:1985:EGB

[Per85] Asher Peres. Einstein, Gödel, Bohr. Foundations of Physics, 15(2): 201–205, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/
10.1007/BF00735292.

Perovic:2006:SIQ

[Per06] Slobodan Perovic. Schrödinger’s interpretation of quantum mechanics and the relevance of Bohr’s experimental critique. Studies in History and Philosophy of Modern Physics, 37(2):275–297, June 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219806000086.

Perovic:2013:ECB

[Per13] Slobodan Perovic. Emergence of complementarity and the Baconian roots of Niels Bohr’s method. Studies in History and Philosophy of Modern Physics, 44(3):162–173, August 2013. ISSN 1355-2198 (print), 1879-

2502 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000427.

Perovic:2021:DQN

[Per21] Slobodan Perović. From Data to Quanta : Niels Bohr’s Vision of Physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2021. ISBN 0-226-79847-X. 251 pp. LCCN QC16.

Petersen:1963:PNB

[Pet63] Aage Petersen. The philosophy of Niels Bohr. Bulletin of the Atomic Scientists, 19(7):8–14, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Petersen:1985:PNB

[Pet85] Aage Petersen. The philosophy of Niels Bohr. In French and Kennedy [FK85], pages 299–310. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Petruccioli:1988:AMP

[Pet88] Sandro Petruccioli. Atomi metafori paradossi: Niels Bohr e la costruzione di una nuova fisica. (Italian) [Atoms as metaphorical paradoxes: Niels Bohr and construction of a new physics]. Edizioni Theoria, Roma, Italia, 1988. ISBN ???? 327 pp. LCCN ????

Petruccioli:2011:CBU

[Pet11] Sandro Petruccioli. Complementarity before uncertainty. Archive for History of Exact Sciences, 65(6):591–624, November 2011. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=article&
issn=0003-9519&volume=65&issue=6&spage=591.

Petruccioli:2014:CPV

[Pet14] Sandro Petruccioli. Correspondence principle versus Planck-type theory of the atom. Archive for History of Exact Sciences, 68(5):599– 639, September 2014. CODEN AHESAN. ISSN 0003-9519 (print),

1432-0657 (electronic). URL http://link.springer.com/article/
10.1007/s00407-014-0137-5.

Pauli:1979:WPW

[PHvMW93] Wolfgang Pauli, Armin Hermann, K. von Meyenn, and Victor F. (Victor Frederick) Weisskopf. Wolfgang Pauli, wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a.. (German) [Wolfgang Pauli, scientific correspondence with Bohr, Einstein, Heisenberg, and others], volume 2, 6, 11 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979–1993. ISBN 0-387-08962-4, 3-540-08962-4, 0-387-13609-6, 3-540-13609-6, 0-387-54911-0, 3-540-54911-0. ???? pp. LCCN QC16.P37 W64.

Piaggio:1923:BRB

[Pia23] H. T. H. Piaggio. Book review: The Theory of Spectra and Atomic Constitution, by Niels Bohr. Mathematical Gazette, 11(164):318, May 1923. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL http://www.jstor.org/stable/3603781.

Piaggio:1924:RAE

[Pia24a] H. T. H. Piaggio. Review: Atomes et Électrons by H. A. Lorentz, E. Rutherford, M. de Broglie, R. A. Millikan, H. Kamerlingh Onnes, P. Weiss, L. Brillouin, W. H. Bragg, W. J. de Haas, N. Bohr, P. Ehrenfest, pp. vii + 271 + i (1923), (Gauthier-Villars). Mathematical Gazette, 12 (170):117–119, May 1924. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL http://www.jstor.org/stable/3604665.

Piaggio:1924:RBE

[Pia24b] H. T. H. Piaggio. Review: Atomes et Électrons by H. A. Lorentz, E. Rutherford, M. de Broglie, R. A. Millikan, H. Kamerlingh Onnes, P.

Weiss, L. Brillouin, W. H. Bragg, W. J. de Haas, N. Bohr, P. Ehrenfest, pp. vii + 271 + i (1923), (Gauthier-Villars). Mathematical Gazette, 12 (170):117–119, May 1924. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL http://www.jstor.org/stable/3604665.

Pihl:1955:KMG

[Pih55] Mogens Pihl. Den klassiske mekanik i geometrisk beskrivelse: dedicated to Professor Niels Bohr on the occasion of his 70th birthday. (Danish) [Classical mechanics from a geometric description], volume 30(12) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 25 pp. LCCN ????

Pihl:1964:NBOb

[Pih64] Mogens Pihl. Niels Bohr og det danske samfund. (Danish) [Niels Bohr and Danish society]. In Bohr and Rozental [BR64], pages 281–291. LCCN QC16.B63 N5. English translation in [Roz67].

Perraud:2008:DMB

[PKWF08] Simon Perraud, Kiyoshi Kanisawa, Zhao-Zhong Wang, and Toshimasa Fujisawa. Direct measurement of the binding energy and Bohr radius of a single hydrogenic defect in a semiconductor quantum well. Physical Review Letters, 100(5):056806, February 8, 2008. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Planck:1923:BAG

[Pla23] Max Planck. Die Bohrsche Atomtheorie. (German) [The Bohr atomic theory]. Die Naturwissenschaften, 11(27):535–537, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Planck:1924:QBAa

[Pla24a] Max Planck. Zur Quantenstatistik des Bohrschen Atommodells. (German) [On quantum statistics of the Bohr atomic model]. Annalen der Physik (Berlin), 75(??):673–684, ???? 1924. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Planck:1924:QBAb

[Pla24b] Max Planck. Zur Quantenstatistik des Bohrschen Atommodells. (German) [On quantum statistics of the Bohr atomic model]. Annalen der Physik (1900), 380(23):673–684, 1924. ISSN 1521-3889.

Plotnitsky:1994:CAE

[Plo94] Arkady Plotnitsky. Complementarity: anti-epistemology after Bohr and Derrida. Duke University Press, Durham, NC, USA, 1994. ISBN 0-8223-1433-9 (hardcover), 0-8223-1437-1 (paperback). 324 pp. LCCN QC174.17.C63 P55 1994.

Plotnitsky:2006:CEQ

[Plo06a] Arkady Plotnitsky. Complementarity, epistemology, and quantum mechanics as an information theory. In Reading Bohr: physics and philosophy [Plo06b], pages 9–47. ISBN 1-4020-5253-7 (cloth), 1-4020-5254-5 (electronic). LCCN QC174.12 .P624 2006. URL http://www.loc.gov/

catdir/enhancements/fy0824/2007295641-d.html;   http://www.
loc.gov/catdir/toc/fy0706/2007295641.html.

Plotnitsky:2006:RBP

[Plo06b] Arkady Plotnitsky. Reading Bohr: physics and philosophy, volume 152 of Fundamental theories of physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-5253-7 (cloth), 1-4020-5254-5 (electronic). xiii + 222 pp. LCCN QC174.12 .P624 2006. URL http://www.loc.gov/

catdir/enhancements/fy0824/2007295641-d.html;   http://www.
loc.gov/catdir/toc/fy0706/2007295641.html.

Plotnitsky:2010:EPB

[Plo10] Arkady Plotnitsky. Epistemology and Probability: Bohr, Heisenberg, Schrödinger, and the Nature of Quantum-Theoretical Thinking. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-387-85334-0. xxviii + 402 pp. LCCN QC173.98 .P56 2010.

Plotnitsky:2011:RUE

[Plo11] Arkady Plotnitsky. On the reasonable and unreasonable effectiveness of mathematics in classical and quantum physics. Foundations of Physics, 41(3):466–491, March 2011. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/
10.1007/s10701-010-9442-2.

Plotnitsky:2012:NBC

[Plo12] Arkady Plotnitsky. Niels Bohr and complementarity: an introduction. SpringerBriefs in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-4516-7, 1-4614-4517-5 (e-book). ISSN 2191-5423. xxi + 199 pp. LCCN QC174.17.C63 P56 2012.

Papenfuss:2002:YWH

[PLS02] Dietrich Papenfuss, D. (Dieter) Lüst, and Wolfgang Schleich, editors. 100 Years Werner Heisenberg: works and impact. Wiley-VCH, Berlin, Germany, Weinheim, Germany, and New York, NY, USA, 2002. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). ix + 299 pp. LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.loc.gov/

catdir/bios/wiley046/2003268171.html;   http://www.loc.gov/
catdir/description/wiley0310/2003268171.html;   http://www.
loc.gov/catdir/toc/wiley031/2003268171.html.

Pinto:2010:BCB

[POC^{+}10] Yaïr Pinto, Marte Otten, Michael A. Cohen, Jeremy M. Wolfe, and Todd S. Horowitz. The boundary conditions for Bohr’s law: when is reacting faster than acting? Attention, Perception, & Psychophysics, 73 (2):613–620, ???? 2010. ISSN 1943-3921 (print), 1943-393X (electronic).

Podgorsak:2010:RPM

[Pod10a] Ervin B. Podgoršak, editor. Radiation Physics for Medical Physicists. Biological and Medical Physics, Biomedical Engineering; SpringerLink: Bücher. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2010. ISBN 3-642-00874-7, 3-642-00875-5 (e-book). xxxiii + 745 pp. LCCN R895 .P632 2010. URL http://link.

springer.com/10.1007/978-3-642-00875-7.

Podgorsak:2010:RBM

[Pod10b] Ervin B. Podgoršak. Rutherford–Bohr model of the atom. In Radiation Physics for Medical Physicists [Pod10a], pages 139–175. ISBN 3-642-00874-7, 3-642-00875-5 (e-book). LCCN R895 .P632 2010. URL

http://link.springer.com/10.1007/978-3-642-00875-7.

Polkinghorne:1958:BRB

[Pol58] J. C. Polkinghorne. Book review: Atomic Physics and Human Knowledge, by Niels Bohr. Physics Today, 11(8):??, August 1958. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Polak:1960:EQA

[Pol60] L. S. Polak. Die Entstehung der Quantentheorie des Atoms (Das Rutherford–Bohrsche Atommodell). (German) [The emergence of the quantum theory of the atom (the Rutherford–Bohr atomic model)]. In Sowjetische Beiträge zur Geschichte der Naturwissenschaft. (German) [Soviet contributions to the history of natural science] [FH60], pages 226–242. LCCN Q125 1960. DM-Ost 17.50.

Popper:1982:CNG

[Pop82] Karl Popper. A critical note on the greatest days of quantum theory. Foundations of Physics, 12(10):971–976, 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). 10.1007/BF01889270.

Powers:2002:LEH

[Pow02] Thomas Powers. Letter to the Editor: Heisenberg and Bohr. New York Times, ??(??):A18, February 16, 2002. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/92195480/13F386CFDB17FA2143.

Powers:2003:LCP

[Pow03] Thomas Powers. A letter from Copenhagen. New York Review of Books, 50(13):55–56, August 14, 2003. ISSN 0028-7504 (print),

1944-7744 (electronic). URL http://werner-heisenberg.unh.edu;
http://www.nybooks.com/articles/archives/2003/aug/14/a-letter-

from-copenhagen/. Previously unpublished letter from German physicist Werner Heisenberg to his wife Elisabeth about his meeting with Niels Bohr in September 1941.

Perez:2016:BET

[PP16] Enric Pérez and Blai Pié i Valls. Bohr and Ehrenfest: transformations and correspondences in the early 1920s. European Physical Journal H, 41(2):93–136, June 2016. CODEN EPJHAD. ISSN 2102-6459 (print),

2102-6467 (electronic). URL http://link.springer.com/article/
10.1140/epjh/e2016-60028-1.

Park:2009:MDE

[PRKH09] Seung H. Park, Rajib Rahman, Gerhard Klimeck, and Lloyd C. L. Hollenberg. Mapping donor electron wave function deformations at a sub-Bohr orbit resolution. Physical Review Letters, 103(10):106802, September 4, 2009. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Pauli:1955:NBD

[PRW55] W. Pauli, L. Rosenfeld, and V. Weisskopf, editors. Niels Bohr and the development of physics: essays dedicated to Niels Bohr on the occasion of his 70th birthday. McGraw-Hill, New York, NY, USA, 1955. vii + 195 pp. LCCN QC71 .P3 1955.

Pantidos:2001:UDS

[PSV01] Panagiotis Pantidos, Kalliopi Spathi, and Evagelos Vitoratos. The use of drama in science education: The case of “Blegdamsvej Faust”. Science & Education (Springer), 10(1–2):107–117, January 2001. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Purrington:2018:HAC

[Pur18] Robert D. Purrington. The Heroic Age: the Creation of Quantum Mechanics, 1925–1940. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-065517-8 (hardcover). xv + 403 pp. LCCN QC173.98 .P87 2018.

Perez:2015:EAH

[PV15] Enric Pérez and Blai Pié Valls. Ehrenfest’s adiabatic hypothesis in Bohr’s quantum theory. In Aaserud and Kragh [AK15], pages 272– 289. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/2015/
06/Bohr-Book.pdf.

Pauli:1993:WBB

[PvM93] Wolfgang Pauli and Karl von Meyenn, editors. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band III: 1940–1949. (German) [Scientific correspondence with Bohr, Einstein, Heisenberg, and others. Volume 3: 1940–1949]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-54911-0 (print), 3-540-78802-6. lxiv + 1076 pp. LCCN ????

Pyenson:1981:BRB

[Pye81] Lewis Pyenson. Book review: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg, u.a. Vol. I: 1919–1929 by Wolfgang Pauli; A. Hermann; K. V. Meyenn; V. F. Weisskopf. Isis, 72 (3):524–525, September 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230307.

Rabi:19xx:RPB

[Rabxx] I. I. (Isidor Isaac) Rabi. I. I. Rabi papers, 1899–1989 (bulk 1945–1968). 19xx. ???? pp. LCCN 0639F; LCA (Unprocessed ac. 22,752); Vault

0202A; Oversize 6:10. URL http://hdl.loc.gov/loc.mss/eadmss.
ms998009.

Radder:1982:ICL

[Rad82] Hans Radder. An immanent criticism of Lakatos’ account of the ‘degenerating phase’ of Bohr’s atomic theory. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):

99–109, ???? 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170611. See

response [Car83].

Rabinowitch:1967:JRO

[RB67] Eugene Rabinowitch and Hans Bethe. J. Robert Oppenheimer, 1904– 1967. Bulletin of the Atomic Scientists, 23(8):2–6, October 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roldan-Charria:1990:LMQ

[RC90] Jairo Roldan-Charria. Langage, mécanique quantique et réalité: un essai sur la pensée de Niels Bohr. (French) [Language, quantum mechanics, and reality: an essay on the thoughts of Niels Bohr]. PhD thesis, Universitè de Lille, Lille, France, 1990. 370 pp. 1 microfiche.

Roldan-Charria:2014:ICO

[RC14] Jairo Roldán-Charria. Indivisibility, complementarity and ontology: A Bohrian interpretation of quantum mechanics. Foundations of Physics, 44(12):1336–1356, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-014-9823-z.

Rechenberg:1986:BRB

[Rec86] H. Rechenberg. Book review: Scientific Correspondence with Bohr, Einstein, Heisenberg and Others. Volume II: 1930–1939 by Wolfgang Pauli; Karl von Meyenn; Armin Hermann; Victor F. Weisskopf. Isis, 77(2):387–388, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232721.

Rechenberg:2009:WHS

[Rec09] Helmut Rechenberg, editor. Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die “fröhliche Wissenschaft” (Jugend bis Nobelpreis). Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-69221-5 (print), 3-540-69222-3 (e-book). xxi + 1001 pp. LCCN QC16.H35 2009. URL http://ebooks.ciando.com/

book/index.cfm/bok\_id/46829;   http://www.ciando.com/img/
books/width167/3540692223\_k.jpg;   http://www.ciando.com/
pictures/bib/3540692223bib\_t\_1.jpg.

Rechenberg:2010:WHS

[Rec10] Helmut Rechenberg. Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die “fröhliche Wissenschaft” (Jugend bis Nobelpreis). (German) [Werner Heisenberg — the language of atoms: Life and Work — a scientific biography: the “Happy Science” (youth to Nobel Prize)]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-540-69221-5 (hardcover), 3-540-69222-3 (e-book). xxi + 1001 pp. LCCN QC16.H35 R434 2010. URL http:

//ebooks.ciando.com/book/index.cfm/bok_id/46829;   http://
www.ciando.com/img/books/width167/3540692223_k.jpg;   http:
//www.ciando.com/pictures/bib/3540692223bib_t_1.jpg.

Redmond:1963:BEA

[Red63] James R. Redmond. Bohr effect: Absence in a molluscan hemocyanin. Science (New Series), 139(3561):1294–1295, March 29, 1963. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-9203 (electronic).

Redhead:1993:BRU

[Red93] Michael Redhead. Book reviews: Unveiling the atom [Abraham Pais, Niels Bohr’s Times, in Physics, Philosophy and Polity. Clarendon Press, Oxford, 1991. Pp. xvii + 565, £25.00. ISBN 0-19-852049-2]. Notes and Records of the Royal Society of London, 47(1):152–154, July 1993. CO-DEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Reed:2005:BMP

[Ree05] B. Cameron Reed. Bohr’s model predicts ionization potential of twoelectron atoms. Physics Education, 40(2):117, 2005. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://stacks.

iop.org/0031-9120/40/i=2/a=F02.

Reed:2009:PMP

[Ree09] Bruce Cameron Reed. Physics of the Manhattan Project. Trafford Publishing, Victoria, BC, Canada, 2009. ISBN 1-4269-0079-1, 1-4269-0081-3 (e-book). 182 pp. LCCN QC790 .R44 2009.

Reed:2011:PMP

[Ree11] Bruce Cameron Reed. The physics of the Manhattan Project. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2011. ISBN 3-642-14708-9 (hardcover), 3-642-14709-7 (e-book). xiii + 170 pp. LCCN QC790 .R44

2011. URL http://d-nb.info/100407090X/04;   http://www.loc.
gov/catdir/enhancements/fy1113/2010937572-d.html;   http://
www.loc.gov/catdir/enhancements/fy1113/2010937572-t.html.

Reed:2014:PMP

[Ree14] Bruce Cameron Reed. The physics of the Manhattan Project. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2014. ISBN 3-662-43532-2 (hardcover), 3-662-43533-0 (ebook). xvii + 222 pp. LCCN QC790 .R44 2015.

Reed:2015:ABS

[Ree15a] Bruce Cameron Reed. The atomic bomb: the story of the Manhattan Project: how nuclear physics became a global geopolitical game-changer. IOP concise physics. Morgan and Claypool Publishers and IOP Publishing, San Rafael, CA, USA and Bristol, UK, 2015. ISBN 1-62705-990-3 (print), 1-62705-991-1 (e-book), 1-62705-993-8 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). 239 (est.) pp. LCCN QC773.3.U5 R443

2015eb. URL http://iopscience.iop.org/book/978-1-6270-5991-
6.

Reed:2015:BS

[Ree15b] Bruce Cameron Reed. The background science. In The atomic bomb: the story of the Manhattan Project: how nuclear physics became a global geopolitical game-changer [Ree15a], pages 2:1–2:30. ISBN 1-62705-990-3 (print), 1-62705-991-1 (e-book), 1-62705-993-8 (mobi). ISSN 2053-2571 (print), 2054-7307 (electronic). LCCN QC773.3.U5 R443 2015eb. URL

http://iopscience.iop.org/book/978-1-6270-5991-6.

Reines:1972:CFOa

[Rei72] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Colorado Associated University Press, Boulder, CO, USA, 1972. ISBN 0-87081-025-1. xiv + 320 pp. LCCN QC780 .C65.

Reingold:1984:MMA

[Rei84] Nathan Reingold. MGM meets the atomic bomb. The Wilson Quarterly (1976–2012), 8(4):154–163, Autumn 1984. ISSN 0363-3276. URL http:

//www.jstor.org/stable/40256804.

Rentetzi:2015:BRF

[Ren15] Maria Rentetzi. Book review: Finn Aaserud; J. L. Heilbron. Love, Literature, and the Quantum Atom: Niels Bohr’s 1913 Trilogy Revisited. Isis, 106(4):972–973, December 2015. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roll-Hansen:2000:ACB

[RH00] Nils Roll-Hansen. The application of complementarity to biology: From Niels Bohr to Max Delbrück. Historical Studies in the Physical and Biological Sciences, 30(2):417–442, ???? 2000. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Rheault:1991:ADP

[Rhe91] Paul Rheault. Autour du principe de realitè physique quantique: Études sur Planck, Bohr et Heisenberg. M.A. dissertation, Université Laval, Sherbrooke, Quebec, Canada, 1991. 104 pp.

Richardson:1914:ETM

[Ric14] Sir Owen Williams Richardson. The Electron Theory of Matter. Cambridge University Press, Cambridge, UK, 1914. viii + 612 pp. LCCN ????

Richardson:1916:ETM

[Ric16] Sir Owen Williams Richardson. The Electron Theory of Matter. Cambridge University Press, Cambridge, UK, second edition, 1916. vi + 631 pp. LCCN ????

Riepe:1960:BRB

[Rie60] Dale Riepe. Book review: Atomic physics and human knowledge, by Niels Bohr. Philosophy and Phenomenological Research, 21(2):276–277,

December 1960. URL http://www.jstor.org/stable/2104340.

Rigden:2002:HEE

[Rig02] John S. Rigden. Hydrogen: the essential element. Harvard University Press, Cambridge, MA, USA, 2002. ISBN 0-674-00738-7. 280 pp.

LCCN QD181.H1 R54 2002. URL http://catdir.loc.gov/catdir/
toc/fy022/2001051708.html.

RiojaNieto:2017:NBP

[Rio17] Ana Rioja Nieto. Niels Bohr and the philosophy of physics: Twentyfirst-century perspectives. International Studies in the Philosophy of Science, 31(4):429–432, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-

9281 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/02698595.2019.1565209.

Rasmussen:1955:SIS

[RM55] Ebbe Rasmussen and Victor Middelboe. Spectroscopic investigations of separated krypton isotopes: dedicated to Professor Niels Bohr on the occasion of his 70th birthday, volume 30(13) of Matematisk-fysiske meddelelser. Munksgaard, København, Danmark, 1955. 21 + 1 pp. LCCN AS281.

Robertson:1979:EYN

[Rob79] Peter Robertson, editor. The early years: the Niels Bohr Institute 1921–1930. Akademisk Forlag, København, Danmark, 1979. ISBN 87-500-1876-0. 175 pp. LCCN QC47.D43 N537. Foreword by Aage Bohr.

Robinson:2012:SED

[Rob12] Andrew Robinson, editor. The scientists: an epic of discovery. Thames and Hudson, New York, NY, USA, 2012. ISBN 0-500-25191-6. 304 pp. LCCN Q141 .S3712 2012.

Robertson:2015:BNP

[Rob15] Peter Robertson. Birthplace of a new physics — the early history of the Niels Bohr Institute. In Aaserud and Kragh [AK15], pages 481–494. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/2015/
06/Bohr-Book.pdf.

Rodgers:2019:TAS

[Rod19] Glen E. Rodgers. Travelling with the Atom: a Scientific Guide to Europe and Beyond. Royal Society of Chemistry, Cambridge, UK, 2019. ISBN 1-78801-528-2 (paperback), 1-78801-702-1 (e-book). xxxii + 551 pp. LCCN QC171.2 .R63 2020.

Rogers:2010:MIS

[Rog10] Kara Rogers, editor. The 100 Most Influential Scientists of All Time. The Britannica guide to the world’s most influential people. Britannica Educational Publishers, in association with Rosen Educational Services, New York, NY, USA, 2010. ISBN 1-61530-002-3 (library binding). 360 pp. LCCN Q162 .A15 2010.

Roque:2017:OHY

[Roq17] Xavier Roqué. One hundred years of the Bohr atom. Proceedings from a conference. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 64(4):384–385, 2017. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00026980.2017.1405897.

Rosenfeld:1935:PDN

[Ros35] Léon Rosenfeld. La plainte du neutrino. (French) [The neutrino’s complaint]. Journal of Jocular Physics, I:35, 1935.

Rosenfeld:1945:NBE

[Ros45] L. (Léon Jacques Henri Constant) Rosenfeld. Niels Bohr: an essay dedicated to him on the occasion of his sixtieth birthday, October 7, 1945. Noord-Hollandsche Uitgevers Mij, Amsterdam, The Netherlands, 1945. 18 pp. LCCN ????

Rosenfeld:1961:NBE

[Ros61] L. (Léon) Rosenfeld. Niels Bohr: an essay dedicated to him on the occasion of his sixtieth birthday, October 7, 1945. North-Holland Publishing Co., Amsterdam, The Netherlands, second corrected edition, 1961. 20 pp. LCCN QC16.B55 R6 1961.

Rosenfeld:1963:NBC

[Ros63] L. Rosenfeld. Niels Bohr’s contribution to epistemology. Physics Today, 16(10):47–54, October 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Rosenfeld:1964:KKO

[Ros64] Léon Rosenfeld. Komplementaritetssynspunktet konsolideres og udbygges. (Danish) [The complementarity view is consolidated and extended]. In Bohr and Rozental [BR64], pages 109–131. LCCN QC16.B63 N5. English translation in [Roz67].

Rosenfeld:1972:NR

[Ros72] Léon Rosenfeld. Nuclear reminiscences. In Reines [Rei72], pages 289– 299. ISBN 0-87081-025-1. LCCN QC780 .C65. URL http://adsabs.

harvard.edu/abs/1972cht..conf..289R.

Rosenfeld:1979:ECB

[Ros79a] Léon Rosenfeld. The epistemological conflict between Einstein and Bohr (dedicated to Max Born on his 80th birthday). In Cohen and Stachel [CS79a], chapter II.10, pages 517–521. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346.

LCCN Q174 .B67 vol. 21 QC7. URL http://link.springer.com/
chapter/10.1007/978-94-009-9349-5_35.

Rosenfeld:1979:NBE

[Ros79b] Léon Rosenfeld. Niels Bohr: An essay dedicated to him on the occasion of his sixtieth birthday, October 7, 1945 (1945; 2nd edition 1961). In Cohen and Stachel [CS79a], chapter I.22, pages 313–326. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21 QC7. URL http://link.springer.

com/chapter/10.1007/978-94-009-9349-5_22.

Rosenfeld:1979:NBC

[Ros79c] Léon Rosenfeld. Niels Bohr’s contribution to epistemology. In Cohen and Stachel [CS79a], chapter II.11, pages 522–535. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21 QC7. URL http://link.springer.

com/chapter/10.1007/978-94-009-9349-5_36.

Rosenfeld:1979:QEA

[Ros79d] Léon Rosenfeld. On quantum electrodynamics (among essays dedicated to Niels Bohr on the occasion of his 70th birthday). In Cohen and Stachel [CS79a], chapter II.3, pages 413–441. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346.

LCCN Q174 .B67 vol. 21 QC7. URL https://link.springer.com/
book/10.1007/978-94-009-9349-5.

Rosenfeld:1979:QTR

[Ros79e] Léon Rosenfeld. Quantum theory in 1929: Recollections from the first Copenhagen conference (1971). In Cohen and Stachel [CS79a], chapter I.21, pages 302–12. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21 QC7.

URL https://link.springer.com/book/10.1007/978-94-009-9349-
5.

Roseberg:1992:NBP

[Rös92] U. Röseberg. Niels Bohr — Physiker und Philosoph. (German) [Niels Bohr — physicist and philosopher]. Physik in unserer Zeit, 23(4):182, ???? 1992. CODEN PHUZAH. ISSN 0031-

9252. URL http://onlinelibrary.wiley.com/doi/10.1002/piuz.
19920230413/abstract.

Roseberg:1994:HHC

[Rös94] Ulrich Röseberg. Hidden historicity: The challenge of Bohr’s philosophical thought. In Faye and Folse [FF94], chapter 15, pages 325–343. ISBN 0-7923-2378-5 (hardcover), 94-015-8106-1 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 153 1994; QC16 1994. URL https://

link.springer.com/chapter/10.1007/978-94-015-8106-6_15.

Roseberg:1995:DTJ

[Rös95] Ulrich Röseberg. Did they just misunderstood each other? Logical empiricists and Bohr’s complementarity argument. In Gavroglu et al. [GSW95], pages 105–124. ISBN 0-7923-2991-0 (set), 0-7923-2988-0. LCCN Q174 .B67 vol. 163. URL http://catdir.

loc.gov/catdir/enhancements/fy0823/94022250-d.html;   http:
//catdir.loc.gov/catdir/enhancements/fy0823/94022250-t.html.

Rose:2000:FCP

[Ros00] Paul Lawrence Rose. Frayn’s Copenhagen plays well, at history’s expense. The Chronicle of Higher Education, ??(??):??, May 5, 2000. ISSN 0009-5982 (print), 1931-1362 (electronic). URL https://www.

chronicle.com/article/Frayns-Copenhagen-Plays/22341.

Rosenkranz:2018:TDA

[Ros18] Ze’ev Rosenkranz, editor. The travel diaries of Albert Einstein. The Far East, Palestine, and Spain, 1922–1923. Princeton University Press, Princeton, NJ, USA, 2018. ISBN 0-691-17441-5. 256 pp. LCCN Q124.6-127.2.

Rothenberg:1987:EBC

[Rot87] Albert Rothenberg. Einstein, Bohr, and creative thinking in science. History of Science (UK), 25(2):147–166, June 1987. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/25/2/147.full.pdf+html.

Rotblat:2000:PJR

[Rot00] Joseph Rotblat. My early years as a physicist in Poland: a talk given to the Group on Monday 8th March 1999. IOP History of Physics Group Newsletter, 13:9–23, Spring 2000. ISSN

1756-168X. URL http://www.iop.org/activity/groups/subject/
hp/newsletter/archive/file_66506.pdf.

Rowlands:2009:LC

[Row09] Peter Rowlands. The Liverpool cyclotrons. IOP History of Physics Newsletter, 25:31–46, February 2009. ISSN 1756-168X. URL http:/

/www.iop.org/activity/groups/subject/hp/newsletter/archive/
file_64509.pdf.

Rozental:1964:FDF

[Roz64a] Stefan Rozental. Forord. (Danish) [Foreword]. In Bohr and Rozental [BR64], pages 7–8. LCCN QC16.B63 N5. English translation in [Roz67].

Rozental:1964:FOH

[Roz64b] Stefan Rozental. Fyrrerne og halvtredserne. (Danish) [Forties and fifties]. In Bohr and Rozental [BR64], pages 145–183. LCCN QC16.B63 N5. English translation in [Roz67].

Rozental:1967:NBH

[Roz67] S. (Stefan) Rozental, editor. Niels Bohr: his life and work as seen by his friends and colleagues. North-Holland Publishing Co., Amsterdam, The Netherlands, 1967. 355 pp. LCCN QC16.B63 N53.

Rozental:1985:NBH

[Roz85] S. (Stefan) Rozental, editor. Niels Bohr: his life and work as seen by his friends and colleagues. North-Holland personal library. Elsevier, Amsterdam, The Netherlands, 1985. ISBN 0-444-86977-8 (US). 355 + 40 pp. LCCN QC16.B63 N53 1985.

Rozental:1998:NB

[Roz98] Stefan Rozental. Niels Bohr: Memoirs of a Working Relationship. Christian Ejlers, Copenhagen, Denmark, 1998. ISBN 87-7241-870-2. 172 pp. LCCN ????

Reeben:2000:GLA

[RR00] M. Reeben and V. Reeben. Is the general law of atomic spectral series by Balmer/Rydberg/Bohr valid for absence epilepsy seizures frequencies as well? European Journal of Neuroscience, 12(Supplement 11):208, ???? 2000. ISSN 0953-816X.

Rosenfeld:1964:NBH

[RRK^{+}64] Léon Rosenfeld, Erik Rüdinger, Oskar Klein, Werner Heisenberg Hendrik G. B. Casimir, Otto Robert Frisch, Stefan Rozental, Aage Bohr, Abraham Pais, Jørgen Kalckar, Christian Møller, Mogens Pihl, Viktor F. Weisskopf, Johannes Pedersen, Viggo Kampmann, Richard Courant, Paul A. M. Dirac, Hans Henrik Koch, William Scharff, Mogens Andersen, Hans Bohr, and Niels Bohr. Niels Bohr: Hans liv or virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: His life and works told by a group of friends and co-workers]. J. H. Schultz Forlag, Copenhagen, DK, 1964. 341 pp. LCCN ????

Rigden:2007:BND

[RS07] John S. Rigden and Roger H. Stuewer. Book notes [David Lindley, Uncertainty: Einstein, Heisenberg, Bohr and the Struggle for the Soul of Science (New York: Doubleday, 2007, 257 pages, $32.00); Stuart Clark, The Sun Kings: The Unexpected Tragedy of Richard Carrington and the Tale of How Modern Astronomy Began (Princeton: Princeton University Press, 2007, xii + 211 pages, $24.95)]. Physics in Perspective (PIP), 9(3):375–377, September 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.com/

article/10.1007/s00016-007-0360-4.

Rubinin:1997:MNB

[Rub97a] P. E. Rubinin. In memory of Niels Bohr (110th anniversary). In Anonymous [Ano97], page ?? ISBN 5-86656-073-9 (hardcover). LCCN ????

Rubinin:1997:NBP

[Rub97b] P. E. Rubinin. Niels Bohr and Petr Léonidovich Kapitza. Physics-Uspekhi, 40(1):95–100, 1997. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL http://stacks.iop.org/1063-

7869/40/i=1/a=A05.

Ruhla:1992:PCB

[Ruh92] Charles Ruhla. The physics of chance: from Blaise Pascal to Niels Bohr. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1992. ISBN 0-19-853960-6 (hardcover), 0-19-853977-0 (paperback). xi + 222 pp. LCCN QC174.8 .R85 1992. US$37.50 (hardcover), US$18.75 (paperback). Translated from the French by G. Barton.

Richardson:1996:RSH

[RW96] W. Mark Richardson and Wesley J. Wildman, editors. Religion and Science: History, Method, Dialogue. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1996. ISBN 0-415-91666-6 (hardcover), 0-415-91667-4 (paperback). xx + 450 pp. LCCN BL240.2 .R43

1996. URL http://catdir.loc.gov/catdir/enhancements/fy0651/
95047045-d.html.

Reiter:2013:ESY

[RY13] Wolfgang Reiter and Jakob Yngvason, editors. Erwin Schrödinger — 50 Years After. European Mathematical Society Publishing House, Zürich, Switzerland, 2013. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (ebook). viii + 185 pp. LCCN QC16.S265 I58 2011.

Rynasiewicz:1983:BRB

[Ryn83] Robert Rynasiewicz. Book review: Ideas of Matter: From Ancient Times to Bohr and Einstein by Mendel Sachs. Isis, 74(3):423–424, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232610.

Rynasiewicz:2015:CP

[Ryn15] Robert Rynasiewicz. The (?) correspondence principle. In Aaserud and Kragh [AK15], pages 175–199. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Soderqvist:2006:VKD

[S^{+}06] Thomas Söderqvist et al., editors. Videnskabernes København. (Danish). [The sciences’ Copenhagen]. Roskilde Universitetsforlag, Roskilde, Danmark, 2006. ISBN 87-7867-038-1, 87-7867-970-2. ???? pp. LCCN Q127.D4 V53 1998. DKR 76,00. URL http:

//elibrary.ebib.dk/Home/html/Gateway.asp?institutionid=22&
ISBN13=9788778679703.

Sachs:1981:IMA

[Sac81] Mendel Sachs. Ideas of matter: from ancient times to Bohr and Einstein. University Press of America, Washington, DC, USA, 1981. ISBN 0-8191-1615-7, 0-8191-1616-5 (paperback). xi + 322 pp. LCCN QC171.2 .S23 1981.

Sachs:1988:EVB

[Sac88] Mendel Sachs. Einstein versus Bohr: the continuing controversies in physics. Open Court, La Salle, IL, USA, 1988. ISBN 0-8126-9064-8, 0-8126-9065-6 (paperback). xxiii + 296 pp. LCCN QC7 .S2 1988.

Salvia:2019:ECP

[Sal19] Stefano Salvia. Embattled cooperation(s): Peaceful atoms, pacifist physicists, and partisans of peace in the Early Cold War (1947–1957). Physics in Perspective (PIP), 21(1):43–62, March 2019. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

SandovalVallarta:1924:BAM

[San24] Manuel Sandoval Vallarta. Bohr’s Atomic Model from the Standpoint of the General Theory of Relativity and of the Calculus Of Perturbations. Ph.D dissertation, Massachusetts Institute of Technology, Cambridge, MA, USA, 1924.

Suwa:1972:BFI

[SB72] K. Suwa and H. H. Bendixen. The Bohr factor: Is it constant? Federation Proceedings, 31(2):355, ???? 1972. CODEN FEPRA7. ISSN 0014-9446.

Scerri:1994:PNH

[Sce94] Eric R. Scerri. Prediction of the nature of hafnium from chemistry, Bohr’s theory and quantum theory. Annals of Science, 51(2):137–150, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Scerri:2016:TSS

[Sce16] Eric R. Scerri. A tale of seven scientists and a new philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-023299-4 (hardcover), 0-19-023300-1 (e-book). xxxiv + 228 pp. LCCN Q175 .S3044 2016.

Schrodinger:1922:DBF

[Sch22] Erwin Schrödinger. Dopplerprinzip und Bohrsche Frequenzbedingung. (German) [Doppler Principle and Bohr’s frequency condition]. Physikalische Zeitschrift, 23(15):301–303, August 1922. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://babel.hathitrust.org/cgi/pt?
id=mdp.39015086723239%3Bview=1up%3Bseq=367.

Schrodinger:1924:BNS

[Sch24] Erwin Schrödinger. Bohrs neue Strahlungshypothese und der Energiesatz (German) [Bohr’s new radiation hypothesis and the energy theorem]. Die Naturwissenschaften, 12(36):720–724, ???? 1924. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Schrodinger:1935:GSQb

[Sch35a] E. Schrödinger. Die gegenwärtige Situation in der Quantenmechanik. (German) [The present situation in quantum mechanics]. Die Naturwissenschaften, 23(49):823–828, December 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/hl6541202ukn4741/fulltext.pdf. En-

glish translation in [Tri80].

Schrodinger:1935:GSQc

[Sch35b] E. Schrödinger. Die gegenwärtige Situation in der Quantenmechanik. (German) [The present situation in quantum mechanics]. Die Naturwissenschaften, 23(50):844–849, December 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/m489741169604131/fulltext.pdf. En-

glish translation in [Tri80].

Schrodinger:1935:GSQa

[Sch35c] Erwin Schrödinger. Die gegenwärtige Situation in der Quantenmechanik. (German) [The present situation in quantum mechanics]. Die Naturwissenschaften, 23(48):807–812, November 1935. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

www.springerlink.com/content/l50v426520375016/fulltext.pdf.

English translation in [Tri80].

Schilpp:1949:AEPa

[Sch49a] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 1 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Schilpp:1949:AEPb

[Sch49b] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 2 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Schmiele:1954:SGS

[Sch54] Walter Schmiele. Skandinavische Geisteswelt von Swedenborg bis Niels Bohr. (German) Scandinavian world of thought from Swedenborg to Niels Bohr. Geist des Abendlandes. Holle, Darmstadt, West Germany, 1954. 340 pp. LCCN PT7095 .S35.

Scharff:1964:MFT

[Sch64a] William Scharff. Minder fra Tisvilde. (Danish) [Memories from Tisvilde]. In Bohr and Rozental [BR64], pages 306–310. LCCN QC16.B63 N5. English translation in [Roz67].

Scheibe:1964:BRAb

[Sch64b] E. Scheibe. Book review: Atomphysik und menschliche Erkenntnis (Atomic Physics and Human Understanding), by Niels Bohr. Series: Die Wissenschaft (Science), Collections of monographs from all fields of natural science. Edited by W. Westphal. Angewandte Chemie, International Edition in English, 3(3):244, March 1964. CODEN ACIEAY. ISSN 0570-

0833. URL http://onlinelibrary.wiley.com/doi/10.1002/anie.
196402442/abstract.

Scheibe:1964:BRAa

[Sch64c] E. Scheibe. Book review: Atomphysik und menschliche Erkenntnis. Von Niels Bohr. Reihe: Die Wissenschaft, Sammlungen v. Einzeldarst. aus allen Gebieten der Naturwissenschaft. Herausgeg. v. W. Westphal. Verlag Friedr. Vieweg & Sohn, Braunschweig 1958. 1. Aufl., XIII, 104 S., 8 Abb., geb. DM 9.40. Angewandte Chemie, 76(7):311, April 7, 1964. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/ange.
19640760729/abstract.

Schmidt:1968:FAT

[Sch68] Egon Schmidt. Fra Arkimedes til Niels Bohr. Træk af fysikkens historie. (Danish) [From Archimedes to Niels Bohr. Traits of the history of physics]. Gyldendal, Oslo, Norway, second edition, 1968. 160 pp. LCCN Q141 .S29 1968. Tegninger af Svend Erik Jensen.

Schweitzer:1977:DIB

[Sch77] P. J. Schweitzer. A definite integral of N. Bohr. SIAM Review, 19(1): 147, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schopf:1988:GBS

[Sch88] Hans-Georg Schöpf. Zur Geschichte der Bohr–Sommerfeldschen Quantentheorie. (German) [On a history of the Bohr–Sommerfeld quantum theory]. Annalen der Physik (1900), 500(8):595–604, 1988. ISSN 1521-3889.

Schweber:1990:YJC

[Sch90] S. S. Schweber. The young John Clarke Slater and the development of quantum chemistry. Historical Studies in the Physical and Biological Sciences, 20(2):339–406, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.jstor.org/stable/

27757647.

Schrag:2000:HHP

[Sch00a] Peter Schrag. History’s Heisenberg Principle. The American Prospect, 11(18):40–??, November 9, 2000. CODEN APROEY. ISSN

1049-7285. URL http://connection.ebscohost.com/c/articles/
3342279/historys-heisenberg-principle;   http://prospect.org/
article/historys-heisenberg-principle.

Schwarzschild:2000:BHS

[Sch00b] Bertram Schwarzschild. Bohr–Heisenberg Symposium marks Broadway opening of Copenhagen. Physics Today, 53(5):51–52, May 2000. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v53/i5/p51/s1.

Schwarzschild:2002:BLC

[Sch02] Bertram Schwarzschild. Bohr letters clarify mystery. Physics Today, 55 (3):??, March 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Schwarz:2011:STN

[Sch11] Stephan Schwarz. Science, technology, and the Niels Bohr Institute in Occupied Denmark. Physics in Perspective (PIP), 13(4):401– 432, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-011-0059-4.

Schwarz:2013:ABM

[Sch13] W. H. Eugen Schwarz. 100th anniversary of Bohr’s model of the atom. Angewandte Chemie, International Edition, 52(47):12228–12238, 2013. CODEN ACIEF5. ISSN 1433-7851 (print), 1521-3773 (electronic). Dedicated to Professor Werner Kutzelnigg on the occasion of his 80th birthday.

Schirrmacher:2015:BGM

[Sch15] Arne Schirrmacher. Bohr’s genuine metaphor: On types, aims and uses of models in the history of quantum theory. In Aaserud and Kragh [AK15], pages 111–140. ISBN 87-7304-387-7. ISSN 1904-

5514. LCCN QC172 .O63 2015. URL http://www.ijqf.org/wps/wp-
content/uploads/2015/06/Bohr-Book.pdf.

Schwarz:2021:ONB

[Sch21] Stephan Schwarz. The occupation of Niels Bohr’s Institute: December 6, 1943–February 3, 1944. Physics in Perspective (PIP), 23(1):49–82, March 2021. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Simoes:2015:SUE

[SDG15] Ana Simões, Maria Paula Diogo, and Kostas Gavroglu. Sciences in the Universities of Europe, Nineteenth and Twentieth Centuries: Academic Landscapes, volume 309 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 94-017-9635-1, 94-017-9636-X (e-book). ISSN 0068-0346. x + 393 pp. LCCN Q183.4.E85 2015. URL https://

link.springer.com/book/10.1007/978-94-017-9636-1.

Sommerfeld:2013:BSA

[SE13] Arnold Sommerfeld and Michael Eckert. Die Bohr–Sommerfeldsche Atomtheorie: Sommerfelds Erweiterung des Bohrschen Atommodells 1915/16. (German) [The Bohr–Sommerfeld atomic theory: Sommerfeld’s expansion of Bohr’s Atommodell 1915/16]. Klassische Texte der Wissenschaft. Springer Spektrum, Berlin, Germany, 2013. ISBN 3-642-35114-X (paperback), 3-642-35115-8 (e-book). xi + 151 pp. LCCN QC173 .S52 2013.

Segre:1976:PSN

[Seg76] Emilio Segrè. Personaggi e scoperte nella fisica contemporanea. (Italian) [Personalities and discoveries in contemporary physics]. Biblioteca della EST. Edizioni scientifiche e tecniche Mondadori, Milano, Italia, 1976. 297 pp. LCCN QC7 .S44. Ciclo di lezioni tenute dal nov. 1972 fino al marzo 1973.

Segre:1980:XRQ

[Seg80] Emilio Segrè. From X-rays to quarks: modern physicists and their discoveries. W. H. Freeman, New York, NY, USA, 1980. ISBN 0-7167-1146-X, 0-7167-1147-8 (paperback). ix + 337 pp. LCCN QC7 .S4413. English translation of [Seg76].

Segre:1983:PSN

[Seg83] Emilio Segrè. Personaggi e scoperte nella fisica contemporanea: dai raggi X ai quark. Biblioteca della EST, 0303-2752. Edizioni scientifiche e tecniche Mondadori (IS), Milano, Italia, second edition, 1983. 297 pp. LCCN ????

Segre:1984:PML

[Seg84] Emilio Segrè. Les physiciens modernes et leurs découvertes: des rayons X aux quarks. (French) [Modern physicists and their discoveries: from X-rays to quarks]. Temps des sciences. Fayard, Paris, France, 1984. 456 pp. LCCN QC7 .S4414 1984. French translation of [Seg76].

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452(1):xv–xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June

Segre:1986:MKR

[Seg86a] Emilio Segre. Mi-karne Ranotgen ove-ad kvarkim: fisikaim modernim ove-tagliyotehe. Keter, Yerushalayim, Israel, 1986. 331 pp. LCCN ???? Hebrew translation of [Seg76].

Segre:1986:WLM

[Seg86b] Emilio Segrè. Wu li ming ren he wu li fa xian. Zhi shi chu ban she, Shanghai, People’s Republic of China, 1986. iii + 366 pp. LCCN QC7 .S4412. Mandarin Chinese translation by Zuwei Liu of [Seg76].

Segre:1987:RXA

[Seg87] Emilio Segre. Dos raios X aos quarks: fisicos modernos e suas descobertas. (Portuguese) [From X-rays to quarks: modern physicists and their discoveries], volume 24 of Pensamento cientifico. Ed. UnB, Brasilia, Brazil, 1987. ISBN 85-230-0078-X. 345 pp. LCCN ????

Segala:1993:BRAa

[Seg93] Marco Segala. Book review: Abraham Pais, Niels Bohr’s Times, in physics, philosophy, and polity, Oxford, Clarendon Press, 1991, xvii + 565 pp. ill. (ISBN 0-19-852049-2). Nuncius, 8(1):364–367, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00343.

Segre:2007:XRQ

[Seg07a] Emilio Segrè. From X-rays to quarks: modern physicists and their discoveries. Dover classics of science and mathematics. Dover, New York, NY, USA, 2007. ISBN 0-486-45783-4. ix + 339 pp. LCCN

QC7 .S4413 2007. URL http://www.loc.gov/catdir/enhancements/
fy0702/2006102450-d.html.

Segre:2007:FCS

[Seg07b] Gino Segrè. Faust in Copenhagen: a struggle for the soul of physics. Viking, New York, NY, USA, 2007. ISBN 0-670-03858-X. x + 310 pp. LCCN QC15 .S427 2007. URL http://www.loc.gov/

catdir/enhancements/fy0743/2006052807-b.html;   http://www.
loc.gov/catdir/enhancements/fy0743/2006052807-d.html.

Sergescu:1958:TDP

[Ser58] P. Sergescu. Review: L’évolution de la notion de phénomène physique des primitifs à Bohr et Louis de Broglie by J. Pelseneer. Revue philosophique de la France et de l’Étranger, 148:102–103, 1958. ISSN 0035-3833 (print), 2104-385X (electronic). URL http://www.jstor.

org/stable/41089669.

Seth:2003:PPC

[Set03] Suman Seth. Principles and problems: Constructions of theoretical physics in Germany, 1890–1918. Ph.D. dissertation, Department of History, Princeton University, Princeton, NJ, USA, November 2003. xiii +

315 pp. URL https://search.proquest.com/pqdtglobal/docview/
288241277/.

Seth:2010:CQA

[Set10] Suman Seth. Crafting the quantum: Arnold Sommerfeld and the practice of theory, 1890–1926. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01373-8 (hardcover). vii + 378 pp. LCCN QC16.S76 2010. US$23.95.

Shull:1959:AU

[SH59] H. Shull and G. G. Hall. Atomic units. Nature, 184(4698):1559–1560, November 14, 1959. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Sharlin:1967:BRB

[Sha67] Harold I. Sharlin. Book review: Niels Bohr: The Man, His Science, & the World They Changed. By Ruth Moore. (New York: Alfred A. Knopf, 1966. xvi + 436 + vii pp. Illustrations and index. $6.95). The Journal of American History, 54(1):193–194, 1967. ISSN 0021-8723 (print), 1945-2314 (electronic). URL http://jah.

oxfordjournals.org/content/54/1/193.full.pdf+html;   http:/
/jah.oxfordjournals.org/content/54/1/193.short.

Shamos:1987:GEPa

[Sha87] Morris H. (Morris Herbert) Shamos, editor. Great experiments in physics: firsthand accounts from Galileo to Einstein. Dover, New York, NY, USA, 1987. ISBN 0-486-25346-5 (paperback). viii + 370 pp. LCCN QC7 .G74 1987.

Sherwin:1972:NBA

[She72] Martin J. Sherwin. Niels Bohr and the atomic bomb: the scientific ideal and international politics, 1943–1944. In Proceedings of the History of the 20th Century Physics conference, Varenna, Italy, 31 July–12 August, 1972, page ?? ????, ????, 1972.

Sherwin:1985:NBF

[She85a] Martin J. Sherwin. Niels Bohr and the first principles of arms control. In ????, editor, Niels Bohr: Physics and the World, Niels Bohr Centennial Symposium, American Academy of Arts and Sciences, Cambridge, Mass., November 12–14, 1985, page ?? ????, ????, 1985.

Sherwin:1985:RHW

[She85b] Martin J. Sherwin. Retrospectives: How well they meant. Bulletin of the Atomic Scientists, 41(7):9–15, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherwin:1986:NBS

[She86] Martin J. Sherwin. Niels Bohr: spurned prophet of arms control. Bulletin of the Atomic Scientists, 42(9):41–45, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shimony:1983:RPB

[Shi83] Abner Shimony. Reflections on the philosophy of Bohr, Heisenberg, and Schrödinger. In Cohen and Laudan [CL83], chapter 11, pages 209–221. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/978-94-
009-7055-7_11.

Shimony:1985:BRB

[Shi85a] Abner Shimony. Book review: The Philosophy of Niels Bohr: The Framework of Complementarity, by Henry Folse. Physics Today, 38(10): ??, October 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [Fol85].

Shimony:1985:BRH

[Shi85b] Abner Shimony. Book review: Henry Folse, The Philosophy of Niels Bohr: The Framework of Complementarity. Physics Today, 38(10):108– 109, October 1, 1985. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.2814739.

Shimony:1986:RPB

[Shi86] Abner Shimony. Reflections on the philosophy of Bohr, Heisenberg, and Schrödinger. In Cohen and Wartofsky [CW85], chapter 20, pages 305–317. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346.

LCCN ???? URL https://link.springer.com/chapter/10.1007/
978-94-009-5345-1_20.

Shomar:2008:BPR

[Sho08] Towfic Shomar. Bohr as a phenomenological realist. Journal for General Philosophy of Science = Zeitschrift für allgemeine Wissenschaftstheorie, 39(2):321–349, December 2008. CODEN JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

40390660;   http://www.jstor.org/stable/i40017446;   http:/
/www.springerlink.com/content/x50856624053n342/.

Shorter:2013:NBQ

[Sho13] R. S. Shorter. Niels Bohr and the Quantum Atom: The Bohr Model of Atomic Structure 1913–1925, by Helge Kragh, Scope: biography, general interest. Level: general readership. Contemporary Physics, 54(2): 133, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Siegfried:2013:WAW

[Sie13] Tom Siegfried. When the atom went quantum: Bohr’s revolutionary atomic theory turns 100. Science News (Washington, DC), 184(1):20–24, July 13, 2013. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930

(electronic). URL http://www.sciencenews.org/view/feature/id/
351277/description/When_the_atom_went_quantum.

Sime:1996:LML

[Sim96] Ruth Lewin Sime. Lise Meitner: a life in physics, volume 13 of California studies in the history of science. University of California Press, Berkeley, CA, USA, 1996. ISBN 0-520-08906-5 (hardcover), 0-520-20860-9 (paperback), 0-585-05524-6 (e-book). xiii + 526 pp.

LCCN QC774.M4 S56 1996. URL http://www.loc.gov/catdir/bios/
ucal051/95035246.html;   http://www.loc.gov/catdir/description/
ucal041/95035246.html.

Sime:2000:STE

[Sim00] Ruth Lewin Sime. The search for transuranium elements and the discovery of nuclear fission. Physics in Perspective (PIP), 2(1):48– 62, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://www.springerlink.com/content/
nr2t13tndn6t9t72/.

Sime:2012:PFO

[Sim12] Ruth Lewin Sime. The politics of forgetting: Otto Hahn and the German Nuclear-Fission Project in World War II. Physics in Perspective (PIP), 14(1):59–94, March 2012. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL

http://link.springer.com/article/10.1007/s00016-011-0065-6;
 http://www.springerlink.com/content/k12202vg92147h68/.

Simoes:2014:SBR

[Sim14] Ana Simões. Scientific biographies revisited: Thomsons’ electrons and Bohr’s quantum atoms. Ambix: Journal of the Society for the History of Alchemy and Chemistry, 61(1):95–99, 2014. CODEN AMBXAO. ISSN 0002-6980 (print), 1745-8234 (electronic). URL http://www.

tandfonline.com/doi/full/10.1179/0002698014Z.00000000047.

Shampo:1975:NB

[SK75] M. A. Shampo and R. A. Kyle. Niels Bohr. JAMA: the journal of the American Medical Association, 233(9):992, September 1, 1975. CODEN JAMAAP. ISSN 0098-7484 (print), 1538-3598 (electronic).

Segre:1969:GMP

[SKST69] Emilio G. Segrè, Joseph Kaplan, Leonard I. Schiff, and Edward Teller. Great men of physics: the humanistic element in scientific work. The University of California letters and science extension series. Tinnon-Brown, Los Angeles, CA, USA, 1969. vi + 110 pp. LCCN ???? With an introduction by Marvin L. Chachere.

Slater:1985:WP

[Sla85] John C. Slater. Waves and particles: 1923–1924. In French and Kennedy [FK85], pages 160–162. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Spangenburg:1995:NBG

[SM95] Ray Spangenburg and Diane Moser. Niels Bohr: gentle genius of Denmark. Makers of modern science. Facts on File, New York, NY, USA, 1995. ISBN 0-8160-2938-5 (hardcover), 1-4381-4897-6 (e-book). xi + 116 pp. LCCN QC16.B63 S63 1995.

Spangenburg:2008:NBA

[SM08] Ray Spangenburg and Diane Kit Moser, editors. Niels Bohr: Atomic Theorist. Makers of modern science. Chelsea House, New York, NY, USA, revised edition, 2008. ISBN 0-8160-6178-5. xvi + 141 pp. LCCN QC16.B63 S63 2008.

Smith:1976:SBH

[Smi76] Brian Clarke Smith. Solution of the Bohr Hamiltonian with Application to High Spin States. Ph.D dissertation, McMaster University, Hamilton, ON, Canada, October 1976.

Smorodinski:1993:BCD

[Smo93] Ya A. Smorodinski. Blegdamsvej 17, Copenhagen, Denmark. Niels Bohr Institute. Physics-Uspekhi, 36(3):197–??, March 31, 1993. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL http:

//stacks.iop.org/1063-7869/36/i=3/a=B09.

Strode:1993:ORD

[SO93] Therkel Strøde and H. Rovsing Olsen, editors. October 1943: the rescue of the Danish Jews from annihilation. Royal Danish Ministry of Foreign Affairs, Copenhagen, Denmark, 1993. 23 pp. LCCN DS135.D4 O28 1993.

Sommerfeld:1915:ADN

[Som15] Arnold Sommerfeld. Die allgemeine Dispersionsformel nach dem Bohrschen Modell. (German) [The general dispersion formula according to Bohr’s model]. In Bergwitz [Ber15], pages 549–584. LCCN QD455

.F47 1915. URL http://catalog.hathitrust.org/api/volumes/
oclc/47733990.html.

Sommerfeld:1918:DDS

[Som18] Arnold Sommerfeld. Die Drudesche Dispersionstheorie vom Standpunkte des Bohrschen Modelles und die Konstitution von H_{2}, O_{2} und N_{2}. (German) [Drudian dispersion theory from the standpoint of Bohr’s model and the constitution of H_{2}, O_{2}, and N_{2}]. Annalen der Physik (1900), 358(15):497–550, 1918. ISSN 0003-3804 (print), 1521-3889 (electronic).

Sommerfeld:1920:GZB

[Som20a] Arnold Sommerfeld. Grundlagen und Ziele der Bohr’schen Theorie von Atomen und Spektren. (German) [Fundamentals and objectives of Bohr’s theory of atoms and spectra]. Export, Berlin, ??(??):258–260, ???? 1920.

Sommerfeld:1920:XHD

[Som20b] Arnold Sommerfeld. XXV. Hauptversammlung der Deutschen Bunsen-Gesellschaft für angewandte physikalische Chemie, E. V. Grundlagen und Ziele der Bohrschen Theorie von Atomen und Spektren. (German) [XXV. General Assembly of the German Bunsen Society for Applied Physical Chemistry, E. V.. Principles and objectives of Bohr’s theory of atomic and spectra]. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 26(13–14):258–260, 1920. CODEN ZEA-PAA. ISSN 0372-8323.

Sommerfeld:1920:VGZ

[Som20c] Arnold Sommerfeld. Zum Vortrag: Grundlagen und Ziele der Bohr’schen Theorie von Atomen und Spektren. — Diskussion und Antworten A. Sommerfelds. (German) [On the lecture: Basics and objectives of Bohr’s theory of atoms and spectra. — Discussion and answers by A. Sommerfeld]. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 26(??):489–490, ???? 1920. CODEN ZEAPAA. ISSN 0372-8323.

Sommerfeld:1920:KBT

[Som20d] Arnold Sommerfeld. Zur Kritik der Bohrschen Theorie der Lichtemission. (German) [Critique of Bohr’s theory of light emission]. Jahrbuch der Radioactivität und Elektronik, Leipzig, 17(??):417–429, ???? 1920.

Sommerfeld:1924:GQBa

[Som24a] Arnold Sommerfeld. Die Grundlagen der Quantentheorie und des Bohr’schen Atommodelles. (German) [The fundamentals of quantum theory and Bohr’s atomic model]. Zeitschrift für Angewandte Chemie, 37(??):787–788, ???? 1924. ISSN 0932-2132.

Sommerfeld:1924:GQBd

[Som24b] Arnold Sommerfeld. Die Grundlagen der Quantentheorie und des Bohr’schen Atommodelles. (German) [The fundamentals of quantum theory and Bohr’s atomic model]. Chemiker-Zeitung, 48(??):721–722, ???? 1924.

Sommerfeld:1924:GQBc

[Som24c] Arnold Sommerfeld. Grundlagen der Quantentheorie und des Bohrschen Atommodelles. Innsbrucker Vortrag. (German) [Fundamentals of quantum theory and Bohr’s atomic model. Innsbruck lecture]. Die Naturwissenschaften, 12(47):1047–1049, November 21, 1924. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01452116.

Sommerfeld:1924:GQBb

[Som24d] Arnold Sommerfeld. Grundlagen der Quantentheorie und des Bohr’schen Atommodells. (German) [Fundamentals of quantum theory and the Bohr atomic model]. Verhandlungen der Gesellschaft deutscher Naturforscher und Ärzte, 88(??):1047–1049, ???? 1924. CO-DEN VGDNAN. ISSN 0172-0651.

Sommerfeld:1931:RBA

[Som31] Arnold Sommerfeld. Rezension: Bohr, Atomtheorie und Naturbeschreibung. (German) [Review: Bohr, Atomic theory and the description of nature]. Deutsche Literaturzeitung, ??(??):2105–??, ???? 1931.

Sommerfeld:2013:WAB

[Som13a] Arnold Sommerfeld. Der weitere Ausbau des Bohr–Sommerfeldschen Atommodells (1916). (German) [The further expansion of Bohr– Sommerfeld’s atomic model (1916)]. In Die Bohr–Sommerfeldsche Atomtheorie: Sommerfelds Erweiterung des Bohrschen Atommodells 1915/16. (German) [The Bohr–Sommerfeld atomic theory: Sommerfeld’s expansion of Bohr’s Atommodell 1915/16] [SE13], pages 41–60. ISBN 3-642-35114-X (paperback), 3-642-35115-8 (e-book). LCCN QC173 .S52 2013.

Sommerfeld:2013:SRB

[Som13b] Arnold Sommerfeld. Sommerfelds Reaktion auf das Bohrsche Atommodell (1913–1914). (German) [Sommerfeld’s reaction to Bohr’s atomic model (1913–1914)]. In Die Bohr–Sommerfeldsche Atomtheorie: Sommerfelds Erweiterung des Bohrschen Atommodells 1915/16. (German) [The Bohr–Sommerfeld atomic theory: Sommerfeld’s expansion of Bohr’s Atommodell 1915/16] [SE13], pages 15–24. ISBN 3-642-35114-X (paperback), 3-642-35115-8 (e-book). LCCN QC173 .S52 2013.

Singh:1999:CVR

[SR99] Rajinder Singh and Falk Riess. C. V. Raman, M. N. Saha and the Nobel Prize for the year 1930. Indian Journal of the History of Science, 34(??):61–75, ???? 1999. CODEN IJHSA4. ISSN 0019-

5235. URL http://insa.nic.in/writereaddata/UpLoadedFiles/
IJHS/Vol34_1_5_RSingh.pdf.

Samuelsson:1998:NLI

[SS98] Bengt Samuelsson and Michael Sohlman, editors. Nobel Lectures: Including Presentation Speeches and Laureates’ Biographies, Physics: 1922– 1941. Nobel lectures, including presentation speeches and laureates’ biographies. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-3402-3. ISSN 0169-006X. x + 456 pp. LCCN ????

Svidzinsky:2005:BMM

[SSH05] Anatoly A. Svidzinsky, Marlan O. Scully, and Dudley R. Herschbach. Bohr’s 1913 molecular model revisited. Proceedings of the National Academy of Sciences of the United States of America, 102(34):11985– 11988, ???? 2005. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Svidzinsky:2014:BMMa

[SSH14a] Anatoly Svidzinsky, Marlan Scully, and Dudley Herschbach. Bohr’s molecular model, a century later. Physics Today, 67(1):33–39, January 2014. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See comments [Amu14, Mlo14, Gru14] and response [SSH14b].

Svidzinsky:2014:BMMb

[SSH14b] Anatoly Svidzinsky, Marlan Scully, and Dudley Herschbach. Bohr’s molecular model and the melding of classical and quantum mechanics. Physics Today, 67(8):10, August 2014. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). See [SSH14a].

Sudoplatov:1995:STM

[SSSS95] Pavel Sudoplatov, Anatolii Pavlovich Sudoplatov, Jerrold L. Schecter, and Leona Schecter. Special Tasks: the Memoirs of an Unwanted Witness, a Soviet Spymaster. Little, Brown and Co., Boston, MA, USA, updated edition, 1995. ISBN 0-316-82115-2. xxxi + 527 pp. LCCN JN6529.I6 S83 1995.

Stapp:1993:MMQ

[Sta93] Henry P. Stapp. Mind, Matter, and Quantum Mechanics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-56289-3 (New York), 3-540-56289-3 (Berlin). xii + 248 pp. LCCN QC174.12 .S8 1993.

Stapp:2004:MMQ

[Sta04] Henry P. Stapp. Mind, matter, and quantum mechanics. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2004. ISBN 3-540-40761-8. ISSN 1612-3018. xii + 297 pp. LCCN QC174.12 .S8 2004. URL http://

www.loc.gov/catdir/enhancements/fy0817/2003061209-d.html;
http://www.loc.gov/catdir/toc/fy043/2003061209.html.

Stachel:2009:BP

[Sta09a] John Stachel. Bohr and the photon. In Quantum Reality, Relativistic Causality, and Closing the Epistemic Circle: essays in honour of Abner Shimony, volume 73(II) of The University of Western Ontario Series in Philosophy of Science, pages 69–83. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-9106-0 (print), 1-4020-9107-9 (electronic).

LCCN QC174.12 .Q365 2009. URL http://www.loc.gov/catdir/
enhancements/fy1102/2008939924-d.htm;   http://www.loc.gov/
catdir/enhancements/fy1102/2008939924-t.htm;   http://www.
springerlink.com/content/v370g7880u258708/.

Stachel:2009:GC

[Sta09b] John J. Stachel. Going Critical: Volume One: The Challenge of Practice, volume 201 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-1308-6 (hardcover). 407 (est.) pp. LCCN ????

Stapp:2009:CP

[Sta09c] Henry Stapp. Complementarity principle. In Greenberger et al. [GHW09], pages 111–113. ISBN 3-540-70626-7, 3-540-70622-4. LCCN C174.12 .C66 2009.

Stapp:2009:EL

[Sta09d] Henry Stapp. Einstein locality. In Greenberger et al. [GHW09], pages 182–188. ISBN 3-540-70626-7, 3-540-70622-4. LCCN C174.12 .C66 2009.

Stapp:2009:MM

[Sta09e] Henry Stapp. Matrix mechanics. In Greenberger et al. [GHW09], pages 368–371. ISBN 3-540-70626-7, 3-540-70622-4. LCCN C174.12 .C66 2009.

Stapp:2009:SC

[Sta09f] Henry Stapp. Schrödinger’s cat. In Greenberger et al. [GHW09], pages 685–689. ISBN 3-540-70626-7, 3-540-70622-4. LCCN C174.12 .C66 2009.

Stapp:2009:MMQ

[Sta09g] Henry P. Stapp. Mind, Matter, and Quantum Mechanics. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2009. ISBN 3-540-89653-8 (hardcover), 3-540-89653-1, 3-540-89654-6 (e-book). ISSN 1612-3018. xv + 300 pp. LCCN QC174.12 .S8 2009. URL http://www.loc.gov/

catdir/enhancements/fy1402/2008942368-d.html;   http://www.
loc.gov/catdir/enhancements/fy1402/2008942368-t.html.

Stewart:1958:BRC

[Ste58] Albert B. Stewart. Book review: Complementarity: Atomic Physics and Human Knowledge, by Niels Bohr. The Antioch Review, 18(3):376–378,

Autumn 1958. URL http://www.jstor.org/stable/4610087.

Stephenson:1986:WJNa

[Ste86a] W. Stephenson. William James: Niels Bohr and complementarity I. Concepts. Psychological Record, 36(4):519–528, ???? 1986. ISSN 0033-2933.

Stephenson:1986:WJNb

[Ste86b] W. Stephenson. William James: Niels Bohr and complementarity II. Pragmatics of a thought. Psychological Record, 36(4):529–544, ???? 1986. ISSN 0033-2933.

Stephenson:1987:WJN

[Ste87] W. Stephenson. William James: Niels Bohr and complementarity III. Schrödinger’s cat. Psychological Record, 37(4):523–544, ???? 1987. ISSN 0033-2933.

Stephenson:1988:WJNa

[Ste88a] W. Stephenson. William James: Niels Bohr and complementarity IV. The significance of time. Psychological Record, 38(1):19–36, ???? 1988. ISSN 0033-2933.

Stephenson:1988:WJNb

[Ste88b] W. Stephenson. William James: Niels Bohr and complementarity V. Phenomenology of subjectivity. Psychological Record, 38(2):203–220, ???? 1988. ISSN 0033-2933.

Stent:1989:LLN

[Ste89] G. S. Stent. Light and life: Niels Bohr’s legacy to contemporary biology. Genome, 31(1):11–15, January 1989. ISSN 0831-2796.

Stenholm:2011:QRB

[Ste11] Stig Stenholm. The quest for reality: Bohr and Wittgenstein, two complementary views. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-960358-8 (hardcover), 0-19-172927-2 (ebook). viii + 222 pp. LCCN BD331 .S747 2011.

Stone:1997:CTE

[Sto97] Richard Stone. Chemistry: Transuranic element names finally final. Science, 277(5332):1601, September 12, 1997. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Strickland:2011:WSC

[Str11] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Stuewer:1979:NPR

[Stu79] Roger H. Stuewer, editor. Nuclear physics in retrospect: proceedings of a symposium on the 1930s. University of Minnesota Press, Minneapolis, MN, USA, 1979. ISBN 0-8166-0869-5. LCCN QC773 .S95 1977.

Stuewer:1984:BRN

[Stu84] Roger H. Stuewer. Book review: Niels Bohr; L. Rosenfeld; J. Rud Nielson; Ulrich Hoyer: Collected Works, Volumes 1–4. Isis, 75(1):234– 235, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232403.

Stuewer:1985:BNF

[Stu85a] Roger H. Stuewer. Bringing the news of fission to America. Physics Today, 38(10):48–56, October 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.org/

resource/1/phtoad/v38/i10/p48_s1.

Stuewer:1985:NBN

[Stu85b] Roger H. Stuewer. Niels Bohr and nuclear physics. In French and Kennedy [FK85], pages 197–220. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Stuewer:1992:BRB

[Stu92] Roger H. Stuewer. Book review: Niels Bohr’s Times, in Physics, Philosophy, and Polity, by Abraham Pais. Physics Today, 45(3):??, March 1992. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Stuewer:1994:OLD

[Stu94] Roger H. Stuewer. The origin of the liquid-drop model and the interpretation of nuclear fission. Perspectives on Science, 2(1):76–129, Spring 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic).

URL http://alsos.wlu.edu/information.aspx?id=1736. See errata

[Ano94a].

Stuewer:2013:ACM

[Stu13] Roger H. Stuewer. An act of creation: the Meitner–Frisch interpretation of nuclear fission. In Katzir et al. [KLR13], pages 231–245. ISBN 3-8442-

5134-0. LCCN QC173.98. URL http://www.edition-open-access.
de/proceedings/5/.

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Sudbery:2011:QEB

[Sud11] Tony Sudbery. Quantum: Einstein, Bohr and the Great Debate About the Nature of Reality, by Manjit Kumar, Scope: general interest. Level: general readership. Contemporary Physics, 52(3):251–254, 2011. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Suvorov:1985:PNB

[Suv85] S. G. Suvorov. On the publication of Niels Bohr’s “Open Letter to the United Nations”. Soviet Physics. Uspekhi, 28(10):935–936, October 31, 1985. CODEN SOPUAP. ISSN 0038-5670. URL http://stacks.iop.

org/0038-5670/28/i=10/a=A05.

Sexl:1985:NBK

[SvMS85] Roman U. Sexl, Karl von Meyenn, and Klaus Stolzenburg, editors. Niels Bohr 1885–1962: Der Kopenhagener Geist in der Physik. (German) [Niels Bohr 1885–1962: The Copenhagen Spirit in Physics]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1985. ISBN 3-528-08922-9. ix + 409 pp. LCCN ????

Sweet:2002:BLN

[Swe02] William Sweet. The Bohr letters: No more uncertainty. Bulletin of the Atomic Scientists, 58(3):20–27, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/3/20.full.pdf+html.

Schleich:1988:APS

[SWW88] W. Schleich, H. Walther, and J. A. Wheeler. Area in phase space as determiner of transition probability: Bohr–Sommerfeld bands, Wigner ripples, and Fresnel zones. Foundations of Physics, 18(10): 953–968, October 1988. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://link.springer.com/article/
10.1007/BF01909932.

Szuromi:2005:RBM

[Szu05] Phillip D. Szuromi. Reviving Bohr molecules. Science (New Series), 309(5740):1459, September 2, 2005. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-9203 (electronic).

Talbot:1979:BRF

[Tal79] Theodore Robert Talbot. Bohr and Rosenfeld’s Foundations for Quantum Electrodynamics. Ph.D dissertation, Columbia University, New York, NY, USA, February 1979. 265 pp.

Tanona:2002:CCE

[Tan02] Scott Daniel Tanona. From correspondence to complementarity: The emergence of Bohr’s Copenhagen interpretation of quantum mechanics. Ph.D. dissertation, Indiana University, Bloomington, IN, USA, 2002. 257 pp.

Tanona:2004:IFB

[Tan04a] Scott Tanona. Idealization and formalism in Bohr’s approach to quantum theory. Philosophy of Science, 71(5):683–695, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/425233.

Tanona:2004:UBR

[Tan04b] Scott Tanona. Uncertainty in Bohr’s response to the Heisenberg microscope. Studies in History and Philosophy of Modern Physics, 35(3): ??, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-

2502 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000371.

Tanona:2009:BRA

[Tan09] Scott Tanona. Book review: Arkady Plotnitsky, Reading Bohr: Physics and Philosophy (2006) Springer, Berlin (242 pp., US $189, Hardcover, ISBN: 978-1-4020-5253-8). Studies in History and Philosophy of Modern Physics, 40(1):90–91, January 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1355219808000567. See [Plo06b].

Teller:1969:NBI

[Tel69] Edward Teller. Niels Bohr and the idea of complementarity. In Great men of physics: the humanistic element in scientific work [SKST69], pages 76–97. LCCN ???? With an introduction by Marvin L. Chachere.

Teller:1981:PPB

[Tel80] Edward Teller. The projection postulate and Bohr’s interpretation of quantum mechanics. In Asquith and Giere [AG81], pages 201–223. ISBN 0-917586-16-6. LCCN ????

Teller:1985:FM

[Tel85] Edward Teller. A few memories. In French and Kennedy [FK85], pages 181–182. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Terletskii:1994:ODN

[Ter94] Yakov Terletskii. Operatsiia “dopros Nil’sa Bora” (Russian) [Operation ‘interrogation of Niels Bohr]. VIET: Voprosy Istorii Estestvoznaniia i Tekhniki, 2(??):21–44, ???? 1994. ISSN 0205-9606.

Terzis:2008:SRB

[Ter08] Andreas F. Terzis. A simple relativistic Bohr atom. European Journal of Physics, 29(4):735, 2008. CODEN EJPHD4. ISSN 0143-0807 (print),

1361-6404 (electronic). URL http://stacks.iop.org/0143-0807/29/
i=4/a=008.

Teuber:2002:HDN

[Teu02] Jan Teuber, editor. Højdepunkter i dansk naturvidenskab. (Danish) [Highlights in Danish natural science]. Gads Forlag, København, Danmark, 2002. ISBN 87-12-03847-4. 191 pp. LCCN Q127.D4 H65 2002.

EST:2000:CC

[The00] The Ensemble Studio Theatre, The Graduate Center of the City University of New York. Creating Copenhagen. 323 minute video, 2000.

Thomas:1926:MSE

[Tho26] L. H. (Llewellyn Hilleth) Thomas. The motion of the spinning electron. Nature, 117(2945):514, April 10, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Tibbetts:2013:HGS

[Tib13] Gary G. Tibbetts. How the Great Scientists Reasoned: the Scientific Method in Action. Elsevier insights. Elsevier, Amsterdam, The Netherlands, 2013. ISBN 0-12-398498-X (hardcover). viii + 148 pp. LCCN Q175 .T547 2013.

Thylwe:2013:BSQ

[TM13] Karl-Erik Thylwe and Patrick McCabe. Bohr–Sommerfeld quantization condition for Dirac states derived from an Ermakov-type invariant. Journal of Mathematical Physics, 54(5):052301, May 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Treiman:1999:OQ

[Tre99] Sam B. Treiman. The Odd Quantum. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-00926-0 (hardcover), 0-691-10300-3 (paperback), 1-4008-2309-9 (e-book). viii + 262 pp. LCCN QC 174.12.T73 1999.

Trimmer:1980:PSQ

[Tri80] John D. Trimmer. The present situation in quantum mechanics: a translation of Schrödinger’s “cat paradox” paper. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 124(5):323–338, October 10, 1980. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor. org/stable/986572. English translation of [Sch35c, Sch35a, Sch35b].

Tucker:1997:SWM

[Tuc97] Robert Edward Tucker. To speak when the mind’s eye fails: Niels Bohr, visualization, and the rhetorical situation of atomic science. Ph.D. dissertation, University of Southern California, Los Angeles, CA, USA, 1997. 251 pp.

Turner:1940:NF

[Tur40] Louis A. Turner. Nuclear fission. Reviews of Modern Physics, 12(1): 1–29, January 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/

10.1103/RevModPhys.12.1;   http://rmp.aps.org/abstract/RMP/

v12/i1/p1_1. This paper surveys the early work from 1934 to 1939 on nuclear fission, including more than 100 from 1939 alone following the news of the Hahn and Strassmann work that Niels Bohr brought to the USA in January 1939.

Turner:1946:AE

[Tur46a] Louis A. Turner. Atomic energy from U^{238}. Physical Review (2), 69(7– 8):366, April 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/PhysRev.

69.366. Publication voluntarily delayed six years because of war-time secrecy.

Turner:1946:AEU

[Tur46b] Louis A. Turner. Atomic energy from U^{238}. Physical Review (2), 69(7– 8):366, April 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/PhysRev.

69.366. Publication voluntarily delayed six years because of war-time secrecy.

Turchetti:2006:BRG

[Tur06] Simone Turchetti. Book review: Gerald Holton, Victory and Vexation in Science. Einstein, Bohr, Heisenberg and Others. Cambridge, Mass.: Harvard University Press, 2005. xiv + 230 pp., ill., ISBN 0-674-01519-3. Nuncius, 21(1):191–192, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00401.

Terhal:2003:QEM

[TWD03] Barbara M. Terhal, Michael M. Wolf, and Andrew C. Doherty. Quantum entanglement: a modern perspective. Physics Today, 54(4):46– 52, April 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://www.physicstoday.org/resource/1/

PHTOAD/v56/i4. This paper discusses the origin of the concept of quantum entanglement, and the famous Schrödinger’s cat experiment, in Schrödinger’s famous three papers [Sch35c, Sch35a, Sch35b], and the later work of John Bell [Bel64]. Those papers were a response to the famous EPR paper [EPR35] that first raised the paradox described by entanglement, and that caused much debate with Bohr, Schrödinger, and others. See also [Eve57, Whe57, Vil86].

Uhlenbeck:1925:EHU

[UG25] George E. Uhlenbeck and Samuel Goudsmit. Ersetzung der Hypothese vom unmechanischen Zwang durch eine Forderung bezüglich des inneren Verhaltens jedes einzelnen Elektrons. (German) [Replacement of the hypothesis of unmechanical force with a requirement for the inner behavior of each electron]. Die Naturwissenschaften, 13(47):953– 954, November 1925. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://link.springer.com/article/
10.1007/BF01558878;   http://www.springerlink.com/content/

u20q85248gp8166q/. See [?] for a computation in 1927 of the radius of a rotating charged particle with the spin of the electron.

Uhlenbeck:1926:SES

[UG26] G. E. Uhlenbeck and S. Goudsmit. Spinning electrons and the structure of spectra. Nature, 117(2938):264–265, February 20, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). The paper is followed by a half-column supportive letter from Niels Bohr. See [Tho26] for a relativistic correction to the Zeeman splitting of spectral lines.

Uhlenbeck:1984:SES

[UGB84] G. E. Uhlenbeck, S. Goudsmit, and Niels Bohr. Spinning electrons and the structure of spectra. In Klaus Stolzenburg, editor, The Emergence of Quantum Mechanics (Mainly 1924–1926), volume 5 of Niels Bohr Collected Works, chapter VI, pages 287–289. Elsevier, Amsterdam, The Netherlands, 1984. ISSN 1876-0503. URL http://www.sciencedirect.

com/science/article/pii/S1876050308705358.

Ullmann-Margalit:1992:SEB

[UM92] Edna Ullmann-Margalit. The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4, volume 146 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346. ix + 302 pp. LCCN Q174 .S354 1992. URL https://link.

springer.com/book/10.1007/978-94-011-2688-5.

Ulbricht:2012:LQC

[UPG^{+}12] Ronald Ulbricht, Joep J. H. Pijpers, Esther Groeneveld, Rolf Koole, Celso de Mello Donega, Daniel Vanmaekelbergh, Christophe Delerue, Guy Allan, and Mischa Bonn. Loosening quantum confinement: observation of real conductivity caused by hole polarons in semiconductor nanocrystals smaller than the Bohr radius. Nano Letters, 12(9):4937– 4942, September 12, 2012. CODEN NALEFD. ISSN 1530-6992.

Urtekin:2006:BMD

[Urt06] Kerim Urtekin. Bohr model and dimensional scaling analysis of atoms and molecules. Ph.D. dissertation, Texas A&M University, College Station, TX, USA, 2006. 133 pp.

Valdevit:2006:OFS

[Val06] Giampaolo Valdevit. Oppenheimer fra scienza e potere. Una storia americana. (Italian) [Oppenheimer between science and power. An American story]. Studi Storici, 47(1):115–142, January/March 2006. ISSN 0039-3037 (print), 2036-458X (electronic). URL http://www.

jstor.org/stable/20567340.

VanDeMark:2003:PKN

[Van03] Brian VanDeMark. Pandora’s keepers: nine men and the atomic bomb. Little, Brown and Co., Boston, MA, USA, 2003. ISBN 0-316-73833-6. xii + 399 + 16 pp. LCCN QC774.A2 V36 2003. URL http://www.loc.

gov/catdir/toc/fy045/2002043646.html.

vanCalmthout:2017:RQSc

[vCK17] Martijn van Calmthout and Leo P. Kouwenhoven. Real Quanta: Simplifying Quantum Physics for Einstein and Bohr. Dundurn Press Limited, Toronto, ON, Canada, 2017. ISBN 1-4597-4049-1 (paperback), 1-4597-4050-5 (PDF), 1-4597-4051-3 (epub). 192 pp. LCCN QC174.123 .C3513 2017.

vanDongen:2015:CHU

[vD15] Jeroen van Dongen. Communicating the Heisenberg uncertainty relations: Niels Bohr, complementarity and the Einstein–Rupp experiments. In Aaserud and Kragh [AK15], pages 310–346. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63 2015. URL http://www.ijqf.

org/wps/wp-content/uploads/2015/06/Bohr-Book.pdf.

vanderWaerden:1967:SQM

[vdW67] B. L. (Bartel Leendert) van der Waerden, editor. Sources of Quantum Mechanics. Classics of science. North-Holland Publishing Co., Amsterdam, The Netherlands, 1967. xi + 430 pp. LCCN QC174.12 S655.

vanderWaerden:1968:SQM

[vdW68] B. L. (Bartel Leendert) van der Waerden, editor. Sources of quantum mechanics, volume 5 of Classics of science. Dover, New York, NY, USA, 1968. ISBN 0-486-61881-1. xi + 430 pp. LCCN QC174.1 .W3 1968.

vanderWaerden:2007:SQM

[vdW07] B. L. (Bartel Leendert) van der Waerden, editor. Sources of quantum mechanics. Dover, New York, NY, USA, 2007. ISBN 0-486-45892-X (paperback). xi + 430 pp. LCCN QC174.12 .W34 2007. URL http://

www.loc.gov/catdir/enhancements/fy0702/2006050791-d.html.

vanFraassen:1976:SAN

[vFH76] B. C. van Fraassen and C. A. Hooker. A semantic analysis of Niels Bohr’s philosophy of quantum theory. In Harper and Hooker [HH76], pages 221–241. ISBN 90-277-0621-2, 94-010-1438-8 (e-book). ISSN 1566-

659X. LCCN B67. URL https://link.springer.com/chapter/10.
1007/978-94-010-1438-0_13.

Vizgin:1987:SNB

[VGEK87] Vladimir P. Vizgin, A. T. Grigor’yan, M. A. El’yashevich, and Ol’ga V. Kuznetsova. Symposium “Niels Bohr and science of the XX century”. Soviet Physics. Uspekhi, 30(3):286–??, 1987. CODEN SOP-

UAP. ISSN 0038-5670. URL http://stacks.iop.org/0038-5670/30/
i=3/a=A07.

vonHevesy:1923:BTR

[vH23] George von Hevesy. Borsche Theorie und Radioaktivität. (German) [Bohr theory and radioactivity]. Die Naturwissenschaften, 11(27):604– 605, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Vickers:2008:BTA

[Vic08] Peter J. Vickers. Bohr’s theory of the atom: Content, closure and con-

sistency, April 2008. URL http://philsci-archive.pitt.edu/4005/
1/Bohr_April_2008.pdf.

Villars:1986:PSC

[Vil86] C. N. Villars. The paradox of Schrödinger’s cat. Physics Education, 21(4):232–237, July 1986. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic).

Vieitez:2008:ORS

[VIR^{+}08] M. O. Vieitez, T. I. Ivanov, E. Reinhold, C. A. de Lange, and W. Ubachs. Observation of a Rydberg series in H + H^{−}: a heavy Bohr atom. Physical Review Letters, 101(16):163001, October 17, 2008. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

vonMeyenn:1993:WPW

[vM93] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band III: 1940–1949. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume III: 1940–1949, volume 11 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-54911-0. ISSN 0172-6315. lxiv + 1070 pp. LCCN QC16.P37 A34. URL http://www.springer.

com/physics/book/978-3-540-54911-6.

vonMeyenn:1996:WPW

[vM96] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil I: 1950– 1952. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume IV, Part I: 1950–1952, volume 14 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-59442-6. ISSN 0172-6315. xxxvii + 968 pp. LCCN QC16.P37 A34. URL http:/

/www.springerlink.com/content/978-3-540-78803-4.

vonMeyenn:1999:WPW

[vM99] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil II: 1953– 1954. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume IV, Part II: 1953–1954, volume 15 of Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-64312-5. ISSN 0172-6315. xxxv + 1100 pp. URL http://www.springer.com/

physics/book/978-3-540-64312-8.

vonMeyenn:2001:WPW

[vM01] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil III: 1955–1956 Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume IV, Part III: 1955–1956, volume 17 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-67591-4, 3-540-78805-0. ISSN 0172-6315. lxv + 994 pp. LCCN QC16.P37 A34.

vonMeyenn:2002:BSA

[vM02] Karl von Meyenn. Vom Bohr-Sommerfeldschen Atom zu Heisenbergs Hexeneinmaleins. Zum 100. Geburtstag von Werner Heisenberg. (German) [From Bohr–Sommerfeld’s atom to Heisenberg’s magic formula. The 100th birthday of Werner Heisenberg]. Naturwissenschaftliche Rundschau, 55(2):69–81, 2002. CODEN NARSAC. ISSN 0028-1050.

vonMeyenn:2005:WPW

[vM05] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil IV: 1957–1958 Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume IV, Part IV: A:1957, B:1958, volume 18 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-40296-9. ISSN 0172-6315. xl + 1585 pp. LCCN QC16.P37 A34.

URL http://www.springerlink.com/content/978-3-540-26832-1.

vonMeyenn:1985:WPW

[vMHW85] Karl von Meyenn, Armin Hermann, and Victor F. (Victor Frederick) Weisskopf, editors. Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band II, 1930–1939. Scientific Correspondence with Bohr, Einstein, Heisenberg a.o. Volume II: 1930–1939, volume 6 of Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-13609-6 (New York), 3-540-13609-6 (Berlin). xxxix + 783 pp. LCCN QC16.P37 A34. URL http://www.

springerlink.com/content/978-3-540-13609-5.

Volkenshtein:1988:CPB

[Vol88] M. V. Vol’kenshteĭn. Complementarity, physics, and biology. Soviet Physics. Uspekhi, 31(2):140, 1988. CODEN SOPUAP. ISSN 0038-5670.

URL http://stacks.iop.org/0038-5670/31/i=2/a=A03.

vonWeizsacker:1971:CI

[vW71] C. F. von Weizsäcker. The Copenhagen interpretation. In Bastin [Bas71], page ?? ISBN 0-521-07956-X. LCCN QC174.1

.Q37. URL http://www.loc.gov/catdir/enhancements/fy0907/
77127237-d.html;   http://www.loc.gov/catdir/enhancements/
fy0907/77127237-t.html.

vonWeizsacker:1985:R

[vW85] C. F. von Weizsäcker. A reminiscence from 1932. In French and Kennedy [FK85], pages 183–190. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Vonborzeszkowski:1986:PMR

[VW86] H. H. Vonborzeszkowski and R. Wahsner. Physical means of recognition and physical reality — a discussion between Niels Bohr and Albert Einstein on the status of quantum mechanics. Deutsche Zeitschrift für Philosophie, 34(12):1098–1106, ???? 1986. ISSN 0012-1045.

vonWeizsacker:1988:BGC

[vW88] Carl Friedrich von Weizsäcker. Bewußtseinswandel. (German) [Change of consciousness]. C. Hanser, Munich, West Germany, second edition, 1988. ISBN 3-446-14649-0. 476 pp. LCCN ????

vonWeizsacker:1991:BGC

[vW91] Carl Friedrich von Weizsäcker. Bewußtseinswandel. (German) [Change of consciousness]. Deutscher Taschenbuch Verlag, Munich, West Germany, 1991. ISBN 3-423-11388-X. 476 pp. LCCN ????

vonWeizsacker:2002:LMH

[vW02a] C. F. von Weizsäcker. Letters to Martin Heidegger, Niels Bohr, Hoimar von Ditfurth, Edward Teller, Joachim Illies, Klaus Scholder, Jürgen Habermas, Jürgen Kuczyński, Hans-Georg Gadamer. Sinn und Form, 54(2):161–176, ???? 2002. ISSN 0037-5756.

vonWeizsacker:2002:GPA

[vW02b] Carl Friedrich von Weizsäcker. Grosse Physiker: von Aristoteles bis Werner Heisenberg. (German) [Great Physicists: from Aristole to Werner Heisenberg], volume 33078 of dtv. Deutscher Taschenbuch-Verlag, München, Germany, unabridged edition, 2002. ISBN 3-423-33078-3. 376 pp. LCCN QC6. EUR-D 12.50.

Walker:2008:BRW

[Wal08] Mark Walker. Book reviews: When physics, politics, and consciousness collide: Faust in Copenhagen: A Struggle for the Soul of Physics, by Gino Segrè. The Mental Aftermath: The Mentality of German Physicists, 1945–1949, by Klaus Hentschel. Quantum Computer Science: An Introduction, by N. David Mermin. Quantum Information: An Overview, by Gregg Jaeger. Physics Today, 61(3):53–54, 2008. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v61/i3/p53/s1.

Walker:2019:CR

[Wal19] Mark Walker. Copenhagen revisited. In Björkman et al. [BLW19], pages 230–247. ISBN 0-8153-9474-8 (hardcover), 1-351-18508-X (mobi), 1-351-18509-8 (e-pub), 1-351-18510-1 (Adobe), 1-351-18511-X (e-book). LCCN DD256.6 .H58 2019.

Wasserman:1981:BKS

[Was81] Neil H. Wasserman. The Bohr–Kramers–Slater Paper and the Development of the Quantum Theory of Radiation in the Work of Niels Bohr. Ph.D dissertation, Harvard University, Cambridge, MA, USA, 1981.

Wasserman:1992:BRB

[Was92] Neil Wasserman. Book review: Redirecting Science: Niels Bohr, Philanthropy, and the Rise of Nuclear Physics by Finn Aaserud. Isis, 83 (2):343–344, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234561.

Breitbarth:1990:BRH

[wB90] Friedrich wilhelm Breitbarth. Book review: Harmony and Unity — The Life of Niels Bohr von N. Blaedel; Madison, Science Tech Publishers, Berlin, Heidelberg, New York, Tokyo, London, Paris, Springer-Verlag, 1988; XII, 323 Seiten mit 154 Bildern; Format 15,5 cm × 23 cm, Pappband zelloph. DM 98,-; ISBN 3-540-19334-0. Zeitschrift f”ur Chemie (Stuttgart, Germany), 30(3):114, March 1990. CODEN ZE-

CEAL. ISSN 0044-2402. URL http://onlinelibrary.wiley.com/
doi/10.1002/zfch.19900300331/abstract.

Wallace:2006:BRL

[WB06] Dorothy I. Wallace and Joseph BelBruno. Bell that rings light: a primer in quantum mechanics and chemical bonding, volume 1 of Mathematics across the curriculum, 1793-3145. World Scientific, Singapore, 2006. ISBN 981-256-705-4. xiv + 137 pp. LCCN QD462.6.M39 W3 2006.

URL http://www.loc.gov/catdir/toc/fy0705/2006287845.html;
http://www.worldscibooks.com/chemistry/6061.html.

Wilczek:2006:FRM

[WD06] Frank Wilczek and Betsy Devine. Fantastic realities: 49 mind journeys and a trip to Stockholm. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2006. ISBN 981-256-649-X (hardcover), 981-256-655-4 (paperback), 981-277-430-0 (e-book). ix + 522 pp. LCCN QC75 .W55 2006.

Weatherburn:1988:TBA

[Wea88] H. Weatherburn. Teaching Bohr atomic theory. Radiography, 54(616): 157, ???? 1988. ISSN 0033-8281.

Weisskopf:1935:KPW

[Wei35] Victor Weisskopf. Komplementäre Philosophie des Witzes. (German) [Complementary philosophy of wit]. Journal of Jocular Physics, I:24–25, 1935.

Weiss:1939:BRB

[Wei39] Paul Weiss. Book review: International Encyclopedia of Unified Science: Vol. I, Foundations of the Unity of Science: No. 1, Encyclopedia and Unified Science, by Otto Neurath, Niels Bohr, John Dewey, Bertrand Russell, Rudolph Carnap, and Charles W. Morris. Ethics, 49

(4):498–500, July 1939. URL http://www.jstor.org/stable/2988923.

Weisskopf:1963:NBM

[Wei63] Victor F. Weisskopf. Niels Bohr: a memorial tribute. Physics Today, 16 (10):58–64, October 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Weisskopf:1964:NBO

[Wei64] Viktor F. Weisskopf. Niels Bohr og internationalt videnskabeligt samarbejde. (Danish) [Niels Bohr and international scientific collaboration]. In Bohr and Rozental [BR64], pages 254–258. LCCN QC16.B63 N5. English translation in [Roz67].

Weiner:1968:IGG

[Wei68] Charles Weiner. Interview with George Gamow at Professor Gamow’s home in Boulder, Colorado April 25, 1968. Niels Bohr Library & Archives, American Institute of Physics, College Park, MD, USA., April

25, 1968. URL http://www.aip.org/history/ohilist/4325.html.

Weisskopf:1970:TSS

[Wei70] Victor F. Weisskopf. Three steps in the structure of matter. Physics Today, 23(8):17–24, August 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.org/

resource/1/phtoad/v23/i8/p17_s1.

Weiner:1977:HTC

[Wei77] Charles Weiner, editor. History of twentieth century physics: Storia della fisica del XX secolo, Proceedings of the International School of Physics “Enrico Fermi” = Rendiconti della Scuola internazionale di fisica “Enrico Fermi”, course 57, July 31–August 12, 1972. Academic Press, New York, USA, 1977. ISBN 0-12-368857-4. LCCN QC7 .V37 1977.

Weisskopf:1985:ENB

[Wei85a] Victor F. Weisskopf. Editorial: Niels Bohr, the quantum, and the world. Physics Today, 38(10):191–192, October 1985. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Weisskopf:1985:NBQ

[Wei85b] Victor F. Weisskopf. Niels Bohr, the quantum, and the world. In French and Kennedy [FK85], pages 19–29. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Weisskopf:1985:PMP

[Wei85c] Victor F. Weisskopf. Personal memories of Pauli. Physics Today, 38(12):36–41, December 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/link/

phtoad/v38/i12/p36/s1.

Weisskopf:1991:LKB

[Wei91] Viktor F. Weisskopf. The lesser known Bohr. Science (New Series), 251 (4994):684–685, February 8, 1991. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Weiner:1994:LFD

[Wei94] Nella Fermi Weiner. Letter: Fermi: Didn’t even tell his wife. Bulletin of the Atomic Scientists, 50(4):3, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberger:2014:NBD

[Wei14] P. Weinberger. Niels Bohr and the dawn of quantum theory. Philosophical Magazine, 94(27):3072–3087, August 2014. CODEN PHMAA4. ISSN 0031-8086.

Wereide:1923:GPR

[Wer23] Th. Wereide. The General Principle of Relativity applied to the Rutherford–Bohr atom-model. Physical Review, 21(4):391–396, April 1923. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.21.391.

Werner:2002:CDR

[Wer02] Emmy E. Werner. A Conspiracy of Decency: the Rescue of the Danish Jews During World War II. Westview Press, Boulder, CO, USA, 2002. ISBN 0-8133-3906-5. 212 pp. LCCN DS135.D4 W47 2002. URL http://

www.loc.gov/catdir/enhancements/fy0831/2003276826-b.html;
http://www.loc.gov/catdir/enhancements/fy0831/2003276826-d.
html.

Wessels:1985:BRI

[Wes85] Linda Wessels. Book review: Issues in quantum physics: The creation of quantum mechanics and the Bohr–Pauli dialogue, John Hendry, Reidel, Boston, 1984. Science (New Series), 229(4717):963–964, September 6, 1985. CODEN SCNCAD. ISSN 0036-8075 (print), 1095-9203 (electronic).

Wheeler:1957:AER

[Whe57] John A. Wheeler. Assessment of Everett’s “relative state” formulation of quantum theory. Reviews of Modern Physics, 29(3):463– 465, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.29.463;   http://rmp.aps.org/abstract/RMP/
v29/i3/p463_1. See [Eve57].

Wheeler:1963:NBN

[Whe63] John Archibald Wheeler. Niels Bohr and nuclear physics. Physics Today, 16(10):36–45, October 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wheaton:1981:BRB

[Whe81] Bruce R. Wheaton. Book review: The Early Years: The Niels Bohr Institute, 1921–1930 by Peter Robertson. Isis, 72(2):329–330, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231033.

Wheeler:1985:PC

[Whe85a] John A. Wheeler. Physics in Copenhagen in 1934 and 1935. In French and Kennedy [FK85], pages 221–226. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Wheeler:1985:NBM

[Whe85b] John Archibald Wheeler. Niels Bohr, the man. Physics Today, 38 (10):66–72, October 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wheeler:1988:WSS

[Whe88] John Archibald Wheeler. World as system self-synthesized by quantum networking. IBM Journal of Research and Development, 32(1):4–15, January 1988. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Wheeler:1992:HU

[Whe92] John Archibald Wheeler. At Home in the Universe, volume 9 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1992. ISBN 0-88318-862-7. ix + 371 pp. LCCN Q158.5 .W44 1992.

Wheeler:2009:MF

[Whe09] John A. Wheeler. Mechanism of fission. Physics Today, 62(4):35– 38, April 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Whitaker:1996:EBQ

[Whi96] Andrew Whitaker. Einstein, Bohr, and the quantum dilemma. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-48220-8 (hardcover), 0-521-48428-6 (paperback). xvii + 349 pp. LCCN

QC174.12 .W48 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http://www.cambridge.org/uk/catalogue/
catalogue.asp?isbn=0521484286;   http://www.loc.gov/catdir/
description/cam027/95018270.html;   http://www.loc.gov/catdir/
toc/cam021/95018270.html.

Whitaker:2004:EPB

[Whi04] M. A. B. Whitaker. The EPR paper and Bohr’s response: a reassessment. Foundations of Physics, 34(9):1305–1340, September 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://www.springerlink.com/content/j032495h61374575/.

Whitaker:2006:EBQ

[Whi06] Andrew Whitaker. Einstein, Bohr and the Quantum Dilemma: From Quantum Theory to Quantum Information. Cambridge University Press, Cambridge, UK, second edition, 2006. ISBN 0-521-67102-7. xviii + 461 pp. LCCN QC174.12 .W48 2006. UK£27.99. URL http://www.

cambridge.org/uk/catalogue/catalogue.asp?isbn=0521671027.

Wilson:1970:CP

[Wil70] Robert Rathbun Wilson. The conscience of a physicist. Bulletin of the Atomic Scientists, 26(6):30–34, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1985:NBY

[Wil85] Robert R. Wilson. Niels Bohr and the young scientists. Bulletin of the Atomic Scientists, 41(7):23–26, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilczek:1992:WDB

[Wil92] Frank Wilczek. What did Bohr do? Book review: Niels Bohr’s Times. Science (New Series), 255(5042):345–347, January 17, 1992. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Winge:2009:SVS

[WS09] Mette Winge and Claus Seidel. Da de store var små — 24 historier fra berømte danskeres barndom. (Danish) [When the great were little — 24 stories from the childhood of famous Danes]. Haase, København, Danmark, 2009. ISBN 87-559-1238-9. 158 pp. LCCN ???? DKR 249,00.

Wyden:1984:DOB

[Wyd84] Peter Wyden. Day One: Before Hiroshima and After. Simon and Schuster, New York, NY, USA, 1984. ISBN 0-671-46142-7 (hardcover). xxxii + 412 pp. LCCN D767.25.H6 W93 1984.

Wyker:2008:RBA

[Wyk08] Brendan Wyker. Realization of the Bohr atom. M.S. dissertation, Rice University, Houston, TX, USA, 2008. 67 pp.

Wheeler:1983:QTM

[WZ83] John Archibald Wheeler and Wojciech Hubert Zurek, editors. Quantum theory and measurement. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-08315-0, 0-691-08316-9 (paperback). xxviii + 811 pp. LCCN QC174.125 .Q38 1983. US$60.00, US$19.50.

Yagi:1964:NSA

[Yag64] Eri Yagi. On Nagaoka’s Saturnian atomic model. Japanese Studies in the History of Science, 3(??):29–47, ???? 1964. CODEN JSHIAE. ISSN 0090-0176.

Yagi:1972:DNS

[Yag72] Eri Yagi. The development of Nagaoka’s Saturnian atomic model II (1904–05). Japanese Studies in the History of Science, 11(??):73–89, ???? 1972. CODEN JSHIAE. ISSN 0090-0176.

Yam:2002:NSN

[Yam02] Philip Yam. News scan: In no uncertain terms [letter of Niels Bohr to Werner Heisenberg]. Scientific American, 286(4):30, April 2002. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/2002/0402issue/0402numbers.html.

York:1975:SAT

[Yor75] Herbert F. York. Sounders of the alarm: Thirty years of attempts to bring the world to grips with the perpetual menace to human security. Bulletin of the Atomic Scientists, 31(10):43–45, December 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yoshida:2009:CNB

[YRBD09] S. Yoshida, C. O. Reinhold, J. Burgdorfer, and F. B. Dunning. Comment on “Nondispersing Bohr wave packets”. Physical Review Letters, 103(14):149302–149301, October 2, 2009. CODEN PRLTAO. ISSN 1079-7114.

Zandonella:2002:UHB

[Zan02] Catherine Zandonella. Uncertainty over Heisenberg’s bomb making ends. New Scientist, ??(??):??, February 7, 2002. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

URL http://www.newscientist.com/article/dn1897-uncertainty-
over-heisenbergs-bomb-making-ends.html.

Zimmer:2011:CPB

[ZDSF11] Karl Günter Zimmer, Max Delbrück, Phillip R. Sloan, and D. Brandon Fogel, editors. Creating a physical biology: the Three-Man Paper and early molecular biology. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-76782-5 (hardcover), 0-226-76783-3 (paperback). 319 pp. LCCN QH506 .C73 2011.

Zinkernagel:2001:PMQ

[Zin01] H. Zinkernagel. Philosophical motivations for quantum gravity – another ‘Bohr/Einstein’ debate? In ????, editor, Proceedings of the III Conference of the Spanish Society for Analytical Philosophy, pages 113– 120. Granada University Press, Granada, Spain, 2001. ISBN ???? LCCN ????

Zinkernagel:2015:WLQ

[Zin15] Henrik Zinkernagel. Are we living in a quantum world? Bohr and quantum fundamentalism. In Aaserud and Kragh [AK15], pages 419– 434. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/2015/
06/Bohr-Book.pdf.

Zinkernagel:2016:NBW

[Zin16] Henrik Zinkernagel. Niels Bohr on the wave function and the classical/quantum divide. Studies in History and Philosophy of Modern Physics, 53(??):9–19, February 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1355219815000969.

Zinkernagel:2019:JFA

[Zin19] Henrik Zinkernagel. Book review: Jan Faye & Henry J. Folse (Eds.): Niels Bohr and the Philosophy of Physics: Twenty-First-Century Perspectives. Journal for General Philosophy of Science = Zeitschrift für allgemeine Wissenschaftstheorie, 50(2):317–322, June 2019. CODEN REFERENCES 305

JGPSE4. ISSN 0925-4560 (print), 1572-8587 (electronic). URL http:/

/link.springer.com/article/10.1007/s10838-019-09448-8.